### v6

# A faster algorithm for Vertex Cover parameterized by solution size

Abstract
We describe a new algorithm for vertex cover with runtime $O^{*}(1.25284^{k})$, where $k$ is the size of the desired solution and $O^{*}$ hides polynomial factors in the input size. This improves over the previous runtime of $O^{*}(1.2738^{k})$ due to Chen, Kanj, & Xia (2010) standing for more than a decade. The key to our algorithm is to use a measure which simultaneously tracks $k$ as well as the optimal value $\operatorname{\lambda}$ of the vertex cover LP relaxation. This allows us to make use of prior algorithms for Maximum Independent Set in bounded-degree graphs and Above-Guarantee Vertex Cover.
The main step in the algorithm is to branch on high-degree vertices, while ensuring that both $k$ and $\mu=k-\operatorname{\lambda}$ are decreased at each step. There can be local obstructions in the graph that prevent $\mu$ from decreasing in this process; we develop a number of novel branching steps to handle these situations.
1 Introduction
For an undirected graph $G=(V,E)$, a subset $S\subseteq V$ is called a vertex cover if every edge has at least one endpoint in $S$. It is closely related to an independent set, since if $S$ is an inclusion-wise minimal vertex cover then $V-S$ is an inclusion-wise maximal independent set, and vice-versa. Finding the size of the smallest vertex cover is a classic NP-complete problem [GJ79]. In particular, $G$ has a vertex cover of size at most $k$ if and only if it has an independent set of size at least $n-k$.
There is natural LP formulation for vertex cover which we denote by $\operatorname{LPVC}(G)$:
minimize $\sum_{v\in V}\theta(v)$
subject to $\theta(u)+\theta(v)\geq 1$ for all edges $e=(u,v)\in E$
$\theta(v)\in[0,1]$ for all vertices $v\in V$
The optimal solution to $\operatorname{LPVC}(G)$, denoted $\lambda(G)$ or just $\lambda$ if $G$ is clear from context, is a lower bound on the size of a minimum vertex cover of $G$. We also define $\mu(G)=k-\lambda(G)$, i.e. the gap between solution size and LP lower bound. This linear program has remarkable properties which have been exploited for a variety of algorithms [NT74, NT75, ${}^{+}$14]. For instance, an optimum basic solution is half-integral and can be found efficiently by a network flow computation.
[…]
1.1 Outline of our results
The main result of this paper is an improved algorithm for vertex cover parameterized by $k$.
Theorem 1.
There is an algorithm for VertexCover with runtime $O^{*}(1.25284^{k})$. Moreover, depending on the maximum vertex degree of the graph $G$, better bounds can be shown:
If $\operatorname{maxdeg}(G)\leq 3$, we get runtime $O^{*}(1.14416^{k})$.
If $\operatorname{maxdeg}(G)\leq 4$, we get runtime $O^{*}(1.21131^{k})$.
If $\operatorname{maxdeg}(G)\leq 5$, we get runtime $O^{*}(1.24394^{k})$.
If $\operatorname{maxdeg}(G)\leq 6$, we get runtime $O^{*}(1.25214^{k})$.
All of these algorithms use polynomial space and are deterministic.
The FPT algorithms for vertex cover, including our new algorithm, are built out of recursively branching on high-degree vertices or other structures, while tracking some “measure” of the graph. Our main new idea is to use a measure which is a piecewise-linear function of $k$ and $\mu$. To illustrate, consider branching on whether some degree-$r$ vertex $u$ is in the cover: we recurse on the subproblem $G_{1}=G-u$ where $k$ is reduced by one and on the subproblem $G_{0}=G-u-N(u)$ where $k$ is reduced by $r$. We must also show that $\mu$ is significantly reduced in the two subproblems.
Suppose that, wishfully speaking, $\vec{\tfrac{1}{2}}$ remains the optimal solution to $\operatorname{LPVC}(G_{0})$ and $\operatorname{LPVC}(G_{1})$. This is what should happen in a “generic” graph. Then $\operatorname{\lambda}(G_{0})=|V(G_{0})|/2=(n-r-1)/2$ and $\operatorname{\lambda}(G_{1})=|V(G_{1})|/2=(n-1)/2$. Since $k_{0}=k-r$ and $k_{1}=k-1$, this implies that $\mu$ is indeed significantly decreased:
$\mu(G_{0})=\mu(G)-(r-1)/2,\qquad\mu(G_{1})=\mu(G)-1/2$
But suppose, on the other hand, that $\vec{\tfrac{1}{2}}$ is not the optimal solution to $\operatorname{LPVC}(G_{0})$ or $\operatorname{LPVC}(G_{1})$. For a concrete example, suppose the neighborhood of $x$ in $G$ is a subset of that of $u$, so $G_{0}$ has an isolated vertex $x$ and an optimal solution $\theta^{*}$ to $\operatorname{LPVC}(G_{0})$ would set $\theta^{*}(x)=0$. In this situation, we develop an alternate branching rule for $G$: rather than branching on $u$ itself, we branch on the two subproblems where (i) both $u,x$ are in the cover and (ii) $x$ is not in the cover. See Figure 1 for an illustration.
Figure 1: In the subgraph $G-N[u]$, the vertex $x$ becomes isolated
We emphasize that this is just one example of our branching analysis. We need to handle more general situations where $\vec{\tfrac{1}{2}}$ is not the optimal solution to $\operatorname{LPVC}(G_{0})$ or $\operatorname{LPVC}(G_{1})$, which can require more complex branching rules. We emphasize that these branching rules are more powerful than simply using conventional branching rules along with the preprocessing rules; the latter may not be able to ensure a good reduction in $\mu$.
In Section 6, we use these ideas for a relatively simple algorithm with runtime $O^{*}(1.2575^{k})$, along with algorithms for bounded-degree graphs. This is already much better than [CKX10]. To improve further, and achieve Theorem 1, we need an additional idea: instead of branching on an arbitrary high-degree vertex, we branch on a targeted vertex to create additional simplifications in the graph. This is much more complex. For example, when we are branching on degree-4 vertices, it is favorable to choose one that has a degree-$3$ neighbor. This will create a degree-$2$ vertex in subproblem $\langle G-u,k-1\rangle$, leading to further simplifications of the graph. See Figure 2.
Figure 2: In the subgraph $G-u$, the vertex $v$ has degree two.
1.2 Related algorithms
Many different kinds of algorithms have been developed for the Vertex Cover problem. We cannot summarize this fully here; we describe some of the most relevant works for our own paper.
As we have mentioned, there is a long line of research into FPT algorithms parametrized by the solution size $k$. Currently, the fastest such algorithm has $O^{*}(1.2738^{k})$ runtime [CKX10]. In addition, [Tsu23] describes algorithms in graphs of maximum degree $3$ and $4$ with runtime respectively $O^{*}(1.1558^{k})$ and $O^{*}(1.2403^{k})$ respectively.
An important variant is Above-Guarantee Vertex Cover (AGVC), where the parameter is the difference between $k$ and various lower bounds on vertex cover size [MR99, MRS09, ${}^{+}$14, GP16, KKK22]. Of these, the parameter $\mu(G)=k-\operatorname{\lambda}(G)$ plays a particularly important role in this paper. We quote the following main result:
Theorem 2 ([${}^{+}$14]).
Vertex cover can be solved in time $O^{*}(2.3146^{\mu})$.
Another natural choice is to measure runtime in terms of the graph size $n$. In this setting, the problem is more commonly referred to as Maximum Independent Set (MaxIS). Xiao and Nagamochi have developed a number of algorithms in this vein. These algorithms, and in particular their performance on bounded-degree graphs, will also play a crucial role in our analysis. We refer to the algorithm targeted for graphs of maximum degree $\Delta$ by the MaxIS-$\Delta$ algorithm.
Theorem 3.
MaxIS-3 can be solved with runtime $O^{*}(1.083506^{n})$ by [XN13]. [^1] This runtime is not claimed directly in [XN13], see [XN17a] instead.
MaxIS-4 can be solved with runtime $O^{*}(1.137595^{n})$ by [XN17a].
MaxIS-5 can be solved with runtime $O^{*}(1.17366^{n})$ by [XN16]. [^2] This runtime is not claimed directly in [XN16], see [XN17b].
MaxIS-6 can be solved with runtime $O^{*}(1.18922^{n})$ by [XN17b].
MaxIS-7 can be solved with runtime $O^{*}(1.19698^{n})$ by [XN17b].
MaxIS in graphs of arbitrary degree can be solved with runtime $O^{*}(1.19951^{n})$ by [XN17b].
1.3 Notation
We consider throughout a simple, undirected, unweighted graph $G=(V,E)$, and we write $n=|V|$. We write an ordered pair $\langle G,k\rangle$ for a graph $G$ where we need to decide if there is a vertex cover of size at most $k$, and we say it is feasible if such a vertex cover exists. We say $C$ is a good cover if it is a vertex cover of size at most $k$.
For a subset $S$ of $V$, the subgraphs of $G$ induced by $S$ and $V\setminus S$ are denoted by $G[S]$ and $G-S$, respectively. We write $u\sim v$ if $(u,v)\in E$ and $u\nsim v$ otherwise. For vertex sets $X,Y$, we write $X\sim Y$ if there exists $x\in X,y\in Y$ with $x\sim y$.
For a vertex $u$, the neighborhood $N_{G}(u)$ is the set $\{u\in V(G):u\sim v\}$ and the closed neighborhood $N_{G}[u]$ is the set $N_{G}(u)\cup\{u\}$. Extending this notation, for a vertex set $S\subseteq V(G)$, we write $N_{G}(S)=\bigl{(}\bigcup_{v\in S}N_{G}(v)\bigr{)}\setminus S$ and $N_{G}[S]=N_{G}(S)\cup S=\bigcup_{v\in S}N_{G}[v]$. For readability, we sometimes write $N(x,y)$ or $N[x,y]$ as shorthand for $N(\{x,y\})$ or $N[\{x,y\}]$.
The degree of vertex $v$, denoted by $\deg_{G}(v)$, is the size of $N_{G}(v)$. We call a vertex of degree $i$ an $i$-vertex; for a vertex $v$, we refer to a neighbor $u$ which has degree $j$ as a $j$-neighbor of $v$. An isolated vertex is a 0-vertex. We say a vertex is subquartic if it has degree in $\{1,2,3,4\}$ and subcubic if it has degree in $\{1,2,3\}$, and subquadratic if it has degree in $\{1,2\}$. (We do not count isolated vertices).
We write $\operatorname{maxdeg}(G)$ and $\operatorname{mindeg}(G)$ for the minimum and maximum vertex degrees in $G$, respectively. The graph is $r$-regular if $\operatorname{maxdeg}(G)=\operatorname{mindeg}(G)=r$. We write $V_{i}$ for the set of degree-$i$ vertices and $n_{i}=|V_{i}|$.
We say that a pair of vertices $u,v$ share neighbor $y$ if $y\in N(u)\cap N(v)$. We denote by $\operatorname{codeg}(u,v)=|N(u)\cap N(v)|$ the number of vertices shared by $u,v$.
An $\ell$-cycle denotes a set of $\ell$ vertices $x_{1},\dots,x_{\ell}$ with a path $x_{1},x_{2},\dots,x_{\ell},x_{1}$.
2 Preliminaries
We review some basic facts about branching and preprocessing rules for vertex cover. Much of this material is standard, see e.g. [${}^{+}$14]. We include proofs in Appendix A for completeness.
Our algorithm will heavily use the LP and its properties. This LP is closely related to properties of independent sets. For brevity, we refer to these as indsets, i.e. sets $X\subseteq V$ where no two vertices in $X$ are adjacent. For an indset $I$ in $G$, we define the surplus by $\operatorname{surp}_{G}(I)=|N_{G}(I)|-|I|$. If $G$ is clear from context, we write just $\operatorname{surp}(I)$ or $N(I)$. We define $\operatorname{minsurp}(G)$ to be the minimum value of $\operatorname{surp}_{G}(I)$ over all non-empty indsets $I$ in $G$, and a min-set to be any non-empty indset $I$ achieving this minimum, i.e. $\operatorname{surp}_{G}(I)=\operatorname{minsurp}(G)$.
Since this comes up in a number of places, we define $\operatorname{minsurp}^{-}(G)=\min\{0,\operatorname{minsurp}(G)\}$.
Proposition 4.
$\lambda(G)=\tfrac{1}{2}(|V|+\operatorname{minsurp}^{-}(G)).$ Moreover, in a basic solution to $\operatorname{LPVC}(G)$, the set $I$ of vertices with value zero forms an indset (possibly empty) with $\operatorname{surp}_{G}(I)=\operatorname{minsurp}^{-}(G)$.
We define a critical-set to be a non-empty indset $I$ with $\operatorname{surp}_{G}(J)\geq\operatorname{surp}_{G}(I)$ for all non-empty subsets $J\subseteq I$. Clearly, any min-set or any singleton set is a critical-set.
Lemma 5.
For any critical-set $I$, there is a good cover $C$ with either $I\subseteq C$ or $I\cap C=\emptyset$.
If $\operatorname{surp}(I)\leq 0$, there is a good cover $C$ with $I\cap C=\emptyset$.
If $\operatorname{surp}(I)=1$, there is a good cover $C$ with either $N[I]\cap C=I$ or $N[I]\cap C=N(I)$.
Proposition 6.
For any non-empty indset $I$ of $G$, there holds $\operatorname{minsurp}(G)\leq\operatorname{minsurp}^{-}(G-N[I])+\operatorname{surp}_{G}(I)$. The equality $\operatorname{minsurp}(G)=\operatorname{minsurp}^{-}(G-N[I])+\operatorname{surp}_{G}(I)$ holds if and only if $I$ is contained in a min-set of $G$.
Proposition 7.
The value $\operatorname{minsurp}(G)$ can be determined in polynomial time. Moreover, for any indset $X$ (possibly $X=\emptyset$), we can efficiently find a min-set of $G$ containing $X$ (if any such exists).
2.1 Preprocessing and branching rules
Our algorithm uses three major preprocessing rules. For this, we need to define the following graph structure: a funnel is a vertex $u$ with a neighbor $x$ such that $G[N(u)\setminus\{x\}]$ forms a clique. Here, we call $x$ the out-neighbor of $u$. For example, if a degree-3 vertex $u$ has a triangle with neighbors $t_{1},t_{2}$, and one other vertex $x$, then $u$ is a funnel with out-neighbor $x$; since this comes up so often, we refer to this as a 3-triangle. A well-known result is that, for a funnel $u$ with out-neighbor $x$, there exists a good cover $C$ with either $u\notin C$ or $x\notin C$.
Preprocessing Rule 1 (P1): Given a critical-set $I$ with $\operatorname{surp}_{G}(I)\leq 0$, form graph $G^{\prime}$ with $k^{\prime}=k-|N(I)|$ by deleting $N[I]$.
Preprocessing Rule 2 (P2): Given a critical-set $I$ with $\operatorname{surp}_{G}(I)=1$, form graph $G^{\prime}$ with $k^{\prime}=k-|I|$ by deleting $N[I]$ and adding a new vertex $y$ adjacent to $N(N(I))$.
Preprocessing Rule 3 (P3): Given a funnel $u$ with out-neighbor $x$, form graph $G^{\prime}$ with $k^{\prime}=k-1-\operatorname{codeg}(u,x)$ by removing vertices $N[u]\cap N[x]$ and adding edges between all vertices in $N(u)\setminus N[x]$ to all vertices in $N(x)\setminus N[u]$.
See Figure 3 for an illustration of rule (P3) applied to a funnel.
Figure 3: A funnel (here, a 3-triangle) $u$ before (left) and after (right) applying P3
These rules can be applied to any eligible sets in any order. Note that if a vertex $v$ has degree zero or one, then applying (P1) to $\{v\}$ removes $v$ and its neighbor (if any). If a vertex $v$ has degree two, then applying (P2) to $\{v\}$ removes $v$ and contracts its neighbors $x,y$ into a new vertex $v^{\prime}$. When no further preprocessing rules can be applied, we say the graph $G$ is simplified; the resulting graph has a number of nice properties, for instance, it has $\operatorname{minsurp}(G)\geq 2$, and the solution $\vec{\tfrac{1}{2}}$ is the unique optimal solution to $\operatorname{LPVC}(G)$, and the minimum degree of any vertex is $3$.
Proposition 8.
After applying any preprocessing rule, we have $\mu(G^{\prime})\leq\mu(G)$, and $\langle G,k\rangle$ is feasible if and only if $\langle G^{\prime},k^{\prime}\rangle$ is feasible.
There is one situation where rule P3 is particularly powerful. Consider a 3-vertex $u$ which has neighbors $x,y,z$ where $x\sim y$ and $y\sim z$. We refer to this as a kite; see Figure 4.
Figure 4: A kite $u,x,y,z$ before (left) and after (right) applying P3
Lemma 9.
Suppose $\operatorname{minsurp}(G)\geq 1$. Then applying (P3) to a kite in $G$ yields a graph $G^{\prime}$ with $k^{\prime}\leq k-2$ and $\mu(G^{\prime})\leq\mu(G)-1/2$.
2.2 Runtime and branching framework
Our algorithm follows the measure-and-conquer approach. Given an input graph $G$, it runs some preprocessing steps and then generates subproblems $G_{1},\dots,G_{t}$ such that $G$ is feasible if and only if at least one $G_{i}$ is feasible. It then runs recursively on each $G_{i}$. Such an algorithm has runtime $O^{*}(e^{\phi(G)})$ as long as the preprocessing steps have runtime $O^{*}(e^{\phi(G)})$ and it satisfies
$\sum_{i=1}^{t}e^{\phi(G_{i})-\phi(G)}\leq 1.$ (1)
(Here and throughout $e=2.718...$ is the base of the natural logarithm.) We refer to $\phi$ as the measure of the algorithm. For the most part, we will use measures of the form
$\phi(G)=a\mu+bk\qquad\qquad\text{for $a,b\geq 0$};$ (2)
We say a subproblem $G^{\prime}$ has drop $(\Delta\mu,\Delta k)$ if $k^{\prime}\leq k-\Delta k$ and $\mu^{\prime}\leq\mu-\Delta\mu$. We say a branching rule has branch sequence (or branch-seq for short) $B=[(\Delta\mu_{1},\Delta k_{1}),\dots,(\Delta\mu_{t},\Delta k_{t})]$ if generates subproblems $G^{\prime}_{1},\dots,G^{\prime}_{t}$ with the given drops. Given values of $a,b$, we define the value of $B$ to be
$\operatorname{val}_{a,b}(B)=\sum_{i=1}^{t}e^{-a\Delta\mu_{i}-b\Delta k_{i}}.$ (3)
We say branch-seq $B^{\prime}$ dominates $B$ if $\operatorname{val}_{a,b}(B^{\prime})\leq\operatorname{val}_{a,b}(B)$ for all $a,b\geq 0$, and $G$ has a branch-seq $B$ available if we can efficiently find a branching rule which has some branch-seq $B^{\prime}$ dominating $B$. In general, a “balanced” branch-seq dominates an “imbalanced” one; formally, if $\Delta k_{1}\leq\Delta k_{2}$ and $\Delta\mu_{1}\leq\Delta\mu_{2}$, then branch-seq $[(\Delta\mu_{1},\Delta k_{1}),(\Delta\mu_{2},\Delta k_{2})]$ dominates $[(\Delta\mu_{1}-f,\Delta k_{1}-g),(\Delta\mu_{1}+f,\Delta k_{2}+g)]$ for any $f,g\geq 0$.
We use the following important compositional property throughout: if we have a branching rule $B$ generating $\ell$ subproblems with drops $(\Delta\mu_{i},\Delta k_{i})$, and we then apply a branching rule with branch-seq $B^{\prime}_{i}$ to each subproblem $i=1,\dots,\ell$, then, overall, we get
$\operatorname{val}_{a,b}(B)=\sum_{i=1}^{\ell}e^{-a\Delta\mu_{i}-b\Delta k_{i}}\operatorname{val}_{a,b}(B_{i}^{\prime}).$
Lemma 10.
Suppose that $\mathcal{G}$ is a class of graphs closed under vertex deletion, and for which there are vertex cover algorithms with runtimes $O^{*}(e^{a\mu+bk})$ and $O(e^{cn})$ for $a,b,c\geq 0$. Then we can solve vertex cover in $\mathcal{G}$ with runtime $O^{*}(e^{dk})$ where $d=\frac{2c(a+b)}{a+2c}$.
To illustrate Lemma 10, consider graphs of maximum degree 3. We can combine the AGVC algorithm with runtime $O^{*}(2.3146^{\mu})$ (corresponding to $a=\log(2.3146),b=0$), and the MaxIS-3 algorithm with runtime $O^{*}(1.083506^{n})$ (corresponding to $c=\log(1.083506)$), to get an algorithm with runtime $O^{*}(1.14416^{k})$; this already gives us one of the results in Theorem 1.
3 Preprocessing rules and reductions in $k$
We define $S(G)$ to be the largest decrease in $k$ obtainable from applying rules (P1) – (P3) exhaustively via some efficiently-computable sequence of operations. [^3] It is not clear how to calculate the absolute largest decrease in $k$ via preprocessing rules, since applying some rules may prevent opportunities for other rules. We assume that we have fixed some polynomial-time computable sequences of potential preprocessing rules in order to reduce $k$ as small as much as possible. At various points in our algorithm, we will describe simplifications available for intermediate graphs. We always assume that our preprocessing rules are chosen to find such simplifications. As a point of notation, we define $R(X)=S(G-X)$ for a vertex set $X$. For vertices $x_{1},\dots,x_{\ell}$ and vertex sets $X_{1},\dots,X_{r}$, we write $R(x_{1},\dots,x_{\ell},X_{1},\dots,X_{r})$ as shorthand for the more cumbersome $R(\{x_{1},\dots,x_{\ell}\}\cup X_{1}\cup\dots\cup X_{r})=S(G-\{x_{1},\dots,x_{\ell}\}-(X_{1}\cup\dots\cup X_{r}))$. Note that we may have $\ell=0$ or $r=0$, e.g. $R(x_{1},\dots,x_{\ell})$ is shorthand for $R(\{x_{1},\dots,x_{\ell}\})$.
We record a few observations on simplifications of various structures.
Observation 11.
Suppose $x,y$ are non-adjacent subquadratic vertices. Then $S(G)\geq 2$ if any of the following three conditions hold: (i) $\operatorname{codeg}(x,y)=0$ or (ii) $\deg(x)+\deg(y)\geq 3$ or (iii) $\operatorname{minsurp}(G)\geq 0$.
Proof.
If $\deg(x)=\deg(y)=1$ and $\operatorname{codeg}(x,y)=0$, then $\{x,y\}$ is a min-set with $\operatorname{surp}(\{x,y\})=0$, and applying P1 reduces $k$ by two. Otherwise, if $\deg(x)\geq 2$, then we can apply (P2) to $x$, and vertex $y$ remains subquadratic, and we can follow up by applying (P1) or (P2) again to $y$. Note that if $\operatorname{minsurp}(G)\geq 0$, then at least one of the conditions (i) or (ii) must hold. ∎
Proposition 12.
If $G$ has 2-vertices $x_{1},\dots,x_{\ell}$ which all have pairwise distance at least $3$, then $S(G)\geq\ell$.
Proof.
We show it by induction on $\ell$. The base case $\ell=0$ is vacuous. For the induction step, we apply (P2) to $x_{\ell}$, forming a graph $G^{\prime}$ where the neighbors $y,z$ get contracted into a single new vertex $t$. By hypothesis, $y,z$ are distinct from $x_{1},\dots,x_{\ell}$. We claim that $\operatorname{dist}_{G^{\prime}}(x_{i},x_{j})\geq 3$ and $\deg_{G^{\prime}}(x_{i})\geq 2$ for any pair $i<j<\ell$. For, clearly $x_{i}\not\sim x_{j}$ in $G^{\prime}$. If $x_{i},x_{j}$ share a neighbor in $G^{\prime}$, it must be vertex $t$ as no other vertices were modified. This is only possible if $\{x_{i},x_{j}\}\sim\{y,z\}$; but this contradicts our hypothesis that $\operatorname{codeg}(x_{i},x_{\ell})=\operatorname{codeg}(x_{j},x_{\ell})=0$. Finally, suppose that $\deg_{G^{\prime}}(x_{i})\leq 1$. This is only possible if the two neighbors of $x_{i}$ got merged together, i.e. $x_{i}\sim y$ and $x_{i}\sim z$. Again, this contradicts that $\operatorname{codeg}(x_{i},x_{\ell})=0$.
By induction hypothesis applied to $G^{\prime}$, we have $S(G^{\prime})\geq\ell-1$ and hence $S(G)\geq\ell$. ∎
Lemma 13.
If $\operatorname{minsurp}(G)\geq 0$ and indset $I$ has $\operatorname{surp}_{G}(I)\leq 1$, then $S(G)\geq|I|$.
Proof.
Let $r=|I|$. If $\operatorname{surp}_{G}(I)=0$, or $I$ is a critical-set, we can apply (P1) or (P2) to $I$. So suppose that $\operatorname{surp}_{G}(I)=1$ and $\operatorname{surp}_{G}(X)=0$ for a non-empty subset $X\subsetneq I$; among all such subsets $X$, choose one of maximum size, and let $s=|X|$. We apply (P1) to $X$, reducing $k$ by $s$ and obtaining a graph $G^{\prime}=G-N[X]$. Now consider the indset $J=I\setminus X$ in $G^{\prime}$, where $|J|=r-s$. We have $|N_{G^{\prime}}(J)|=|N_{G}(I)\setminus N_{G}(X)|=r+1-s$ and so $\operatorname{surp}_{G^{\prime}}(J)=1$.
We claim that $J$ is a critical-set in $G^{\prime}$. For, if some non-empty subset $X^{\prime}\subseteq J$ has $|N_{G^{\prime}}(X^{\prime})|\leq|X^{\prime}|$, then indset $I^{\prime}=X\cup X^{\prime}\subseteq I$ would have $|N_{G}(I^{\prime})|\leq|N_{G}(X)|+|N_{G^{\prime}}(X^{\prime})|\leq|X|+|X^{\prime}|=|I^{\prime}|$. Hence $\operatorname{surp}_{G}(I^{\prime})\leq 0$, contradicting maximality of $X$. So, we can apply (P2) to $J$ in $G^{\prime}$, getting a further drop of $r-s$. Overall, we get a net drop of $s+(r-s)=r$ as desired. ∎
4 Branching rules
Most of our branching rules can be viewed in terms of guessing that a certain set of vertices $X_{1}$ is in the cover. In the graph $G-X_{1}$, a set of vertices $X_{0}$ may become isolated. We apply rule (P1) to remove $X_{0}$; the resulting subproblem $\langle G-(X_{0}\cup X_{1}),k-|X_{1}|\rangle$ is called principal, and we define the excess to be $|X_{0}|$. Note that the graph $G^{\prime}=G-(X_{0}\cup X_{1})$ may still contain isolated vertices, or have additional simplifications available. We say $\langle G^{\prime},k-|X_{1}|\rangle$ has drop $(\Delta\mu^{\prime},|X_{1}|)$ directly, and the final subproblem $G^{\prime\prime}=\langle G^{\prime\prime},k^{\prime\prime}\rangle$ has drop $(\Delta\mu^{\prime\prime},\Delta k^{\prime\prime})$ after simplification.
Since this plays a central role in our analysis, we define the shadow of a vertex set $X$, denoted $\operatorname{shad}(X)$, to be $\operatorname{minsurp}(G-X)$. For readability, we write $\operatorname{shad}(x_{1},\dots,x_{\ell},X_{1},\dots,X_{r})$ as shorthand for $\operatorname{shad}(\{x_{1},\dots,x_{\ell}\}\cup X_{1}\cup\dots\cup X_{r})=\operatorname{minsurp}(G-\{x_{1},\dots,x_{\ell}\}-(X_{1}\cup\dots\cup X_{r}))$.
Proposition 14.
If $\operatorname{minsurp}(G)\geq 2$, then a principal subproblem $\langle G-X,k^{\prime}\rangle$ with excess $s$ has $\Delta\mu=\tfrac{1}{2}(\Delta k-s+\operatorname{minsurp}^{-}(G^{\prime}))=\tfrac{1}{2}(\Delta k-s+\min\{0,\operatorname{shad}(X)\})$.
Proof.
Here Proposition 4 gives $\lambda(G)=n/2$. Let $G^{\prime}=G-X$, where $|X|=\Delta k+s$. Then $\mu(G^{\prime})=k^{\prime}-\operatorname{\lambda}(G^{\prime})=k^{\prime}-\tfrac{1}{2}(n^{\prime}+\operatorname{minsurp}^{-}(G^{\prime}))$, where $n^{\prime}=n-X=2\lambda(G)-(s+\Delta k)$, which simplifies to $\mu(G^{\prime})=\mu(G)-\tfrac{1}{2}(\Delta k-s+\operatorname{minsurp}^{-}(G^{\prime}))$ as desired. ∎
Observation 15.
Suppose $\operatorname{minsurp}(G)\geq 2$. Then for any indset $I$, there holds $\operatorname{shad}(I)\geq 2-|I|$ and $\operatorname{shad}(N[I])\geq 2-\operatorname{surp}_{G}(I)$.
In particular, for a vertex $u$, there holds $\operatorname{shad}(u)\geq 1$ and $\operatorname{shad}(N[u])\geq 3-\deg(u)$.
Proof.
If $J$ is a min-set of $G-I$, then $2\leq\operatorname{surp}_{G}(J)\leq\operatorname{surp}_{G-I}(J)+|I|$. So $\operatorname{surp}_{G-I}(J)\geq 2-|I|$. Similarly, if $J$ is a min-set of $G-N[I]$, then $2\leq\operatorname{surp}_{G}(I\cup J)=\operatorname{surp}_{G-N[I]}(J)+\operatorname{surp}_{G}(I)$. ∎
By combining Observation 11 and Observation 15, we immediately get the following simplification rule (which is used ubiquitously):
Observation 16.
If $G$ is simplified and $G-\{u,v\}$ has two non-adjacent subquadratic vertices, then $R(u,v)\geq 2$.
Our bread-and-butter branching rule is to choose some vertex $u$, and branch on whether it is in the cover. We refer to this as splitting on $u$. This generates two principal subproblems $\langle G-u,k-1\rangle$ and $\langle G-N[u],k-\deg(u)\rangle$. More generally, in light of Lemma 5, we know that for any critical-set $I$, there is either a good cover containing $I$ or omitting $I$. We can branch on subproblems $\langle G-I,k-|I|\rangle$ and $\langle G-N[I],k-N[I]\rangle$, and we refer to this as splitting on $I$. Splitting on a vertex is a special case where $I$ is a singleton.
Consider the effect of splitting on $u$. By Observation 15, the subproblem $\langle G-u,k-1\rangle$ has drop $(0.5,1)$ and the subproblem $\langle G-N[u],k-\deg(u)\rangle$ has drop $(1,\deg(u))$. The latter bound, however, is very loose and should be improved to get an optimized analysis. In light of Proposition 14, it revolves around the value of $\operatorname{shad}(N[u])$. If $\operatorname{shad}(N[u])\geq 0$, then $\lambda$ drops by $\frac{\deg(u)+1}{2}$ and so the subproblem has drop $(\tfrac{\deg(u)-1}{2},\deg(u))$. This is the “generic” situation for splitting on $u$.
When $\operatorname{shad}(N[u])\leq 0$, we say that $u$ is blocked. If $x$ is a vertex in a min-set of $G-N[u]$, we say that $x$ is a blocker of $u$. Necessarily, $x\nsim u$. This motivates the following powerful branching rule:
(B) If $x$ is a blocker of vertices $u_{1},\dots,u_{\ell}$, then branch on subproblems $\langle G-\{u_{1},\dots,u_{\ell},x\},k-\ell-1\rangle$ and $\langle G-N[x],k-\deg(x)\rangle$.
Proposition 17.
Rule (B) is a valid branching rule.
Proof.
The subproblem $\langle G-N[x],k-\deg(x)\rangle$ is feasible if and only if $G$ has a good cover omitting $x$. The subproblem $\langle G-\{u_{1},\dots,u_{\ell},x\},k-\ell-1\rangle$ is feasible if and only if $G$ has a good cover which includes all vertices $u_{1},\dots,u_{\ell},x$. We claim that, if $G$ is feasible, at least one of these cases holds. For, suppose a good cover of $G$ omits vertex $u_{i}$. Then $G-N[u_{i}]$ has a cover of size at most $k-\deg(u_{i})$. By Lemma 5, this implies that $G-N[u_{i}]$ has such a cover $C^{\prime}$ which omits any min-set $I$ of $G-N[u_{i}]$, and in particular $x\notin C^{\prime}$. Then $G$ has a good cover $C\cup N(u_{i})$ omitting $x$. ∎
In the vast majority of cases where we use (B), we have $\ell=1$, that is, $x$ is a blocker of a single vertex $u$. To provide intuition, keep in mind the picture that if $G-N[u]$ has $s$ isolated vertices $v_{1},\dots,v_{s}$, then $\{v_{1},\dots,v_{s}\}$ would be an indset in $G-N[u]$ with zero neighbors and surplus $-s$, in particular, $\operatorname{shad}(N[u])\leq-s$. These vertices were hidden in the “shadow” of $N[u]$. See Figure 5.
Figure 5: Here, vertices $v_{1},v_{2}$ are both blockers of $u$; in particular, we have $\operatorname{shad}(N[u])\leq-2$.
Also, note that if $\operatorname{shad}(N[u])=0$ exactly, then we could still split on $u$ if desired and get the ideal drop in $\mu$. However, we can also apply rule (B) if desired, and this is usually more profitable. Vertices with $\operatorname{shad}(N[u])=0$ share many properties with vertices with $\operatorname{shad}(N[u])<0$ strictly. This is the reason for the somewhat unintuitive definition of blocked vertex.
5 Analysis of blockers and splitting
We will develop a series of branching rules (typically using rule (B)) to handle blocked vertices, along with a number of other related cases. The most significant consequence of these rules is that we are able to get near-ideal branch sequences for high-degree vertices. Specifically, we will show the following main result which is the heart of our branching algorithm:
{restatable}
theoremthmA Suppose $G$ is simplified, and let $r=\operatorname{maxdeg}(G)$. Depending on $r$, then following branch-seqs are available:
If $r\geq 4$: $[(1,3),(1,5)]$ or $[(0.5,1),(1.5,4)]$.
If $r\geq 5$: $[(1,3),(1,5)]$ or $[(0.5,1),(2,5)]$.
If $r\geq 6$: $[(1,3),(1,5)]$, $[(0.5,2),(2,5)]$, or $[(0.5,1),(2.5,r)]$.
The proof of Section 5 has many cases which will require us to build up slowly from analysis of low-degree vertices. For Section 5 only, we always assume that the starting graph $\boldsymbol{G}$ is simplified. We begin with a few elementary observations.
Observation 18.
Suppose $u$ is blocked and let $I$ be a min-set of $G-N[u]$. There is at least one vertex $x\in I$ with $\operatorname{codeg}(x,u)\geq 1$.
Proof.
If not, we would have $\operatorname{surp}_{G}(I)=\operatorname{surp}_{G-N[u]}(I)\leq 0$, contradicting that $G$ is simplified. ∎
Proposition 19.
Suppose vertex $u$ has $\operatorname{shad}(N[u])\leq 4-\deg(u)$ and $x$ is a blocker of $u$. If we apply (B) to vertices $u,x$, then subproblem $G^{\prime}=\langle G-N[x],k-\deg(x)\rangle$ has drop $(1,4)$. Moreover, if $\deg(x)=3$ and $x$ is contained in a non-singleton min-set of $G-N[u]$, then $G^{\prime}$ has drop $(1,5)$.
Proof.
Since $G$ is simplified, we have $\deg(x)\geq 3$ and $\operatorname{minsurp}(G)\geq 2$. If $\deg(x)=4$, then $G^{\prime}$ has drop $(1,4)$ directly. So, suppose $\deg(x)=3$ exactly. Consider any min-set $I$ of $G-N[u]$ with $x\in I$. The indset $J=I\cup\{u\}\setminus\{x\}$ has $\operatorname{surp}_{G-N[x]}(J)\leq\operatorname{surp}_{G-N[u]}(I)+(\deg(u)-\deg(x))\leq 1$. On the other hand, $\operatorname{shad}(N[x])\geq\operatorname{minsurp}(G)-\deg(x)+1\geq 0$. So Lemma 13 gives $R(N[x])\geq|J|=|I|$.
By Proposition 14 we see that $G^{\prime}$ has drop $(1,3)$ directly. Since $|I|\geq 1$ trivially, we have $R(N[x])\geq 1$ and $G^{\prime}$ has drop $(1,4)$ after simplification. If $|I|\geq 2$, then $R(N[x])\geq 2$ and likewise $G^{\prime}$ has drop $(1,5)$ after simplification. ∎
5.1 Branching rules for indsets with surplus two.
It will be very important to develop special branching rules for indsets with surplus two. Since $G$ is assumed to be simplified, note that any such set $I$ will be a critical set, and we can split on $I$ to get subproblems $\langle G-I,k-|I|\rangle$ and $\langle G-N[I],k-|I|-2\rangle$. However, this branching rule is not powerful enough for us; we develop specialized branching rules for smaller indsets (indsets with cardinality two or three).
Proposition 20.
Suppose $G$ has an indset $I$ with $\operatorname{surp}_{G}(I)=2$, and let $z\in N(I)$. If we split on $z$, the subproblem $\langle G-z,k-1\rangle$ has drop $(0.5,1+|I|)$. If $z$ has a blocker $x\notin I$ and we apply (B) to $z,x$, the subproblem $\langle G-\{z,x\},k-2\rangle$ has drop $(1,2+|I|)$.
Proof.
Subproblem $\langle G-z,k-1\rangle$ has drop $(0.5,1)$ directly. We have $\operatorname{surp}_{G-z}(I)=\operatorname{surp}_{G}(I)-1=1$, and $\operatorname{shad}(G-z)\geq\operatorname{minsurp}(G)-1\geq 1$. So (P2) applied to $I$ gives $R(z)\geq|I|$. Likewise, subproblem $\langle G-\{z,x\},k-2\rangle$ has drop $(1,2)$ directly, and $\operatorname{shad}(z,x)\geq\operatorname{minsurp}(G)-2\geq 0$ and $\operatorname{surp}_{G-\{z,x\}}(I)\leq\operatorname{surp}_{G}(I)-1\leq 1$. So Lemma 13 gives $R(z,x)\geq|I|$. ∎
Lemma 21.
Suppose $G$ has an indset $I$ with $\operatorname{surp}_{G}(I)=2,|I|\geq 3$. Then $G$ has available branch-seq $[(1,4),(1,5)]$ or $[(0.5,4),(2,5)]$.
Proof.
If $|I|\geq 4$, then splitting on $I$ gives branch-seq $[(1,4),(1,6)]$. So suppose $|I|=3$. We first claim that there is some vertex $z$ with $|N(z)\cap I|=2$. For, consider the set of vertices $J=\{z:|N(z)\cap I|\geq 2\}$. The vertices in $I$ are independent and have degree at least three, so there are at least 9 edges from $I$ to $N(I)$. Since $|N(I)|=5$, by the pigeonhole principle we must have $|J|\geq 2$. If $N(z)\not\subseteq I$ for all $z\in J$, then $J$ would be an indset with surplus at most one, contradicting that $G$ is simplified. Hence there is $z\in J$ with $|N(z)\cap I|=2$ exactly.
Let $z$ be any such vertex, and let $I=\{x_{1},x_{2},y\}$ where $z\sim x_{1},z\sim x_{2},z\not\sim y$ and $\deg(x_{1})\leq\deg(x_{2})$. If $\deg(z)\leq 4$, then we split on $I$, generating subproblems $\langle G-I,k-3\rangle$ and $\langle G-N[I],k-5\rangle$ with drops $(1,3)$ and $(1,5)$ directly, where $G-I$ has a subquadratic vertex $z$ so it has drop $(1,4)$ after simplification. If $\deg(z)\geq 5$ and $\operatorname{shad}(N[z])\geq 5-\deg(z)$, then we split on $z$; by Proposition 20, subproblem $\langle G-z,k-1\rangle$ has drop $(0.5,4)$, and subproblem $\langle G-N[z],k-\deg(z)\rangle$ has drop $(2,5)$ directly. If $\deg(z)\geq 5$ and $z$ has a blocker $t\neq y$, then we apply (B) to $z,t$; by Proposition 19, subproblem $\langle G-N[t],k-\deg(t)\rangle$ has drop $(1,4)$ and by Proposition 20, subproblem $\langle G-\{z,t\},k-2\rangle$ has drop $(1,5)$.
The only remaining possibility is if $\deg(z)\geq 5$ and $y$ is the only blocker of $z$, i.e. $\deg(z)=5,\deg(y)=3$ and $N(y)\subseteq N(z)$. Each vertex $x_{i}$ has $\deg(x_{i})+\deg(y)-\operatorname{codeg}(x_{i},y)\leq|N(I)|$, i.e. $\operatorname{codeg}(x_{i},y)\geq\deg(x_{i})-2$. If $\deg(x_{1})=3$, this implies that $x_{1},y$ share some neighbor $u$; but in this case $G$ has a 3-triangle $x_{1},u,z$, contradicting that $G$ is simplified.
So we suppose $\deg(x_{2})\geq\deg(x_{1})\geq 4$ and then $\operatorname{codeg}(x_{1},y)\geq 2$ and $\operatorname{codeg}(x_{2},y)\geq 2$. Equivalently, $\deg_{G-N[x_{1}]}(y)\leq 1$ and $\deg_{G-N[x_{2}]}(y)\leq 1$. So $y$ is a blocker of all three vertices $x_{1},x_{2},z$. We apply (B) giving subproblems $\langle G-N[y],k-\deg(y)\rangle$ and $\langle G-\{x_{1},x_{2},z,y\},k-4\rangle$, with drops $(1,3)$ and $(1,4)$ directly. Furthermore, $\operatorname{surp}_{G-N[y]}(\{x_{1},x_{2}\})\leq 2-|N_{G-N[y]}(I)|=2-2=0$ and neither $x_{1}$ or $x_{2}$ is isolated in $G-N[y]$ (they retain neighbor $z$). So we can apply (P1) to indset $\{x_{1},x_{2}\}$ in $G-N[y]$, giving $R(N[y])\geq 2$. So $G-N[y]$ has drop $(1,5)$ after simplification. ∎
Lemma 22.
Suppose $G$ has an indset $I$ with $\operatorname{surp}_{G}(I)=2,|I|\geq 2$. Then $G$ has available branch-seq $[(1,4),(1,4)]$ or $[(0.5,3),(2,5)]$.
Proof.
If $|I|\geq 3$, then we apply Lemma 21. So let $I=\{x,y\}$, where $\deg(x)\leq\deg(y)$ and let $A=N(x,y)$. Note that $|A|=\deg(x)+\deg(y)-4\geq\deg(x)-1$. The vertices in $A$ cannot form a clique, as then $x$ would have a funnel. So consider vertices $z_{1},z_{2}\in A$ with $z_{1}\not\sim z_{2}$.
If $z_{1}$ and $z_{2}$ are both subquartic, then we split on $I$; the subproblem $\langle G-N[I],k-|N[I]|-2$ has drop $(1,4)$ directly. The subproblem $G-I=G-\{x,y\}$ has non-adjacent subquadratic vertices $z_{1},z_{2}$, so by Observation 16, we have $R(x,y)\geq 2$ and subproblem $\langle G-I,k-|I|\rangle$ has drop $(1,4)$ after simplification.
So suppose $\deg(z_{1})\geq 5$. If $\operatorname{shad}(N[z_{1}])\geq 0$, then split on $z_{1}$; by Proposition 20, the subproblem $\langle G-z_{1},k-1\rangle$ has drop $(0.5,3)$, while the subproblem $\langle G-N[z_{1}],k-\deg(z_{1})\rangle$ has drop $(2,5)$ directly. Otherwise, suppose $\operatorname{shad}(N[z_{1}])\leq-1$ and $t$ is a blocker of $z_{1}$. Note that $t\notin I$ since $t\not\sim z_{1}$ whereas $z_{1}\in N(x)\cap N(y)$. We then apply (B) to $t,z_{1}$; the subproblem $\langle G-N[z_{1}],k-\deg(z_{1})\rangle$ has drop $(1,5)$ directly, and by Proposition 20, the subproblem $\langle G-\{t,z\},k-2\rangle$ has drop $(1,4)$. ∎
5.2 Branching rules for blocked vertices.
We now turn to developing branching rules for blocked vertices of various types. We need to work our way from low to high degree.
Proposition 23.
Suppose vertex $u$ has $\operatorname{shad}(N[u])\leq 5-\deg(u)$ and $u$ has a blocker $x$ of degree at least 4. Then $G$ has available branch-seq $[(1,4),(1,4)]$ or $[(0.5,2),(2,5)]$.
Proof.
If $\operatorname{shad}(N[x])\leq 3-\deg(x)$ and $J$ is a min-set of $G-N[x]$, then $\operatorname{surp}_{G}(J\cup\{x\})\leq(3-\deg(x))+(\deg(x)-1)\leq 2$ and we can apply Lemma 22 to indset $J\cup\{x\}$.
So suppose $\operatorname{shad}(N[x])\geq 4-\deg(x)$. In this case, we apply (B) to $u,x$. Subproblem $\langle G-\{u,x\},k-2\rangle$ has drop $(1,2)$ directly. If $\deg(x)\geq 5$, then subproblem $\langle G-N[x],k-\deg(x)\rangle$ has drop $(1.5,5)$ directly. If $\deg(x)=4$, then consider a min-set $I$ of $G-N[u]$ with $x\in I$. We have $\operatorname{surp}_{G-N[x]}(I\cup\{u\}\setminus\{x\})\leq\operatorname{surp}_{G-N[u]}(I)+\deg(u)-\deg(x)\leq 1$. So $R(N[x])\geq 1$ and $G^{\prime}$ has drop $(1.5,5)$ after simplification. In either case, $\langle G-\{u,x\},k-2\rangle$ and $\langle G-N[x],k-\deg(x)\rangle$ give drops $(1.2),(1.5,5)$ respectively. This branch-seq $[(1,2),(1.5,5)]$ dominates $[(0.5,2),(2,5)]$. ∎
Lemma 24.
Suppose $G$ has a vertex of degree at least 5 which has a 3-neighbor. Then $G$ has available branch-seq $[(1,3),(1,5)]$ or $[(0.5,2),(2.5)]$.
Proof.
Let $u$ be a vertex of degree at least $5$ and let $z$ be a 3-neighbor of $u$. There are a number of cases to consider.
Case I: $\boldsymbol{\operatorname{shad}(N[u])\geq 5-\deg(u)}$. Splitting on $u$ gives subproblems $G-u,G-N[u]$ with drops $(0.5,1),(2,5)$ directly; the former has drop $(0.5,2)$ after simplification due to its 2-neighbor $z$.
Case II: $\boldsymbol{u}$ has a blocker $\boldsymbol{x}$ with $\boldsymbol{\deg(x)\geq 4}$. We suppose that $\operatorname{shad}(N[u])\leq 4-\deg(u)$ since otherwise it would be covered in Case I. Then the result follows from Proposition 23, noting that $[(1,4),(1,4)]$ dominates $[(1,3),(1,5)]$.
Case III: $\boldsymbol{u}$ has a blocker $\boldsymbol{x}$ with $\boldsymbol{R(u,x)\geq 2}$. We apply (B) to vertices $u,x$. Subproblem $\langle G-\{u,x\},k-2\rangle$ has drop $(1,4)$ after simplification, and by Proposition 19 subproblem $\langle G-N[x],k-\deg(x)\rangle$ has drop $(1,4)$. We get branch-seq $[(1,4),(1,4)]$, which dominates $[(1,3),(1,5)]$.
Case IV: $\boldsymbol{G-N[u]}$ has some non-singleton min-set $\boldsymbol{I}$. Let $x\in I$. Necessarily $\deg(x)=3$, else it would be covered in Case II. We apply (B) to $u,x$; subproblem $\langle G-\{u,x\},k-2\rangle$ has drop $(1,3)$ after simplification due to subquadratic vertex $z$. By Proposition 19 subproblem $\langle G-N[x],k-\deg(x)\rangle$ has drop $(1,5)$ after simplification.
Case V: No previous cases apply. Suppose that no vertices in the graph are covered by any of the previous cases. Let $U$ denote the non-empty set of vertices $u$ with $\deg(u)\geq 5$ and $u$ having a 3-neighbor. We claim that every vertex $u\in U$ has degree exactly 5. For, suppose $\deg(u)\geq 6$, and since we are not in Case I we have $\operatorname{shad}(N[u])\leq 4-\deg(u)\leq-2$. Then necessarily a min-set of $G-N[u]$ would have size at least two, and it would be covered in Case IV.
So every $u\in U$ has degree exactly 5, and is blocked by a singleton 3-vertex $x$, i.e. $N(x)\subseteq N(u)$. Let us say in this case that $x$ is linked to $u$; we denote by $X$ the set of all such 3-vertices linked to any vertex in $U$. We claim that each $x\in X$ has at least two neighbors of degree at least 5, which must be in $U$ due to their 3-neighbor $x$. For, suppose that $x$ has two subquartic neighbors $z_{1},z_{2}$; necessarily $z_{1}\not\sim z_{2}$ since $G$ is simplified and $x$ has degree 3. Then $R(u,x)\geq 2$ by Observation 16, which would have been covered in Case III.
On the other hand, we claim that each vertex $u\in U$ has at most one neighbor in $X$. For, suppose $u$ is linked to $x$ and $u$ has two 3-neighbors $x_{1},x_{2}\in X$, which must be non-adjacent since $G$ is simplified. Then Observation 16 gives $R(u,x)\geq 2$ due to subquadratic vertices $x_{1},x_{2}$. This would have been covered in Case III.
Thus, we see that each vertex in $X$ has at least two neighbors in $U$ and each vertex in $U$ has at most one neighbor in $X$. Hence $|U|\geq 2|X|$, and so by the pigeonhole principle, there must be some vertex $x\in X$ which is linked to two vertices $u_{1},u_{2}\in U$. We apply (B) to $u_{1},u_{2},x$, getting subproblems $\langle G-\{u_{1},u_{2},x\},k-3\rangle$ and $\langle G-N[x],k-\deg(x)\rangle$. By Proposition 19, the latter subproblem has drop $(1,4)$. Furthermore, if $z_{1},z_{2}$ are two 5-neighbors of $x$, then $z_{1}\not\sim z_{2}$ else $x$ would have a 3-triangle. So $G-\{u_{1},u_{2},x\}$ has non-adjacent 2-vertices $z_{1},z_{2}$, and subproblem $\langle G-\{u_{1},u_{2},x\},k-3\rangle$ has drop $(1,5)$ after simplification. ∎
Proposition 25.
Suppose a vertex $u$ has $\deg(u)\geq 4$ and $\operatorname{shad}(N[u])\leq 4-\deg(u)$. Then $G$ has available branch-seq $[(1,3),(1,5)]$ or $[(0.5,2),(2,5)]$.
Proof.
Let $I$ be a min-set of $G-N[u]$. By Observation 18, there is a vertex $x\in I$ which shares some neighbor $t$ with $u$. If $\deg(x)\geq 4$, then we apply Proposition 23, noting that $[(1,4),(1,4)]$ dominates $[(1,3),(1,5)]$. If $\deg(x)=3$ and $\deg(t)\geq 5$, we apply Lemma 24 to $t$. If $\deg(x)=3,\deg(t)\leq 4,|I|\geq 2$, we apply (B) to $u,x$; subproblem $\langle G-\{u,x\},k-2\rangle$ has drop $(1,3)$ after simplification (due to subquadratic vertex $t$) and by Proposition 19 subproblem $\langle G-N[x],k-\deg(x)\rangle$ has drop $(1,5)$.
So finally take $I=\{x\}$ where $\deg(x)=3$ and $x$ shares two neighbors $t_{1},t_{2}$ with $u$. Necessarily $t_{1}\not\sim t_{2}$ since $G$ has no 3-triangles. If either $t_{1}$ or $t_{2}$ has degree 5 or more, we can apply Lemma 24. If they are both subquartic, then we apply (B) to $u,x$; by Proposition 19, subproblem $\langle G-N[x],k-\deg(x)\rangle$ has drop $(1,4)$ and by Observation 16 there holds $R(u,x)\geq 2$ and subproblem $\langle G-\{u,x\},k-2\rangle$ has drop $(1,4)$ after simplification. Here, $[(1,4),(1,4)]$ dominates $[(1,3),(1,5)]$. ∎
Lemma 26.
Suppose vertex $u$ has $\deg(u)\geq 5$ and $\operatorname{shad}(N[u])\leq 5-\deg(u)$.
  • If $\deg(u)=5$, then $G$ has available branch-seq $[(1,3),(1,4)]$ or $[(0.5,2),(2,5)]$.
  • If $\deg(u)\geq 6$, then $G$ has available branch-seq $[(1,3),(1,5)]$ or $[(0.5,2),(2,5)]$.
Proof.
Let $I$ be a min-set of $G-N[u]$ of smallest size. If any vertex in $I$ has degree at least 4, the result follows from Proposition 23 (noting that $[(1,4),(1,4)]$ dominates $[(1,3),(1,5)]$. So suppose that all vertices in $I$ have degree 3. By Observation 18, we may select some $x\in I$ with $\operatorname{codeg}(x,u)\geq 1$. For this vertex $x$, define $Z=N(x)\cap N(u)$ and $Y=N_{G-N[u]}(x)$, where $|Z|+|Y|=\deg(x)=3$ and $|Z|\geq 1$.
If any vertex $z\in Z$ has degree at least 5, we apply Lemma 24 to $z$ with its 3-neighbor $x$. If any vertex $z\in Z$ has degree 3, we apply Lemma 24 to $u$ with its 3-neighbor $z$. If $\operatorname{codeg}(x,x^{\prime})\geq 2$ for any vertex $x^{\prime}\in I\setminus\{x\}$, then we apply Lemma 22 to the surplus-two indset $\{x,x^{\prime}\}$. So we suppose that $\operatorname{codeg}(x,x^{\prime})\leq 1$ for all $x^{\prime}\in I\setminus\{x\}$ and that all vertices in $Z$ have degree exactly four.
The vertices in $Z$ must be independent, as otherwise $G$ would have a 3-triangle with $x$. If $\operatorname{codeg}(z,z^{\prime})\geq 3$ for any vertices $z,z^{\prime}\in Z$, then $\operatorname{shad}(N[z])\leq 0$ (since $G-N[z]$ has a 1-vertex $z^{\prime}$), and we can apply Proposition 25. So we suppose the vertices in $Z$ share no neighbors besides $u$ and $x$. In this case, in the graph $G-\{u,x\}$, the vertices in $Z$ become subquadratic and share no common neighbors. By Proposition 12, we thus have $R(u,x)\geq|Z|$.
We next claim that if $|I|\geq 2$, then $R(u,x)\geq|Y|$. For, every vertex $y\in Y$ must have a neighbor $\sigma(y)\in I\setminus\{x\}$, as otherwise we would have $\operatorname{surp}_{G-N[u]}(I^{\prime})\leq\operatorname{surp}_{G-N[u]}(I)$ for non-empty ind-set $I^{\prime}=I\setminus\{x\}$, contradicting minimality of $I$. The vertices $\sigma(y):y\in Y$ must be distinct, as a common vertex $x^{\prime}=\sigma(y)=\sigma(y^{\prime})$ would have $\operatorname{codeg}(x,x^{\prime})\geq 2$ which we have already ruled out. So, in the graph $G-N[x]$, each vertex $\sigma(y)\in I$ loses neighbor $y$. By Observation 15, we have $\operatorname{shad}(N[x])\geq 3-\deg(x)=0$. So $G-N[x]$ has $|Y|\leq 2$ non-adjacent subquadratic vertices. Observation 16 implies that $R(N[x])\geq|Y|$.
At this point, our strategy is to apply (B) to $u,x$. If $|I|=1$, then either $\deg(u)=5,|Z|\geq 2$ or $\deg(u)\geq 6,|Z|=3$. Subproblem $\langle G-N[x],k-\deg(x)\rangle$ has drop $(1,3)$ directly and subproblem $\langle G-\{u,x\},k-2\rangle$ has drop $(1,2+|Z|)$ after simplification; that is, it has drop $(1,4)$ if $\deg(u)=5$ and drop $(1,5)$ otherwise.
Otherwise, if $|I|\geq 2$, then subproblem $\langle G-\{u,x\},k-2\rangle$ has drop $(1,2+|Z|)$ after simplification and subproblem $\langle G-N[x],k-\deg(x)\rangle$ has drop $(1,3+|Y|)$ after simplification. Since $|Z|+|Y|=3$ and $|Z|\geq 1$, this always gives drops $(1,3),(1,5)$ or $(1,4),(1,4)$. ∎
[…]
Proof.
Consider an $r$-vertex $u$. If $r=4$ and $\operatorname{shad}(N[u])\geq 0$, then split on $u$ with branch-seq $[(0.5,1),(1.5,4)]$. If $r=4$ and $\operatorname{shad}(N[u])\leq-1$, then use Proposition 25. If $r=5$ and $\operatorname{shad}(N[u])\geq 0$, then split on $u$ with branch-seq $[(0.5,1),(2,5)]$. If $r=5$ and $\operatorname{shad}(N[u])\leq-1$, then apply Proposition 25. If $r\geq 6$ and $\operatorname{shad}(N[u])\geq 6-r$, then split on $u$, getting branch-seq $[(0.5,1),(2.5,r)]$. If $r\geq 6$ and $\operatorname{shad}(N[u])\leq 5-r$, then apply Lemma 26. ∎
6 A simple branching algorithm
We now describe our first branching algorithm for vertex cover. As discussed earlier, the measure is a piecewise-linear function of $k$ and $\mu$. It is more convenient to describe it in terms of multiple self-contained “mini-algorithms,” with linear measure functions. As a starting point, we can describe the first algorithm:
Simplify $G$. If $\operatorname{maxdeg}(G)\leq 3$, then run either the algorithm of Theorem 2 or the MaxIS-3 algorithm, whichever is cheaper, and return. Otherwise, apply Section 5 to an arbitrary vertex of degree at least $4$, and run Branch4Simple on the two resulting subproblems.
Algorithm 1 Function Branch4Simple$(G,k)$
To clarify, despite the name Branch4Simple, the graph is allowed to have vertices of either higher or lower degrees. What we mean is that, depending on the current values of $k$ and $\mu$, it is advantageous to branch on a vertex of degree 4 or higher. Our algorithm is looking for such a vertex; if it is not present (i.e. the graph has maximum degree 3) then an alternate algorithm should be used instead. Specifically, here we use either the AGVC algorithm with runtime depending on $\mu$ or the MaxIS-3 algorithm with runtime depending on $n=2(k-\mu)$.
Henceforth we describe our branching algorithms more concisely. Whenever we apply a branching rule, we assume that we recursively use the algorithm under consideration (here, Branch4Simple) on the resulting subproblems and return. Likewise, when we list some other algorithms for the problem, we assume they are dovetailed (effectively running the cheapest of them).
Proposition 27.
Branch4Simple has measure $a\mu+bk$ for $a=0.71808,b=0.019442$.
Proof.
Simplifying the graph in Line 1 can only reduce $\phi(G)=a\mu+bk$. If Section 5 is used, it gives a branch-seq $B$ with drops dominated by $[(1,3),(1,5)]$ or $[(0.5,1),(1.5,4)]$. To satisfy Eq. (3), we thus need:
$\operatorname{val}_{a,b}(B)\leq\max\{e^{-a-3b}+e^{-a-5b},e^{-0.5a-b}+e^{-1.5a-4b}\}\leq 1$ (4)
Otherwise, suppose $G$ has maximum degree $3$. Since it is simplified, it has $n=2\lambda=2(k-\mu)$, so the MaxIS-3 algorithm has runtime $O^{*}(1.083506^{2(k-\mu)})$. Then, it suffices to show that
$\min\{\mu\log(2.3146),2(k-\mu)\log(1.083506)\}\leq a\mu+bk.$ (5)
This can be verified mechanically. ∎
By Lemma 10, this combined with the MaxIS-4 algorithm (with $a=0.71808,b=0.019442,c=\log 1.137595$) immediately gives runtime $O^{*}(1.2152^{k})$ for degree-4 graphs.
Although Proposition 27 is easy to check directly, the choices for the parameters $a$ and $b$ may seem mysterious. The explanation is that the most constraining part of the algorithm is dealing with the lower-degree graphs, i.e. solving the problem when the graph has maximum degree $3$. The inequality (5), which governs this case, should be completely tight: there should be a “triple point” with respect to Branch4Simple, the MaxIS-3 algorithm, and the algorithm of Theorem 2. That is, for chosen parameters $a,b$, we should have
$\mu\log(2.3146)=2(k-\mu)\log(1.083506)=a\mu+bk$
This allows us to determine $b$ in terms of $a$. Then our goal is to minimize $a$ whilst respecting the branching constraints of Eq. (4). This same reasoning will be used for parameters in all our algorithms. For example, we can define algorithms for branching on degree 5 and degree 6 vertices.
Simplify $G$ If $\operatorname{maxdeg}(G)\leq 4$, then run either the MaxIS-4 algorithm or Branch4Simple Otherwise, apply Section 5 to an arbitrary vertex of degree at least $5$.
Algorithm 2 Function Branch5Simple$(G,k)$
Simplify $G$ If $\operatorname{maxdeg}(G)\leq 5$, then run either the MaxIS-5 algorithm or Branch5Simple Otherwise, apply Section 5 to an arbitrary vertex of degree at least $6$.
Algorithm 3 Function Branch6Simple$(G,k)$
Along similar lines, we immediately obtain the results:
Proposition 28.
Branch5Simple has measure $a\mu+bk$ for $a=0.44849,b=0.085297$.
Branch6Simple has measure $a\mu+bk$ for $a=0.20199,b=0.160637$.
Combined with the MaxIS-5 algorithm using Lemma 10, Branch5Simple immediately gives an algorithm for degree-5 graphs with runtime $O^{*}(1.2491^{k})$. The final algorithm is very similar, but we only keep track of runtime in terms of $k$, not $\mu$ or $n$.
If $\operatorname{maxdeg}(G)\leq 6$, then use Lemma 10 with algorithms of Branch6Simple and MaxIS-6 Otherwise, split on an arbitrary vertex of degree at least $7$.
Algorithm 4 Function Branch7Simple$(G,k)$
Theorem 29.
Algorithm Branch7Simple$(G,k)$ runs in time $O^{*}(1.2575^{k})$.
Proof.
If we split on a vertex of degree at least $7$, then we generate subproblems with vertex covers of size at most $k-1$ and $k-7$. These have cost $1.2575^{k-1}$ and $1.2575^{k-7}$ by induction. Since $11.2575^{-1}+1.2575^{-7}<$ 1, the overall cost is $O^{*}(1.2575^{k})$. Otherwise the two algorithms in question have measure $a\mu+bk$ and $cn$ respectively; where $a=0.20199,b=0.160637,c=\log(1.1893)$; by applying Lemma 10, we get a combined algorithm with cost $O^{*}(1.2575^{k})$ as desired. ∎
7 The advanced branching algorithm: overview
In the next sections, we consider more elaborate graph structures and branching rules. This improves the runtime to $O^{*}(1.25284^{k})$, along with improved algorithms for lower-degree graphs. The outline is similar to Section 6: we develop a series of branching algorithms targeted to graphs with different degrees, building our way up from degree-4 graphs to degree-7 graphs. Here we highlight some of the main changes in our analysis.
Multiway branching. So far, we have only analyzed two-way branching rules. The advanced algorithms will branch on multiple vertices at a time. Specifically, we say that we split on a sequence of vertices $x_{1},\dots,x_{\ell}$ if we first split on $x_{1}$; in the subproblem $\langle G-x_{1},k-1\rangle$ we split on $x_{2}$; in the subproblem $\langle G-\{x_{1},x_{2}\},k-2\rangle$ we split on $x_{3}$; and so on. Assuming that $x_{1}\dots,x_{\ell}$ are independent, this generates $\ell+1$ subproblems $\langle G-N[x_{1}],k-\deg(x_{1})\rangle,\langle G-x_{1}-N[x_{2}],k-1-\deg(x_{2})\rangle,\langle G-\{x_{1},x_{2}\}-N[x_{3}],k-2-\deg(x_{3})\rangle,\dots,\langle G-\{x_{1},\dots,x_{\ell-1}\}-N[x_{\ell}],k-(\ell-1)-\deg(x_{\ell})\rangle,\langle G-\{x_{1},\dots,x_{\ell}\},k-\ell\rangle$. We emphasize that the subproblem $G-N[x_{1}]$ does not have any further branching, such as guessing the status of $x_{2}$. Splitting on $x_{1},x_{2}$ is just a 3-way branch.
Miscellaneous branching rules. We occasionally use some other branching rules; some of these can be applied to graphs which are not simplified. The following rules are used repeatedly:
(A1): Branch on 2-vertex. For a 2-vertex $v$ with neighbors $x$, $y$, Lemma 5 implies there is either a good cover omitting $v$, or omitting both $x,y$. Thus, we can branch on subproblems $\langle G-N[v],k-|N(v)|\rangle$ and $\langle G-N[x,y],k-|N(x,y)|\rangle$.
(A2): Branch on a funnel. For a funnel $u$ with out-neighbor $x$ and any other neighbor $v$ of $u$, we branch on whether $v\in C$ or whether $\{v,x\}\cap C=\emptyset$; this generates subproblems $\langle G-v,k-1\rangle$ and $\langle G-N[v,x],k-|N(v,x)|\rangle$.
(A3): Branch on shared neighbors. For vertices $u,v$ branch on whether $\{u,v\}\subseteq C$, or $N(u)\cap N(v)\subseteq C$; this generates subproblems $\langle G-\{u,v\},k-2\rangle$ and $\langle G-N(u)\cap N(v),k-\operatorname{codeg}(u,v)\rangle$.
We use a few specialized branching rules in various places; these will be discussed as needed.
Creation of high-degree vertices and surplus-two indsets. One common tactic in our analysis is to branch in a targeted way to allow us the use a particularly profitable “follow-up” branching rule in the resulting graph. Two of the most common situations are to use preprocessing rules P2 or P3 in order to create new high-degree vertices, or branching to create a large surplus-two indset for which we can use brancing rules in Section 5.1.
We say graph $G$ is $r$-vulnerable (or $r$-vul for short) if either (i) $S(G)\geq 1$ or (ii) $G$ has an indset with surplus at most two and size at least $r$.
Proposition 30.
Suppose that $G$ has one of the following two structures:
(i) a 2-vertex $u$ with neighbors $x,y$ or
(ii) a funnel $u$ with out-neighbor $x$ and a vertex $y\in N(u)\setminus N(x)$.
In either case, let $r=\deg(x)+\deg(y)-\operatorname{codeg}(x,y)-1$.
If $r\geq 5$, then $G$ has available branch-seq $[(0,2)]$, $[(1,4),(1,6)]$ or $[(0.5,2),(2,6)]$. If $r\geq 6$ then $G$ has available branch-seq $[(0,2)]$, $[(1,4),(1,6)]$, $[(0.5,3),(2,6)]$, or $[(0.5,2),(2.5,r+1)]$.
Proof.
In case (i), we apply (P2) to merge $x,y$ into a new $r$-vertex $z$. In case (ii), we apply (P3); if $\operatorname{codeg}(u,x)\geq 0$ this directly gives drop $[(0,2)]$. Otherwise, it produces an $r$-vertex $y$. If the resulting graph $G^{\prime}$ can be simplified further (aside from isolated vertices), then we get net drop $(0,2)$. Otherwise, we apply Section 5 to $G^{\prime}$. ∎
Proposition 31.
Suppose $\operatorname{mindeg}(G)\geq 1$.
If $G$ is $2$-vulnerable it has available branch-seq $[(0,1)]$, $[(1,4),(1,4)]$, or $[(0.5,3),(2,5)]$.
If $G$ is $3$-vulnerable it has available branch-seq $[(0,1)]$, $[(1,4),(1,5)]$, or $[(0.5,4),(2,5)]$.
If $G$ is $r$-vulnerable for $r\geq 4$, it has available branch-seq $[(0,1)]$ or $[(1,r),(1,r+2)]$.
Proof.
If $G$ is not simplified, then $S(G)\geq 1$ giving branch-seq $(0,1)$. Otherwise, let $I$ be a surplus-two min-set of size $r$. If $r\geq 4$ then splitting on $I$ gives subproblems $G-I,G-N[I]$ with drops $(1,r)$ and $(1,r+2)$ respectively. If $r=2$ or $r=3$, we apply Lemma 22 or Lemma 21 respectively. ∎
For Propositions 30 and 31, we use the following notation for $\operatorname{val}_{a,b}(B)$ of the given branch-seq $B$.
[…]
Notational convention: omitting the value $k$. For the remainder of this paper, when analyzing branching subproblems, we will usually omit the solution size $k^{\prime}$. It should be determined from context, from the following mnemonic: for a subproblem $G-\{x_{1},\dots,x_{\ell}\}-N[y]$, we take $k^{\prime}=k-\ell-\deg(y)$. In using this notation, we must be careful to ensure that $\{x_{1},\dots,x_{\ell}\}\not\sim y$, as otherwise some $x_{i}$ may be double-counted.
8 Branching on 4-vertices
We now develop an improved algorithm Branch4. At a high level, the improvement over the baseline algorithm Branch4Simple comes from tracking the presence of 3-vertices. If a 4-vertex $u$ has a 3-neighbor $v$, then we can split on $u$ and then, in the branch $G-u$, the degree of $v$ gets reduced to two. We then use (P2) to reduce $k$ by one – a “bonus” reduction. We can also ensure that this creates new 3-vertices; thus, the process is a self-sustaining chain reaction.
We sketch an outline of the algorithm here; many additional details will be provided later.
Apply branching rules to ensure $G$ has nice structure, e.g. simplified and no 4-cycles.                  (We will call this 4-base-position (4BP) later.) If $V_{3}\neq\emptyset$ and $V_{3}\not\sim V_{4}$, then solve vertex cover exactly on $V_{3}$. Split on a 4-vertex with a 3-neighbor, chosen according to tie-breaking rules to be described later. Branch on a triangle. Branch on a vertex $u$ according to tie-breaking rule to be described later.
Algorithm 5 Function Branch4$(G,k)$
Theorem 32.
Algorithm Branch4 has measure $a_{4}\mu+b_{4}k$ for $a_{4}=0.59394,b_{4}=0.039361$.
Via Lemma 10, in combination with the MaxIS-4 algorithm, this will give our algorithm with runtime $O^{*}(1.21131^{k})$ for degree-4 graphs. Theorem 32 is very length to prove. One major complication is that it requires us to use a measure for $G$ tracking additional information beyond just the values $\mu$ and $k$. Specifically, our strategy is to show that Branch4 has the measure of the form
$\phi(G)=a\mu+bk-\operatorname{bonus}(G);$
where we define function $\operatorname{bonus}(G)$ as follows. First, we define a special 4-vertex to be a 4-vertex with at least two 3-neighbors, and we say $G$ is special if it has a special 4-vertex. We then define:
$\operatorname{bonus}(G)$ =$=\begin{cases}\beta&\text{if $\operatorname{mindeg}(G)=3$ and $\operatorname{maxdeg}(G)=4$ and $G$ is special}\\ \alpha&\text{if $G$ has a subcubic vertex and the previous case does not hold}\\ 0&\text{otherwise}\end{cases}$
for parameters $\alpha=0.03894,\beta=0.05478$
Note that $\operatorname{bonus}(G)$ is bounded by a constant and so this in turn will show Theorem 32. Although the precise value of the parameters is not critical, we often use the important inequalities
$0\leq\alpha\leq b\leq\beta\leq 2b$ (6)
Motivated by Eq. (6), we write $S(G)\geq\ell^{*}$ if either (i) $S(G)\geq\ell+1$, or (ii) $S(G)=\ell$ and there is a simplified graph $\langle G^{\prime},k-\ell\rangle$ with a 3-vertex. We also write $R(X)\geq\ell^{*}$ if $S(G-X)\geq\ell^{*}$, etc.
Similarly, if we transform a graph $\langle G,k\rangle$ into a graph $\langle G^{\prime},k^{\prime}\rangle$, we say $G^{\prime}$ has “extended drop” $(\Delta\mu,\Delta k^{*})$ if $\mu^{\prime}\leq\mu-\Delta\mu$ and either (i) $G^{\prime}$ has a subcubic vertex and $k^{\prime}=k-\Delta k$ or (ii) $k^{\prime}<k-\Delta k$ strictly. To explain, note that since $\alpha\leq b$, we have $\phi(G^{\prime})\leq\phi(G)-a\Delta\mu-b\Delta k-\alpha+\operatorname{bonus}(G)$. So the contribution of $G^{\prime}$ to the sum in Eq. (3) is at most $e^{-a\Delta\mu-b\Delta k-\alpha}$.
8.1 Some observations about simplification and degree-3 vertices
We record a few general observations about simplification in graphs.
Proposition 33.
Let $s\geq 1$. If $G$ has at least $2(s+1)$ subcubic vertices, and it has a critical-set $I$ with $\operatorname{surp}(I)=1,|I|\geq s$, then $S(G)\geq s^{*}$.
Proof.
We apply (P2) to $I$. If $|I|>s$, this gives $S(G)\geq s+1$ directly. Otherwise, all subcubic vertices outside $N[I]$ remain subcubic, leaving at least $12(s+1)-|N[I]|=$ 1 subcubic vertices.
Proposition 34.
Suppose $G$ has no 4-cycles, has at least 7 subcubic vertices, and has two non-adjacent 2-vertices. Then $S(G)\geq 2^{*}$.
Proof.
Let $u,v$ be non-adjacent 2-vertices. If we apply (P2) to $v$, then all the subcubic vertices in $G$ outside from $N[v]$ remain subcubic in the resulting graph $G^{\prime}$. Since $G$ has no 4-cycle, we have $\deg_{G^{\prime}}(u)=2$. So $G^{\prime}$ has at least 4 subcubic vertices and has a 2-vertex $u$. By Proposition 33, this implies $S(G^{\prime})\geq 1^{*}$ and hence $S(G)\geq 2^{*}$. ∎
Lemma 35.
Let $X$ be a vertex set. If $\operatorname{maxdeg}(G)\leq 4$ and $G$ has no components of size at most $8|X|$, then $G-X$ has a subcubic vertex.
Proof.
There must be some vertex $z$ at distance two from $X$, since otherwise the component(s) containing $X$ would have size at most $(4+\operatorname{maxdeg}(G))|X|\leq 8|X|$. So suppose $z$ has a neighbor $y\sim X$. Here $\deg_{G-X}(y)\geq 1$ (due to neighbor $z$) and also $\deg_{G-X}(y)\leq 3$ (it loses some neighbor in $X$). Hence, $G-X$ has a subcubic vertex. ∎
Proposition 36.
Let $r\geq 2$ be an arbitrary constant. Suppose that $G$ has at least 4 subcubic vertices and $\operatorname{maxdeg}(G)\leq 4$ and $\operatorname{minsurp}(G)\geq 1$ and $G$ has an indset $I$ with $|I|\geq r$ and $\operatorname{surp}_{G}(I)\leq 2$. Then $G$ has available branch-seq $[(0,1^{*})]$ or $[(1,r^{*}),(1,(r+2)^{*})]$.
Proof.
If $G$ has any components of size less than $8(r+2)$, we solve them exactly in $O(1)$ time and remove them from $G$; since $\operatorname{minsurp}(G)\geq 1$, this reduces $k$ by at least two and hence gives drop $[(0,2)]$. If $\operatorname{surp}_{G}(I)=1$, then Proposition 33 with $s=1$ gives $S(G)\geq 1^{*}$. Otherwise, we split on min-set $I$; if $|I|>r$ strictly, the subproblem $G-I$ and $G-N[I]$ have drop $(1,(r+1))$ and $(1,(r+3))$ directly. If $|I|=r$, these subproblems have drops $(1,r)$ and $(1,r+2)$ and, by Lemma 35, both of them have subcubic vertices.
We say that $G$ is $r^{*}$-vul if either $S(G)\geq 1^{*}$ or the preconditions of Proposition 36 hold, and we correspondingly define
$\gamma^{*}_{r}=\max\{e^{-b-\alpha},e^{-a-rb-\alpha}+e^{-a-(r+2)b-\alpha}\}$
8.2 Regularizing the graph and basic analysis
We will first ensure that $G$ has nice properties, before the “heavy-duty” branching rules. Specifically, we say $G$ is in 4-base position (or 4BP for short) if the following conditions hold:
  • $G$ is simplified.
  • $\operatorname{maxdeg}(G)\leq 4$
  • $\operatorname{shad}(N[u])\geq 1$ for all vertices $u$.
  • $G$ has no 4-cycle
  • $G$ has no connected-components of size larger than some constant $K$.
Note: Let us explain the last clause in the definition of 4BP, and the choice of constant $K$ (which we will not specify explicitly). In light of Lemma 35, our branching algorithms will generate principal subproblems $G-(U_{0}\cup U_{1})$ where $\Delta k=|U_{1}|$ is lower-bounded by some constant value $c$. Such subproblems may be produced directly (e.g. by guessing a vertex is in the cover), or may be obtained by applying (P1) to a principal subproblem $G-X$. If the constant $K$ is large enough (in particular, larger than $8c$), then any such principal subproblem $G-(U_{0}\cup U_{1})$ with drop $(\Delta\mu,c)$ automatically has drop $(\Delta\mu,c^{*})$ “for free”. For example, this principle was used in Proposition 36; we will use it implicitly through the remainder of the algorithm analysis.
Proposition 37.
If $\operatorname{maxdeg}(G)\leq 4$ and $\operatorname{minsurp}(G)\geq 2$, but $G$ is not in 4BP, then it has available branch-seq $[(0,2)]$, $[(1,3),(1,5)]$, $[(0.5,2),(2,5)]$, or $[(1,3^{*}),(1,3^{*})]$
Proof.
If $G$ has a connected-component $X$ of constant size, we may solve vertex cover on it exactly in constant time, and update $k^{\prime}=k-|C_{X}|$ where $C_{X}$ is a good cover of $X$. Since $\operatorname{minsurp}(G)\geq 2$, we have $|C_{X}|\geq 2$, and so removing this component gives drop $[(0,2)]$.
If $G$ has a funnel $u$ with out-neighbor $x$, then we branch on subproblems $G-N[u],G-N[x]$; these both have drop $(1,3^{*})$ by Lemma 35. Since $\operatorname{minsurp}(G)\geq 2$, rules (P1) and (P2) cannot be applied; so for the remainder of the analysis we may suppose that $G$ is simplified.
If $G$ has a blocked 4-vertex, then we apply Proposition 25.
Finally, if $G$ has a 4-cycle $u_{1},u_{2},u_{3},u_{4}$, we apply (A3) to get subproblems $G-\{u_{1},u_{3}\}$ and $G-N(u_{1},u_{3})$. We claim these subproblems both have drops $(1,3^{*})$ after simplification. Let us first show this holds for subproblem $G^{\prime}=G-\{u_{1},u_{3}\}$, which has drop $(1,2)$ directly. If $u_{2}\not\sim u_{4}$, then Observation 16 immediately gives $S(G^{\prime})\geq 2$. So suppose $u_{2}\sim u_{4}$; since $G$ is simplified, this is only possible if $u_{1},u_{2},u_{3},u_{4}$ all have degree 4 (else they would have a 3-triangle). Then $u_{1}\not\sim u_{3}$ as otherwise $u_{1}$ would have a funnel with its neighbors $u_{2},u_{3},u_{4}$. So $G^{\prime}$ has at least 4 subcubic vertices (the neighbors of $u_{1}$) and it has a 2-vertex $u_{2}$, so Proposition 33 with $s=1$ gives $S(G^{\prime})\geq 1^{*}$.
We next argue that $G^{\prime\prime}=G-N(u_{1},u_{3})$ has drop $(1,3^{*})$. If $\operatorname{codeg}(u_{1},u_{3})\geq 3$, then $G^{\prime\prime}$ has drop $(1,3^{*})$ directly by Lemma 35. Otherwise, if $\operatorname{codeg}(u_{1},u_{3})=2$, then $G^{\prime\prime}=G-\{u_{2},u_{4}\}$; this has drop $(1,3^{*})$ by precisely the same reasoning as for the graph $G^{\prime}=G-\{u_{1},u_{3}\}$. ∎
Proposition 38.
If $G$ is in 4BP and $u$ is a 4-vertex, then $G-N[u]$ has at least 8 subcubic vertices.
Proof.
Suppose $u$ has 3-neighbors $v_{1},\dots,v_{s}$ and 4-neighbors $v_{s+1},\dots,v_{s}$ We have $\operatorname{codeg}(u,v_{i})=0$ for $i\leq s$, since otherwise $v_{i}$ would have a 3-triangle. Likewise, we have $\operatorname{codeg}(u,v_{i})\leq 1$ for $i>s$, as otherwise $G$ would have a 4-cycle. Thus, the neighborhood of $u$ has at most $\lfloor(4-s)/2\rfloor$ edges, and hence at least $3s+4(4-s)-2\lfloor(4-s)/2\rfloor-4\geq 8$ edges from $N(u)$ to vertices $t$ at distance two from $u$. Since $G$ has no 4-cycles, each such vertex $t$ has an edge to precisely one vertex in $N(u)$. ∎
We now analyze the easier steps of the algorithm line-by-line. Note that, starting at Line 4, the graph $G$ is 4-regular. Lines 3 and 5, which are more involved, will be considered later.
Line 1. We may remove any isolated vertices, as this can only increase $\operatorname{bonus}(G)$.
First suppose that $\operatorname{bonus}(G)=0$. If $G$ is not simplified, then rules (P1)–(P3) give drop $[(0,1)]$. If $\operatorname{maxdeg}(G)\geq 5$, then we apply Section 5; otherwise, we apply Proposition 37. We need:
$\max\{e^{-b},e^{-0.5a-b}+e^{-2a-5b},e^{-a-3b-\alpha}+e^{-a-3b-\alpha},e^{-a-3b}+e^{-a-5b}\}$ $\leq 1$ (4-1)
Next suppose $\operatorname{bonus}(G)=\beta$. Since $\operatorname{mindeg}(G)\geq 3$, applying (P1) or (P2) reduces $k$ by at least two. Otherwise, if $\operatorname{minsurp}(G)\geq 2$, then we apply Proposition 37. Overall, we need:
$\max\{e^{-2b},e^{-0.5a-2b}+e^{-2a-5b},e^{-a-3b-\alpha}+e^{-a-3b-\alpha},e^{-a-3b}+e^{-a-5b}\}$ $\leq e^{-\beta}$ (4-2)
Finally suppose $\operatorname{bonus}(G)=\alpha$. If $G$ is not simplified, then rules (P1)–(P3) give drop $[(0,1)]$. If $\operatorname{maxdeg}(G)\leq 4$, then we can apply Proposition 37. Finally, suppose $G$ is simplified and has a vertex $u$ of degree at least $5$. If $\operatorname{shad}(N[u])\leq 4-\deg(u)$, then we apply Lemma 26. If $\operatorname{shad}(N[u])\geq 5-\deg(u)$, then we split on $u$; subproblem $G-N[u]$ has drop $(2,5)$, while subproblem $G-u$ retains the subcubic vertex in the original graph and so has drop $(0.5,1^{*})$. Over all cases, we need:
$\max\{e^{-b},e^{-0.5a-b-\alpha}+e^{-2a-5b},e^{-a-3b-\alpha}+e^{-a-3b-\alpha},e^{-a-3b}+e^{-a-5b}\}$ $\leq e^{-\alpha}$ (4-3)
Line 2. We run either the MaxIS-3 algorithm or Theorem 2 on $V_{3}$. Let $k_{3}$ be the size of the minimum vertex cover for $V_{3}$ and let $\mu_{3}=k_{3}-n_{3}/2$. Removing $V_{3}$ reduces $\phi(G)$ by at least $bk_{3}-\beta\geq 2b-\beta\geq 0$. Running the algorithms on $V_{3}$ has cost $O^{*}(2.3146^{\mu_{3}})$ or $O^{*}(1.085306^{n_{3}})$ respectively. Since $n_{3}\leq n=2(k-\mu)$ and $\mu_{3}\leq\mu$, it suffices to show:
$\min\{\mu\log(2.3146),2(k-\mu)\log(1.085306)\}\leq a\mu+bk$ (4-4)
Line 4. Let $N(u)=\{v_{1},\dots,v_{4}\}$ where $v_{1}\sim v_{2}$. Since $G$ has no 4-cycles, $v_{1}\not\sim\{v_{3},v_{4}\}$. We split on $v_{3}$, getting subproblems $G-N[v_{3}],G-v_{3}$ with drops $(1.5,4^{*})$ and $(0.5,1^{*})$. In the latter subproblem, we apply Proposition 30 to the funnel $u$ with out-neighbor $v_{4}$ and vertex $v_{1}\in N(u)\setminus N(v_{4})$, where $r=\deg_{G-v_{3}}(v_{1})+\deg_{G-v_{3}}(v_{4})-\operatorname{codeg}_{G-v_{3}}(v_{1},v_{4})-1\geq 4+3-1-1=5$. We need:
$e^{-1.5a-4b-\alpha}+e^{-0.5a-b}\psi_{5}\leq 1$ (4-5)
8.3 Line 3: Branching on a 4-vertex with a 3-neighbor
Suppose Line 3 selects a 4-vertex $u$, with 3-neighbors $v_{1},\dots,v_{s}$ and $4$-neighbors $v_{s+1},\dots,v_{4}$. Let $x_{1},x_{2}$ be the other two neighbors of $v_{1}$, sorted so $\deg(x_{1})\geq\deg(x_{2})$. Since $G$ has no 4-cycles, and the 3-vertices $v_{1},\dots,v_{s}$ have no triangles, all vertices $u,v_{i},x_{i}$ are distinct, and $\{x_{1},x_{2}\}\not\sim\{u,v_{2},v_{3},v_{4}\}$. Note that there may be multiple ways of ordering vertices in this way; e.g., we can permute the vertices $v_{1},\dots,v_{s}$ or $v_{s+1},\dots,v_{4}$, or if if $\deg(x_{1})=\deg(x_{2})$ we can swap $x_{1}/x_{2}$.
Let us fix some ordering and consider the graphs $G_{0}=G-N[u]$ and $G_{1}=G-u$ produced by splitting on $u$. There are two strategies we can use for further branching. First, we can apply (P2) to the graph $G_{1}$, contracting its 2-vertices $v_{1},\dots,v_{s}$ into vertices $z_{1},\dots,z_{s}$; we let $G_{1}^{\prime}$ denote the resulting graph, and in this case we overall generate two subproblems $G_{0},G_{1}^{\prime}$. Second, we can apply (A1) to the 2-vertex $v_{1}$ in $G-u$, generating further subproblems $G_{2}=G-N[v_{1}],G_{3}=G-v_{1}-N[x_{1},x_{2}]$; so in this case we overall generate three subproblems $G_{0},G_{2},G_{3}$.
Throughout, we write $r=\deg(x_{1})+\deg(x_{2})-2=\deg_{G_{1}^{\prime}}(z_{1})$.
Proposition 39.
If $s\geq 2$, then $S(G_{2})\geq 1^{*}$. If $r=4$, then $S(G_{0})\geq 2^{*}$. If $r=5$, then $S(G_{0})\geq 1^{*}$.
Proof.
For the first result, if $s\leq 3$, then $G-N[v_{1}]$ has a 3-vertex $v_{4}$ and a 2-vertex $v_{2}\not\sim v_{4}$; applying (P2) to $v_{2}$ reduces $k$ by one, and still leaves a subcubic vertex $v_{4}$. If $s=4$, then $G-N[v_{1}]$ has non-adjacent 2-vertices $v_{2},v_{3}$, so $S(G_{2})\geq 2$ by Observation 11.
For the next two results, recall from Proposition 38 that $G_{0}$ has at least 8 subcubic vertices. If $r=4$, then $\deg_{G_{0}}(x_{1})=\deg_{G_{0}}(x_{2})=2$ and $x_{1}\not\sim x_{2}$, so by Proposition 34 we have $S(G_{0})\geq 2^{*}$. Similarly, if $r=5$, then $\deg_{G_{0}}(x_{2})=2$ so by Proposition 33 we have $S(G_{0})\geq 1^{*}$. ∎
Proposition 40.
Suppose $r\geq 5$. Then $G_{3}$ is a principal subproblem with excess two and $\Delta k=r+2$ and $\operatorname{minsurp}(G_{3})\geq 4-r$. Furthermore, the following bounds all hold:
  • If $r=5$ and $\operatorname{minsurp}(G_{3})=-1$, then $R(N[x_{1}])\geq 2^{*}$.
  • If $r=5$ and $\operatorname{minsurp}(G_{3})=-1$ and $s=1$, then $R(N[x_{1}])\geq 3^{*}$.
  • If $r=6$ and $\operatorname{minsurp}(G_{3})=-2$, then $R(N[x_{1}])\geq 3^{*}$.
  • If $r=6$ and $\operatorname{minsurp}(G_{3})=-1$, then $G-N[x_{1}]$ is $3^{*}$-vul.
Proof.
Let $H=G-u-N[x_{1},x_{2}]$. Since $G$ has no 4-cycles, we have $|N[x_{1},x_{2}]|=r+1$ and so $H$ is principal with excess two and $\Delta k=r+2$. Since $\deg(x_{1})\geq\deg(x_{2})$, we have $\deg(x_{1})=4$ and $\deg(x_{2})=r-2$. Note that if $s=1$, then the sole 3-neighbor of $u$ gets removed from $H$, and so $\operatorname{mindeg}(H)\geq 1$. Also note that, by Proposition 38, the graph $G-N[x_{1}]$ has at least 8 subcubic vertices, and since $G$ is in 4BP we have $\operatorname{shad}(N[x_{1}])\geq 1$.
Let $I$ be a min-set in $H$ and $\ell=\operatorname{surp}_{H}(I)$. If $\ell\leq-2$ or $s=1$ then necessarily $|I|\geq 2$. Then $\operatorname{surp}_{G-N[x_{1}]}(I\cup\{x_{2}\})\leq\ell+(\deg_{G}(x_{2})-1)=\ell+r-3$. Since $G$ is in 4BP, we have $\operatorname{shad}(N[x_{1}])\geq 1$ so $\ell\geq 4-r$. Furthermore, if $\ell=4-r$ exactly, then $\operatorname{surp}_{G-N[x_{1}]}(I\cup\{x_{2}\})=1$, and then Proposition 33 applied to indset $I\cup\{x_{2}\}$ gives $R(N[x_{1}])\geq 2^{*}$ for $|I|=1$ and $R(N[x_{1}])\geq 3^{*}$ for $|I|\geq 2$.
Finally, suppose $r=6$ and $\ell=-1$. If $|I|=1$, then $H$ has an isolated vertex $t$; since $G$ has no 4-cycles, this is only possible if $\deg_{G}(t)=3$ and $t\sim u$ and $\operatorname{codeg}(t,x_{1})=\operatorname{codeg}(t,x_{2})=1$. In this case, $\deg_{G-N[x_{1}]}(t)=2$ and by Proposition 33 we have $R(N[x_{1}])\geq 1^{*}$. Otherwise, we have $\operatorname{surp}_{G-N[x_{1}]}(I\cup\{x_{2}\})\leq 2$, which by Proposition 36 implies that $G-N[x_{1}]$ is $3^{*}$-vul. ∎
There are a number of cases, which implicitly define our tie-breaking rule for $u$: when considering a given case, we assume that no other 4-vertex in the graph and no other ordering of its neighbors would be covered by any of the previous cases.
Case I: $\boldsymbol{s\geq 2}$ and $\boldsymbol{S(G_{0})\geq 2^{*}}$. We generate subproblems $G_{0},G_{1}^{\prime}$. If $s\in\{2,3\}$, then $G_{1}^{\prime}$ has drop $(0.5,3)$ or $(0.5,4)$ and also has a 3-vertex $v_{4}$; if $s=4$, then $G_{1}^{\prime}$ has drop $(0.5,5)$ directly. In any case, $G_{1}^{\prime}$ has drop $(0.5,3^{*})$. We then simply need:
$e^{-1.5a-6b-\alpha}+e^{-0.5a-3b-\alpha}\leq e^{-\beta}$ (4-6)
Case II: $\boldsymbol{G_{0}}$ is $\boldsymbol{3^{*}}$-vul and $\boldsymbol{S(G_{2})\geq 1^{*}}$ and $\boldsymbol{r\geq 5}$ and $\boldsymbol{\operatorname{minsurp}(G_{3})\geq 5-r}$. We generate subproblems $G_{0},G_{2},G_{3}$, where $G_{2}$ has drop $(1,4^{*})$ after simplification, and $G_{3}$ has drop $(2.5,7^{*})$. We need:
$e^{-1.5a-4b}\gamma_{3}^{*}+e^{-a-4b-\alpha}+e^{-2.5a-7b-\alpha}\leq e^{-\beta}$ (4-7)
Case III: $\boldsymbol{s\geq 2}$. By Proposition 39, we have $S(G_{2})\geq 1^{*}$; also, we must have $r\geq 5$, since if $r=4$ then $S(G_{0})\geq 2^{*}$ and it would covered in Case I. In particular, $\deg(x_{1})=4$.
We claim that $G-N[x_{1}]$ is not $3^{*}$-vul. For, suppose it were, and consider $u^{\prime}=x_{1}$ and its 3-neighbor $v_{1}^{\prime}=v_{1}$ and its two neighbors $x_{1}^{\prime}=u,x_{2}^{\prime}=x_{2}$. We have $r^{\prime}=\deg(x_{1}^{\prime})+\deg(x_{2}^{\prime})-2=4+\deg(x_{2})=r$. If $\operatorname{shad}(u^{\prime},N[x_{1}^{\prime},x_{2}^{\prime}])\leq 4-r^{\prime}$, then Proposition 40 applied to $u^{\prime}$ would give $R(N[x_{1}^{\prime}])=S(G_{0})\geq 2^{*}$ so $u$ would be covered in Case I. On the other hand, if $\operatorname{shad}(u^{\prime},N[x_{1}^{\prime},x_{2}^{\prime}])=\operatorname{minsurp}(G_{3}^{\prime})\geq 5-r^{\prime}$, then vertex $u^{\prime}$, with $v_{1}^{\prime},x_{1}^{\prime},x_{2}^{\prime}$, would be covered by Case II.
We claim next that $r=6$. For, suppose $r=5$, so $S(G_{0})\geq 1^{*}$. If $\operatorname{minsurp}(G_{3})\geq 0$, then we would be in Case II. On the other hand, if $\operatorname{minsurp}(G_{3})=-1$, then Proposition 40 gives $R(N[x_{1}])\geq 2^{*}$ and in particular $G-N[x_{1}]$ would be $3^{*}$-vul.
Now our branching rule is to generate subproblems $G_{0},G_{2},G_{3}$. By Proposition 40, we have $\operatorname{minsurp}(G_{3})\geq 0$ since $G-N[x_{1}]$ is not $3^{*}$-vul. So $G_{2}$ has drop $(3,8^{*})$ and we need:
$e^{-1.5a-4b-\alpha}+e^{-a-4b-\alpha}+e^{-3a-8b-\alpha}\leq e^{-\beta}$ (4-8)
Note: henceforth, we know the graph has no special 4-vertex (it would be covered in Case III or earlier). In particular, we have $s=1$ and $\operatorname{bonus}(G)=\alpha$.
Case IV: $\boldsymbol{S(G_{0})\geq 3^{*}}$. Here, we generate subproblems $G_{0},G_{1}^{\prime}$. Since $s=1$, the graph $G_{1}^{\prime}$ has a subcubic vertex $v_{2}$ and so it has drop $(0.5,2^{*})$. We need:
$e^{-1.5a-7b-\alpha}+e^{-0.5a-2b-\alpha}\leq e^{-\alpha}$ (4-9)
Case V: $\boldsymbol{r\geq 5}$ and $\boldsymbol{G_{0}}$ is $\boldsymbol{3^{*}}$-vul. Here we generate subproblems $G_{0},G_{2},G_{3}$, where $G_{2}$ has drop $(1,3^{*})$. We have $\operatorname{minsurp}(G_{3})\geq 5-r$, as otherwise Proposition 40 would give $R(N[x_{1}])\geq 3^{*}$ (bearing in mind that $s=1$); then $u^{\prime}=x_{1}$ would be covered in Case IV. So subproblem $G_{3}$ has drop $(2.5,7^{*})$. Overall, we need:
$e^{-1.5a-4b}\gamma_{3}^{*}+e^{-a-3b-\alpha}+e^{-2.5a-7b-\alpha}\leq e^{-\alpha}$ (4-10)
Case VI: $\boldsymbol{r=6}$. Again we generate subproblems $G_{0},G_{2},G_{3}$, where $G_{2}$ has drop $(1,3^{*})$ as before. If $G-N[x_{1}]$ is $3^{*}$-vul, then vertex $u^{\prime}=x_{1}$ would be covered in Case V earlier (it has a 3-neighbor $v_{1}$ with a 4-neighbor $u$). So assume $G-N[x_{1}]$ is not $3^{*}$-vul; by Proposition 40, this implies $\operatorname{minsurp}(G_{3})\geq 0$ and so $G_{3}$ has drop $(3,8^{*})$. We need:
$e^{-1.5a-4b-\alpha}+e^{-a-3b-\alpha}+e^{-3a-8b-\alpha}\leq e^{-\alpha}$ (4-11)
Case VII: $\boldsymbol{r=4}$. Here, $G$ cannot have any 3-vertex $v$ with two 4-neighbors $y_{1},y_{2}$, as then vertex $u^{\prime}=y_{1}$ would have $r^{\prime}\geq 5$ and be covered in previous cases. So we assume that all 3-vertices in the graph have at least two 3-neighbors. In particular, we have $x_{1}\sim t_{1}$ and $x_{2}\sim t_{2}$ for 3-vertices $t_{1},t_{2}$ distinct from $v_{1}$. Then in $G_{1}^{\prime}$, the merged vertex $z_{1}$ has degree 4 and has 3-neighbors $t_{1},t_{2}$, that is, it is a special vertex (see Figure 6). Since $\operatorname{mindeg}(G_{1}^{\prime})\geq 3$ and $\operatorname{maxdeg}(G_{1}^{\prime})\leq 4$, we get $\operatorname{bonus}(G_{1}^{\prime})=\beta$. Also, $S(G_{0})\geq 2^{*}$ since $r=4$. We need:
$e^{-1.5a-6b-\alpha}+e^{-0.5a-2b-\beta}\leq e^{-\alpha}$ (4-12)
Figure 6: We show the degrees of each vertex. In $G-u$, the 2-vertex $v_{1}$ has neighbors $x_{1},x_{2}$, which get merged into $z_{1}$. The new vertex has 4 neighbors in total and at least two 3-neighbors.
8.4 Line 5: Branching on a vertex in a 4-regular graph
We select a vertex $u$ and an ordering of its neighbors $v_{1},v_{2},v_{3},v_{4}$; our general strategy is to split on $v_{1},v_{2}$. The graph $H=G-\{v_{1},v_{2}\}-N[v_{3},v_{4}]$ plays a central role in the analysis. We let $I$ be a min-set of $H$; for brevity, we define $A=\{v_{1},v_{2}\}$ in this section.
Proposition 41.
Suppose $G$ has a 6-cycle $x_{1},\dots,x_{6}$. Then $R(x_{1},N[x_{4}])\geq 2^{*}$; furthermore, either $\operatorname{codeg}(x_{1},x_{4})=1$ or $x_{4}$ is a special 4-vertex in $G-N[x_{1}]$.
Proof.
At this point, since $\operatorname{girth}(G)\geq 5$ and $G$ is 4-regular, each vertex $v$ has twelve vertices at distance two which all have degree 3 in $G-N[v]$. The graph $G^{\prime}=G-x_{1}-N[x_{4}]$ has two 2-vertices $x_{2},x_{6}$. We have $x_{2}\not\sim x_{6}$, as otherwise would have a triangle $x_{1},x_{2},x_{6}$. Since $G^{\prime}$ has at least 11 subcubic vertices, Proposition 34 gives $S(G^{\prime})\geq 2^{*}$.
Now suppose $\operatorname{codeg}(x_{1},x_{4})=0$. Then the graph $G-N[x_{1}]$ retains vertex $x_{4}$ and its neighbors, while $x_{3},x_{5}$ lose their neighbors $x_{2},x_{6}$ respectively. Since $\operatorname{girth}(G)\geq 5$, we have $x_{3}\neq x_{5}$ and the degrees of $x_{3},x_{5}$ are not reduced by two. So $x_{4}$ has two 3-neighbors $x_{3},x_{5}$ in $G-N[x_{1}]$. ∎
Proposition 42.
We have $\operatorname{mindeg}(H)\geq 1$. Moreover, for any vertex $x\in H$, there holds:
  • If $\deg_{H}(x)=1$, then $x\sim A$ and $R(u,N[x])\geq 2^{*}$.
  • If $\deg_{H}(x)=2$ and $x\not\sim A$, then $x$ is a special 4-vertex in $G-N[u]$.
Proof.
Because $\operatorname{girth}(G)\geq 5$, we have $\operatorname{codeg}(x,v_{3})\leq 1,\operatorname{codeg}(x,v_{4})\leq 1$ and $|N(x)\cap A|\leq 1$. If $\deg_{H}(x)=1$, then necessarily $x$ has a neighbor $v_{i}\in A$ and a neighbor $r_{3}\sim v_{3}$ and a neighbor $r_{4}\sim v_{4}$ and thus $u,v_{3},r_{3},x,r_{4},v_{4}$ is a 6-cycle; by Proposition 41 we have $R(u,N[x])\geq 2^{*}$. Likewise, if $\deg_{H}(x)=2$ and $x\not\sim A$, then $x$ has neighbors $r_{3}\sim v_{3},r_{4}\sim v_{4}$, and then $G$ has a 6-cycle $u,v_{3},r_{3},x,r_{4},v_{4}$. Furthermore, $x\not\sim A$ by hypothesis and $x\not\sim\{v_{3},v_{4}\}$ since $x\in H$. Hence $\operatorname{codeg}(u,x)=0$ and so by Proposition 41 the vertex $x$ is special in $G-N[u]$. ∎
Proposition 43.
If $\operatorname{minsurp}(H)<0$ and $\deg_{H}(x)\leq 2$ for all $x\in I$, then $G-N[u]$ is special.
Proof.
We assume that $x\sim A$ for all $x\in I$, as otherwise the result follows from Proposition 42. Since $\operatorname{mindeg}(H)\geq 1$, and $|N(I)|<|I|$ strictly, there must be some vertex $y\in H$ with two neighbors $x_{1},x_{2}\in I$. The neighbors of $x_{1},x_{2}$ in $A$ must be distinct else $G$ would have a 4-cycle. So suppose without loss of generality $x_{1}\sim v_{1},x_{2}\sim v_{2}$ and $G$ has a 6-cycle $u,v_{1},x_{1},y,x_{2},v_{2}$. Note that $y\not\sim\{v_{3},v_{4}\}$ since $y\in H$ and $y\not\sim\{v_{1},v_{2}\}$ since then $G$ would have a 3-cycle $v_{1},x_{1},y$ or $v_{2},x_{2},y$. So $\operatorname{codeg}(u,y)=0$ and hence $G-N[u]$ is special by Proposition 41. ∎
For our tie-breaking rule, we assume that if we are in a given case that ordering of neighbors $v_{1}^{\prime},\dots,v_{4}^{\prime}$ would fall into a previous case.
Case I: $\boldsymbol{\operatorname{shad}(N[v_{3}])\leq 1}$. Let $J$ be a min-set of $G-N[v_{3}]$; since $\operatorname{mindeg}(G-N[v_{3}])\geq 3$, necessarily $|J|\geq 2$, and since $G$ is in 4BP we have $\operatorname{surp}_{G-N[v_{3}]}(J)=1$. Here, we split on $v_{3}$. By Proposition 38, $G-N[v_{3}]$ has at least 8 subcubic vertices, so $R(N[v_{3}])\geq 2^{*}$ by Proposition 33. Overall, we need:
$e^{-0.5a-b-\alpha}+e^{-1.5a-6b-\alpha}\leq 1$ (4-13)
Case II: $\boldsymbol{\operatorname{minsurp}(H)\leq-2}$. Since $\operatorname{mindeg}(H)\geq 1$, we must have $|I|\geq 3$. Let $J=I\cup\{v_{4}\}$. Here, $\operatorname{surp}_{G-N[v_{3}]}(J)\leq 2$, and $|J|=|I|+1\geq 4$. Since we are not in Case I, we have $\operatorname{shad}(N[v_{3}])\geq 2$ and so $J$ is a min-set in $G-N[v_{3}]$.
Here again we split on $v_{3}$ and then split on $J$ in the subproblem $G-N[v_{3}]$, generating subproblems $G-v_{3},G-N[v_{3}]-J,G-N[v_{3},J]$. Since we are not in Case II, we have $\operatorname{shad}(N[v_{3}],J)\geq\operatorname{shad}(N[v_{3}])-|J|\geq 2-|J|$ and likewise $\operatorname{shad}(N[v_{3},J])\geq\operatorname{shad}(N[v_{3}]-|N[J]|\geq 2-|N[J]|$. So the latter two subproblems have drops $(2.5,8^{*}),(2.5,10^{*})$ respectively. Here we need:
$e^{-0.5a-b-\alpha}+e^{-2.5a-8b-\alpha}+e^{-2.5a-10b-\alpha}\leq 1$ (4-14)
Note: In all subsequent cases, our strategy is to split on $v_{1},v_{2}$, and then apply (A1) to vertex $u$ in subproblem $G-\{v_{1},v_{2}\}$. This gives four subproblems $G_{1}=G-N[u],G_{2}=G-N[v_{1}],G_{3}=G-v_{1}-N[v_{2}],H$. Here, $G_{1},G_{2},G_{3}$ have drop $(1.5,4^{*}),(1.5,4^{*}),(2,5^{*})$ directly.
Case III: $\boldsymbol{\operatorname{minsurp}(H)=0}$. Here $H$ has drop $(3.5,9^{*})$, so we need:
$e^{-1.5a-4b-\alpha}+e^{-1.5a-4b-\alpha}+e^{-2a-5b-\alpha}+e^{-3.5a-9b-\alpha}\leq 1$ (4-15)
Note: For the remainder of the analysis, we assume $\operatorname{minsurp}(H)=-1$, and we let $r=|I|$; we have $r\geq 2$ since $\operatorname{mindeg}(H)\geq 1$. For $r\leq O(1)$, the graph $H$ has drop $(3,(8+r)^{*})$ after applying (P1).
Case IV: $\boldsymbol{S(G_{3})\geq 2^{*}}$. There are a few different subcases depending on $r$. If $r\geq 4$, then $H$ has drop $(3,12^{*})$ and we need:
$e^{-1.5a-4b-\alpha}+e^{-1.5a-4b-\alpha}+e^{-2a-7b-\alpha}+e^{-3a-12b-\alpha}\leq 1$ (4-16)
Next suppose that $r=3$. Clearly $\deg_{H}(x)\leq 2$ for all $x\in I$, so $G-N[u]$ is special by Proposition 43. Since in this case $H$ has drop $(3,11^{*})$, we then need:
$e^{-1.5a-4b-\beta}+e^{-1.5a-4b-\alpha}+e^{-2a-7b-\alpha}+e^{-3a-11b-\alpha}\leq 1$ (4-17)
Finally suppose $r=2$. Again, $G-N[u]$ is special by Proposition 43. Let $I=\{x_{1},x_{2}\},N_{H}(I)=\{y\}$. Since $\deg_{H}(x_{1})=\deg_{H}(x_{2})=1$, Proposition 42 implies that $x_{1}\sim A$ and $x_{2}\sim A$; since they already have common neighbor $y$, we assume without loss of generality that $x_{1}\sim v_{1}$ and $x_{2}\sim v_{2}$. Now consider the 6-cycle $u,v_{1},x_{1},y,x_{2},v_{2}$. Here, $x_{2}$ has neighbors $v_{2},y,r_{3}\sim y_{3},r_{4}\sim y_{4}$. We must have $\operatorname{codeg}(v_{1},x_{2})=0$; for, if $v_{1}\sim y$, then $G$ would have a 3-cycle $v_{1},x_{1},y$, and if $v_{1}\sim v_{2}$, then $G$ would have a 3-cycle $u,v_{1},v_{2}$, and if $v_{1}\sim r_{i}$ for $i\in\{3,4\}$, then $G$ would have a 4-cycle $r_{i},v_{i},u,v_{1}$. So by Proposition 41, the graph $G_{2}=G-N[v_{1}]$ is also special.
Accordingly, in this case we need:
$e^{-1.5a-4b-\beta}+e^{-1.5a-4b-\beta}+e^{-2a-7b-\alpha}+e^{-3a-10b-\alpha}\leq 1$ (4-18)
Case V: $\boldsymbol{G_{3}}$ is $\boldsymbol{3^{*}}$-vul. If $\deg_{H}(x)=1$ for any $x\in I$, then by Proposition 42, $u,x$ share a neighbor $v\in A$ and $R(u,N[x])\geq 2^{*}$. But then vertex $u^{\prime}=v$ and its two neighbors $v_{1}^{\prime}=u,v_{2}^{\prime}=x$ would be covered in Case IV, or an earlier case.
Accordingly, we must have $\deg_{H}(x)\geq 2$ for all $x\in I$. At this point, consider counting the two-element subsets of $N_{H}(I)$. Each vertex $x\in I$ contributes $\binom{\deg_{H}(x)}{2}$ such subsets and since $G$ has no four-cycles these are all distinct. So $\sum_{x\in I}\binom{\deg_{H}(x)}{2}\leq\binom{|N_{H}(I)|}{2}=\binom{r-1}{2}$. Since $\deg_{H}(x)\geq 2$ for all $x\in I$, this implies that $r\geq 5$ and, furthermore, if $r=5$ then $\deg_{H}(x)=2$ for all $x\in I$; in the latter case, by Proposition 43 the graph $G-N[u]$ is special. So depending on whether $r=5$ or $r\geq 6$, we need respectively:
$e^{-1.5a-4b-\beta}+e^{-1.5a-4b-\alpha}+e^{-2a-5b}\gamma_{3}^{*}+e^{-3a-13b-\alpha}\leq 1$ (4-19)
$e^{-1.5a-4b-\alpha}+e^{-1.5a-4b-\alpha}+e^{-2a-5b}\gamma_{3}^{*}+e^{-3a-14b-\alpha}\leq 1$ (4-20)
Case VI: Everything else. Since $G$ is in 4BP, we have $\operatorname{shad}(v_{1},N[v_{3}])\geq\operatorname{shad}(N[v_{3}])-1\geq 0$. On the other hand, $\operatorname{surp}_{G-v_{1}-N[v_{3}]}(I\cup\{v_{4}\})\leq\operatorname{minsurp}(H)+1+(\deg_{G-N[v_{3}]}(v_{4})-1)=2$. So $G-v_{1}-N[v_{3}]$ is $3^{*}$-vul via indset $I\cup\{v_{4}\}$; it would be covered by Case V. So Case VI is impossible.
9 Branching on 5-vertices
We will show for the following main result in this section:
Theorem 44.
Branch5$(G,k)$ has measure $a_{5}\mu+b_{5}k$ for $a_{5}=0.379406,b_{5}=0.097471$.
By Lemma 10, Theorem 44 combined with the MaxIS-5 algorithm immediately gives the algorithm with runtime $O^{*}(1.24394^{k})$ for degree-5 graphs.
There is a major difference between the degree-4 and degree-5 settings. In the former, we can ensure that the graph has many 3-vertices, and so branching on a single 4-vertex can create new 2-vertices. This structure cannot be maintained in the degree-5 case: at least two 5-vertices are jointly needed for the desired simplifications. The process is not self-sustaining.
Our strategy depends on the relative density of 5-vertices. When 5-vertices are very common, we can find closely-clustered groups of them for more advanced branching rules; the denser the 5-vertices, the more favorable the branching. We will actually develop three separate degree-5 algorithms with different “intensity” levels.
The first algorithm is very simple and is targeted to the regime where 5-vertices are rare:
Simplify $G$ If $\operatorname{maxdeg}(G)\leq 4$, then run either the MaxIS-4 algorithm or Branch4 Apply Section 5 to an arbitrary vertex of degree at least $5$.
Algorithm 6 Function Branch5-1$(G,k)$
Proposition 45.
Branch5-1 has measure $a_{51}\mu+b_{51}k$ for $a_{51}=0.496708,b_{51}=0.064301$.
Proof.
Given the potential branch-seqs from Section 5, we need the following constraints:
$\max\{e^{-a-3b}+e^{-a-5b},e^{-0.5a-b}+e^{-2a-5b}\}\leq 1$ (5-1)
$\min\{a_{4}\mu+b_{4}k,2(k-\mu)\log(1.137595)\}\leq a\mu+bk$ (5-2)
The latter constraint can be routinely checked by LP algorithms. ∎
The next algorithms Branch5-2 and Branch5-3 are targeted to medium and high density of 5-vertices respectively. They share some branching rules, so we describe them in a single listing.
1 Apply miscellaneous branching rules to ensure $G$ has nice structure.                                (We call this 5-base-position (5BP).) Branch on non-adjacent 5-vertices $u,v$ with $R(u,v)\geq 1$ and $R(u,N[v])\geq 1$. Branch on vertices $u,v$ where $\deg(u)=5,\deg_{G-u}(v)\geq 4$, and $G-\{u,v\}$ has a kite. Branch on a 4-vertex with two non-adjacent 5-neighbors Branch on a 5-vertex $u$ with neighbors $v_{1},v_{2},v_{3},v_{4},v_{5}$ where $v_{1}\sim v_{2}$ and $\deg(v_{1})=\deg(v_{3})=\deg(v_{4})=\deg(v_{5})=5$. Branch on a $5$-vertex with five $5$-neighbors. if $L=2$ then
2        Branch on a 5-vertex $u$ where $G-N[u]$ is 3-vul. Branch on non-adjacent 5-vertices $u,v$ with $R(u,v)\geq 1$. Branch on a $4$-vertex which has a triangle involving a 5-vertex. Branch on a 5-vertex with fewer than three 4-neighbors.
Run the cheaper of Branch5-$(L-1)$$(G,k)$ or the MaxIS-5 algorithm
Algorithm 7 Function Branch5-$L$$(G,k)$ for $L=2,3$
We will show the following results for these algorithms:
Theorem 46.
Branch5-2$(G,k)$ has measure $a_{52}\mu+b_{52}k$ for $a_{52}=0.437086,b_{52}=0.080751$.
Branch5-3$(G,k)$ has measure $a_{53}\mu+b_{53}k$ for $a_{53}=0.379406,b_{53}=0.097471$.
Thus, algorithm Branch5-3 will achieve the claim of Theorem 44.
9.1 Regularizing the graph and structural properties
We say that $G$ is in 5-base position (or 5BP for short) if the following conditions hold:
  • $G$ is simplified.
  • $\operatorname{maxdeg}(G)\leq 5$.
  • Every 5-vertex $u$ has $\operatorname{shad}(N[u])\geq 1$ and $R(u)=0$.
  • Every 4-vertex $u$ has $\operatorname{shad}(N[u])\geq 1$.
  • There are no vertices $u,v$ with $\operatorname{codeg}(u,v)\geq 2$ and $R(u,v)+R(N(u)\cap N(v))+\operatorname{codeg}(u,v)\geq 5$.
Note that, if $G$ is in 5BP, then any 5-vertex $u$ only has neighbors of degree 4 and 5 (else we would have $R(u)\geq 1$). Unlike in the degree-3 case, we cannot ensure that $G$ has no four-cycles. However, we can show that the four-cycles are heavily restricted.
Proposition 47.
If $G$ is in 5BP, then any pair of non-adjacent vertices $x,y$ have $\operatorname{codeg}(x,y)\leq 2$.
Proof.
Suppose $\deg(x)\geq\deg(y)$ and $\operatorname{codeg}(x,y)\geq 3$. If $\deg(x)=3$, then $\{x,y\}$ would be an indset with surplus one, contradicting that $G$ is simplified. If $\deg(y)=4$, then $\operatorname{shad}(N[y])\leq 0$ (due to vertex $x$), contradicting that $G$ is in 5BP. So we suppose $\deg(x)=\deg(y)=5$. Again, if $\operatorname{codeg}(x,y)\geq 4$ then $\operatorname{shad}(N[y])\leq 0$ due to vertex $x$, so we suppose that $\operatorname{codeg}(x,y)=3$ exactly. Then $R(N(x)\cap N(y))\geq 2$ due to non-adjacent 2-vertices $x,y$. So $R(x,y)+R(N(x)\cap N(y))+\operatorname{codeg}(x,y)\geq 5$ contradicting that $G$ is in 5BP. ∎
Proposition 48.
If $G$ is in 5BP, then any 4-cycle $x_{1},x_{2},x_{3},x_{4}$ satisfies the following properties:
  • All the vertices $x_{i}$ have degree at least four.
  • If $\deg(x_{1})=4$ then $x_{1}\not\sim x_{3}$
  • At least two of the vertices $x_{i}$ have degree five.
Proof.
For the first result, suppose that $\deg(x_{1})=3$. If $x_{1}\sim x_{3}$ or $x_{2}\sim x_{4}$, then $x_{1}$ would have a 3-triangle. So suppose $x_{1}\not\sim x_{3}$ and $x_{2}\not\sim x_{4}$. By Proposition 47, this means that $x_{2},x_{4}$ are the only shared neighbors of $x_{1},x_{3}$. If $\deg(x_{3})=4$, then we would have $\operatorname{shad}(N[x_{3}])\leq 0$ (due to 1-vertex $x_{1}$), contradicting that $G$ is simplified. So, suppose that $\deg(x_{3})=5$. Since $x_{2},x_{4}$ are neighbors of 5-vertex $x_{3}$ and 3-vertex $x_{1}$, they must have degree 4 (as 5-vertices cannot be neighbors of 3-vertices). At this point, observe that $R(x_{1},x_{3})\geq 2$ (due to non-adjacent 2-vertices $x_{2},x_{4}$) and $R(N(x_{1})\cap N(x_{3}))=R(x_{2},x_{4})\geq 1$ (due to 1-vertex $x_{1}$), contradicting that $G$ is in 5BP.
For the second result, suppose $\deg(x_{1})=4$ and $x_{1}\sim x_{3}$ and $\deg(x_{2})\geq\deg(x_{4})$; we have already shown that $\deg(x_{4})\geq 4$. We have $x_{2}\not\sim x_{4}$, as otherwise $x_{1}$ has a funnel (due to clique $x_{2},x_{3},x_{4}$). If $\deg(x_{2})=5$, then $R(x_{2})\geq 1$ (due to 3-triangle $x_{1},x_{3},x_{4}$). Otherwise, if $\deg(x_{2})=\deg(x_{4})=4$, then $R(N(x_{1})\cap N(x_{3}))\geq 1$ (due to subquadratic vertex $x_{2}$) and $R(x_{1},x_{3})\geq 2$ (due to non-adjacent 2-vertices $x_{2},x_{4}$); this contradicts that $G$ is in 5BP.
For the final result, suppose that $\deg(x_{1})=\deg(x_{2})=\deg(x_{3})=4$. We have already shown then that $x_{1}\not\sim x_{3}$, and Proposition 47 gives $\operatorname{codeg}(x_{1},x_{3})\leq 2$. So $G-\{x_{2},x_{4}\}$ has two non-adjacent subquadratic vertices $x_{1},x_{3}$. Thus $R(N(x_{1})\cap N(x_{3}))=R(x_{2},x_{4})\geq 2$ by Observation 16, and $R(x_{1},x_{3})\geq 1$ due to subquadratic vertex $x_{2}$. This contradicts that $G$ is in 5BP. ∎
Proposition 49.
Suppose $G$ is in 5BP. If a 4-vertex $u$ has a 5-neighbor, then the neighbors of $u$ can be enumerated as $v_{1},v_{2},v_{3},v_{4}$, where $\deg(v_{1})=5$ and $\{v_{3},v_{4}\}\not\sim\{v_{1},v_{2},v_{3},v_{4}\}$. Moreover, after Line 4, we may further suppose that $v_{3},v_{4}$ are subquartic.
Proof.
Let $v_{1}$ be a 5-neighbor of $u$. By Proposition 48, we must have $\operatorname{codeg}(u,v_{1})\leq 1$. Also, if there are any edges among the other neighbors of $u$, then $R(v_{1})\geq 1$ due to the resulting 3-triangle on $u$. Finally, if either $v_{3}$ or $v_{4}$ has degree 5, then $u$ would have two non-adjacent 5-neighbors and we would branch at Line 4. ∎
Proposition 50.
After Line 4, if a 5-vertex $u$ has at least four 5-neighbors, then the induced neighborhood $G[N(u)]$ is a matching.
Proof.
Suppose $x\sim u$ and $u,x$ share neighbors $y_{1},y_{2}$; let $v_{1},v_{2}$ be the two remaining neighbors of $u$, where $\deg(v_{1})=5$. Since we did not branch on $v_{1},v_{2}$ at Line 3, we must have $\deg(v_{2})=4,v_{1}\sim v_{2}$ and $\deg(y_{3})=\deg(y_{2})=\deg(x)=5$. By Proposition 49, we have $\{y_{1},y_{2},x\}\not\sim v_{2}$. Then $y_{1}\not\sim v_{1}$, as otherwise $G-\{y_{2},x\}$ would have a kite, where $\deg_{G-y_{2}}(v_{2})=\deg_{G}(v_{2})=4$, and we would branch at Line 3. So $R(v_{1},y_{1})\geq 1$ (due to the 3-triangle $u,x,y_{2}$) and $R(v_{1},N[y_{1}])\geq 1$ (due to 2-vertex $v_{2}$), and we would branch on non-adjacent 5-vertices $v_{1},y_{1}$ at Line 2. ∎
Lemma 51.
After Line 6, the graph has $n_{4}\geq n_{5}$.
Proof.
Let $A_{51},A_{5+}$ be the set of 5-vertices with exactly one 4-neighbor or more than one 4-neighbor, respectively, and let $B$ be the set of 4-vertices with exactly one 5-neighbor. Note that $A_{51},A_{5+}$ partition $V_{5}$ at this point; for, if a 5-vertex $u$ has five independent 5-neighbors, we would branch on it at Line 6; if it has a triangle in its neighborhood, we would branch on $u$ at Line 5.
We claim that, for any vertex $u\in A_{51}$ with 4-neighbor $x$, we have $x\in B$. For, suppose that $x$ has a second 5-neighbor $v$; we must have $u\sim v$ else we would branch on $x$ at Line 4. Then $u$ has three other 5-neighbors $v_{3},v_{4},v_{5}$; we would branch on $u$ at Line 5.
By counting edges between $B$ and $A_{51}$, we thus see that $|B|\geq|A_{51}|$. Next let us count the edges between $V_{4}$ and $V_{5}$. Clearly there are least $|A_{51}|+2|A_{5+}|$ such edges. Furthermore, each vertex $v\in B$ has exactly one 5-neighbor, and by Proposition 49 any 4-vertex $v$ has at most two 5-neighbors. So $|B|+2(|V_{4}|-|B|)\leq|A_{51}|+2|A_{5+}|$. Combining this inequality with $|B|\geq|A_{51}|$, we conclude that $|V_{4}|=|B|+(|V_{4}|-|B|)\geq|A_{51}|+|A_{5+}|=|V_{5}|.$
Observation 52.
After Line 8 of algorithm Branch5-2, any 5-vertex $x$ has $\operatorname{shad}(N[x])\geq 2$, and has no 3-vertex within distance two. Any vertex $y\not\sim x$ has $\operatorname{codeg}(x,y)\leq\deg(y)-3$.
Proof.
If $\operatorname{shad}(N[x])\leq 1$, then since $G$ is in 5BP we would have $R(N[x])\geq 1$ and we would branch on $x$ at Line 2. So $\deg_{G-N[x]}(y)\geq 3$, i.e. $\operatorname{codeg}(x,y)\leq\deg(y)-3$. We have already seen that $x$ has no 3-neighbor since $G$ is in 5BP; this also shows it cannot have any 3-vertex at distance two. ∎
Lemma 53.
After Line 9 of algorithm Branch5-2, any non-adjacent vertices $x,y$ with $\deg(x)=5,\deg(y)=4,\operatorname{codeg}(x,y)=1$ satisfy $\operatorname{shad}(N[x,y])\geq 1$.
Proof.
Let $G^{\prime}=G-N[x,y]$. We first claim that $\operatorname{mindeg}(G^{\prime})\geq 2$. For, consider a vertex $t\in G^{\prime}$. If $\deg_{G}(t)=3$, then by Observation 52 we have $\operatorname{codeg}(x,t)=0$ and by Proposition 48 we have $\operatorname{codeg}(y,t)\leq 1$. So $\deg_{G^{\prime}}(t)\geq 2$. Likewise, if $\deg_{G}(t)=5$, then $\operatorname{codeg}(y,t)\leq 1$ and $\operatorname{codeg}(x,t)\leq 2$. So again $\deg_{G^{\prime}}(t)\geq 2$.
Next suppose $\deg_{G}(t)=4$; by Proposition 47 and Observation 52 we have $\operatorname{codeg}(y,t)\leq 2$ and $\operatorname{codeg}(x,t)\leq 1$. To get $\deg_{G^{\prime}}(t)=1$, vertex $t$ must share two neighbors $z_{1},z_{2}$ with $y$ and share a distinct neighbor $z_{3}$ with $x$. By Proposition 48 we must have $\deg(z_{1})=\deg(z_{2})=5$. If $z_{1}\not\sim z_{2}$, then $R(z_{1},z_{2})\geq 1$ (due to subquadratic vertex $y$), and we would branch on $z_{1},z_{2}$ at Line 9. If $z_{1}\sim z_{2}$, then in the graph $G-N[x]$, vertex $t$ loses its neighbor $z_{3}$, and so it has a 3-triangle $z_{1},z_{2}$; in particular, $R(N[x])\geq 1$ and we would branch on $x$ at Line 8.
Thus, in all cases, $\deg_{G^{\prime}}(t)\geq 2$. Now, suppose $G^{\prime}$ has an indset $I$ with $\operatorname{surp}_{G^{\prime}}(I)\leq 0$. Since $\operatorname{mindeg}(G^{\prime})\geq 2$, necessarily $|I|\geq 2$. Then $\operatorname{surp}_{G-N[x]}(I\cup\{y\})\leq 0+\deg_{G-N[x]}(y)-1=2$. So $G-N[x]$ would be 3-vul and we would branch on $x$ at Line 8. ∎
Observation 54.
After Line 10 of algorithm Branch5-2, any 4-vertex has at most one 5-neighbor and is not part of any 4-cycle.
Proof.
Consider a 4-vertex $v$. If $v$ has two 5-neighbors we would branch at Line 6 or Line 10, depending if they were adjacent or not. Now consider a four-cycle $v,x_{1},x_{2},x_{3}$. By Observation 52, we have $\deg(x_{2})\leq 4$. Thus, by Proposition 48, both $x_{1}$ and $x_{3}$ have degree 5; that is, $v$ has two 5-neighbors, which is what we have showed is impossible.
9.2 Basic branching rules
We analyze it line by line; Lines 6 and 11, which are more involved, will be handled later.
Line 1: We can first simplify $G$. If a vertex $u$ has degree at least 6, we apply Section 5. If a 5-vertex $u$ has $\operatorname{shad}(N[u])\leq 0$ we apply Lemma 26. Otherwise if a 5-vertex $u$ has $R(u)\geq 1$, then we split on $u$ giving branch-seq $[(0.5,2),(2,5)]$. If a 4-vertex $u$ has $\operatorname{shad}(N[u])\leq 0$, then we apply Proposition 25. If vertices $u,v$ have $R(u,v)+R(N(u)\cap N(v)))+\operatorname{codeg}(u,v)\geq 5$, we apply (A3).
Over all cases, we get branch-seq $[(1,3),(1,4)]$, $[(1,2),(1,5)]$, $[(0.5,2),(2,5)]$, or $[(0.5,1),(2.5,6)]$. Accordingly, we need:
$\max\{e^{-a-3b}+e^{-a-4b},e^{-a-2b}+e^{-a-5b},e^{-0.5a-2b}+e^{-2a-5b},e^{-0.5a-b}+e^{-2.5a-6b}\}\leq 1$ (5-3)
Line 2: We split on $u,v$, obtaining subproblems $G_{1}=G-u,G_{2}=G-u-N[v],G_{3}=G-\{u,v\}$. Since $G$ is in 5BP, they have drop $(2,5),(2.5,6),(1,2)$ directly. After simplification, $G_{2}$ has drop $(2.5,7)$ and $G_{3}$ has drop $(1,3)$. So we need:
$e^{-2a-5b}+e^{-2.5a-7b}+e^{-a-3b}\leq 1$ (5-4)
Line 3: We split on $u,v$ and apply (P3) to $G-\{u,v\}$. Since $G$ is in 5BP, subproblems $G-N[u]$ and $G-u-N[v]$ both have drops $(2,5)$. Also, $\operatorname{shad}(u,v)\geq\operatorname{shad}(u)-1\geq 1$ so by Lemma 9 the subproblem $G-\{u,v\}$ has drop $(1.5,4)$ after simplification. We need:
$e^{-2a-5b}+e^{-2a-5b}+e^{-1.5a-4b}\leq 1$ (5-5)
Line 4: Consider a 4-vertex $u$ with non-adjacent 5-neighbors $y_{1},y_{2}$, and let $x_{1},x_{2}$ be the other two neighbors of $u$ where $\deg(x_{1})\leq\deg(x_{2})$. Our plan is to split on $y_{1},y_{2}$, generating subproblems $G_{1}=G-N[y_{1}],G_{2}=G-y_{1}-N[y_{2}],G_{3}=G-\{y_{1},y_{2}\}$. Since $G$ is in 5BP, the subproblems have drops $(2,5),(2.5,6),(1,2)$ directly.
By Proposition 49, the neighbors of $u$ must be independent. If $\deg(x_{1})\leq 3$, then $R(y_{1},y_{2})\geq 1$ (due to 2-vertex $u$) and $R(y_{1},N[y_{2}])\geq 1$ (due to 2-vertex $x_{1}$); in this case we would branch on $y_{1},y_{2}$ at Line 2. So we may assume that $\deg(x_{1})\geq 4$. We can apply Proposition 30 to 2-vertex $u$ in $G_{3}$ with $r=\deg(x_{1})+\deg(x_{2})-\operatorname{codeg}(x_{1},x_{2})-1$. If $\deg(x_{1})=\deg(x_{2})=4$, then Proposition 48 gives $\operatorname{codeg}(x_{1},x_{2})=1$ (they share 4-neighbor $u$). If $\deg(x_{2})=5$, then Proposition 47 gives $\operatorname{codeg}(x_{1},x_{2})\leq 2$. In either case $r\geq 6$ and so we need:
$e^{-2a-5b}+e^{-2.5a-6b}+e^{-a-2b}\psi_{6}\leq 1$ (5-6)
Line 5: Vertices $v_{3},v_{4},v_{5}$ cannot form a clique (else $R(v_{1})\geq 1$ due to the funnel $u$), so assume without loss of generality $v_{3}\not\sim v_{4}$. Then we split on $v_{3},v_{4}$, generating subproblems $G_{1}=G-N[v_{3}],G_{2}=G-v_{3}-N[v_{4}],G_{3}=G-\{v_{3},v_{4}\}$, which have drops $(2,5),(2.5,6),(1,2)$ directly.
In $G_{3}$, we also apply Proposition 30 to vertices $v_{1},v_{5}$ for the 3-triangle $u,v_{1},v_{2}$. By Proposition 50, at this point $\operatorname{codeg}(v_{1},u)\leq 1$ and $\operatorname{codeg}(v_{5},u)\leq 1$. So $v_{1}\not\sim v_{5}$ and $\deg_{G_{3}}(v_{1})=5,\deg_{G_{3}}(v_{5})\geq 4$ and, by Proposition 47, we have $\operatorname{codeg}(v_{1},v_{5})\leq 2$. So $\deg_{G_{3}}(v_{1})+\deg_{G_{3}}(v_{5})-\operatorname{codeg}_{G_{3}}(v_{1},v_{5})-1\geq 5+4-2-1=6$. Overall we need
$e^{-2a-5b}+e^{-2.5a-6b}+e^{-a-2b}\psi_{6}\leq 1$
which is just a repetition of (5-6).
Line 8: Since $G$ is in 5BP, subproblem $G-N[u]$ has drop $(2,5)$ directly, and subproblem $G-u$ has drop $(0.5,1)$. We apply Proposition 31 to $G-N[u]$, so we need:
$e^{-2a-5b}\gamma_{3}+e^{-0.5a-b}\leq 1\qquad\text{(when $L=2$)}$ (5-7)
Line 9: We get subproblems $G-N[u],G-u-N[v],G-\{u,v\}$. Since $G$ is in 5BP, these have drops $(2,5),(2.5,6),(1,2)$ directly, and the latter has drop $(1,3)$ after simplification. We need:
$e^{-2a-5b}+e^{-2.5a-6b}+e^{-a-3b}\leq 1\qquad\text{(when $L=2$)}$ (5-8)
Line 10: Suppose we are branching on a 4-vertex $u$ with neighbors $x_{1},x_{2},y_{1},y_{2}$ where $\deg(x_{1})=5$ and $x_{1}\sim x_{2}$. By Proposition 49, $u$ has no other edges in its neighborhood and $y_{1},y_{2}$ are subquartic. Since $x_{2},y_{1},y_{2}$ are within distance two of 5-vertex $x_{1}$ we have $\deg(y_{1})=\deg(y_{2})=4$ and so by Observation 52 we have $\operatorname{codeg}(x_{1},y_{2})=\operatorname{codeg}(x_{1},y_{1})=1$.
Here, we split on $y_{1}$, and then apply (A2) to $x_{1}$ in subproblem $G-y_{1}$. This generates subproblems $G_{1}=G-N[y_{1}],G_{2}=G-y_{1}-N[y_{2},x_{1}]$, and $G_{3}=G-\{y_{1},x_{1}\}$. Since $G$ is in 5BP, subproblem $G_{1}$ has drop $(1.5,4)$ directly. Also, by Lemma 53, we have $\operatorname{minsurp}(G_{2})\geq\operatorname{shad}(N[y_{2},x_{1}])-1\geq 0$. So $G_{2}$ has drop $(3.5,9)$. Subproblem $G_{3}$ has drop $(1,2)$ directly, and has a 2-vertex $u$ so it has drop $(1,3)$ after simplification. We need:
$e^{-1.5a-4b}+e^{-3.5a-9b}+e^{-a-3b}\leq 1\qquad\text{(when $L=2$)}$ (5-9)
Line 12: We use a more refined property of the MaxIS-5 algorithm of [XN16]; specifically, their algorithm has runtime $O^{*}(1.17366^{w_{3}n_{3}+w_{4}n_{4}+n_{5}})$ for constants $w_{3}=0.5093,w_{4}=0.8243$. For Branch5-3, Lemma 51 implies that $n_{4}\geq n_{5}$, and so we need:
$\min\{a_{52}\mu+b_{52}k,2(k-\mu)\cdot(w_{4}\cdot 1/2+1/2)\log 1.17366\}\leq a\mu+bk$ (5-10)
which can be checked routinely by LP algorithms.
For Branch5-2, each 5-vertex has at least three 4-neighbors (else we would branch at Line 12), and each 4-vertex has at most one 5-neighbor. So $n_{4}\geq 3n_{5}$, and we need
$\min\{a_{51}\mu+b_{51}k,2(k-\mu)\cdot(w_{4}\cdot 3/4+1/4)\log 1.17366\}\leq a\mu+bk$ (5-11)
which can also be checked mechanically.
9.3 Line 6: A 5-vertex with five independent 5-neighbors
Our basic plan is to choose an ordering of the neighbors of $u$ as $v_{1},\dots,v_{5}$, and then split on $v_{1},v_{2},v_{3}$ to create a 2-vertex $u$. The graph $H=G-\{v_{1},v_{2},v_{3}\}-N[v_{4},v_{5}]$ is central to this analysis; note that $\operatorname{codeg}(v_{4},v_{5})\in\{1,2\}$ by Proposition 47.
Let $I$ be a min-set of $H$, and define parameters
$s=|N(I)|,\qquad\ell=\operatorname{minsurp}(H),\qquad d_{45}=\operatorname{codeg}(v_{4},v_{5})$
For this section, we define $A=\{v_{1},v_{2},v_{3}\}$ and $B=\{v_{4},v_{5}\}$, so that $H=G-(A\cup N[B])$.
Proposition 55.
If there is any vertex $x\in H$ with $\deg_{H}(x)\leq 5-\deg_{G}(x)$, then $u$ has neighbors $v_{i},v_{j}$ with $R(v_{i},N[v_{j}])\geq 1$.
Proof.
If $\deg_{G}(x)=3$, then $x\not\sim A$; to get $\deg_{H}(x)\leq 2$, we must have $\operatorname{codeg}(x,v_{j})\geq 1$ for $v_{j}\in B$, so $R(v_{1},N[v_{j}])\geq 1$. If $\deg_{G}(x)=4$, then $|N(x)\cap A|\leq 1$ (else we would branch on $x$ at Line 3); so to get $\deg_{H}(x)\leq 1$, we have either $\operatorname{codeg}(x,v_{j})\geq 2$ for $v_{j}\in B$, or $|N(x)\cap A|\geq 1$. If $\operatorname{codeg}(x,v_{j})\geq 2$ for $v_{j}\in B$, then $R(v_{1},N[v_{j}])\geq 1$ (due to subquadratic vertex $x$), while if $\operatorname{codeg}(x,v_{4})=\operatorname{codeg}(x,v_{5})=1,x\sim v_{i}\in A$, then $R(v_{i},N[v_{4}])\geq 1$ (due to subquadratic vertex $x$).
Finally, suppose $\deg_{G}(x)=5$ and $x$ is isolated in $H$. By Proposition 47 we have $\operatorname{codeg}(x,u)\leq 2$ and $\operatorname{codeg}(x,v_{4})\leq 2$ and $\operatorname{codeg}(x,v_{5})\leq 2$. If $\operatorname{codeg}(x,v_{4})=\operatorname{codeg}(x,v_{5})=2$, then $x$ has a neighbor $v_{i}\in A$, so $R(v_{i},N[v_{4}])\geq 1$ (due to subquadratic vertex $x$). Otherwise, if $\operatorname{codeg}(x,v_{j})=2$ for $v_{j}\in B$ and $x\sim v_{i}\in A$, then $R(v_{i},N[v_{j}])\geq 1$ (due to subquadratic vertex $x$) ∎
Proposition 56.
If $s=1$ and $\ell=-1$, then one of the following holds:
  • There are distinct neighbors $v_{i},v_{j}$ where $G-v_{i}-N[v_{j}]$ is 3-vul.
  • $d_{45}=1$ and there are neighbors $v_{i},v_{i^{\prime}},v_{j}$ with $R(v_{i},v_{i^{\prime}},N[v_{j}])\geq 1$.
  • $d_{45}=2$ and the graph $G-N[u]$ has a 5-vertex with a 3-neighbor.
Proof.
We assume that $I=\{x_{1},x_{2}\}$ for 5-vertices $x_{1},x_{2}$ where $\deg_{H}(x_{1})=\deg_{H}(x_{2})=1$, as otherwise the result follows from Proposition 55. Let $y$ be the single shared neighbor of $x_{1},x_{2}$ in $H$. Since $y$ has non-adjacent 5-neighbors $x_{1},x_{2}$, we have $\deg_{G}(y)=5$. By Proposition 47, each $x\in I$ has at most two neighbors in $A$ and shares at most two neighbors with each $v_{j}\in B$.
If $I\not\sim A$, then $\operatorname{surp}_{G-N[v_{4},v_{5}]}(I)=-1$, and hence $\operatorname{surp}_{G-v_{1}-N[v_{4}]}(I\cup\{v_{5}\})\leq 2$, and $G-v_{1}-N[v_{4}]$ is 3-vul via in the indset $J=I\cup\{v_{5}\}$.
If some $x\in I$ has exactly one neighbor $v_{i}\in A$, then $\operatorname{codeg}(x,v_{j})\geq 2$ for some $v_{j}\in B$ and $R(v_{i},N[v_{j}])\geq 1$ (due to subquadratic vertex $x$).
If $d_{45}=1$, the only remaining case is that there is some $x\in I$ with two neighbors $v_{i},v_{i^{\prime}}$ in $A$. Then $\operatorname{codeg}(x,v_{j})\geq 1$ for some $v_{j}\in B$, and so $R(v_{i},v_{i^{\prime}},N[v_{j}])\geq 1$ as claimed.
Finally suppose $d_{45}=2$. If $|N(I)\cap A|<3$, say without loss of generality $v_{1}\not\sim I$, then $\operatorname{surp}_{G-\{v_{2},v_{3}\}-N[v_{4},v_{5}]}(I)=-1$ and hence $\operatorname{surp}_{G-v_{2}-N[v_{4}]}(I\cup\{v_{5}\})\leq 2$; so $G-v_{2}-N[v_{4}]$ is 3-vul (via indset $I\cup\{v_{5}\})$. So let us suppose that every vertex in $A$ has at least one neighbor in $I$, and no $x\in I$ has exactly one neighbor in $A$. Without loss of generality, we may suppose that $v_{1}\sim x_{1}$ and $v_{2}\sim x_{2}$, and $v_{3}\sim x_{1}$ and $v_{3}\sim x_{2}$, with no other edges between $A$ and $I$.
Now $v_{3}$ has four 5-neighbors $u,x_{1},x_{2},y$ and $y,v_{3}$ share neighbors $x_{1},x_{2}$. So by Proposition 50 we have $y\not\sim v_{3}$. If $y\sim v_{i}$ for $v_{i}\in\{v_{1},v_{2}\}$, then $R(v_{i},N[v_{3}])\geq 1$ due to subquadratic vertex $y$. Otherwise, if $y\not\sim A$, then in the graph $G-N[u]$, vertices $x_{1},x_{2}$ remain with degree 3 (losing neighbors $v_{1},v_{2},v_{3}$), while $y$ remains with degree 5. ∎
We now analyze the subproblems. When describing each case, we assume that no other vertex ordering could go into an earlier case.
Case I: $\boldsymbol{\operatorname{shad}(v_{1},v_{2},N[v_{3}])\leq-1.}$ Let $G^{\prime}=G-\{v_{1},v_{2}\}-N[v_{3}]$. We claim that $\operatorname{mindeg}(G^{\prime})\geq 1$. For, consider a vertex $x\in G^{\prime}$. By Proposition 47, we have $\operatorname{codeg}(x,v_{3})\leq 2$. If $\deg(x)=5$, this implies $\deg_{G^{\prime}}(x)\geq 1$. If $\deg(x)=4$, then $|N(x)\cap\{v_{1},v_{2}\}|\leq 1$ (else we would branch on $x$ at Line 4), so again $\deg_{G^{\prime}}(x)\geq 1$. Finally, if $\deg(x)=3$, then $x\not\sim\{v_{1},v_{2}\}$ and $\deg_{G^{\prime}}(x)\geq 1$.
So $\operatorname{surp}_{G^{\prime}}(J)\leq-1$ for an indset $J$, where necessarily $|J|\geq 2$. Then $\operatorname{surp}_{G-N[v_{3}]}(J)\leq 1$; since $G$ is in 5BP, this implies $R(N[v_{3}])\geq|J|\geq 2$. Our strategy is to split on $v_{3}$, generating subproblems $G-v_{3},G-N[v_{3}]$; these have drops $(0.5,1),(2,7)$ after simplification. We need:
$e^{-0.5a-b}+e^{-2a-7b}\leq 1$ (5-12)
Case II: $\boldsymbol{R(v_{1},v_{2},N[v_{3}])\geq 3.}$ Split on $v_{1},v_{2},v_{3}$, generating subproblems $G-N[v_{1}],G-v_{1}-N[v_{2}],G-\{v_{1},v_{2}\}-N[v_{3}],G-\{v_{1},v_{2},v_{3}\}$. Since we are not in Case I, the first three have drops $(2,5),(2.5,6),(3,7)$ directly. Since $G$ is in 5BP, subproblem $G-\{v_{1},v_{2},v_{3}\}$ has drop $(1.5,3)$ directly; we apply Proposition 30 to its 2-vertex $u$ where $\deg(v_{4})+\deg(v_{5})-d_{45}-1\geq 7$. So we need:
$e^{-2a-5b}+e^{-2.5a-6b}+e^{-3a-10b}+e^{-1.5a-3b}\psi_{7}\leq 1$ (5-13)
Note: After Case II, we suppose that $\ell\geq-3+d_{45}$. For, suppose $\operatorname{surp}_{H}(I)\leq-4+d_{45}\leq-2$, where necessarily $|I|\geq 2$. Then $\operatorname{surp}_{G-\{v_{1},v_{2}\}-N[v_{4}]}(I\cup\{v_{5}\})\leq 1$. So either $\operatorname{shad}(v_{1},v_{2},N[v_{4}])\leq-1$ (which would put us into Case I with alternate vertex ordering $v_{1},v_{2},v_{4}$) or, by Lemma 13 we have $R(v_{1},v_{2},N[v_{4}])\geq|I\cup\{v_{5}\}|\geq 3$, putting us into Case II.
Our strategy in all subsequent cases is to split on $v_{1},v_{2},v_{3}$, generating subproblems $G_{1}=G-N[v_{1}],G_{2}=G-v_{1}-N[v_{2}],G_{3}=G-\{v_{1},v_{2}\}-N[v_{3}]$, and $G-\{v_{1},v_{2},v_{3}\}$. We then apply (A1) to the 2-vertex $u$ in $G-\{v_{1},v_{2},v_{3}\}$, obtaining subproblems $G_{0}=G-N[u]$ and $H=G-\{v_{1},v_{2},v_{3}\}-N[v_{4},v_{5}]$. Since we are not in Case I, subproblems $G_{0},G_{1},G_{2},G_{3}$ have drops $(2,5),(2,5),(2.5,6)$ and $(3,7)$ respectively, and the final subproblem $H$ has drop
$(5.5-d_{45}/2,13-d_{45})\qquad\text{ or }\qquad(5.5-(d_{45}+\ell)/2,13-d_{45}+s)$
where the latter is obtained if we apply P1 to $I$.
Case III: $\boldsymbol{G_{3}}$ is 2-vul, and either $\boldsymbol{G_{1}}$ or $\boldsymbol{G_{2}}$ is 3-vul. Here, $H$ has drop $(4,11)$. We can apply Proposition 31 to $G_{3}$ and $G_{1}$ or $G_{2}$; overall, in the two cases, we need:
$e^{-2a-5b}+\max\{e^{-2a-5b}\gamma_{3}+e^{-2.5a-6b},e^{-2a-5b}+e^{-2.5a-6b}\gamma_{3}\}+e^{-3a-7b}\gamma_{2}+e^{-4a-11b}\leq 1$ (5-14)
Case IV: Either $\boldsymbol{G_{1}}$ or $\boldsymbol{G_{2}}$ is 3-vul. If $\ell=-3+d_{45}$, then $\operatorname{surp}_{G-\{v_{1},v_{2}\}-N[v_{4}]}(I\cup\{v_{5}\})\leq 2$ and so $G-\{v_{1},v_{2}\}-N[v_{4}]$ is 2-vul; this would put us into Case III via vertex ordering $v_{1}^{\prime}=v_{1},v_{2}^{\prime}=v_{2},v_{3}^{\prime}=v_{4}$. So we suppose $\ell\geq-2+d_{45}$ and hence $H$ has drop $(4.5,11)$. We need:
$e^{-2a-5b}+\max\{e^{-2a-5b}\gamma_{3}+e^{-2.5a-6b},e^{-2a-5b}+e^{-2.5a-6b}\gamma_{3}\}+e^{-3a-7b}+e^{-4.5a-11b}\leq 1$ (5-15)
Note: After Case IV, we must have $s\geq 1$; for, if $s=0$, then $H$ has an isolated vertex; by Proposition 55 there are then neighbors $v_{i},v_{j}$ with $R(v_{i},N[v_{j}])\geq 1$, putting us into Case IV.
Case V: $\boldsymbol{G_{3}}$ is 4-vul. We know $\ell\geq-3+d_{45}$. First, suppose $d_{45}=2,\ell=-1,s=1$ exactly. So $H$ has drop $(4,12)$. By Proposition 56, there are either vertices $v_{i},v_{j}$ where $G-v_{i}-N[v_{j}]$ is 3-vul (putting us into Case IV), or $G_{0}$ has a 5-vertex with a 3-neighbor. In the latter case, we can either simplify $G_{0}$ further, or apply Lemma 24 to $G_{0}$, and we then need:
$e^{-2a-5b}\max\{e^{-b},e^{-a-3b}+e^{-a-5b},e^{-0.5a-2b}+e^{-2a-5b}\}$
$\qquad\qquad+e^{-2a-5b}+e^{-2.5a-6b}+e^{-3a-7b}\gamma_{4}+e^{-4a-12b}\leq 1$ (5-16)
Otherwise, if $\ell\geq-2+d_{45}$ or $s\geq 2$ or $d_{45}=1$, then $H$ has drop $(4.5,11)$ or $(4,13)$, and we need:
$e^{-2a-5b}+e^{-2a-5b}+e^{-2.5a-6b}+e^{-3a-7b}\gamma_{4}+\max\{e^{-4a-13b},e^{-4.5a-11b}\}\leq 1$ (5-17)
Case VI: $\boldsymbol{d_{45}=1}$. If $\ell\leq-2$, then $\operatorname{surp}_{G-\{v_{1},v_{2}\}-N[v_{4}]}(I\cup\{v_{5}\})\leq 2$ and $|I|=s-\ell\geq 3$, so $G-\{v_{1},v_{2}\}-N[v_{4}]$ is 4-vul; this would put us into Case V. If $s=1,\ell=-1$, then by Proposition 56, there are either neighbors $v_{i},v_{i^{\prime}},v_{j}$ with $R(v_{i},v_{i^{\prime}},N[v_{j}])\geq 1$ (putting us into Case V) or $G-v_{i}-N[v_{j}]$ is 3-vul (putting us into Case IV). So either we haver $\ell\geq 0$ or we have $\ell=-1,s\geq 2$, and thus $H$ has drop $(4.5,14)$ or $(5,12)$. We need:
$e^{-2a-5b}+e^{-2a-5b}+e^{-2.5a-6b}+e^{-3a-7b}+\max\{e^{-4.5a-14b},e^{-5a-12b}\}\leq 1$ (5-18)
Case VII: Everything else. Since no ordering of the vertices $v_{1},\dots,v_{5}$ puts us into Case V, we suppose that every pair $v_{i},v_{j}$ shares some neighbor $x_{ij}\neq u$. The vertices $x_{ij}$ are distinct, since otherwise $\operatorname{codeg}(x_{ij},u)\geq 3$ for some $i,j$, contradicting Proposition 47. So the distance-two vertices from $u$ are precisely $\{x_{ij}:1\leq i<j\leq 5\}$. Then indset $I=\{v_{2},\dots,v_{5}\}$ has precisely $6$ neighbors in $G-N[v_{1}]$, namely vertices $x_{ij}$ for $2\leq i<j\leq 5$. So $G-N[v_{1}]$ is 4-vul via indset $I$; this would be covered in Case IV. So Case VII is impossible.
9.4 Line 11: A 5-vertex with fewer than three 4-neighbors
We suppose we are in Branch5-2 here. Let $U$ be the non-empty set of 5-vertices which have fewer than three 4-neighbors. For any vertex $u\in U$, define the induced neighborhood $H_{u}=G[N(u)]$. Note that any 4-vertex is isolated in $H_{u}$, as otherwise we would apply Line 10 to it with its 5-neighbor $u$. The following theorem characterizes the possible neighborhood structure (see Figure 7).
Figure 7: The four cases for the neighborhood of $u$. Dashed edges may or may not be present.
Theorem 57.
There is some vertex $u\in U$ and an enumeration of its neighbors $v_{1},\dots,v_{5}$ to satisfy one of the following three conditions:
  1. (I)
    $v_{1},\dots,v_{5}$ are independent and $\deg(v_{1})=\deg(v_{2})=\deg(v_{3})=5$.
  2. (II)
    $H_{u}$ is a matching and $v_{1}\sim v_{2},\deg(v_{1})=\deg(v_{2})=\deg(v_{3})=5,\deg(v_{5})=4,\deg_{G-v_{3}}(v_{4})=4$ and $\operatorname{codeg}(v_{1},v_{4})=\operatorname{codeg}(v_{2},v_{4})=\operatorname{codeg}(v_{1},v_{5})=\operatorname{codeg}(v_{2},v_{5})=\operatorname{codeg}(v_{3},v_{4})=1$.
  3. (III)
    $v_{1}\sim v_{2},v_{1}\sim v_{3}$, and $\deg(v_{1})=5,\deg(v_{4})=\deg(v_{5})=4$ and $\{v_{4},v_{5}\}\not\sim v_{1}$.
  4. (IV)
    $v_{1}\sim v_{2},v_{3}\sim v_{4},\deg(v_{1})=\deg(v_{2})=\deg(v_{3})=\deg(v_{4})=5$. Furthermore, there are 5-vertices $x_{13},x_{24},y_{2},y_{4}$ and 4-vertices $z_{2},z_{4}$, all distinct from each other and from $N[u]$, such that $v_{2},x_{24},y_{2}$ and $v_{4},x_{24},y_{4}$ are both triangles in $G$ and $x_{2}\sim x_{24}\sim x_{4}$ and $x_{13}\not\sim\{x_{24},v_{2},v_{4}\}$ and $v_{2}\sim z_{2},v_{4}\sim z_{4}$.
Proof.
If $H_{u}$ has no edges, it falls into the first category (choose $v_{1},v_{2},v_{3}$ to be 5-neighbors of $u$).
Suppose $H_{u}$ is not a matching, i.e. there is a neighbor $v_{1}$ with $\operatorname{codeg}(v_{1},u)\geq 2$. Let $v_{2},v_{3}$ be two shared neighbors of $v_{1},u$ and let $v_{4},v_{5}$ be any other neighbors of $u$. Note that $G-\{v_{4},v_{5}\}$ has a kite. Since 4-vertices are isolated in $H_{u}$, the only possible structure where we would not use the branching rule at Line 3 must be that $\deg(v_{1})=5,\deg(v_{4})=\deg(v_{5})=4$; since 4-vertices are isolated in $H_{u}$ necessarily $\{v_{1},v_{2},v_{3}\}\not\sim\{v_{4},v_{5}\}$. This gives Case III.
Suppose $H_{u}$ is a matching with exactly one edge $(v_{1},v_{2})$, and an additional 5-vertex $v_{3}$. If it has any other 5-neighbor $t$, then $R(v_{3},t)\geq 1$ (due to 3-triangle $u,v_{1},v_{2}$) and we would branch at Line 9. So $u$ must have two 4-neighbors $v_{4},v_{5}$. Then Observation 54 implies $\operatorname{codeg}(v_{1},v_{4})=\operatorname{codeg}(v_{2},v_{4})=\operatorname{codeg}(v_{1},v_{5})=\operatorname{codeg}(v_{2},v_{5})=\operatorname{codeg}(v_{3},v_{4})=1$. This gives Case II.
Thus, for the remainder of the proof, we suppose that, for every vertex $u\in U$, the graph $H_{u}$ is a matching with exactly two edges, where the one unmatched neighbor has degree 4.
Consider some $u\in U$, where $H_{u}$ has edges $(v_{1},v_{2}),(v_{3},v_{4})$ and has a 4-vertex $v_{5}$. By Observation 52 we automatically have $\operatorname{codeg}(v_{1},v_{5})=\operatorname{codeg}(v_{2},v_{5})=1$. We claim that $\operatorname{codeg}(v_{3},v_{4})=1$. For, suppose $v_{3},v_{4}$ share neighbor $t\neq u$; since $t$ has two 5-neighbors necessarily $\deg(t)=5$. Then $v_{3}\in U$ since it has three 5-neighbors $u,v_{4},t$, but $\deg_{H_{v_{3}}}(v_{4})\geq 2$, which contradicts our assumption on structure of $U$. By completely symmetric arguments also $\operatorname{codeg}(v_{1},v_{2})=1$.
Now consider forming a bipartite graph $J$ with left-nodes $v_{1},v_{2}$ and right-nodes $v_{3},v_{4}$, with an edge $(v_{i},v_{j})$ if $\operatorname{codeg}_{G}(v_{i},v_{j})=2$. If any node has degree zero in $J$, we may suppose without loss of generality that $\deg_{J}(v_{4})=0$ in which case $\operatorname{codeg}(v_{1},v_{4})=\operatorname{codeg}(v_{2},v_{4})=1$ giving us Case II.
Otherwise, if every node in $J$ has at least one neighbor, then by Hall’s theorem $J$ has a perfect matching; by swapping vertices as needed we assume this matching is $(v_{1},v_{3}),(v_{2},v_{4})$, that is, $v_{1},v_{3}$ share a neighbor $x_{13}\neq u$ and $v_{2},v_{4}$ share a neighbor $x_{24}\neq u$. Here, $\deg(x_{13})=\deg(x_{24})=5$ since they each have two 5-neighbors. Clearly $u\not\sim\{x_{13},x_{24}\}$. By Observation 52, we have $\operatorname{codeg}(x_{13},u)\leq 2$ and $\operatorname{codeg}(x_{24},u)\leq 2$, so $x_{13}\not\sim\{v_{2},v_{4}\}$ and $x_{24}\not\sim\{v_{1},v_{3}\}$. In particular, $x_{13}\neq x_{24}$. Also, $x_{13}\not\sim x_{24}$, as otherwise $x_{13}$ would have three independent 5-neighbors $x_{24},v_{1},v_{3}$, contradicting our assumption of structure of $U$.
Now $v_{2}\in U$ as it has three 5-neighbors $v_{1},u,x_{24}$ and similarly $v_{4}\in U$ as it has three 5-neighbors $v_{3},u,x_{24}$. By our hypothesis on the structure of $U$, it must be that $v_{2}$ has an edge $(x_{24},y_{2})$ in its neighborhood and has a 4-neighbor $z_{2}$, and likewise $v_{4}$ has an edge $(x_{24},y_{4})$ in its neighborhood and has a 4-neighbor $z_{4}$. By Observation 52, we have $\operatorname{codeg}(x_{24},u)\leq 2$, and so $y_{2},y_{4}\notin N[u]$. Also $y_{2}\neq y_{4}$, as otherwise the vertex $y=y_{2}=y_{4}$ would have $\operatorname{codeg}(v_{2},v_{4})\geq 3$, sharing neighbors $u,x_{24},y$. Since $x_{13}\not\sim x_{24}$, it must be that $y_{2},y_{4}$ are distinct from $x_{13}$. Likewise, we have $z_{2}\neq z_{4}$ since otherwise the vertex $z=z_{2}=z_{4}$ would have two 5-neighbors $v_{2},v_{4}$. This gives Case IV. ∎
We thus choose a vertex $u$ and an enumeration of its neighbors $v_{1},\dots,v_{5}$ to satisfy Theorem 57. We then use the following branching rules:
Case I. Split on $v_{1},v_{2},v_{3}$, getting subproblems $G_{1}=G-N[v_{1}],G_{2}=G-v_{1}-N[v_{2}],G_{3}=G-\{v_{1},v_{2}\}-N[v_{3}],G_{4}=G-\{v_{1},v_{2},v_{3}\}$. By Observation 52, these have drops $(2,5),(2.5,6),(3,7),(1,3)$ directly. In $G_{4}$, we apply Proposition 30 to 2-vertex $u$, where by Observation 52 and Observation 54 we have $r=\deg_{G_{4}}(v_{4})+\deg_{G_{4}}(v_{5})-\operatorname{codeg}_{G_{4}}(v_{4},v_{5})-1\geq 6$ irrespective of the degrees of $v_{4},v_{5}$. We need:
$e^{-2a-5b}+e^{-2.5a-6b}+e^{-3a-7b}+e^{-1.5a-3b}\psi_{6}\leq 1$ (5-19)
Case IIa: $\boldsymbol{\operatorname{shad}(v_{3},N[v_{1},v_{4}])\geq 1}$. Split on $v_{3},v_{5}$ and apply (A2) to vertex $v_{1}$ in $G-\{v_{3},v_{5}\}$. This generates subproblems $G_{1}=G-N[v_{3}],G_{2}=G-v_{3}-N[v_{5}],G_{3}=G-\{v_{3},v_{5}\}-N[v_{1},v_{4}],G_{4}=G-\{v_{1},v_{3},v_{5}\}$. Since $G$ is in 5BP, and $\deg_{G}(v_{3})=5,\deg_{G}(v_{5})=4,v_{3}\not\sim v_{5}$, subproblems $G_{1},G_{2}$ both have drop $(2,5)$ directly. Case IIa gives $\operatorname{minsurp}(G_{3})\geq 0$; since $\operatorname{codeg}(v_{1},v_{4})=1$, it has drop $(4,10)$. Subproblem $G_{4}$ has drop $(1.5,3)$ directly, and we also apply Proposition 30 to 2-vertex $u$ in $G_{4}$, noting that $\deg_{G_{4}}(v_{2})+\deg_{G_{4}}(v_{4})-1-\operatorname{codeg}_{G_{4}}(v_{2},v_{4})=4+4-1-1=6$.
Overall, we need:
$e^{-2a-5b}+e^{-2a-5b}+e^{-4a-10b}+e^{-1.5a-3b}\psi_{6}\leq 1$ (5-20)
Case IIb: $\boldsymbol{\operatorname{shad}(v_{3},N[v_{1},v_{4}])\leq 0}$. Split on $v_{3},v_{4}$, then apply (A2) to vertex $v_{1}$ in $G-\{v_{3},v_{4}\}$. We get subproblems $G_{1}=G-N[v_{3}],G_{2}=G-v_{3}-N[v_{4}],G_{3}=G-\{v_{3},v_{4}\}-N[v_{1},v_{5}],G_{4}=G-\{v_{1},v_{3},v_{4}\}$. Since $G$ is in 5BP, subproblems $G_{1},G_{2}$ both have drop $(2,5)$.
In subproblem $G_{2}$, consider vertex $v_{1}$; it has degree 4 since $\operatorname{codeg}(v_{1},v_{4})=1$ and $v_{1}\not\sim v_{3}$. Case IIb gives $\operatorname{shad}_{G_{2}}(v_{1})=\operatorname{shad}(v_{3},N[v_{1},v_{4}])\leq 0$, that is, $v_{1}$ is a blocked 4-vertex in $G_{2}$. Also, $\operatorname{minsurp}(G_{2})\geq\operatorname{shad}(N[v_{5}])-1\geq 0$. So, either $S(G_{2})\geq 1$ or we can apply Proposition 25 to $G_{2}$.
For subproblem $G_{3}$, Lemma 53 gives $\operatorname{minsurp}(G_{3})\geq\operatorname{shad}(N[v_{1},v_{5}])-2\geq-1$. Also, since $\operatorname{codeg}(v_{3},v_{4})=1$, we have $\operatorname{mindeg}(G_{3})\geq\operatorname{mindeg}(G-N[v_{1},v_{5}])-1\geq 1$. Now $G_{3}$ is principal with $\Delta k=10$ and excess two; if $\operatorname{minsurp}(G_{3})\geq 0$, it has drop $(4,10)$ and if $\operatorname{minsurp}(G_{3})=-1$ it has drop $(3.5,11)$ after applying (P1).
Finally, subproblem $G_{4}$ has drop $(1.5,3)$ directly. We apply Proposition 30 to 2-vertex $u$ in $G_{4}$, noting that $\deg_{G_{4}}(v_{2})+\deg_{G_{4}}(v_{5})-1-\operatorname{codeg}(v_{2},v_{5})=4+4-2=6$. Overall we need:
$e^{-2a-5b}+e^{-2a-5b}\max\{e^{-b},e^{-a-3b}+e^{-a-5b},e^{-0.5a-2b}+e^{-2a-5b}\}$
$\qquad\qquad+\max\{e^{-3.5a-11b},e^{-4a-10b}\}+e^{-1.5a-3b}\psi_{6}\leq 1$ (5-21)
Case III: We claim that there is some good cover $C$ satisfying one of the following three cases: (i) $v_{1}\in C$; or (ii) $v_{1}\notin C,v_{4}\notin C$; or (iii) $v_{1}\notin C,v_{4}\in C,v_{5}\notin C$. For, if not, then consider a good cover $C$ where $v_{1}\notin C,v_{4}\in C,v_{5}\in C$. Then $C^{\prime}=C\cup\{v_{1}\}\setminus\{u\}$ would be another cover, of the same size, which would be in case (i).
So our strategy is to branch on subproblems $G_{1}=G-v_{1},G_{2}=G-N[v_{1},v_{4}],G_{3}=G-v_{4}-N[v_{1},v_{5}]$. Here, $G_{1}$ has drop $(0.5,1)$. By Lemma 53, $\operatorname{minsurp}(G_{2})=\operatorname{shad}(N[v_{1},v_{4}])\geq 1$ and $\operatorname{minsurp}(G_{3})\geq\operatorname{shad}(N[v_{1},v_{5}])-1\geq 0$, so $G_{2},G_{3}$ have drops $(3,8),(3.5,9)$ respectively. We need:
$e^{-0.5a-b}+e^{-3a-8b}+e^{-3.5a-9b}\leq 1$ (5-22)
Case IV: Split on the independent 5-vertices $u,x_{13},x_{24}$, generating subproblems $G_{1}=G-N[u],G_{2}=G-u-N[x_{13}],G_{3}=G-\{u,x_{13}\}-N[x_{24}],G_{4}=G-\{x_{14},x_{23},u\}$. Note that $\operatorname{minsurp}(G_{3})\geq\operatorname{shad}(N[x_{24}])-2\geq 0$ and $\operatorname{minsurp}(G_{4})\geq\operatorname{shad}(x_{14})-2\geq 0$. So these have drops $(2,5),(2.5,6),(3,7),(1.5,3)$ directly.
Consider graph $G_{2}$. By Observation 52, we have $\operatorname{minsurp}(G_{2})\geq\operatorname{shad}(N[x_{13}])-1\geq 1$, in particular $\operatorname{mindeg}(G_{2})\geq 2$. Also note that $v_{4}\not\sim v_{4}$ and both $z_{2},z_{4}$ remain in the graph $G_{2}$, since $z_{2},z_{4}$ are 4-vertices which already have a 5-neighbors $v_{2}$ or $v_{4}$ respectively.
We claim that $S(G_{2})\geq 2$; there are a few different cases depending on $|N(x_{13})\cap\{y_{2},y_{4}\}|$. First, suppose $x_{13}\sim y_{2}$ and $x_{13}\sim y_{4}$; then, in the graph $G_{2}$, vertices $v_{2}$ and $v_{4}$ have degree two (losing neighbors $v_{1},u,y_{2}$ and $v_{3},u,y_{4}$ respectively). So by Observation 11 we immediately have $S(G_{2})\geq 2$. Next, suppose $|N(x_{13})\cap\{y_{2},y_{4}\}|=1$, say without loss of generality $x_{13}\sim y_{2},x_{13}\not\sim y_{4}$. In this case, we apply (P2) to the 2-vertex $v_{2}$ in $G_{2}$; in the resulting graph $G_{2}^{\prime}$, the vertices $z_{2},x_{24}$ get merged into a new vertex $z^{\prime}$. In $G_{2}^{\prime}$, the vertex $v_{4}$ has neighbors $z_{4},z^{\prime},y_{4}$ where $z^{\prime}\sim y_{4}$ in $G_{2}^{\prime}$ since $x_{24}\sim y_{4}$ in $G_{2}$. So we can apply (P3) to 3-triangle $v_{2}^{\prime}$, yielding $S(G_{2}^{\prime})\geq 1$ and hence $S(G_{2})\geq 2$. Finally suppose $x_{13}\not\sim\{y_{2},y_{4}\}$; then in $G_{2}$, the vertex $v_{2}$ has degree 3 with a triangle $v_{1},y_{2}$. We apply (P3) to $v_{2}$; let $G_{2}^{\prime}$ denote the resulting graph, where we remove $z_{2}$ and add edges from $v_{1},y_{2}$ to the neighbors of $z_{2}$. Since $v_{4}\not\sim v_{2}$, the degree of $v_{4}$ does not change in $G_{2}^{\prime}$, and in particular $v_{4}$ still has degree 3 with a triangle $v_{3},y_{4}$. So $S(G_{2}^{\prime})\geq 1$ and $S(G_{2})\geq 2$.
Consider graph $G_{3}$; here, vertex $v_{1}$ loses neighbors $u,x_{13},v_{2}$ and likewise $v_{3}$ loses neighbors $u,x_{13},v_{4}$). Moreover $\operatorname{minsurp}(G_{3})\geq\operatorname{minsurp}(G_{24})-2\geq 0$. Thus, by Observation 11 we have $S(G_{3})\geq 2$.
So $G_{2}$ has drop $(2.5,8)$ after simplification and $G_{3}$ has drop $(3,9)$ after simplification. We need:
$e^{-2a-5b}+e^{-2.5a-8b}+e^{-3a-9b}+e^{-1.5a-3b}\leq 1$ (5-23)
10 Branching on 6-vertices
Similar to the degree-5 setting, we will have a few algorithms targeted for different densities of degree-6 vertices in the graph. We will show the following result:
Theorem 58.
Branch6$(G,k)$ has measure $a_{6}\mu+b_{6}k$ for $a_{6}=0.166944,b_{6}=0.166214$.
Combined with Lemma 10, this gives the $O^{*}(1.25214^{k})$ algorithm for degree-6 graphs. The first algorithm is very simple, and is targeted toward the regime where degree-6 vertices are sparse.
Simplify $G$ If $\operatorname{maxdeg}(G)\leq 5$, then run the cheaper of Branch5-3 or the MaxIS-5 algorithm. Apply Section 5 to a vertex of degree at least $6$
Algorithm 8 Function Branch6-1$(G,k)$
Proposition 59.
Branch6-1 has measure $a_{61}\mu+b_{61}k$ for $a_{61}=0.254135,b_{61}=0.137360$.
Proof.
Given the potential branch-seqs from Section 5, we have the following constraints:
$\max\{e^{-a-3b}+e^{-a-5b},e^{-0.5a-2b}+e^{-2a-5b},e^{-0.5a-b}+e^{-2.5a-6b}\}\leq 1$ (6-1)
$\min\{a_{53}\mu+b_{53}k,2(k-\mu)\log(1.17366)\}\leq a\mu+bk$ (6-2)
The latter constraint can be routinely checked by LP algorithms. ∎
The more advanced algorithms try to branch on multiple 6-vertices in order to get bonus simplifications. These algorithms differ only in a few small steps, so we we describe them jointly.
1 Apply miscellaneous branching rules to ensure $G$ has nice structure e.g. max degree 6. Split on a 6-vertex $u$ where $G-N[u]$ is 2-vul. Branch on vertices $u,v$ with $\deg_{G}(u)=6,\deg_{G-u}(v)\geq 5$, and either $R(u,v)\geq 1$ or $\operatorname{shad}(u,N[v])\leq 0$. Branch on non-adjacent 6-vertices $x_{1},x_{2},x_{3}$ with $R(x_{1},x_{2},x_{3})\geq 1$. Branch on a 6-vertex $u$ with six 6-neighbors, or five 6-neighbors and one 5-neighbor. if $L=2$ then
2        Branch on non-adjacent 6-vertices $x_{1},x_{2}$ with $R(x_{1},N[x_{2}])\geq 1$. Branch on a 5-vertex $u$ which has two 6-neighbors and three 5-neighbors,            UNLESS each of these 6-neighbors itself has at least four 5-neighbors.
Run the cheaper of Branch6-$(L-1)(G,k)$ or the MaxIS algorithm
Algorithm 9 Function Branch6-$L$$(G,k)$ for $L=2,3$
We will show the following:
Theorem 60.
Branch6-2 has measure $a_{62}\mu+b_{62}k$ for $a_{62}=0.202348,b_{62}=0.154382$.
Branch6-3 has measure $a_{63}\mu+b_{63}k$ for $a_{63}=0.166944,b_{63}=0.166214$
10.1 Regularizing the graph and structural properties
There is a grab bag of isolated local obstructions that we will remove. We say that $G$ is in 6-base position (or 6BP for short) if the following conditions hold:
  • $G$ is simplified.
  • $\operatorname{maxdeg}(G)\leq 6$.
  • Every 6-vertex $u$ has $\operatorname{shad}(N[u])\geq 2$ and $R(u)=0$.
  • Every 5-vertex $u$ has $\operatorname{shad}(N[u])\geq 1$ and $R(u)=0$.
  • There are no vertices $u,v$ with $\operatorname{codeg}(u,v)\geq 2$ and $R(u,v)+R(N(u)\cap N(v))+\operatorname{codeg}(u,v)\geq 5$.
Observation 61.
  • If $G$ is in 5BP, then any 6-vertex $u$ and vertex $v\not\sim u$ have $\operatorname{codeg}(u,v)\leq\deg(v)-3$.
  • If $G$ is in 5BP, then any 5-vertex $u$ and vertex $v\not\sim u$ have $\operatorname{codeg}(u,v)\leq 2$.
  • After Line 3, any vertices $u,v$ with $\deg(u)=6$ and $\deg_{G-u}(v)\geq 5$ have $\operatorname{shad}(u,v)\geq 2$ and $\operatorname{shad}(u,N[v])\geq 1$.
  • After Line 4, any non-adjacent 6-vertices $x_{1},x_{2},x_{3}$ have $\operatorname{shad}(x_{1},x_{2},x_{3})\geq 2$.
Proof.
In the first case, if $\operatorname{codeg}(u,v)\geq\deg(v)-2$, then $\operatorname{shad}(N[u])\leq 1$. In the second case, we may suppose that also $\deg(v)\leq 5$ (else it is covered by the first case); then the analysis is the same as Proposition 47. In the third case, if $\operatorname{shad}(u,v)\leq 1$ or $\operatorname{shad}(u,N[v])\leq 0$, we would branch at Line 3. In the final case, we have $\operatorname{shad}(x_{1},x_{2},x_{3})\geq\operatorname{shad}(x_{1},x_{2})-1\geq 1$. If $\operatorname{shad}(x_{1},x_{2},x_{3})\leq 1$, then we would branch at Line 4. ∎
Proposition 62.
After Line 4, any 5-vertex $u$ has at most three 6-neighbors; moreover, if $u$ has three 6-neighbors, its other two neighbors have degree 4.
Proof.
Suppose $u$ has 6-neighbors $v_{1},v_{2},v_{3}$. Let $v_{4},v_{5}$ be the other two neighbors of $u$. They must have degree at least 4, else $R(u)\geq 1$ contradicting that $G$ is in 6BP. If $v_{1},v_{2},v_{3}$ are independent, then $R(v_{1},v_{2},v_{3})\geq 1$ due to 2-vertex $u$, and we would apply Line 4 to $v_{1},v_{2},v_{3}$. So we may suppose without loss of generality that $v_{1}\sim v_{2}$.
Now suppose for contradiction that $\deg(v_{4})\geq 5$. If $v_{3}\sim v_{4}$, then $R(v_{1},v_{2})\geq 1$ due to 3-triangle $u,v_{3},v_{4}$, and we would branch on $v_{1},v_{2}$ at Line 3. Otherwise, if $v_{3}\not\sim v_{4}$, then $R(v_{3},v_{4})\geq 1$ due to 3-triangle $u,v_{1},v_{2}$, and we would branch on $v_{3},v_{4}$ at Line 3. ∎
Proposition 63.
Before Line 9 in the algorithm Branch6-2, the graph has $5n_{4}+n_{5}\geq 2n_{6}$.
Before Line 9 in the algorithm
Branch6-3, the graph has $5n_{4}+2n_{5}\geq 2n_{6}$.
Proof.
For $i\geq 0$, let $A_{6i}$ be the set of 6-vertices with exactly $i$ 5-neighbors and let $A_{5i}$ be the set of 5-vertices with exactly $i$ 6-neighbors. By Proposition 62, $A_{54}=A_{55}=\emptyset$ and any vertex in $A_{53}$ has at least two 4-neighbors. Also let $A_{52}^{\prime}\subseteq A_{52}$ be the set of 5-vertices with two 6-neighbors and at least one 4-neighbor. By double-counting edges between $V_{5}$ and $V_{6}$, we have:
$|A_{51}|+2|A_{52}|+3|A_{53}|=|A_{61}|+2|A_{62}|+3|A_{63}|+4|A_{64}|+5|A_{65}|+6|A_{66}|$ (7)
Due to the branching rule at Line 5, any vertex in $A_{60}\cup A_{61}$ has at least one 4-neighbor. Each 4-vertex has at most one 6-neighbor (else we would branch on the 6-neighbors at Line 3), and so:
$|V_{4}|\geq|A_{60}|+|A_{61}|$ (8)
Now consider the edges between $V_{4}$ and $V_{5}\cup V_{6}$. Each vertex in $A_{60}\cup A_{61}$ contributes one such edge, each vertex in $A_{52}^{\prime}$ contributes at least one edge, each vertex in $A_{53}$ contributes at least two edges by Proposition 62, and each vertex in $V_{4}$ has at most four such edges. So we get:
$4|V_{4}|$ $\geq|A_{60}|+|A_{61}|+2|A_{53}|+|A_{52}^{\prime}|$ (9)
Adding inequalities (7), (8), (9) and collecting terms gives:
$5|V_{4}|+|A_{51}|+2|A_{52}|+|A_{53}|-|A_{52}^{\prime}|\geq 2|A_{60}|+3|A_{61}|+3|A_{63}|+4|A_{64}|+5|A_{65}|+6|A_{66}|$ (10)
Since $|V_{5}|=|A_{51}|+|A_{52}|+|A_{53}|$ and $|V_{6}|=|A_{60}|+|A_{61}|+\dots+|A_{66}|$, this implies $5n_{4}+2n_{5}\geq 2n_{6}$.
In the algorithm Branch5-2, further consider any vertex $u\in A_{52}\setminus A_{52}^{\prime}$. Both the 6-neighbors of $u$ must be in $A_{64}\cup A_{65}\cup A_{66}$ (else we would have branched on $u$ at Line 8). By double-counting the edges between $A_{52}\setminus A_{52}^{\prime}$ and $V_{6}$, we get:
$4|A_{64}|+5|A_{65}|+6|A_{66}|\geq 2(|A_{52}|-|A_{52}^{\prime}|)$ (11)
Now adding inequality (10) to $\tfrac{1}{2}$ times inequality (11), and collecting terms again, gives:
$5|V_{4}|+|A_{51}|+|A_{52}|+|A_{53}|\geq 2|A_{60}|+3|A_{61}|+3|A_{63}|+2|A_{64}|+2.5|A_{65}|+3|A_{66}|$ (12)
which in turn implies $5n_{4}+n_{5}\geq 2n_{6}$ as desired. ∎
10.2 Basic analysis
We begin by analyzing the easier steps of the algorithm; Lines 5 and 8 will be handled later.
Line 1. We first simplify $G$. If a vertex $u$ of degree 5 or 6 has $\operatorname{shad}(N[u])\leq 5-\deg(u)$, we apply Lemma 26. If a vertex $u$ of degree 5 or 6 has $R(u)\geq 1$, we split on $u$ with branch-seq $[(0.5,2),(2,5)]$. If a vertex has degree at least 7, we apply Section 5. If there are vertices $u,v$ with $R(u,v)+R(N(u)\cap N(v))+\operatorname{codeg}(u,v)\geq 5$, we apply (A3).
Over all cases, we get branch-seq $[(1,3),(1,4)]$, $[(1,2),(1,5)]$, or $[(0.5,2),(2,5)]$. We need:
$\max\{e^{-a-3b}+e^{-a-4b},e^{-a-2b}+e^{-a-5b},e^{-0.5a-2b}+e^{-2a-5b}\}\leq 1$ (6-3)
Line 2. Since $G$ is in 6BP, subproblems $G-u$ and $G-N[u]$ have drop $(0.5,1)$ and $(2.5,6)$ directly, and we can apply Proposition 31 to the latter problem. We need:
$e^{-0.5a-b}+e^{-2.5a-6b}\gamma_{2}\leq 1$ (6-4)
Line 3. Split on $u$, and in subproblem $G-u$, either apply Lemma 26 or split on $v$. Since $G$ is in 6BP, subproblems $G-N[u],G-u$ have drops $(2.5,6),(0.5,1)$ directly and are simplified. We need:
$e^{-2.5a-6b}+e^{-0.5a-b}\max\{e^{-b},e^{-a-3b}+e^{-a-4b},e^{-0.5a-2b}+e^{-2a-5b}\}\leq 1$ (6-5)
Line 4. Split on $x_{1},x_{2},x_{3}$, getting subproblems $G_{1}=G-N[x_{1}],G_{2}=G-x_{1}-N[x_{2}],G_{3}=G-\{x_{1},x_{2}\}-N[x_{3}],G_{4}=G-\{x_{1},x_{2},x_{3}\}$. By Observation 61 we have $\operatorname{minsurp}(G_{1})\geq 2,\operatorname{minsurp}(G_{2})\geq 1,\operatorname{minsurp}(G_{3})\geq\operatorname{shad}(N[x_{3}])-2\geq 0$ and $\operatorname{minsurp}(G_{4})\geq\operatorname{shad}(x_{1},x_{2})-1\geq 1$. Because $S(G_{4})\geq 1$, these subproblems $G_{1},G_{2},G_{3},G_{4}$ have drops $(2.5,6),(3,7),(3.5,8),(2,5)$. We need:
$e^{-2.5a-6b}+e^{-3a-7b}+e^{-3.5a-8b}+e^{-1.5a-4b}\leq 1$ (6-6)
Line 7. Since $G$ is in 6BP, subproblems $G-N[x_{1}]$ and $G-x_{1}-N[x_{2}]$ and $G-\{x_{1},x_{2}\}$ have drops $(2.5,6),(3,7),(1,2)$ directly, and $G-x_{1}-N[x_{2}]$ has drop $(3,8)$ after simplification. We need:
$e^{-2.5a-6b}+e^{-3a-8b}+e^{-a-2b}\leq 1\qquad\text{(only for $L=2$)}$ (6-7)
Line 9. We use a more refined property of the MaxIS-6 algorithm of [XN17b]; specifically, it has runtime $O^{*}(1.18922^{w_{3}n_{3}+w_{4}n_{4}+w_{5}n_{5}+n_{6}})$ for constants $w_{3}=0.49969,w_{4}=0.76163,w_{5}=0.92401$.
For Branch6-2 we have $5n_{4}+n_{5}\geq 2n_{6}$. It can be checked that the maximum value of $w_{3}n_{3}+w_{4}n_{4}+w_{5}n_{5}+n_{6}$, subject to constraints $n_{3}+n_{4}+n_{5}+n_{6}=n,n_{i}\geq 0,5n_{4}+n_{5}\geq 2n_{6}$, occurs at $n_{5}=2n/3,n_{6}=n/3,n_{3}=n_{4}=0$. So the runtime is at most $1.18922^{n(w_{5}(2/3)+(1/3))}$ and we need:
$\min\{a_{61}\mu+b_{61}k,2(k-\mu)\cdot(w_{5}\cdot 2/3+1/3)\log(1.18922)\}\leq a\mu+bk$ (6-8)
which can also be checked mechanically.
For Branch6-3 we have $5n_{4}+2n_{5}\geq 2n_{6}$. Again, the maximum value of $w_{3}n_{3}+w_{4}n_{4}+w_{5}n_{5}+n_{6}$, subject to constraints $n_{3}+n_{4}+n_{5}+n_{6}=n,n_{i}\geq 0,5n_{4}+2n_{5}\geq 2n_{6}$, occurs at $n_{6}=n_{5}=n/2,n_{3}=n_{4}=0$. So the overall runtime is at most $1.18922^{n(w_{5}(1/2)+(1/2))}$ and we need:
$\min\{a_{62}\mu+b_{62}k,2(k-\mu)\cdot(w_{5}/2+1/2)\log(1.18922)\}\leq a\mu+bk$ (6-9)
which can also be checked mechanically. This shows Theorem 60.
10.3 Line 5: branching on a 6-vertex with high-degree neighbors
Consider a vertex $u$ here; we denote the induced neighborhood graph by $H=G[N(u)]$.
Proposition 64.
The neighbors of $u$ can be ordered as $v_{1},\dots,v_{6}$ to satisfy one of the following conditions:
  1. (I)
    $v_{1},\dots,v_{5}$ are independent and $\deg(v_{1})=\deg(v_{2})=\deg(v_{3})=\deg(v_{4})=\deg(v_{5})=6$.
  2. (II)
    $H$ is a matching and $v_{1}\sim v_{2},v_{3}\sim v_{4},\deg_{G-v_{6}}(v_{5})=5,\deg(v_{1})=\deg(v_{3})=\deg(v_{4})=\deg(v_{6})=6$.
  3. (III)
    $v_{4}\sim v_{5},v_{4}\sim v_{6}$ and $\deg_{G}(v_{1})=6,\deg_{G-v_{1}}(v_{2})\geq 5,\deg_{G-\{v_{1},v_{2}\}}(v_{3})\geq 5$.
Proof.
First observe that $H$ cannot have any triangle $v_{1},v_{2},v_{3}$; for, in that case, we would have $R(v_{4},v_{5})\geq 1$ where $v_{4},v_{5}$ are any other 6-neighbors of $u$; we would branch on $v_{4},v_{5}$ at Line 3.
If $H$ has no edges, then let $v_{1},\dots,v_{5}$ be any five 6-neighbors of $u$, and we get Case I.
If $H$ is a matching with exactly one edge $(v_{1},v_{2})$, with three other 6-neighbors $v_{3},v_{4},v_{5}$, we would have $R(v_{3},v_{4},v_{5})\geq 1$ due to the 3-triangle $u,v_{1},v_{2}$ and we would branch Line 4.
If $H$ is a matching with exactly two edges, suppose without loss of generality they are $(v_{1},v_{2}),(v_{3},v_{4})$ where $\deg(v_{4})=6$. If $\deg(v_{5})=\deg(v_{6})=6$, then again have $R(v_{4},v_{5},v_{6})\geq 1$ due to 3-triangle $u,v_{1},v_{2}$, and we would branch at Line 4. So suppose without loss of generality that $\deg(v_{5})=5$, in which case the ordering $v_{1},\dots,v_{6}$ satisfies Case II.
If $H$ is a matching with exactly three edges, we can arrange them as $(v_{1},v_{2}),(v_{3},v_{4}),(v_{5},v_{6})$, where all vertices except $v_{2}$ have degree 6. This also satisfies Case II.
Next suppose $H$ is not a matching. First, suppose some $v_{4}\in H$ has neighbors $v_{5},v_{6}\in H$ where one of the vertices $v_{4},v_{5},v_{6}$ has degree 5. Since $H$ has no triangle, we order the other vertices as $v_{1},v_{2},v_{3}$ where $v_{2}\not\sim v_{3}$, satisfying Case III.
Finally, suppose $v_{4}\in H$ has neighbors $v_{5},v_{6}\in H$ where $\deg(v_{4})=\deg(v_{5})=\deg(v_{6})=6$. The induced subgraph $H^{\prime}=G[\{v_{1},v_{2},v_{3}\}]$ must be a matching, as otherwise it would be covered by the argument in the preceding paragraph. So let $x$ be a vertex which is isolated in $H^{\prime}$. If $\deg(x)=6$, we set $v_{1}=x$ and $v_{2}$ to be the other 5-vertex (if any) and $v_{3}$ to be the 6-vertex. If $\deg(x)=5$, we set $v_{1},v_{2}$ to be the 6-neighbors of $u$ and $v_{3}=x$. ∎
See Figure 8 for an illustration.
Figure 8: Possible cases for the neighborhood of $u$. Dashed edges may or may not be present.
In light of Proposition 64, we use the following branching rules for $u$. When analyzing a given case, we assume no earlier case was possible for any another vertex ordering.
Case Ia: $\boldsymbol{G-\{v_{1},v_{2}\}-N[v_{3}]}$ is 3-vul. Split on $v_{1},\dots,v_{4}$, generating subproblems $G_{1}=G-N[v_{1}],G_{2}=G-v_{1}-N[v_{2}],G_{3}=G-\{v_{1},v_{2}\}-N[v_{3}],G_{4}=G-\{v_{1},v_{2},v_{3}\}-N[v_{4}],G_{5}=G-\{v_{1},v_{2},v_{3},v_{4}\}$. By Observation 61, we have $\operatorname{minsurp}(G_{1})\geq 2$, $\operatorname{minsurp}(G_{2})\geq 1$, $\operatorname{minsurp}(G_{3})\geq\operatorname{shad}(N[v_{3}])-2\geq 0$, $\operatorname{minsurp}(G_{4})\geq\operatorname{shad}(N[v_{4}])-3\geq-1$ and $\operatorname{minsurp}(G_{5})\geq\operatorname{shad}(v_{1},v_{2},v_{3})-1\geq 1$. So $G_{1},G_{2},G_{3},G_{4},G_{5}$ have drop $(2.5,6),(3,7),(3.5,8),(3.5,9),(2,4)$ directly. We can apply Proposition 31 to $G_{3}$, and have $S(G_{5})\geq 1$ due to 2-vertex $u$. We need:
$e^{-2.5a-6b}+e^{-3a-7b}+e^{-3.5a-8b}\gamma_{3}+e^{-3.5a-9b}+e^{-2a-5b}\leq 1$ (6-10)
Case Ib: Split on $v_{1},v_{2},v_{3},v_{4}$, and apply (A1) to vertex $u$ in the subproblem $G-\{v_{1},v_{2},v_{3},v_{4}\}$. This generates subproblems $G_{1}=G-N[v_{1}],G_{2}=G-v_{1}-N[v_{2}],G_{3}=G-\{v_{1},v_{2}\}-N[v_{3}],G_{4}=G-\{v_{1},v_{2},v_{3}\}-N[v_{4}],G_{0}=G-N[u]$ and $G_{5}=G-\{v_{1},v_{2},v_{3},v_{4}\}-N[v_{5},v_{6}]$.
Since $G$ is 6BP, $G_{0},G_{1},G_{2},G_{3}$ have drops $(2.5,6),(2.5,6),(3,7),(3.5,8)$ directly. We must have $\operatorname{minsurp}(G_{4})\geq 0$, as otherwise $\operatorname{shad}(v_{1},v_{2},N[v_{4}])\leq 0$, which would put us into Case Ia (with swapping the order of $v_{3}/v_{4}$). So $G_{4}$ has drop $(4,9)$.
For subproblem $G_{5}$, let $d=\deg_{G-N[v_{5}]}(v_{6})$; by Observation 61 we have $d\geq 3$. We claim that $\operatorname{minsurp}(G_{5})\geq 2-d$. For, consider an indset $I$ with $\operatorname{surp}_{G_{5}}(I)\leq 1-d\leq-2$; necessarily $|I|\geq 2$. Then $\operatorname{surp}_{G-\{v_{1},v_{2}\}-N[v_{5}]}(I\cup\{v_{6}\})\leq(1-d)+(d-1)+1=2$. So $G-\{v_{1},v_{2}\}-N[v_{5}]$ is 3-vul, putting us into Case Ia with alternate vertex ordering $v_{1}^{\prime}=v_{1},v_{2}^{\prime}=v_{2},v_{3}^{\prime}=v_{5}$. So $G_{5}$ is principal with excess two and $\Delta k=10+d\geq 13$ and $\Delta\mu\geq 5$.
Putting all the subproblems together, we need:
$e^{-2.5a-6b}+e^{-2.5a-6b}+e^{-3a-7b}+e^{-3.5a-8b}+e^{-4a-9b}+e^{-5a-13b}\leq 1$ (6-11)
Case IIa: $\boldsymbol{G-v_{6}-N[v_{1}]}$ is 3-vul. Split on $v_{6},v_{1},v_{5}$, obtaining subproblems $G_{1}=G-N[v_{6}],G_{2}=G-v_{6}-N[v_{1}],G_{3}=G-\{v_{1},v_{6}\}-N[v_{5}],G_{4}=G-\{v_{1},v_{6},v_{5}\}$. By Observation 61, we have $\operatorname{minsurp}(G_{1})\geq 2,\operatorname{minsurp}(G_{2})\geq 1,\operatorname{minsurp}(G_{3})\geq\operatorname{shad}(v_{6},N[v_{5}])-1\geq 0,\operatorname{minsurp}(G_{4})\geq\operatorname{shad}(v_{1},v_{6})-1\geq 1$. So $G_{1},G_{2},G_{3},G_{4}$ have drops $(2.5,6),(3,7),(3,7),(1.5,3)$ directly. We apply Proposition 31 to $G_{3}$ and we have $S(G_{4})\geq 1$ (due to 3-triangle $u,v_{3},v_{4}$), so overall we need:
$e^{-2.5a-6b}+e^{-3a-7b}\gamma_{3}+e^{-3a-7b}+e^{-1.5a-4b}\leq 1$ (6-12)
Case IIb: Split on $v_{6},v_{4},v_{3}$ and apply (A2) to vertex $v_{1}$ in subproblem $G-\{v_{6},v_{4},v_{3}\}$. Overall, this generates subproblems $G_{1}=G-N[v_{6}],G_{2}=G-v_{6}-N[v_{4}],G_{3}=G-v_{6}-N[v_{3}],G_{4}=G-\{v_{3},v_{4},v_{6},v_{1}\}$ and $G_{5}=G-\{v_{3},v_{4},v_{6}\}-N[v_{1},v_{5}]$. By Observation 61, we have $\operatorname{minsurp}(G_{1})\geq 2$ and $\operatorname{minsurp}(G_{2})\geq 1$ and $\operatorname{minsurp}(G_{3})\geq 1$ and $\operatorname{minsurp}(G_{4})\geq\operatorname{shad}(v_{4},v_{6})-2\geq 0$. So the subproblems have drops $(2.5,6),(3,7),(3,7),(2,4)$ directly; subproblem $G_{4}$ has a 2-vertex $u$, so it has drop $(2,5)$ after simplification.
For subproblem $G_{5}$, let $d=\deg_{G-v_{6}-N[v_{1}]}(v_{5})$. We have $d\geq 3$, as otherwise $R(v_{6},N[v_{1}])\geq 1$ and it would be covered in Case IIa. We must have $\operatorname{minsurp}(G_{5})\geq 2-d$; for, if we have an indset $I$ with $\operatorname{surp}_{G_{5}}(I)\leq 1-d\leq-2$, where necessarily $|I|\geq 2$, then $\operatorname{surp}_{G-v_{6}-N[v_{1}]}(I\cup\{v_{5}\})\leq 2$ and $G-v_{6}-N[v_{1}]$ is 3-vul; it would be covered in Case IIa.
Thus $G_{5}$ has $\Delta k=9+d\geq 12$ and $\Delta\mu\geq 4.5$. Putting all the subproblems together, we need:
$e^{-2.5a-6b}+e^{-3a-7b}+e^{-3a-7b}+e^{-2a-5b}+e^{-4.5a-12b}\leq 1$ (6-13)
Case III. Split on $v_{1},v_{2},v_{3}$, obtaining subproblems $G_{1}=G-N[v_{1}],G_{2}=G-v_{1}-N[v_{2}],G_{3}=G-\{v_{1},v_{2}\}-N[v_{3}],G_{4}=G-\{v_{1},v_{2},v_{3}\}$. By Observation 61, we have $\operatorname{minsurp}(G_{1})\geq 2$ and $\operatorname{minsurp}(G_{2})\geq 1$ and $\operatorname{minsurp}(G_{3})\geq\operatorname{shad}(v_{1},N[v_{3}])-1\geq 0$ and $\operatorname{minsurp}(G_{4})\geq\operatorname{shad}(v_{1},v_{2})-1\geq 1$. So the subproblems have drops $(2.5,6),(2.5,6),(3,7),(1.5,3)$ directly; furthermore, $G_{4}$ has a kite $u,v_{4},v_{5},v_{6}$ and so it has net drop $(2,5)$ by Lemma 9. We need:
$e^{-2.5a-6b}+e^{-2.5a-6b}+e^{-3a-7b}+e^{-2a-5b}\leq 1$ (6-14)
10.4 Line 8: Branching on a 5-vertex with two 6-neighbors
In the algorithm Branch6-2, we consider a 5-vertex $u$ with 6-neighbors $y_{1},y_{2}$ and 5-neighbors $v_{1},v_{2},v_{3}$, where either $y_{1}$ or $y_{2}$ has at most three 5-neighbors. See Figure 9 for an illustration.
Figure 9: The neighborhood of vertex $u$. The edge $(y_{1},y_{2})$ may or may not be present.
Proposition 65.
  • The vertices $v_{1},v_{2},v_{3}$ are independent and $\{y_{1},y_{2}\}\not\sim\{v_{1},v_{2},v_{3}\}$.
  • If any pair $y_{i},v_{j}$ share a vertex $x\neq u$, then $\deg(x)=5$ and $x\not\sim u$.
  • There is some pair $y_{i},v_{j}$ with $\operatorname{codeg}(y_{i},v_{j})=1$.
Proof.
  • Vertices $v_{1},v_{2},v_{3}$ are independent, as otherwise $R(y_{1},y_{2})\geq 1$ (due to 3-triangle $u$) and we would branch at Line 3. There must be some pair $y_{i},v_{j}$ with $y_{i}\not\sim v_{j}$ (else both $y_{1},y_{2}$ would have four 5-neighbors $u,v_{1},v_{2},v_{3}$); say without loss of generality that $y_{1}\not\sim v_{1}$. We have $y_{2}\not\sim\{v_{1},v_{2},v_{3}\}$, else $R(y_{1},v_{1})\geq 1$ due to the 3-triangle involving $u,y_{2},v_{j}$, and we would apply Line 3 to $y_{1},v_{1}$. This in turn implies that $y_{1}\not\sim\{v_{2},v_{3}\}$, else $R(y_{2},v_{1})\geq 1$ due to the resulting 3-triangle and we would apply Line 3 to $y_{2},v_{1}$.
  • Say without loss of generality that $y_{1},v_{1}$ share a vertex $x\neq u$. Since none of the neighbors of $u$ are adjacent to $v_{1}$, we have $x\not\sim u$. If $\deg(x)\leq 4$, then $R(y_{1},v_{1})\geq 1$ (due to subquadratic vertex $x$), and we would branch on $y_{1},v_{1}$ at Line 3. If $\deg(x)=6$ and $x\sim y_{2}$, then $R(N[x])\geq 1$ (due to vertex $u$ losing neighbors $y_{1},v_{1},y_{2}$) and we would branch at Line 2. If $\deg(x)=6$ and $x\not\sim y_{2}$, then $R(y_{2},N[x])\geq 1$ (again due to vertex $u$ losing neighbors $y_{1},v_{1},y_{2}$), and we would branch on $y_{2},x$ at Line 7.
  • Suppose $y_{i}$ shares neighbor $t_{ij}\neq u$ with $v_{j}$ for $j=1,2,3$; as we have just seen, $\deg(t_{ij})=5$ and $t_{ij}\not\sim u$ for all $j$. Vertices $t_{i1},t_{i2},t_{i3}$ must be distinct; for, if $t_{ij}=t_{ij^{\prime}}=t$, then $\operatorname{codeg}(t,u)\geq 3$ (sharing vertices $v_{j},v_{j^{\prime}},y_{i}$), contradicting Observation 61. So $y_{i}$ has four distinct 5-neighbors $t_{i1},t_{i2},t_{i3},u$. By hypothesis, this does not hold for both $y_{1}$ and $y_{2}$. ∎
Proposition 66.
Suppose $y_{1}\sim y_{2}$, and let $i\in\{1,2\},j\in\{1,2,3\}$. For the graph $G^{\prime}_{ij}=G-\{v_{1},v_{2},v_{3}\}-N[y_{i},v_{j}]$, we have $\operatorname{minsurp}(G^{\prime}_{ij})\geq-3+\operatorname{codeg}(y_{i},v_{j})$. Moreover, if $\operatorname{codeg}(y_{i},v_{j})=1$ and $\operatorname{minsurp}(G^{\prime}_{ij})=-2$, then either $S(G^{\prime}_{ij})\geq 1$ or $R(v_{j^{\prime}},N[v_{j}])\geq 1$ for some $j^{\prime}\in\{1,2,3\}$.
Proof.
Suppose without loss of generality $i=j=1$ and write $G^{\prime}=G^{\prime}_{ij}=G-\{v_{2},v_{3}\}-N[y_{1},v_{1}]$ for brevity. Let $I$ be a min-set of $G^{\prime}$. If $\operatorname{surp}_{G^{\prime}}(I)\leq-4+\operatorname{codeg}(y_{1},v_{1})$, then $\operatorname{surp}_{G-N[y_{1}]}(I\cup\{v_{1}\})\leq 2$. So $G-N[y_{1}]$ is 2-vul (via indset $I\cup\{v_{1}\}$) and we would branch on $y_{2}$ at Line 2.
Now suppose $\operatorname{codeg}(y_{1},v_{1})=1$ and $\operatorname{surp}_{G^{\prime}}(I)=-2$. If $|I|\geq 3$ then $S(G^{\prime})\geq 1$ as desired, so suppose $|I|=2$. By Observation 61, we have $\operatorname{codeg}(v_{2},v_{3})\leq 2$, so there must be some vertex $x\in I$ and $v_{j}\in\{v_{2},v_{3}\}$ with $x\not\sim v_{j}$; say without loss of generality $x\not\sim v_{3}$ and $x$ is isolated in $G^{\prime}$.
Let $d=\deg_{G}(x)$. By Observation 61 we have $\operatorname{codeg}(x,y_{1})\leq d-3$ and $\operatorname{codeg}(x,v_{1})\leq 2$. The only way for $x$ to be isolated in $G^{\prime}$ is to have $x\sim v_{2},\operatorname{codeg}(x,v_{1})=2,\operatorname{codeg}(x,y_{1})=d-3$. If $d\leq 5$, we thus have $R(v_{2},N[v_{1}])\geq 1$ due to subquadratic vertex $x$. If $d=6$, then by Proposition 66 we have $x\not\sim y_{2}$ (else the 6-vertex $x$ would be shared by $y_{2},v_{2}$); in this case, $G-N[x]$ has a 3-vertex $y_{1}$ with a triangle $y_{1},y_{2},u$, and we would branch on $x$ at Line 2. ∎
There are now a few different cases for the analysis.
Case I: $\boldsymbol{y_{1}\not\sim y_{2}}$. Split on $y_{1},y_{2},v_{1}$, getting subproblems $G_{1}=G-N[y_{1}],G_{2}=G-y_{1}-N[y_{2}],G_{3}=G-\{y_{1},y_{2}\}-N[v_{1}],G_{4}=G-\{y_{1},y_{2},v_{1}\}$. By Observation 61, we have $\operatorname{minsurp}(G_{1})\geq 2$ and $\operatorname{minsurp}(G_{2})\geq 1$ and $\operatorname{minsurp}(G_{3})\geq\operatorname{shad}(y_{1},v_{1})-1\geq 1$ and $\operatorname{minsurp}(G_{4})\geq\operatorname{shad}(y_{1},y_{2})-1\geq 1$. So $G_{1},G_{2},G_{3},G_{4}$ have drops $(2.5,6),(3,7),(3,7),(1.5,3)$ directly. We also apply Proposition 30 to 2-vertex $u$ in $G_{4}$, where $\deg(v_{2})+\deg(v_{3})-1-\operatorname{codeg}(v_{2},v_{3})\geq 7$ by Observation 61. So we need:
$e^{-2.5a-6b}+e^{-3a-7b}+e^{-3a-7b}+e^{1.5a-3b}\psi_{7}\leq 1$ (6-15)
Case II: $\boldsymbol{y_{1}\sim y_{2}}$. We choose some enumeration $v_{1},v_{2},v_{3}$, and split on $v_{1},v_{2}$; in the subproblem $G-\{v_{1},v_{2}\}$, we apply (A2) to $u,y_{1}$. This generates subproblems $G_{1}=G-N[v_{1}],G_{2}=G-v_{1}-N[v_{2}],G_{3}=G\{v_{1},v_{2}\}-N[v_{3},y_{1}],G_{4}=G-\{v_{1},v_{2},y_{1}\}$.
By Observation 61, $G_{1},G_{2},G_{4}$ have drops $(2,5),(2.5,6),(1.5,3)$ directly; furthermore, we can apply Proposition 30 to the 2-vertex $u$ in $G_{4}$, where $r=\deg(v_{1})+(\deg(y_{2})-1)-\operatorname{codeg}(y_{2},v_{1})-1\geq 5+5-2-1=7$ by Observation 61.
Proposition 66 and Observation 61 give $\operatorname{minsurp}(G_{3})\geq-3+\operatorname{codeg}(y_{1},v_{3})$ and $\operatorname{codeg}(y_{1},v_{3})\leq 2$. In light of Proposition 65, there are two cases to select the ordering:
Case IIa: $\boldsymbol{R(v_{1},N[v_{2}])\geq 1.}$ In this case $G_{2}$ has drop $(2.5,7)$ after simplification. Since $G_{3}$ has drop $(4,11)$, we need:
$e^{-2a-5b}+e^{-2.5a-7b}+e^{-4a-11b}+e^{-1.5a-3b}\psi_{7}\leq 1$ (6-16)
Case IIb: $\boldsymbol{\operatorname{codeg}(y_{1},v_{3})=1.}$ By Proposition 66, either $R(v_{j},N[v_{3}])\geq 1$ for some $j$ (which would put us into Case IIa, with some other vertex ordering), or $S(G_{3})\geq 1$, or $\operatorname{minsurp}(G_{3})\geq-1$. In the latter two cases, $G_{3}$ has drop $(4,13)$ or $(4.5,12)$. So we need:
$e^{-2a-5b}+e^{-2.5a-6b}+\max\{e^{-4a-13b},e^{-4.5a-12b}\}+e^{-1.5a-3b}\psi_{7}\leq 1$ (6-17)
11 Branching on 7-vertices and finishing up.
We first describe the degree-7 branching algorithm.
Simplify $G$ Branch on a vertex $u$ with degree larger than 7 Split on a 7-vertex $u$ with $\operatorname{shad}(N[u])\geq 0$. Branch on a 7-vertex $u$ which has a blocker $x$ with $\deg(x)\geq 5$. Branch on a vertex $x$ which is a blocker of three 7-vertices $u_{1},u_{2},u_{3}$. Run either Branch6-3 or the MaxIS-7 algorithm.
Algorithm 10 Function Branch7$(G,k)$
Theorem 67.
Branch7 has measure $a_{7}\mu+b_{7}k$ for $a_{7}=0.01266,b_{7}=0.221723$.
To show this, let us examine the algorithm line by line.
Line 2. Here, we apply Section 5; we need:
$\max\{e^{-a-3b}+e^{-a-4b},e^{-0.5a-2b}+e^{-2a-5b},e^{-0.5a-b}+e^{-2.5a-8b}\}\leq 1$ (7-1)
Line 3. In this case $G-N[u]$ has drop $(3,7)$. So we need:
$e^{-0.5a-b}+e^{-3a-7b}\leq 1$ (7-2)
Line 4. Apply (B) to $u,x$, getting subproblems $G-N[x],G-\{u,x\}$ with drops $(1,5)$ and $(1,2)$ respectively. We need:
$e^{-a-2b}+e^{-a-5b}\leq 1$ (7-3)
Line 5. Apply (B) to vertices $u_{1},u_{2},u_{3},x$. The resulting subproblems $G-\{u_{1},u_{2},u_{3},x\}$ and $G-N[x]$ have drops $(1,4),(1,3)$ directly; the branch-seq $[(1,3),(1,4)]$ is already covered in (7-1).
Line 6. At this point, every 7-vertex $u$ has a blocker $x$ (else we would branch on $u$ at Line 3), where $\deg(x)\leq 4$ (else we would branch on $u,x$ at Line 4). Due to Line 5, each vertex $x$ serves as a blocker to at most two 7-vertices $u,u^{\prime}$. Thus, at this stage, the graph satisfies
$n_{3}+n_{4}\geq n_{7}/2$
We now use a more refined property of the MaxIS-7 algorithm of [XN17b]; specifically, it has runtime $O^{*}(1.19698^{w_{3}n_{3}+w_{4}n_{4}+w_{5}n_{5}+w_{6}n_{6}+n_{7}})$ for constants $w_{3}=0.65077,w_{4}=0.78229,w_{5}=0.89060,w_{6}=0.96384$. Since $n_{3}+n_{4}\geq n_{7}/2$, it can be mechanically checked that $w_{3}n_{3}+w_{4}n_{4}+w_{5}n_{5}+w_{6}n_{6}+n_{7}$ is maximized at $n_{6}=n,n_{3}=n_{4}=n_{7}=0$. Accordingly, we need to show
$\min\{a_{63}\mu+b_{63}k,2(k-\mu)\cdot(w_{6}\log(1.19698))\}\leq a\mu+bk$ (7-4)
which can also be checked mechanically. This shows Theorem 67.
To finish up, we use a simple degree-8 branching algorithm.
If $\operatorname{maxdeg}(G)\leq 7$, then use Lemma 10 with the algorithms Branch7 or MaxIS-7. Otherwise, split on a vertex of degree at least $8$.
Algorithm 11 Function Branch8$(G,k)$
Theorem 68.
Branch8 has runtime $O^{*}(1.25284^{k})$.
Proof.
Splitting on a vertex of degree $r\geq 8$ generates subproblems with vertex covers of size at most $k-1$ and $k-8$. These have cost $1.25284^{k-1}$ and $1.25284^{k-8}$ by induction. Since $11.25284^{-1}+1.25284^{-8}<$ 1, the overall cost is $O^{*}(1.25284^{k})$. Otherwise, Lemma 10 (with parameters $a_{7},b_{7}$ and $c=\log(1.19698)$) gives the desired algorithm with runtime $O^{*}(1.25284^{k})$. ∎
Appendix A Proofs for basic results
Proof of Proposition 4.
Consider any basic solution $\theta\in\{0,\tfrac{1}{2},1\}^{V}$ to $\operatorname{LPVC}(G)$. Let $I_{0}=\theta^{-1}(0)$ and $I_{1}=\theta^{-1}(1)$; note that $I_{0}$ is an indset (possibly empty) and $N(I_{0})\subseteq I_{1}$. So $\sum_{v}\theta(v)=|I_{1}|+\tfrac{1}{2}(n-|I_{0}|-|I_{1}|)\geq(n+|N(I_{0})|-|I_{0}|)/2=(n+\operatorname{surp}(I_{0}))/2\geq(n+\operatorname{minsurp}^{-}(G))/2$. On the other hand, if $I_{0}$ is a min-set of $G$, then setting $\theta(v)=1$ for $v\in N(I_{0})$ and $\theta(v)=0$ for $v\in I_{0}$ and $\theta(v)=1/2$ for all other vertices $v$ gives a fractional vertex cover with $\sum_{v}\theta(v)=(n+\operatorname{minsurp}^{-}(G))/2$. ∎
Proof of Lemma 5.
Let $I$ be a critical-set; we first claim there is a good cover $C$ with $I\subseteq C$ or $I\cap C=\emptyset$. For, suppose that $I\cap C=J$ where $\emptyset\neq J\subsetneq I$. In this case, we must also have $N(I\setminus J)\subseteq C$, so overall $|C\cap N[I]|\geq|J|+|N(I\setminus J)|=|J|+|I\setminus J|+\operatorname{surp}(I\setminus J)$; since $I\setminus J$ is a non-empty subset of $I$, we have $\operatorname{surp}(I\setminus J)\geq\operatorname{surp}(I)$, hence this is at least $|J|+|I\setminus J|+\operatorname{surp}(I)=|I|+\operatorname{surp}(I)\geq N(I)$. Thus, replacing $C\cap N[I]$ by $N(I)$ would give a vertex cover $C^{\prime}$ with $C^{\prime}\cap I=\emptyset$ and $|C^{\prime}|\leq|C|$.
Furthermore, if $\operatorname{surp}(I)\leq 0$, then consider a good cover $C$ with $I\subseteq C$. Then replacing $I$ with $N(I)$ would give another cover $C^{\prime}$ with $|C^{\prime}|\leq|C|+|N(I)|-|I|\leq|C|$ and $C^{\prime}\cap I=\emptyset$.
Finally, suppose $\operatorname{surp}(I)=1$ and let $J=N(I)$. If $G$ has a good cover with $I\cap C=\emptyset$, then $J\subseteq C$ as desired. Otherwise, suppose $G$ has a good cover with $I\subseteq C$. Then, $C^{\prime}=(C\setminus I)\cup J$ would be another cover of size $|C|-|I|+(|J|-|C\cap J|)$; since $|J|=|I|+1$, this is $|C|+1-|C\cap J|$. If $C\cap J\neq\emptyset$, this would be a good cover with $I\cap C=\emptyset$. ∎
Proof of Proposition 6.
Let $J$ be an indset with $\operatorname{surp}_{G-N[I]}(J)=\operatorname{minsurp}^{-}(G-N[I])$. Then $J\cup I$ is a non-empty indset of $G$ with surplus $\operatorname{surp}_{G-N[I]}(J)+\operatorname{surp}_{G}(I)=\operatorname{minsurp}^{-}(G-N[I])+\operatorname{surp}_{G}(I)$; in particular, if $\operatorname{minsurp}^{-}(G-N[I])+\operatorname{surp}_{G}(I)=\operatorname{minsurp}(G)$, then $I$ is a min-set of $G$. On ther other hand, if $I$ is contained in a min-set $J$ of $G$, then $\operatorname{surp}_{G-N[I]}(J\setminus I)=\operatorname{surp}_{G}(J)+|I|-|N(I)|=\operatorname{surp}_{G}(J)-\operatorname{surp}_{G}(I)$. So $\operatorname{minsurp}^{-}(G-N[I])\leq\operatorname{minsurp}(G)-\operatorname{surp}_{G}(I)$. ∎
Proof of Proposition 7.
By applying Proposition 6 to indsets $I=\{x\}$, we get $\operatorname{minsurp}(G)=\min_{x\in V}\bigl{(}\operatorname{minsurp}^{-}(G-N[x])+\deg(x)-1\bigr{)}$. We can use Proposition 4 to compute each value $\operatorname{minsurp}^{-}(G-N[x])$ in polynomial time (by solving the LP). This gives an algorithm to compute $\operatorname{minsurp}(G)$. Likewise, given any vertex-set $X$, we can use Proposition 6 to determine if $X$ is contained in a min-set of $G$ and, if so, we can use Proposition 4 to find a vertex set $J$ attaining $\operatorname{surp}_{G-N[X]}(J)=\operatorname{minsurp}^{-}(G-N[X])$; then $J\cup X$ is a min-set of $G$ containing $X$. ∎
Proof of Proposition 8.
Let us first consider P1. If $G^{\prime}$ has a good cover $C^{\prime}$, then $C^{\prime}\cup N(I)$ is a cover for $G$ of size $k$. Conversely, by Lemma 5, there is a good cover $C$ of $G$ with $C\cap I=\emptyset$, and then then $C\setminus N(I)$ is a cover of $G^{\prime}$ of size $k-|N(I)|=k^{\prime}$. Similarly, if $G^{\prime}$ has a fractional cover $\theta$, then we can extend to $G$ by setting $\theta(v)=0$ for $v\in I$ and $\theta(v)=1$ for $v\in N(I)$, with total weight $\lambda(G^{\prime})+(k-k^{\prime})$. So $\mu(G^{\prime})=k^{\prime}-\lambda(G^{\prime})\leq k^{\prime}-(\lambda(G)-(k-k^{\prime}))=k-\lambda(G)=\mu(G)$.
Next, for rule P2, let $J=N(I)$. Since $I$ is a critical-set with surplus one, it cannot contain any isolated vertex. Given any good cover $C$ of $G$ with $J\subseteq C$, observe that $(C\setminus J)\cup\{y\}$ is a cover of $G^{\prime}$ of size $k-|J|+1=k^{\prime}$. Likewise, given a good cover $C$ of $G$ with $J\cap C=\emptyset$, we have $N(J)\subseteq C$; in particular, $I\subseteq C$ since $I$ has no isolated vertices, so $C\setminus I$ is a cover of $G^{\prime}$ of size $k-|I|=k^{\prime}$. Conversely, consider a good cover $C^{\prime}$ of $G^{\prime}$; if $y\in C^{\prime}$, then $(C^{\prime}\cup J)\setminus\{y\}$ is a cover of $G$, of size $k^{\prime}+|J|-1=k$. If $y\notin C^{\prime}$, then $C^{\prime}\cup I$ is a cover of $G$, of size $k^{\prime}+|I|=k$. Similarly, if $G$ has an optimal fractional cover $\theta$, it can be extended to $G$ by setting $\theta(v)=\theta(y)$ for $v\in J$, and $\theta(v)=1-\theta(y)$ for $v\in I$. This has weight $\lambda(G^{\prime})+|I|(1-\theta(y))+|J|\theta(y)-\theta(y)=\lambda(G^{\prime})+|I|\geq\lambda(G)$. So $\mu(G^{\prime})=k^{\prime}-\lambda(G^{\prime})\leq(k-|I|)-(\lambda(G)-|I|)=\mu(G)$.
Now for P3, let $A=N_{G}(u)\cap N_{G}(x)$ and $B_{x}=N_{G}(x)\setminus N_{G}[u]$ and $B_{u}=N_{G}(u)\setminus N_{G}[x]$. To show the validity of the preprocessing rule, suppose $\langle G,k\rangle$ is feasible, and let $C$ be a good cover of $G$ with either $|C\cap\{u,x\}|=1$; then $C\setminus(\{u,x\}\cup A)$ is a cover of $G^{\prime}$ of size $k^{\prime}$. Conversely, suppose $\langle G^{\prime},k^{\prime}\rangle$ is feasible with a good cover $C^{\prime}$. So $|B_{u}\setminus C^{\prime}|\leq 1$. if $B_{u}\setminus C^{\prime}=\emptyset$, then $C=C^{\prime}\cup A\cup\{x\}$ is a cover of $G$. If $|B_{u}\setminus C^{\prime}|=1$, then necessarily $B_{x}\subseteq C^{\prime}$ and $C=C^{\prime}\cup A\cup\{u\}$ is a cover of $G$.
Finally, we claim that (P3) satisfies $\mu(G^{\prime})\leq\mu(G)$; since $k^{\prime}=k-1-|A|$, it suffices to show that $\lambda(G)\leq\lambda(G^{\prime})+1+|A|$. For, take an optimal fractional cover $\theta$ for $G^{\prime}$. To extend it to a fractional cover for $G$, we set $\theta(v)=1$ for $v\in A$ and there are a few cases to determine the values for $\theta(u)$ and $\theta(v)$. If $B_{u}=\emptyset$, we set $\theta(u)=0,\theta(x)=1$. Otherwise, if $B_{u}\neq\emptyset$, let $b_{u}=\min_{v\in B_{u}}\theta^{\prime}(v)$ and then set $\theta(u)=1-b_{u},\theta(x)=b_{u}$; since $G^{\prime}$ has a biclique between $B_{u}$ and $B_{x}$, this covers any edge between $x$ and $v\in B_{x}$ with $\theta(x)+\theta(v)=b_{u}+\theta(v)\geq 1$. Overall, $\theta$ is a fractional vertex cover of $G$ of weight $\sum_{v\in G^{\prime}}\theta(v)+\theta(u)+\theta(x)+|A|=\lambda(G^{\prime})+1+\operatorname{codeg}(u,x)$. ∎
Proof of Lemma 9.
Consider a kite $u,x,y,z$. We can view the application of (P3) to funnel $u$ as a two-part process. First, we remove the shared neighbor $y\in N(u)\cap N(x)$; then, we merge vertices $z,x$ into a new vertex $u^{\prime}$. In the first step, we obtain a graph $G^{\prime\prime}=G-y$ with $k^{\prime\prime}=k-1$ and $n^{\prime\prime}=n-1$. So, by Proposition 4, we have $\lambda(G^{\prime\prime})=\lambda(G)-1/2$ and hence $\mu(G^{\prime\prime})\leq\mu(G)-1/2$ and $k^{\prime\prime}=k-1$. By Proposition 8, the second step gives $\mu(G^{\prime})\leq\mu(G^{\prime\prime})$ and $k^{\prime}=k^{\prime\prime}-1$. ∎
Proof of Lemma 10.
First, exhaustively apply rule (P1) to $G$; the resulting graph $G^{\prime}$ has $n^{\prime}\leq n$ and $k^{\prime}\leq k$, and $\operatorname{minsurp}(G^{\prime})\geq 0$. In particular, $\mu(G^{\prime})=k^{\prime}-\lambda(G^{\prime})=k^{\prime}-n^{\prime}/2$. Note that $G^{\prime}\in\mathcal{G}$ since (P1) just deletes vertices. Now dovetail the two algorithms for $G^{\prime}$, running both simultaneously and returning the output of whichever terminates first. This has runtime $O^{*}(\min\{e^{cn^{\prime}},e^{a(k^{\prime}-n^{\prime}/2)+bk^{\prime}}\})$. For fixed $k^{\prime}$, this is maximized at $n^{\prime}=\frac{2k^{\prime}(a+b)}{a+2c}$, at which point it takes on value $e^{dk^{\prime}}\leq e^{dk}$. ∎

### v7

@@
-A faster algorithm for Vertex Cover parameterized by solution size
+A faster algorithm for Vertex Cover parameterized by solution size [^1] This is an extended version of a paper appearing in the 41st international Symposium of Theoretical Aspects of Computer Science (STACS) 2024
 Abstract
@@
 1 Introduction
-For an undirected graph $G=(V,E)$, a subset $S\subseteq V$ is called a vertex cover if every edge has at least one endpoint in $S$.
-It is closely related to an independent set, since if $S$ is an inclusion-wise minimal vertex cover then $V-S$ is an inclusion-wise maximal independent set, and vice-versa.
+For an undirected graph $G=(V,E)$, a subset $C\subseteq V$ is called a vertex cover if every edge has at least one endpoint in $C$.
+It is closely related to an independent set, since if $C$ is an inclusion-wise minimal vertex cover then $V\setminus C$ is an inclusion-wise maximal independent set, and vice-versa.
 Finding the size of the smallest vertex cover is a classic NP-complete problem [GJ79].
-In particular, $G$ has a vertex cover of size at most $k$ if and only if it has an independent set of size at least $n-k$.
+In particular, $G$ has a vertex cover of size at most $k$ if and only if it has an independent set of size at least nkn-k.
 There is natural LP formulation for vertex cover which we denote by $\operatorname{LPVC}(G)$:
 minimize $\sum_{v\in V}\theta(v)$
-subject to $\theta(u)+\theta(v)\geq 1$ for all edges $e=(u,v)\in E$
+subject to $\theta(u)+\theta(v)\geq 1$ for all edges $(u,v)\in E$
 $\theta(v)\in[0,1]$ for all vertices $v\in V$
-The optimal solution to $\operatorname{LPVC}(G)$, denoted $\lambda(G)$ or just $\lambda$ if $G$ is clear from context, is a lower bound on the size of a minimum vertex cover of $G$.
-We also define $\mu(G)=k-\lambda(G)$, i.e.
+The optimal solution to $\operatorname{LPVC}(G)$, denoted $\lambda(G)$ or just $\lambda$ if $G$ is clear from context, is a lower bound on the size of a minimum vertex cover of $G$. We also define
+$\mu(G)=k-\lambda(G),$
+i.e.
 the gap between solution size and LP lower bound.
-This linear program has remarkable properties which have been exploited for a variety of algorithms [NT74, NT75, ${}^{+}$14].
+This linear program has remarkable algorithmic properties [NT74, NT75, LNR+14].
 For instance, an optimum basic solution is half-integral and can be found efficiently by a network flow computation.
-[…]
+Fixed-parameter tractable (FPT) and exact algorithms explore a landscape of parameters to understand the complexity of different problems [Nie06, FG06, DF99, CFK+15, FK10].
+Vertex Cover was one of the first studied FPT problems with respect to the parameter $k$, the optimal solution size.
+Building on a long line of research, this culminated in an algorithm with $O^{*}(1.2738^{k})$ runtime [CKX10].
+[^2] Throughout, we write $O^{*}(T)$ as shorthand for $T\cdot{\mathsf{poly}}(n)$.
+This record has been standing for more than a decade.
+Assuming the Exponential Time Hypothesis, no algorithm with runtime $O^{*}(2^{o(k)})$ is possible [CFK+15].
+Many different kinds of algorithms have been developed for the Vertex Cover problem. We cannot summarize them fully here; we describe some of the most relevant works for our own paper.
+An important variant is Above-Guarantee Vertex Cover (AGVC), where the parameter is the difference between $k$ and various lower bounds on vertex cover size [MR99, MRS09, LNR+14, GP16, KKK22].
+Of these, the parameter $\mu$ plays a particularly important role in this paper.
+We quote the following main result:
+Theorem 1 ([LNR+14]).
+Vertex cover can be solved in time $O^{*}(2.3146^{\mu(G)})$.
+Another natural choice is to measure runtime in terms of the graph size $n$.
+In this setting, the problem is more commonly referred to as Maximum Independent Set (MaxIS).
+Xiao and Nagamochi have developed a number of algorithms in this vein.
+These algorithms, and in particular their performance on bounded-degree graphs, will also play a crucial role in our analysis.
+We refer to the algorithm targeted for graphs of maximum degree $\Delta$ by the MaxIS-$\Delta$ algorithm.
+Theorem 2.
+MaxIS-3 can be solved with runtime $O^{*}(1.083506^{n})$ by [XN13]. [^3] This runtime is not claimed directly in [XN13], see [XN17a] instead.
+MaxIS-4 can be solved with runtime $O^{*}(1.137595^{n})$ by [XN17a].
+MaxIS-5 can be solved with runtime $O^{*}(1.17366^{n})$ by [XN16]. [^4] This runtime is not claimed directly in [XN16], see [XN17b] instead.
+MaxIS-6 can be solved with runtime $O^{*}(1.18922^{n})$ by [XN17b].
+MaxIS-7 can be solved with runtime $O^{*}(1.19698^{n})$ by [XN17b].
+MaxIS in graphs of arbitrary degree can be solved with runtime $O^{*}(1.19951^{n})$ by [XN17b].
+In addition, [Tsu23] describes algorithms in graphs of maximum degree $3$ and $4$ with runtime respectively $O^{*}(1.1558^{k})$ and $O^{*}(1.2403^{k})$ respectively.
 1.1 Outline of our results
@@
 All of these algorithms use polynomial space and are deterministic.
-The FPT algorithms for vertex cover, including our new algorithm, are built out of recursively branching on high-degree vertices or other structures, while tracking some “measure” of the graph.
+The FPT algorithms for vertex cover, including our new algorithm, are built out of recursive branching while tracking some “measure” of the graph.
 Our main new idea is to use a measure which is a piecewise-linear function of $k$ and $\mu$.
-To illustrate, consider branching on whether some degree-$r$ vertex $u$ is in the cover: we recurse on the subproblem $G_{1}=G-u$ where $k$ is reduced by one and on the subproblem $G_{0}=G-u-N(u)$ where $k$ is reduced by $r$.
+To illustrate, consider branching on whether some degree-$r$ vertex $u$ is in the cover: we recurse on the subproblem $G_{1}=G-u$ with a solution of size $k_{1}=k-1$ and on the subproblem $G_{0}=G-u-N(u)$ with the solution size $k_{0}=k-r$.
 We must also show that $\mu$ is significantly reduced in the two subproblems.
-Suppose that, wishfully speaking, $\vec{\tfrac{1}{2}}$ remains the optimal solution to $\operatorname{LPVC}(G_{0})$ and $\operatorname{LPVC}(G_{1})$.
+Suppose that, wishfully speaking, $\theta=\vec{\tfrac{1}{2}}$ remains the optimal solution to $\operatorname{LPVC}(G_{0})$ and $\operatorname{LPVC}(G_{1})$.
 This is what should happen in a “generic” graph.
-Then $\operatorname{\lambda}(G_{0})=|V(G_{0})|/2=(n-r-1)/2$ and $\operatorname{\lambda}(G_{1})=|V(G_{1})|/2=(n-1)/2$.
-Since $k_{0}=k-r$ and $k_{1}=k-1$, this implies that $\mu$ is indeed significantly decreased:
+Then $\operatorname{\lambda}(G_{0})=\frac{|V(G_{0})|}{2}=\frac{n-r-1}{2}$ and $\operatorname{\lambda}(G_{1})=\frac{|V(G_{1})|}{2}=\frac{n-1}{2}$.
+This implies that $\mu$ is indeed significantly decreased:
 $\mu(G_{0})=\mu(G)-(r-1)/2,\qquad\mu(G_{1})=\mu(G)-1/2$
@@
 In this situation, we develop an alternate branching rule for $G$: rather than branching on $u$ itself, we branch on the two subproblems where (i) both $u,x$ are in the cover and (ii) $x$ is not in the cover.
-See Figure 1 for an illustration.
-Figure 1: In the subgraph $G-N[u]$, the vertex $x$ becomes isolated
+(See Figure 1.)
+Figure 1: In the subgraph $G_{0}$, the vertex $x$ becomes isolated.
 We emphasize that this is just one example of our branching analysis.
-We need to handle more general situations where $\vec{\tfrac{1}{2}}$ is not the optimal solution to $\operatorname{LPVC}(G_{0})$ or $\operatorname{LPVC}(G_{1})$, which can require more complex branching rules.
-We emphasize that these branching rules are more powerful than simply using conventional branching rules along with the preprocessing rules; the latter may not be able to ensure a good reduction in $\mu$.
+We need to handle more general situations where $\vec{\tfrac{1}{2}}$ is not the optimal solution to $\operatorname{LPVC}(G_{0})$ or $\operatorname{LPVC}(G_{1})$, which can require more complex and fundamentally new branching rules.
 In Section 4, we use these ideas for a relatively simple algorithm with runtime $O^{*}(1.2575^{k})$, along with algorithms for bounded-degree graphs.
@@
 For example, when we are branching on degree-4 vertices, it is favorable to choose one that has a degree-$3$ neighbor.
-This will create a degree-$2$ vertex in subproblem $\langle G-u,k-1\rangle$, leading to further simplifications of the graph.
+This will create a degree-$2$ vertex in subproblem GuG-u, leading to further simplifications of the graph.
 See Figure 2.
-Figure 2: In the subgraph $G-u$, the vertex $v$ has degree two.
-1.2 Related algorithms
-Many different kinds of algorithms have been developed for the Vertex Cover problem. We cannot summarize this fully here; we describe some of the most relevant works for our own paper.
-As we have mentioned, there is a long line of research into FPT algorithms parametrized by the solution size $k$.
-Currently, the fastest such algorithm has $O^{*}(1.2738^{k})$ runtime [CKX10].
-In addition, [Tsu23] describes algorithms in graphs of maximum degree $3$ and $4$ with runtime respectively $O^{*}(1.1558^{k})$ and $O^{*}(1.2403^{k})$ respectively.
-An important variant is Above-Guarantee Vertex Cover (AGVC), where the parameter is the difference between $k$ and various lower bounds on vertex cover size [MR99, MRS09, ${}^{+}$14, GP16, KKK22].
-Of these, the parameter $\mu(G)=k-\operatorname{\lambda}(G)$ plays a particularly important role in this paper.
-We quote the following main result:
-Theorem 2 ([${}^{+}$14]).
-Vertex cover can be solved in time $O^{*}(2.3146^{\mu})$.
-Another natural choice is to measure runtime in terms of the graph size $n$.
-In this setting, the problem is more commonly referred to as Maximum Independent Set (MaxIS).
-Xiao and Nagamochi have developed a number of algorithms in this vein.
-These algorithms, and in particular their performance on bounded-degree graphs, will also play a crucial role in our analysis.
-We refer to the algorithm targeted for graphs of maximum degree $\Delta$ by the MaxIS-$\Delta$ algorithm.
-Theorem 3.
-MaxIS-3 can be solved with runtime $O^{*}(1.083506^{n})$ by [XN13]. [^1] This runtime is not claimed directly in [XN13], see [XN17a] instead.
-MaxIS-4 can be solved with runtime $O^{*}(1.137595^{n})$ by [XN17a].
-MaxIS-5 can be solved with runtime $O^{*}(1.17366^{n})$ by [XN16]. [^2] This runtime is not claimed directly in [XN16], see [XN17b].
-MaxIS-6 can be solved with runtime $O^{*}(1.18922^{n})$ by [XN17b].
-MaxIS-7 can be solved with runtime $O^{*}(1.19698^{n})$ by [XN17b].
-MaxIS in graphs of arbitrary degree can be solved with runtime $O^{*}(1.19951^{n})$ by [XN17b].
+Figure 2: In the subgraph GuG-u, the vertex $v$ has degree two.
 1.2 Notation
@@
 We say $C$ is a good cover if it is a vertex cover of size at most $k$.
-For a subset $S$ of $V$, the subgraphs of $G$ induced by $S$ and $V\setminus S$ are denoted by $G[S]$ and $G-S$, respectively.
+For a subset $S$ of $V$, the subgraphs of $G$ induced by $S$ and $V\setminus S$ are denoted by $G[S]$ and GSG-S, respectively.
 We write $u\sim v$ if $(u,v)\in E$ and $u\nsim v$ otherwise.
 For vertex sets $X,Y$, we write $X\sim Y$ if there exists $x\in X,y\in Y$ with $x\sim y$.
-For a vertex $u$, the neighborhood $N_{G}(u)$ is the set $\{u\in V(G):u\sim v\}$ and the closed neighborhood $N_{G}[u]$ is the set $N_{G}(u)\cup\{u\}$.
+For a vertex $u$, the neighborhood $N_{G}(u)$ is the set $\{v\in V(G):u\sim v\}$ and the closed neighborhood $N_{G}[u]$ is the set $N_{G}(u)\cup\{u\}$.
 Extending this notation, for a vertex set $S\subseteq V(G)$, we write $N_{G}(S)=\bigl(\bigcup_{v\in S}N_{G}(v)\bigr)\setminus S$ and $N_{G}[S]=N_{G}(S)\cup S=\bigcup_{v\in S}N_{G}[v]$.
@@
 We say a vertex is subquartic if it has degree in $\{1,2,3,4\}$ and subcubic if it has degree in $\{1,2,3\}$, and subquadratic if it has degree in $\{1,2\}$.
-(We do not count isolated vertices).
 We write $\operatorname{maxdeg}(G)$ and $\operatorname{mindeg}(G)$ for the minimum and maximum vertex degrees in $G$, respectively.
@@
 We say that a pair of vertices $u,v$ share neighbor $y$ if $y\in N(u)\cap N(v)$. We denote by $\operatorname{codeg}(u,v)=|N(u)\cap N(v)|$ the number of vertices shared by $u,v$.
+An independent set (or, for brevity, an indset) denotes a set $X\subseteq V$ where no two vertices in $X$ are adjacent, i.e. $X\not\sim X$.
 An $\ell$-cycle denotes a set of $\ell$ vertices $x_{1},\dots,x_{\ell}$ with a path $x_{1},x_{2},\dots,x_{\ell},x_{1}$.
 2 Preliminaries
-We review some basic facts about branching and preprocessing rules for vertex cover. Much of this material is standard, see e.g. [${}^{+}$14]. We include proofs in Appendix A for completeness.
-Our algorithm will heavily use the LP and its properties.
-This LP is closely related to properties of independent sets.
-For brevity, we refer to these as indsets, i.e.
-sets $X\subseteq V$ where no two vertices in $X$ are adjacent.
-For an indset $I$ in $G$, we define the surplus by $\operatorname{surp}_{G}(I)=|N_{G}(I)|-|I|$.
-If $G$ is clear from context, we write just $\operatorname{surp}(I)$ or $N(I)$.
-We define $\operatorname{minsurp}(G)$ to be the minimum value of $\operatorname{surp}_{G}(I)$ over all non-empty indsets $I$ in $G$, and a min-set to be any non-empty indset $I$ achieving this minimum, i.e.
-$\operatorname{surp}_{G}(I)=\operatorname{minsurp}(G)$.
-Since this comes up in a number of places, we define $\operatorname{minsurp}^{-}(G)=\min\{0,\operatorname{minsurp}(G)\}$.
+Our algorithm will heavily use the LP and its properties, which are closely related to properties of indsets.
+We review some basic facts here.
+Much of this material is standard, see e.g.
+[LNR+14]; we defer some proofs to Appendix A for completeness.
+For an indset $I$ in $G$, we define the surplus by
+$\operatorname{surp}_{G}(I)=|N_{G}(I)|-|I|.$
+We define $\operatorname{minsurp}(G)$ to be the minimum value of $\operatorname{surp}_{G}(I)$ over non-empty indsets $I$ in $G$, and a min-set to be any non-empty indset $I$ with $\operatorname{surp}_{G}(I)=\operatorname{minsurp}(G)$.
+Since this plays a central role in our analysis, we define the shadow of a vertex set $X$, denoted $\operatorname{shad}_{G}(X)$, to be
+$\operatorname{shad}_{G}(X)=\operatorname{minsurp}(G-X).$
+We write $\operatorname{shad}_{G}(x_{1},\dots,x_{\ell},X_{1},\dots,X_{r})$ as shorthand for $\operatorname{shad}_{G}(\{x_{1},\dots,x_{\ell}\}\cup X_{1}\cup\dots\cup X_{r})=\operatorname{minsurp}(G-\{x_{1},\dots,x_{\ell}\}-(X_{1}\cup\dots\cup X_{r}))$.
+Since this comes up in a number of places, we define $\operatorname{minsurp}^{0}(G)=\min\{0,\operatorname{minsurp}(G)\}$.
+If $G$ is clear from context, we write just $\operatorname{surp}(I),N(I)$, etc.
+To explain the notation, consider the following diagram:
+Figure 3: In the subgraph GXG-X, there is an isolated vertex, hidden in the “shadow” of $X$. Here $\operatorname{shad}(X)=-1$.
 Proposition 4.
-$\lambda(G)=\tfrac{1}{2}(|V|+\operatorname{minsurp}^{-}(G)).$ Moreover, in a basic solution to $\operatorname{LPVC}(G)$, the set $I$ of vertices with value zero forms an indset (possibly empty) with $\operatorname{surp}_{G}(I)=\operatorname{minsurp}^{-}(G)$.
+There holds $\lambda(G)=\tfrac{1}{2}(|V|+\operatorname{minsurp}^{0}(G)).$
+Proposition 5.
+Given any indset $I$ and vertex set $X\supseteq I$, one can efficiently find an indset which attains the minimum value of $\operatorname{surp}_{G}(J)$, over all indsets $J$ such that $I\subseteq J\subseteq X$.
+Using Proposition 5, we can efficiently find min-sets of various types.
+For example, we can determine $\operatorname{minsurp}(G)$ by applying Proposition 5 with $X=V,I=\{v\}$ for all choices of vertex $v$; the minimum over all these values gives a (non-empty) min-set II of $G$.
+Similarly, we can find inclusion-wise minimal or maximal min-sets by repeatedly applying Proposition 5 to try to include or exclude each vertex from a given min-set.
+For the rest of the paper, we will not discuss the computational aspects of finding min-sets that are needed in our algorithm.
 We define a critical-set to be a non-empty indset $I$ with $\operatorname{surp}_{G}(J)\geq\operatorname{surp}_{G}(I)$ for all non-empty subsets $J\subseteq I$.
@@
 Lemma 6.
-For any critical-set $I$, there is a good cover $C$ with either $I\subseteq C$ or $I\cap C=\emptyset$.
+For a critical-set $I$, there is a good cover $C$ with either $I\subseteq C$ or $I\cap C=\emptyset$. Moreover:
 If $\operatorname{surp}(I)\leq 0$, there is a good cover $C$ with $I\cap C=\emptyset$.
 If $\operatorname{surp}(I)=1$, there is a good cover $C$ with either $N[I]\cap C=I$ or $N[I]\cap C=N(I)$.
-Proposition 6.
-For any non-empty indset $I$ of $G$, there holds $\operatorname{minsurp}(G)\leq\operatorname{minsurp}^{-}(G-N[I])+\operatorname{surp}_{G}(I)$.
-The equality $\operatorname{minsurp}(G)=\operatorname{minsurp}^{-}(G-N[I])+\operatorname{surp}_{G}(I)$ holds if and only if $I$ is contained in a min-set of $G$.
-Proposition 7.
-The value $\operatorname{minsurp}(G)$ can be determined in polynomial time.
-Moreover, for any indset $X$ (possibly $X=\emptyset$), we can efficiently find a min-set of $G$ containing $X$ (if any such exists).
-2.1 Preprocessing and branching rules
-Our algorithm uses three major preprocessing rules.
-For this, we need to define the following graph structure: a funnel is a vertex $u$ with a neighbor $x$ such that $G[N(u)\setminus\{x\}]$ forms a clique.
-Here, we call $x$ the out-neighbor of $u$.
-For example, if a degree-3 vertex $u$ has a triangle with neighbors $t_{1},t_{2}$, and one other vertex $x$, then $u$ is a funnel with out-neighbor $x$; since this comes up so often, we refer to this as a 3-triangle.
-A well-known result is that, for a funnel $u$ with out-neighbor $x$, there exists a good cover $C$ with either $u\notin C$ or $x\notin C$.
+2.1 Preprocessing rules and reductions in $k$
+Our algorithm uses three major preprocessing rules. The first two are defined in terms of critical sets of small surplus.
 Preprocessing Rule 1 (P1): Given a critical-set $I$ with $\operatorname{surp}_{G}(I)\leq 0$, form graph $G^{\prime}$ with $k^{\prime}=k-|N(I)|$ by deleting $N[I]$.
 Preprocessing Rule 2 (P2): Given a critical-set $I$ with $\operatorname{surp}_{G}(I)=1$, form graph $G^{\prime}$ with $k^{\prime}=k-|I|$ by deleting $N[I]$ and adding a new vertex $y$ adjacent to $N(N(I))$.
-Preprocessing Rule 3 (P3): Given a funnel $u$ with out-neighbor $x$, form graph $G^{\prime}$ with $k^{\prime}=k-1-\operatorname{codeg}(u,x)$ by removing vertices $N[u]\cap N[x]$ and adding edges between all vertices in $N(u)\setminus N[x]$ to all vertices in $N(x)\setminus N[u]$.
+The final rule is based on a graph structure called a funnel.
+This structure consists of a vertex $u$ with a neighbor $x$ such that $G[N(u)\setminus\{x\}]$ forms a clique.
+We call $x$ the out-neighbor of $u$.
+For example, if a degree-3 vertex $u$ has a triangle with neighbors $t_{1},t_{2}$, and one other vertex $x$, then $u$ is a funnel with out-neighbor $x$; since this comes up so often, we refer to this as a 3-triangle.
+The following result is well-known:
+Observation 7.
+For a funnel $u$ with out-neighbor $x$, there exists a good cover $C$ satisfying either (i) $u\notin C$ or (ii) $x\notin C$ and $|C\setminus N(u)|=1$.
+This leads to the following simplification rule:
+Preprocessing Rule 3 (P3): Given a funnel $u$ with out-neighbor $x$, form graph $G^{\prime}$ with $k^{\prime}=k-1-\operatorname{codeg}(u,x)$ by removing vertices the vertices in $N[u]\cap N[x]$ and adding edges between all vertices in $N(u)\setminus N[x]$ to all vertices in $N(x)\setminus N[u]$.
 See Figure 4 for an illustration of rule (P3) applied to a funnel.
@@
 If a vertex $v$ has degree two, then applying (P2) to $\{v\}$ removes $v$ and contracts its neighbors $x,y$ into a new vertex $v^{\prime}$.
-When no further preprocessing rules can be applied, we say the graph $G$ is simplified; the resulting graph has a number of nice properties, for instance, it has $\operatorname{minsurp}(G)\geq 2$, and the solution $\vec{\tfrac{1}{2}}$ is the unique optimal solution to $\operatorname{LPVC}(G)$, and the minimum degree of any vertex is $3$.
+When no further preprocessing rules can be applied, we say $G$ is simplified; the graph then has a number of nice properties, for instance, $\operatorname{minsurp}(G)\geq 2,\operatorname{mindeg}(G)\geq 3$, and the solution $\vec{\tfrac{1}{2}}$ is the unique optimal solution to $\operatorname{LPVC}(G)$.
 Proposition 8.
 After applying any preprocessing rule, we have $\mu(G^{\prime})\leq\mu(G)$, and $\langle G,k\rangle$ is feasible if and only if $\langle G^{\prime},k^{\prime}\rangle$ is feasible.
-There is one situation where rule P3 is particularly powerful. Consider a 3-vertex $u$ which has neighbors $x,y,z$ where $x\sim y$ and $y\sim z$. We refer to this as a kite; see Figure 4.
+Observation 9.
+Suppose $\operatorname{minsurp}(G)\geq 2$.
+Then any vertex set $X$ satisfies $\operatorname{shad}(X)\geq 2-|X|$, and any indset $I$ satisfies $\operatorname{shad}(N[I])\geq 2-\operatorname{surp}_{G}(I).$
+In particular, a vertex $u$ satisfies $\operatorname{shad}(u)\geq 1$ and $\operatorname{shad}(N[u])\geq 3-\deg(u)$.
+We define $S(G)$ to be the largest decrease in $k$ obtainable from applying rules (P1) – (P3) in some efficiently-computable sequence.
+[^5] It is not clear how to calculate the largest decrease in $k$ via preprocessing rules, since applying some rules may prevent other rules.
+We assume that we have fixed some polynomial-time computable sequence of potential preprocessing rules to reduce $k$ as much as possible.
+At various points in our algorithm, we will describe simplifications available for intermediate graphs.
+We always assume that our preprocessing rules can find such simplifications.
+For vertices $x_{1},\dots,x_{\ell}$ and vertex sets $X_{1},\dots,X_{r}$, we write $S(x_{1},\dots,x_{\ell},X_{1},\dots,X_{r})$ as shorthand for $S(G-\{x_{1},\dots,x_{\ell}\}-X_{1}-\dots-X_{r})$.
+We record a few observations on simplifications of various structures.
+Proposition 10.
+Let $x,y$ be non-adjacent subquadratic vertices.
+Then $S(G)\geq 2$ if any of the three conditions hold: (i) $\operatorname{codeg}(x,y)=0$ or (ii) $\deg(x)+\deg(y)\geq 3$ or (iii) $\operatorname{minsurp}(G)\geq 0$.
+Proposition 11.
+If $G$ has 2-vertices $x_{1},\dots,x_{\ell}$ with pairwise distance at least $3$, then $S(G)\geq\ell$.
+Since a simplified graph $G$ has $\operatorname{minsurp}(G)\geq 2$, we have the following immediate observation from Proposition 10:
+Observation 12.
+If $G$ is simplified and $G-\{u,v\}$ has two non-adjacent subquadratic vertices, then $S(u,v)\geq 2$.
+We now consider a more involved simplification rules.
+Lemma 13.
+If $\operatorname{minsurp}(G)\geq 0$ and we have an indset $I$ with $\operatorname{surp}_{G}(I)\leq 1$, then $S(G)\geq|I|$.
+Proof.
+Let $r=|I|$.
+If $\operatorname{surp}_{G}(I)=0$, or $I$ is a critical-set, we can apply (P1) or (P2) to II.
+So suppose that $\operatorname{surp}_{G}(I)=1$ and $\operatorname{surp}_{G}(K)=0$ for a non-empty subset $K\subsetneq I$.
+Choose $K$ to be inclusion-wise-maximal with this property, and let $s=|K|$.
+We apply (P1) to $K$, reducing $k$ by $s$ and obtaining a graph $G^{\prime}=G-N[K]$.
+Now consider the indset $J=I\setminus K$ in $G^{\prime}$, where |J|=rs|J|=r-s.
+We have $|N_{G^{\prime}}(J)|=|N_{G}(I)\setminus N_{G}(K)|=r+1-s$ and so $\operatorname{surp}_{G^{\prime}}(J)=1$.
+We claim that $J$ is a critical-set in $G^{\prime}$.
+For, if some non-empty subset $K^{\prime}\subseteq J$ has $|N_{G^{\prime}}(K^{\prime})|\leq|K^{\prime}|$, then indset $I^{\prime}=K\cup K^{\prime}\subseteq I$ would have $|N_{G}(I^{\prime})|\leq|N_{G}(K)|+|N_{G^{\prime}}(K^{\prime})|\leq|K|+|K^{\prime}|=|I^{\prime}|$.
+Hence $\operatorname{surp}_{G}(I^{\prime})\leq 0$, contradicting maximality of $K$.
+So, we can apply (P2) to $J$ in $G^{\prime}$, getting a further drop of rsr-s.
+Overall, we get a net drop of s+(rs)=rs+(r-s)=r as desired.
+∎
+Lemma 14.
+We define a kite to be a 3-vertex $u$ with neighbors $x,y,z$ where $x\sim y$ and $y\sim z$.
+If $\operatorname{minsurp}(G)\geq 1$, then applying (P3) to a kite $u$ yields a graph $G^{\prime}$ with $k^{\prime}\leq k-2$ and $\mu(G^{\prime})\leq\mu(G)-1/2$. (See Figure 5.)
 Figure 5: A kite $u,x,y,z$ before (left) and after (right) applying P3
+Proof.
+We view the application of (P3) to $u$ as a two-part process.
+First, we remove the shared neighbor $y\in N(u)\cap N(x)$; then, we merge $z,x$ into a new vertex $u^{\prime}$.
+The first step gives graph $G^{\prime\prime}=G-y$ with $k^{\prime\prime}=k-1,n^{\prime\prime}=n-1$.
+So, by Proposition 4, we have $\lambda(G^{\prime\prime})=\lambda(G)-1/2$ and so $\mu(G^{\prime\prime})\leq\mu(G)-1/2$.
+By Proposition 8, the second step gives $\mu(G^{\prime})\leq\mu(G^{\prime\prime})$ and $k^{\prime}=k^{\prime\prime}-1$.
+∎
 Lemma 15.
-Suppose $\operatorname{minsurp}(G)\geq 1$. Then applying (P3) to a kite in $G$ yields a graph $G^{\prime}$ with $k^{\prime}\leq k-2$ and $\mu(G^{\prime})\leq\mu(G)-1/2$.
-2.2 Runtime and branching framework
+Suppose that $\mathcal{G}$ is a class of graphs closed under vertex deletion, and for which there are vertex cover algorithms with runtimes $O^{*}(e^{a\mu+bk})$ and $O(e^{cn})$ for $a,b,c\geq 0$.
+Then we can solve vertex cover in $\mathcal{G}$ with runtime $O^{*}(e^{dk})$ where $d=\frac{2c(a+b)}{a+2c}$.
+Proof.
+First, exhaustively apply (P1) to $G$; the resulting graph $G^{\prime}$ has $n^{\prime}\leq n,k^{\prime}\leq k$, and $\operatorname{minsurp}(G^{\prime})\geq 1$.
+In particular, $\mu(G^{\prime})=k^{\prime}-\lambda(G^{\prime})=k^{\prime}-n^{\prime}/2$.
+Note that $G^{\prime}\in\mathcal{G}$ since (P1) just deletes vertices.
+Now dovetail the two algorithms for $G^{\prime}$, running both simultaneously and returning the output of whichever terminates first.
+This has runtime $O^{*}(\min\{e^{cn^{\prime}},e^{a(k^{\prime}-n^{\prime}/2)+bk^{\prime}}\})$.
+For fixed $k^{\prime}$, this is maximized at $n^{\prime}=\frac{2k^{\prime}(a+b)}{a+2c}$, at which point it takes on value $e^{dk^{\prime}}\leq e^{dk}$.
+∎
+To illustrate Lemma 15, consider graphs of maximum degree 3.
+We can combine the AGVC algorithm with runtime $O^{*}(2.3146^{\mu})$ (corresponding to $a=\log(2.3146),b=0$), and the MaxIS-3 algorithm with runtime $O^{*}(1.083506^{n})$ (corresponding to $c=\log(1.083506)$), to get an algorithm with runtime $O^{*}(1.14416^{k})$; this already gives us one of the results in Theorem 3.
+2.2 Branching framework
 Our algorithm follows the measure-and-conquer approach.
@@
 It then runs recursively on each $G_{i}$.
-Such an algorithm has runtime $O^{*}(e^{\phi(G)})$ as long as the preprocessing steps have runtime $O^{*}(e^{\phi(G)})$ and it satisfies
-$\sum_{i=1}^{t}e^{\phi(G_{i})-\phi(G)}\leq 1.$ (1)
-(Here and throughout $e=2.718...$ is the base of the natural logarithm.) We refer to $\phi$ as the measure of the algorithm. For the most part, we will use measures of the form
+Such an algorithm has runtime $O^{*}(e^{\phi(G)})$ as long as the preprocessing steps have runtime $O^{*}(e^{\phi(G)})$ and it satisfies [^6] Here and throughout $e=2.718$...
+is the base of the natural logarithm.
+$\sum_{i=1}^{t}e^{\phi(G_{i})-\phi(G)}\leq 1$ (Branching Inequality)
+We refer to $\phi$ as the measure of the algorithm. For the most part, we will use measures of the form
 $\phi(G)=a\mu+bk\qquad\qquad\text{for $a,b\geq 0$};$ (1)
@@
 $\operatorname{val}_{a,b}(B)=\sum_{i=1}^{t}e^{-a\Delta\mu_{i}-b\Delta k_{i}}.$ (2)
-We say branch-seq $B^{\prime}$ dominates $B$ if $\operatorname{val}_{a,b}(B^{\prime})\leq\operatorname{val}_{a,b}(B)$ for all $a,b\geq 0$, and $G$ has a branch-seq $B$ available if we can efficiently find a branching rule which has some branch-seq $B^{\prime}$ dominating $B$.
-In general, a “balanced” branch-seq dominates an “imbalanced” one; formally, if $\Delta k_{1}\leq\Delta k_{2}$ and $\Delta\mu_{1}\leq\Delta\mu_{2}$, then branch-seq $[(\Delta\mu_{1},\Delta k_{1}),(\Delta\mu_{2},\Delta k_{2})]$ dominates $[(\Delta\mu_{1}-f,\Delta k_{1}-g),(\Delta\mu_{1}+f,\Delta k_{2}+g)]$ for any $f,g\geq 0$.
-We use the following important compositional property throughout: if we have a branching rule $B$ generating $\ell$ subproblems with drops $(\Delta\mu_{i},\Delta k_{i})$, and we then apply a branching rule with branch-seq $B^{\prime}_{i}$ to each subproblem $i=1,\dots,\ell$, then, overall, we get
-$\operatorname{val}_{a,b}(B)=\sum_{i=1}^{\ell}e^{-a\Delta\mu_{i}-b\Delta k_{i}}\operatorname{val}_{a,b}(B_{i}^{\prime}).$
-Lemma 10.
-Suppose that $\mathcal{G}$ is a class of graphs closed under vertex deletion, and for which there are vertex cover algorithms with runtimes $O^{*}(e^{a\mu+bk})$ and $O(e^{cn})$ for $a,b,c\geq 0$.
-Then we can solve vertex cover in $\mathcal{G}$ with runtime $O^{*}(e^{dk})$ where $d=\frac{2c(a+b)}{a+2c}$.
-To illustrate Lemma 10, consider graphs of maximum degree 3.
-We can combine the AGVC algorithm with runtime $O^{*}(2.3146^{\mu})$ (corresponding to $a=\log(2.3146),b=0$), and the MaxIS-3 algorithm with runtime $O^{*}(1.083506^{n})$ (corresponding to $c=\log(1.083506)$), to get an algorithm with runtime $O^{*}(1.14416^{k})$; this already gives us one of the results in Theorem 1.
-3 Preprocessing rules and reductions in $k$
-We define $S(G)$ to be the largest decrease in $k$ obtainable from applying rules (P1) – (P3) exhaustively via some efficiently-computable sequence of operations.
-[^3] It is not clear how to calculate the absolute largest decrease in $k$ via preprocessing rules, since applying some rules may prevent opportunities for other rules.
-We assume that we have fixed some polynomial-time computable sequences of potential preprocessing rules in order to reduce $k$ as small as much as possible.
-At various points in our algorithm, we will describe simplifications available for intermediate graphs.
-We always assume that our preprocessing rules are chosen to find such simplifications.
-As a point of notation, we define $R(X)=S(G-X)$ for a vertex set $X$.
-For vertices $x_{1},\dots,x_{\ell}$ and vertex sets $X_{1},\dots,X_{r}$, we write $R(x_{1},\dots,x_{\ell},X_{1},\dots,X_{r})$ as shorthand for the more cumbersome $R(\{x_{1},\dots,x_{\ell}\}\cup X_{1}\cup\dots\cup X_{r})=S(G-\{x_{1},\dots,x_{\ell}\}-(X_{1}\cup\dots\cup X_{r}))$.
-Note that we may have $\ell=0$ or $r=0$, e.g.
-$R(x_{1},\dots,x_{\ell})$ is shorthand for $R(\{x_{1},\dots,x_{\ell}\})$.
-We record a few observations on simplifications of various structures.
-Observation 11.
-Suppose $x,y$ are non-adjacent subquadratic vertices.
-Then $S(G)\geq 2$ if any of the following three conditions hold: (i) $\operatorname{codeg}(x,y)=0$ or (ii) $\deg(x)+\deg(y)\geq 3$ or (iii) $\operatorname{minsurp}(G)\geq 0$.
+We say branch-seq $B^{\prime}$ dominates $B$ if $\operatorname{val}_{a,b}(B^{\prime})\leq\operatorname{val}_{a,b}(B)$ for all $a,b\geq 0$, and $G$ has a branch-seq $B$ available if we can efficiently find a branching rule with a branch-seq dominating $B$.
+Most of our branching rules come from guessing that a certain vertex set $X$ is in the cover, and then removing a set of vertices $Y$ which are isolated in GXG-X.
+The resulting subproblem $\langle G^{\prime},k^{\prime}\rangle$ with $G^{\prime}=G-(X\cup Y),k^{\prime}=k-|X|$ is called principal with $\Delta k=|X|$ and excess $s=|Y|$.
+The graph $G^{\prime}$ may have other simplifications available.
+We say $\langle G^{\prime},k^{\prime}\rangle$ has drop $(\Delta\mu^{\prime},|X|)$ directly, and the final subproblem $G^{\prime\prime}=\langle G^{\prime\prime},k^{\prime\prime}\rangle$ has drop $(\Delta\mu^{\prime\prime},\Delta k^{\prime\prime})$ after simplification.
+Note that if we apply (P1) to $G^{\prime}$, then the subproblem $G^{\prime\prime}=\langle G-(X\cup Y\cup N[Z]),k-|X|-|N(Z)|\rangle$ is again principal, with excess $|Y|+|Z|$; however, if we apply (P2) or (P3) to $G^{\prime}$, then $G^{\prime\prime}$ is not principal as it requires other operations such as adding edges.
+Proposition 16.
+If $\operatorname{minsurp}(G)\geq 0$, then a principal subproblem $\langle G-Z,k-\Delta k\rangle$ with excess $s$ has $\Delta\mu=\tfrac{1}{2}(\Delta k-s+\operatorname{minsurp}^{0}(G-X))$.
 Proof.
-If $\deg(x)=\deg(y)=1$ and $\operatorname{codeg}(x,y)=0$, then $\{x,y\}$ is a min-set with $\operatorname{surp}(\{x,y\})=0$, and applying P1 reduces $k$ by two.
-Otherwise, if $\deg(x)\geq 2$, then we can apply (P2) to $x$, and vertex $y$ remains subquadratic, and we can follow up by applying (P1) or (P2) again to $y$.
-Note that if $\operatorname{minsurp}(G)\geq 0$, then at least one of the conditions (i) or (ii) must hold.
+Proposition 4 gives $\lambda(G)=n/2$.
+Let $G^{\prime}=G-Z$, where $|Z|=\Delta k+s$.
+Then $\mu(G^{\prime})=k^{\prime}-\operatorname{\lambda}(G^{\prime})=k^{\prime}-\tfrac{1}{2}(n^{\prime}+\operatorname{minsurp}^{0}(G^{\prime}))$, where $n^{\prime}=n-|Z|=2\lambda(G)-(s+\Delta k)$, which simplifies to $\mu(G^{\prime})=\mu(G)-\tfrac{1}{2}(\Delta k-s+\operatorname{minsurp}^{0}(G^{\prime}))$.
 ∎
+Consider a critical-set II of $G$.
+By Lemma 6, we know that there is either a good cover containing $I$, or disjoint to $I$.
+In the latter case, all the neighbors of $I$ must go into the cover.
+There is a natural branching rule to guess which of these possibilities holds:
+(B-Crit): For a critical-set $I$ branch on subproblems $\langle G-I,k-|I|\rangle$ and $\langle G-N[I],k-|N(I)|\rangle$.
 Proposition 17.
-If $G$ has 2-vertices $x_{1},\dots,x_{\ell}$ which all have pairwise distance at least $3$, then $S(G)\geq\ell$.
+Suppose that $\operatorname{minsurp}(G)\geq 2$.
+Let $I$ be a critical-set with $\operatorname{surp}_{G}(I)=\ell,|I|=r$.
+Then (B-Crit) generates principal subproblems $G_{1}=\langle G-I,k-r\rangle$ and $G_{0}=\langle G-N[I],k-r$-$\ell\rangle$ with excess 0 and $r$ respectively.
+Furthermore:
+  • If $r=1$, then the subproblem $G_{1}$ has $\Delta\mu=\tfrac{1}{2}$.
+  • If $r>1$, then the subproblem $G_{1}$ has $\Delta\mu=\tfrac{1}{2}(r+\operatorname{minsurp}^{0}(G-I))\geq 1$.
+  • The subproblem $G_{0}$ has $\Delta\mu=\tfrac{1}{2}(\ell+\operatorname{minsurp}^{0}(G-N[I]))\geq 1$.
 Proof.
-We show it by induction on $\ell$.
-The base case $\ell=0$ is vacuous.
-For the induction step, we apply (P2) to $x_{\ell}$, forming a graph $G^{\prime}$ where the neighbors $y,z$ get contracted into a single new vertex $t$.
-By hypothesis, $y,z$ are distinct from $x_{1},\dots,x_{\ell}$.
-We claim that $\operatorname{dist}_{G^{\prime}}(x_{i},x_{j})\geq 3$ and $\deg_{G^{\prime}}(x_{i})\geq 2$ for any pair $i<j<\ell$.
-For, clearly $x_{i}\not\sim x_{j}$ in $G^{\prime}$.
-If $x_{i},x_{j}$ share a neighbor in $G^{\prime}$, it must be vertex $t$ as no other vertices were modified.
-This is only possible if $\{x_{i},x_{j}\}\sim\{y,z\}$; but this contradicts our hypothesis that $\operatorname{codeg}(x_{i},x_{\ell})=\operatorname{codeg}(x_{j},x_{\ell})=0$.
-Finally, suppose that $\deg_{G^{\prime}}(x_{i})\leq 1$.
-This is only possible if the two neighbors of $x_{i}$ got merged together, i.e.
-$x_{i}\sim y$ and $x_{i}\sim z$.
-Again, this contradicts that $\operatorname{codeg}(x_{i},x_{\ell})=0$.
-By induction hypothesis applied to $G^{\prime}$, we have $S(G^{\prime})\geq\ell-1$ and hence $S(G)\geq\ell$. ∎
-Lemma 13.
-If $\operatorname{minsurp}(G)\geq 0$ and indset $I$ has $\operatorname{surp}_{G}(I)\leq 1$, then $S(G)\geq|I|$.
+The bound $\Delta\mu=\tfrac{1}{2}(r+\operatorname{minsurp}^{0}(G-I))$ for $G_{1}$ follows from Proposition 16.
+Also, by Observation 9, we have $\operatorname{minsurp}(G-I)\geq\operatorname{minsurp}(G)-|I|\geq 2-r$.
+Thus, if $r=1$, we have $\operatorname{minsurp}^{0}(G-I)=0$.
+Similarly, by Proposition 16, we have $\Delta\mu=\tfrac{1}{2}(\ell+\operatorname{minsurp}^{0}(G-N[I]))$ and by Observation 9, we have $\operatorname{minsurp}(G-N[I])\geq 2-\operatorname{surp}_{G}(I)=2$-$\ell$.
+∎
+3 Analysis of vertex branching
+Our bread-and-butter branching rule is to choose a vertex $u$ and branch on whether it is in the cover.
+We refer to this as splitting on $u$.
+It is a special case of (B-Crit) with a singleton set $I=\{u\}$, generating principal subproblems $\langle G-u,k-1\rangle$ and $\langle G-N[u],k-\deg(u)\rangle$.
+By Proposition 16, if we start with a simplified graph, then the subproblems have drops $(0.5,1)$ and $(1,\deg(u))$ respectively.
+Let us examine the subproblem $\langle G-N[u],k-\deg(u)\rangle$ more carefully: if $\operatorname{shad}(N[u])\geq 0$, then the subproblem has drop $(\tfrac{\deg(u)-1}{2},\deg(u))$.
+This is the “generic” and desired situation.
+Otherwise, when $\operatorname{shad}(N[u])\leq 0$, we say that $u$ is blocked.
+If $x$ is a vertex in a min-set of GN[u]G-N[u], we say that $x$ is a blocker of $u$.
+This motivates the following powerful branching rule:
+(B-Block): Given a vertex $x$ which is a blocker of each of the vertices $u_{1},\dots,u_{\ell}$, branch on subproblems $G_{1}=\langle G-\{u_{1},\dots,u_{\ell},x\},k-\ell-1\rangle$ and $G_{0}=\langle G-N[x],k-\deg(x)\rangle$.
+Proposition 18.
+(B-Block) is a valid branching rule.
 Proof.
-Let $r=|I|$.
-If $\operatorname{surp}_{G}(I)=0$, or $I$ is a critical-set, we can apply (P1) or (P2) to $I$.
-So suppose that $\operatorname{surp}_{G}(I)=1$ and $\operatorname{surp}_{G}(X)=0$ for a non-empty subset $X\subsetneq I$; among all such subsets $X$, choose one of maximum size, and let $s=|X|$.
-We apply (P1) to $X$, reducing $k$ by $s$ and obtaining a graph $G^{\prime}=G-N[X]$.
-Now consider the indset $J=I\setminus X$ in $G^{\prime}$, where $|J|=r-s$.
-We have $|N_{G^{\prime}}(J)|=|N_{G}(I)\setminus N_{G}(X)|=r+1-s$ and so $\operatorname{surp}_{G^{\prime}}(J)=1$.
-We claim that $J$ is a critical-set in $G^{\prime}$.
-For, if some non-empty subset $X^{\prime}\subseteq J$ has $|N_{G^{\prime}}(X^{\prime})|\leq|X^{\prime}|$, then indset $I^{\prime}=X\cup X^{\prime}\subseteq I$ would have $|N_{G}(I^{\prime})|\leq|N_{G}(X)|+|N_{G^{\prime}}(X^{\prime})|\leq|X|+|X^{\prime}|=|I^{\prime}|$.
-Hence $\operatorname{surp}_{G}(I^{\prime})\leq 0$, contradicting maximality of $X$.
-So, we can apply (P2) to $J$ in $G^{\prime}$, getting a further drop of $r-s$.
-Overall, we get a net drop of $s+(r-s)=r$ as desired.
-∎
-4 Branching rules
-Most of our branching rules can be viewed in terms of guessing that a certain set of vertices $X_{1}$ is in the cover.
-In the graph $G-X_{1}$, a set of vertices $X_{0}$ may become isolated.
-We apply rule (P1) to remove $X_{0}$; the resulting subproblem $\langle G-(X_{0}\cup X_{1}),k-|X_{1}|\rangle$ is called principal, and we define the excess to be $|X_{0}|$.
-Note that the graph $G^{\prime}=G-(X_{0}\cup X_{1})$ may still contain isolated vertices, or have additional simplifications available.
-We say $\langle G^{\prime},k-|X_{1}|\rangle$ has drop $(\Delta\mu^{\prime},|X_{1}|)$ directly, and the final subproblem $G^{\prime\prime}=\langle G^{\prime\prime},k^{\prime\prime}\rangle$ has drop $(\Delta\mu^{\prime\prime},\Delta k^{\prime\prime})$ after simplification.
-Since this plays a central role in our analysis, we define the shadow of a vertex set $X$, denoted $\operatorname{shad}(X)$, to be $\operatorname{minsurp}(G-X)$.
-For readability, we write $\operatorname{shad}(x_{1},\dots,x_{\ell},X_{1},\dots,X_{r})$ as shorthand for $\operatorname{shad}(\{x_{1},\dots,x_{\ell}\}\cup X_{1}\cup\dots\cup X_{r})=\operatorname{minsurp}(G-\{x_{1},\dots,x_{\ell}\}-(X_{1}\cup\dots\cup X_{r}))$.
-Proposition 14.
-If $\operatorname{minsurp}(G)\geq 2$, then a principal subproblem $\langle G-X,k^{\prime}\rangle$ with excess $s$ has $\Delta\mu=\tfrac{1}{2}(\Delta k-s+\operatorname{minsurp}^{-}(G^{\prime}))=\tfrac{1}{2}(\Delta k-s+\min\{0,\operatorname{shad}(X)\})$.
-Proof.
-Here Proposition 4 gives $\lambda(G)=n/2$.
-Let $G^{\prime}=G-X$, where $|X|=\Delta k+s$.
-Then $\mu(G^{\prime})=k^{\prime}-\operatorname{\lambda}(G^{\prime})=k^{\prime}-\tfrac{1}{2}(n^{\prime}+\operatorname{minsurp}^{-}(G^{\prime}))$, where $n^{\prime}=n-X=2\lambda(G)-(s+\Delta k)$, which simplifies to $\mu(G^{\prime})=\mu(G)-\tfrac{1}{2}(\Delta k-s+\operatorname{minsurp}^{-}(G^{\prime}))$ as desired.
-∎
-Observation 15.
-Suppose $\operatorname{minsurp}(G)\geq 2$. Then for any indset $I$, there holds $\operatorname{shad}(I)\geq 2-|I|$ and $\operatorname{shad}(N[I])\geq 2-\operatorname{surp}_{G}(I)$.
-In particular, for a vertex $u$, there holds $\operatorname{shad}(u)\geq 1$ and $\operatorname{shad}(N[u])\geq 3-\deg(u)$.
-Proof.
-If $J$ is a min-set of $G-I$, then $2\leq\operatorname{surp}_{G}(J)\leq\operatorname{surp}_{G-I}(J)+|I|$.
-So $\operatorname{surp}_{G-I}(J)\geq 2-|I|$.
-Similarly, if $J$ is a min-set of $G-N[I]$, then $2\leq\operatorname{surp}_{G}(I\cup J)=\operatorname{surp}_{G-N[I]}(J)+\operatorname{surp}_{G}(I)$.
-∎
-By combining Observation 11 and Observation 15, we immediately get the following simplification rule (which is used ubiquitously):
-Observation 16.
-If $G$ is simplified and $G-\{u,v\}$ has two non-adjacent subquadratic vertices, then $R(u,v)\geq 2$.
-Our bread-and-butter branching rule is to choose some vertex $u$, and branch on whether it is in the cover.
-We refer to this as splitting on $u$.
-This generates two principal subproblems $\langle G-u,k-1\rangle$ and $\langle G-N[u],k-\deg(u)\rangle$.
-More generally, in light of Lemma 5, we know that for any critical-set $I$, there is either a good cover containing $I$ or omitting $I$.
-We can branch on subproblems $\langle G-I,k-|I|\rangle$ and $\langle G-N[I],k-N[I]\rangle$, and we refer to this as splitting on $I$.
-Splitting on a vertex is a special case where $I$ is a singleton.
-Consider the effect of splitting on $u$.
-By Observation 15, the subproblem $\langle G-u,k-1\rangle$ has drop $(0.5,1)$ and the subproblem $\langle G-N[u],k-\deg(u)\rangle$ has drop $(1,\deg(u))$.
-The latter bound, however, is very loose and should be improved to get an optimized analysis.
-In light of Proposition 14, it revolves around the value of $\operatorname{shad}(N[u])$.
-If $\operatorname{shad}(N[u])\geq 0$, then $\lambda$ drops by $\frac{\deg(u)+1}{2}$ and so the subproblem has drop $(\tfrac{\deg(u)-1}{2},\deg(u))$.
-This is the “generic” situation for splitting on $u$.
-When $\operatorname{shad}(N[u])\leq 0$, we say that $u$ is blocked.
-If $x$ is a vertex in a min-set of $G-N[u]$, we say that $x$ is a blocker of $u$.
-Necessarily, $x\nsim u$.
-This motivates the following powerful branching rule:
-(B) If $x$ is a blocker of vertices $u_{1},\dots,u_{\ell}$, then branch on subproblems $\langle G-\{u_{1},\dots,u_{\ell},x\},k-\ell-1\rangle$ and $\langle G-N[x],k-\deg(x)\rangle$.
-Proposition 17.
-Rule (B) is a valid branching rule.
-Proof.
-The subproblem $\langle G-N[x],k-\deg(x)\rangle$ is feasible if and only if $G$ has a good cover omitting $x$.
-The subproblem $\langle G-\{u_{1},\dots,u_{\ell},x\},k-\ell-1\rangle$ is feasible if and only if $G$ has a good cover which includes all vertices $u_{1},\dots,u_{\ell},x$.
-We claim that, if $G$ is feasible, at least one of these cases holds.
+The subproblem $G_{0}$ is feasible if and only if $G$ has a good cover omitting $x$, and the subproblem $G_{1}$ is feasible if and only if $G$ has a good cover which includes all vertices $u_{1},\dots,u_{\ell},x$.
+We claim that, if $\langle G,k\rangle$ is feasible, at least one of these cases holds.
 For, suppose a good cover of $G$ omits vertex $u_{i}$.
 Then $G-N[u_{i}]$ has a cover of size at most $k-\deg(u_{i})$.
-By Lemma 5, this implies that $G-N[u_{i}]$ has such a cover $C^{\prime}$ which omits any min-set $I$ of $G-N[u_{i}]$, and in particular $x\notin C^{\prime}$.
-Then $G$ has a good cover $C\cup N(u_{i})$ omitting $x$.
+By Lemma 6, then $G-N[u_{i}]$ has such a cover $C^{\prime}$ which omits any min-set II of $G-N[u_{i}]$, and in particular $x\notin C^{\prime}$.
+Then $G$ has a good cover $C=C^{\prime}\cup N(u_{i})$ omitting $x$.
 ∎
-In the vast majority of cases where we use (B), we have $\ell=1$, that is, $x$ is a blocker of a single vertex $u$.
-To provide intuition, keep in mind the picture that if $G-N[u]$ has $s$ isolated vertices $v_{1},\dots,v_{s}$, then $\{v_{1},\dots,v_{s}\}$ would be an indset in $G-N[u]$ with zero neighbors and surplus $-s$, in particular, $\operatorname{shad}(N[u])\leq-s$.
-These vertices were hidden in the “shadow” of $N[u]$.
+In the vast majority of cases, we use (B-Block) with $\ell=1$, that is, $x$ is a blocker of a single vertex $u$.
+For intuition, keep in mind the picture where GN[u]G-N[u] has $s$ isolated vertices $v_{1},\dots,v_{s}$, so that $\{v_{1},\dots,v_{s}\}$ is an indset in GN[u]G-N[u] with zero neighbors and surplus s-s, in particular, $\operatorname{shad}(N[u])\leq-s$.
 See Figure 6.
-Figure 5: Here, vertices $v_{1},v_{2}$ are both blockers of $u$; in particular, we have $\operatorname{shad}(N[u])\leq-2$.
+Figure 6: Here, vertices $v_{1},v_{2}$ are both blockers of $u$; in particular, $\operatorname{shad}(N[u])\leq-2$.
 Also, note that if $\operatorname{shad}(N[u])=0$ exactly, then we could still split on $u$ if desired and get the ideal drop in $\mu$.
-However, we can also apply rule (B) if desired, and this is usually more profitable.
+However, we can also apply rule (B-Block), and this is usually more profitable.
 Vertices with $\operatorname{shad}(N[u])=0$ share many properties with vertices with $\operatorname{shad}(N[u])<0$ strictly.
 This is the reason for the somewhat unintuitive definition of blocked vertex.
-5 Analysis of blockers and splitting
-We will develop a series of branching rules (typically using rule (B)) to handle blocked vertices, along with a number of other related cases.
-The most significant consequence of these rules is that we are able to get near-ideal branch sequences for high-degree vertices.
-Specifically, we will show the following main result which is the heart of our branching algorithm:
+We will develop a series of branching rules to handle blocked vertices and other related cases.
+The most significant consequence of these rules is to give near-ideal branch sequences for high-degree vertices.
+Specifically, we have the following main result which is the heart of our algorithm:
 {restatable}
-theoremthmA Suppose $G$ is simplified, and let $r=\operatorname{maxdeg}(G)$. Depending on $r$, then following branch-seqs are available:
+theoremthmA Suppose $G$ is simplified, and let $r=\operatorname{maxdeg}(G)$. Depending on $r$, the following branch-seqs are available:
 If $r\geq 4$: $[(1,3),(1,5)]$ or $[(0.5,1),(1.5,4)]$.
@@
 If $r\geq 6$: $[(1,3),(1,5)]$, $[(0.5,2),(2,5)]$, or $[(0.5,1),(2.5,r)]$.
-The proof of Section 5 has many cases which will require us to build up slowly from analysis of low-degree vertices.
-For Section 5 only, we always assume that the starting graph $\boldsymbol{G}$ is simplified.
-We begin with a few elementary observations.
-Observation 18.
-Suppose $u$ is blocked and let $I$ be a min-set of $G-N[u]$. There is at least one vertex $x\in I$ with $\operatorname{codeg}(x,u)\geq 1$.
+The proof of Figure 6 has many cases which we will build up slowly. For Section 3 only, we always assume that the starting graph $\boldsymbol{G}$ is simplified.
+3.1 Branching rules for indsets with surplus two
+We first develop branching rules for non-singleton indsets with surplus two.
+As a starting point, note that since $G$ is assumed to be simplified, we have $\operatorname{minsurp}(G)\geq 2$ hence such a set $I$ would be a critical-set.
+We can apply (B-Crit) to obtain subproblems $\langle G-I,k-|I|\rangle$ and $\langle G-N[I],k-r-2\rangle$.
+These are principal subproblems, and they have drops $(1,r),(1,r+2)$ respectively.
+This is already a powerful branching rule, but it can be improved when $|I|=2$ or $|I|=3$ by finding additional simplifications in the subproblems. We start with a few preliminary observations.
+Observation 19.
+Suppose $G$ has an indset $I$ with $\operatorname{surp}_{G}(I)=2$, and let $z\in N(I)$.
+Then $S(z)\geq|I|$, and for any vertex $x\notin I$, there holds $S(z,x)\geq|I|$.
 Proof.
-If not, we would have $\operatorname{surp}_{G}(I)=\operatorname{surp}_{G-N[u]}(I)\leq 0$, contradicting that $G$ is simplified. ∎
+We have $\operatorname{surp}_{G-z}(I)=\operatorname{surp}_{G}(I)-1=1$, and $\operatorname{shad}(z)\geq\operatorname{minsurp}(G)-1\geq 1$.
+So (P2) applied to $I$ gives $S(z)\geq|I|$.
+Likewise, $\operatorname{shad}(z,x)\geq\operatorname{minsurp}(G)-2\geq 0$ and $\operatorname{surp}_{G-\{z,x\}}(I)\leq\operatorname{surp}_{G}(I)-1\leq 1$.
+So Lemma 13 gives $S(z,x)\geq|I|$.
+∎
 Proposition 20.
-Suppose vertex $u$ has $\operatorname{shad}(N[u])\leq 4-\deg(u)$ and $x$ is a blocker of $u$.
-If we apply (B) to vertices $u,x$, then subproblem $G^{\prime}=\langle G-N[x],k-\deg(x)\rangle$ has drop $(1,4)$.
-Moreover, if $\deg(x)=3$ and $x$ is contained in a non-singleton min-set of $G-N[u]$, then $G^{\prime}$ has drop $(1,5)$.
+Suppose $\operatorname{shad}(N[u])\leq 4-\deg(u)$ and $x$ is a blocker of $u$.
+If we apply (B-Block) to vertices $u,x$, then subproblem $G^{\prime}=\langle G-N[x],k-\deg(x)\rangle$ has drop $(1,4)$.
+Moreover, if $\deg(x)=3$ and $x$ is contained in a non-singleton min-set of GN[u]G-N[u], then $G^{\prime}$ has drop $(1,5)$.
 Proof.
-Since $G$ is simplified, we have $\deg(x)\geq 3$ and $\operatorname{minsurp}(G)\geq 2$.
 If $\deg(x)=4$, then $G^{\prime}$ has drop $(1,4)$ directly.
-So, suppose $\deg(x)=3$ exactly.
-Consider any min-set $I$ of $G-N[u]$ with $x\in I$.
+So, suppose $\deg(x)=3$; by Proposition 16, $G^{\prime}$ has drop $(1,3)$ directly.
+Let $I\ni x$ be a min-set of GN[u]G-N[u].
 The indset $J=I\cup\{u\}\setminus\{x\}$ has $\operatorname{surp}_{G-N[x]}(J)\leq\operatorname{surp}_{G-N[u]}(I)+(\deg(u)-\deg(x))\leq 1$.
 On the other hand, $\operatorname{shad}(N[x])\geq\operatorname{minsurp}(G)-\deg(x)+1\geq 0$.
-So Lemma 13 gives $R(N[x])\geq|J|=|I|$.
-By Proposition 14 we see that $G^{\prime}$ has drop $(1,3)$ directly.
-Since $|I|\geq 1$ trivially, we have $R(N[x])\geq 1$ and $G^{\prime}$ has drop $(1,4)$ after simplification.
-If $|I|\geq 2$, then $R(N[x])\geq 2$ and likewise $G^{\prime}$ has drop $(1,5)$ after simplification.
-∎
-5.1 Branching rules for indsets with surplus two.
-It will be very important to develop special branching rules for indsets with surplus two.
-Since $G$ is assumed to be simplified, note that any such set $I$ will be a critical set, and we can split on $I$ to get subproblems $\langle G-I,k-|I|\rangle$ and $\langle G-N[I],k-|I|-2\rangle$.
-However, this branching rule is not powerful enough for us; we develop specialized branching rules for smaller indsets (indsets with cardinality two or three).
-Proposition 20.
-Suppose $G$ has an indset $I$ with $\operatorname{surp}_{G}(I)=2$, and let $z\in N(I)$.
-If we split on $z$, the subproblem $\langle G-z,k-1\rangle$ has drop $(0.5,1+|I|)$.
-If $z$ has a blocker $x\notin I$ and we apply (B) to $z,x$, the subproblem $\langle G-\{z,x\},k-2\rangle$ has drop $(1,2+|I|)$.
-Proof.
-Subproblem $\langle G-z,k-1\rangle$ has drop $(0.5,1)$ directly.
-We have $\operatorname{surp}_{G-z}(I)=\operatorname{surp}_{G}(I)-1=1$, and $\operatorname{shad}(G-z)\geq\operatorname{minsurp}(G)-1\geq 1$.
-So (P2) applied to $I$ gives $R(z)\geq|I|$.
-Likewise, subproblem $\langle G-\{z,x\},k-2\rangle$ has drop $(1,2)$ directly, and $\operatorname{shad}(z,x)\geq\operatorname{minsurp}(G)-2\geq 0$ and $\operatorname{surp}_{G-\{z,x\}}(I)\leq\operatorname{surp}_{G}(I)-1\leq 1$.
-So Lemma 13 gives $R(z,x)\geq|I|$.
+So Lemma 13 gives $S(N[x])\geq|J|=|I|$.
 ∎
@@
 Proof.
-If $|I|\geq 4$, then splitting on $I$ gives branch-seq $[(1,4),(1,6)]$.
+If $|I|\geq 4$, then we simply apply (B-Crit).
 So suppose $|I|=3$.
-We first claim that there is some vertex $z$ with $|N(z)\cap I|=2$.
-For, consider the set of vertices $J=\{z:|N(z)\cap I|\geq 2\}$.
-The vertices in $I$ are independent and have degree at least three, so there are at least 9 edges from $I$ to $N(I)$.
-Since $|N(I)|=5$, by the pigeonhole principle we must have $|J|\geq 2$.
-If $N(z)\not\subseteq I$ for all $z\in J$, then $J$ would be an indset with surplus at most one, contradicting that $G$ is simplified.
-Hence there is $z\in J$ with $|N(z)\cap I|=2$ exactly.
-Let $z$ be any such vertex, and let $I=\{x_{1},x_{2},y\}$ where $z\sim x_{1},z\sim x_{2},z\not\sim y$ and $\deg(x_{1})\leq\deg(x_{2})$.
-If $\deg(z)\leq 4$, then we split on $I$, generating subproblems $\langle G-I,k-3\rangle$ and $\langle G-N[I],k-5\rangle$ with drops $(1,3)$ and $(1,5)$ directly, where $G-I$ has a subquadratic vertex $z$ so it has drop $(1,4)$ after simplification.
-If $\deg(z)\geq 5$ and $\operatorname{shad}(N[z])\geq 5-\deg(z)$, then we split on $z$; by Proposition 20, subproblem $\langle G-z,k-1\rangle$ has drop $(0.5,4)$, and subproblem $\langle G-N[z],k-\deg(z)\rangle$ has drop $(2,5)$ directly.
-If $\deg(z)\geq 5$ and $z$ has a blocker $t\neq y$, then we apply (B) to $z,t$; by Proposition 19, subproblem $\langle G-N[t],k-\deg(t)\rangle$ has drop $(1,4)$ and by Proposition 20, subproblem $\langle G-\{z,t\},k-2\rangle$ has drop $(1,5)$.
-The only remaining possibility is if $\deg(z)\geq 5$ and $y$ is the only blocker of $z$, i.e.
-$\deg(z)=5,\deg(y)=3$ and $N(y)\subseteq N(z)$.
-Each vertex $x_{i}$ has $\deg(x_{i})+\deg(y)-\operatorname{codeg}(x_{i},y)\leq|N(I)|$, i.e.
-$\operatorname{codeg}(x_{i},y)\geq\deg(x_{i})-2$.
-If $\deg(x_{1})=3$, this implies that $x_{1},y$ share some neighbor $u$; but in this case $G$ has a 3-triangle $x_{1},u,z$, contradicting that $G$ is simplified.
-So we suppose $\deg(x_{2})\geq\deg(x_{1})\geq 4$ and then $\operatorname{codeg}(x_{1},y)\geq 2$ and $\operatorname{codeg}(x_{2},y)\geq 2$.
-Equivalently, $\deg_{G-N[x_{1}]}(y)\leq 1$ and $\deg_{G-N[x_{2}]}(y)\leq 1$.
-So $y$ is a blocker of all three vertices $x_{1},x_{2},z$.
-We apply (B) giving subproblems $\langle G-N[y],k-\deg(y)\rangle$ and $\langle G-\{x_{1},x_{2},z,y\},k-4\rangle$, with drops $(1,3)$ and $(1,4)$ directly.
-Furthermore, $\operatorname{surp}_{G-N[y]}(\{x_{1},x_{2}\})\leq 2-|N_{G-N[y]}(I)|=2-2=0$ and neither $x_{1}$ or $x_{2}$ is isolated in $G-N[y]$ (they retain neighbor $z$).
-So we can apply (P1) to indset $\{x_{1},x_{2}\}$ in $G-N[y]$, giving $R(N[y])\geq 2$.
-So $G-N[y]$ has drop $(1,5)$ after simplification.
+The vertices in $I$ are independent, so there are at least $\sum_{x\in I}\deg(x)\geq 9$ edges from $I$ to $N(I)$.
+Since $|N(I)|=5$, by the pigeonhole principle there must be at least one vertex $z$ with $|N(z)\cap I|\geq 2$.
+Our strategy will revolve around a chosen vertex $z$; as a tie-breaking rule, we assume that in a given case analysis there were no other vertices in a previous case.
+Case I: $\boldsymbol{\deg(z)\leq 4}$ and $\boldsymbol{N(z)\not\subseteq I}$.
+We apply (B-Crit) to $I$, generating subproblems $\langle G-I,k-3\rangle$ and $\langle G-N[I],k-5\rangle$ with drops $(1,3)$ and $(1,5)$ directly.
+The former subproblem has a subquadratic vertex $z$ so it has drop $(1,4)$ after simplification.
+Case II: $\boldsymbol{\deg(z)\geq 5}$ and $\boldsymbol{\operatorname{shad}(N[z])\geq 5-\deg(z)}$.
+We split on $z$; by Observation 19, subproblem $\langle G-z,k-1\rangle$ has drop $(0.5,4)$ after simplification, and subproblem $\langle G-N[z],k-\deg(z)\rangle$ has drop $(2,5)$ directly.
+Case III: $\boldsymbol{\deg(z)\geq 5}$ and $\boldsymbol{z}$ has a blocker $\boldsymbol{t\notin I}$.
+We apply (B-Block) to $z,t$; by Proposition 20, subproblem $\langle G-N[t],k-\deg(t)\rangle$ has drop $(1,4)$ and by Observation 19, subproblem $\langle G-\{z,t\},k-2\rangle$ has drop $(1,5)$ after simplification.
+Case IV: $\boldsymbol{\deg(z)\geq 5}$.
+Let $I=\{x_{1},x_{2},y\}$ where $z\sim x_{1},z\sim x_{2}$.
+Since we are not in Case II or Case III, it must be that $y$ is the sole blocker of $z$: that is, we have $N(y)\subseteq N(z)$.
+Each vertex $x_{i}$ has $|N(x_{i},y)|=\deg(x_{i})+\deg(y)-\operatorname{codeg}(x_{i},y)\leq|N(I)|$, i.e.
+$\operatorname{codeg}(x_{i},y)\geq\deg(x_{i})-2\geq 1$.
+So we must have $\deg(x_{1}),\deg(x_{2})\geq 4$, as otherwise some $x_{i}$ would have a 3-triangle with $u,z$, contradicting that $G$ is simplified.
+As a result, there are at least $11$ edges going from $I$ to $N(I)$; since $|N(I)|=5$ the pigeonhole principle implies there must be some vertex $z^{\prime}$ with $|N(z^{\prime})\cap I|=3$.
+Since $z^{\prime}\in N(y)\subseteq N(z)$, we must have $z^{\prime}\sim z$ as well.
+So $N(z^{\prime})\not\subseteq I$.
+Necessarily, any blocker of this vertex $z^{\prime}$ would be outside $I$.
+Depending on the degree of $z^{\prime}$, it would fall into one of the earlier cases.
+This contradicts our tie-breaking rule for $z$.
+So Case IV is impossible.
+Figure 7: Illustration of Case IV
+Case V: $\boldsymbol{N(z)\subseteq I}$.
+Necessarily $\deg(z)=3$.
+There are at least $6$ edges from $I$ to $N(I)\setminus\{z\}$, so by the pigeonhole principle there is another vertex $z^{\prime}\neq z$ with $|N(z^{\prime})\cap I|\geq 2$.
+If $N(z^{\prime})\subseteq I$, then necessarily $z\not\sim z^{\prime}$.
+So $\{z,z^{\prime}\}$ is an indset of surplus one, contradicting that $G$ is simplified.
+If $N(z^{\prime})\not\subseteq I$, it would fall into one of the earlier cases.
+In either event, Case V is also impossible.
 ∎
@@
 If $|I|\geq 3$, then we apply Lemma 21.
-So let $I=\{x,y\}$, where $\deg(x)\leq\deg(y)$ and let $A=N(x,y)$.
-Note that $|A|=\deg(x)+\deg(y)-4\geq\deg(x)-1$.
-The vertices in $A$ cannot form a clique, as then $x$ would have a funnel.
+So let $I=\{x_{1},x_{2}\}$ and let $A=N(x_{1})\cap N(x_{2})$.
+Note that $|A|=\deg(x_{1})+\deg(x_{2})-|N(I)|\geq\deg(x_{1})-1\geq 2$.
+The vertices in $A$ cannot form a clique, as then $x_{1}$ would have a funnel.
 So consider vertices $z_{1},z_{2}\in A$ with $z_{1}\not\sim z_{2}$.
-If $z_{1}$ and $z_{2}$ are both subquartic, then we split on $I$; the subproblem $\langle G-N[I],k-|N[I]|-2$ has drop $(1,4)$ directly.
-The subproblem $G-I=G-\{x,y\}$ has non-adjacent subquadratic vertices $z_{1},z_{2}$, so by Observation 16, we have $R(x,y)\geq 2$ and subproblem $\langle G-I,k-|I|\rangle$ has drop $(1,4)$ after simplification.
+If $z_{1}$ and $z_{2}$ are both subquartic, then we apply (B-Crit) to $I$; the subproblem $\langle G-N[I],k-4\rangle$ has drop $(1,4)$ directly.
+The subproblem $\langle G-I,k-2\rangle$ has non-adjacent subquadratic vertices $z_{1},z_{2}$, so by Observation 12 it has drop $(1,4)$ after simplification.
 So suppose $\deg(z_{1})\geq 5$.
-If $\operatorname{shad}(N[z_{1}])\geq 0$, then split on $z_{1}$; by Proposition 20, the subproblem $\langle G-z_{1},k-1\rangle$ has drop $(0.5,3)$, while the subproblem $\langle G-N[z_{1}],k-\deg(z_{1})\rangle$ has drop $(2,5)$ directly.
+If $\operatorname{shad}(N[z_{1}])\geq 0$, we split on $z_{1}$; by Observation 19, the subproblem $\langle G-z_{1},k-1\rangle$ has drop $(0.5,3)$ after simplification, while the subproblem $\langle G-N[z_{1}],k-\deg(z_{1})\rangle$ has drop $(2,5)$ directly.
 Otherwise, suppose $\operatorname{shad}(N[z_{1}])\leq-1$ and $t$ is a blocker of $z_{1}$.
-Note that $t\notin I$ since $t\not\sim z_{1}$ whereas $z_{1}\in N(x)\cap N(y)$.
-We then apply (B) to $t,z_{1}$; the subproblem $\langle G-N[z_{1}],k-\deg(z_{1})\rangle$ has drop $(1,5)$ directly, and by Proposition 20, the subproblem $\langle G-\{t,z\},k-2\rangle$ has drop $(1,4)$.
+Note that $t\notin I$ since $t\not\sim z_{1}$ whereas $z_{1}\in N(x_{1})\cap N(x_{2})$.
+We then apply (B-Block) to $t,z_{1}$; the subproblem $\langle G-N[z_{1}],k-\deg(z_{1})\rangle$ has drop $(1,5)$ directly, and by Observation 19, the subproblem $\langle G-\{t,z\},k-2\rangle$ has drop $(1,4)$ after simplification.
 ∎
-5.2 Branching rules for blocked vertices.
-We now turn to developing branching rules for blocked vertices of various types. We need to work our way from low to high degree.
+3.2 Branching rules for blocked vertices
+We now turn to developing branching rules for blocked vertices of various types.
 Proposition 23.
-Suppose vertex $u$ has $\operatorname{shad}(N[u])\leq 5-\deg(u)$ and $u$ has a blocker $x$ of degree at least 4. Then $G$ has available branch-seq $[(1,4),(1,4)]$ or $[(0.5,2),(2,5)]$.
+Suppose $\operatorname{shad}(N[u])\leq 5-\deg(u)$ and $u$ has a blocker $x$ of degree at least 4. Then $G$ has available branch-seq $[(1,4),(1,4)]$ or $[(0.5,2),(2,5)]$.
 Proof.
-If $\operatorname{shad}(N[x])\leq 3-\deg(x)$ and $J$ is a min-set of $G-N[x]$, then $\operatorname{surp}_{G}(J\cup\{x\})\leq(3-\deg(x))+(\deg(x)-1)\leq 2$ and we can apply Lemma 22 to indset $J\cup\{x\}$.
+If $\operatorname{shad}(N[x])\leq 3-\deg(x)$ and $J$ is a min-set of GN[x]G-N[x], then $\operatorname{surp}_{G}(J\cup\{x\})\leq(3-\deg(x))+(\deg(x)-1)\leq 2$ and we can apply Lemma 22 to indset $J\cup\{x\}$.
 So suppose $\operatorname{shad}(N[x])\geq 4-\deg(x)$.
-In this case, we apply (B) to $u,x$.
+In this case, we apply (B-Block) to $u,x$.
 Subproblem $\langle G-\{u,x\},k-2\rangle$ has drop $(1,2)$ directly.
 If $\deg(x)\geq 5$, then subproblem $\langle G-N[x],k-\deg(x)\rangle$ has drop $(1.5,5)$ directly.
-If $\deg(x)=4$, then consider a min-set $I$ of $G-N[u]$ with $x\in I$.
+If $\deg(x)=4$, then consider a min-set II of GN[u]G-N[u] with $x\in I$.
 We have $\operatorname{surp}_{G-N[x]}(I\cup\{u\}\setminus\{x\})\leq\operatorname{surp}_{G-N[u]}(I)+\deg(u)-\deg(x)\leq 1$.
-So $R(N[x])\geq 1$ and $G^{\prime}$ has drop $(1.5,5)$ after simplification.
+So $S(N[x])\geq 1$ and $G^{\prime}$ has drop $(1.5,5)$ after simplification.
 In either case, $\langle G-\{u,x\},k-2\rangle$ and $\langle G-N[x],k-\deg(x)\rangle$ give drops $(1.2),(1.5,5)$ respectively.
@@
 Case I: $\boldsymbol{\operatorname{shad}(N[u])\geq 5-\deg(u)}$.
-Splitting on $u$ gives subproblems $G-u,G-N[u]$ with drops $(0.5,1),(2,5)$ directly; the former has drop $(0.5,2)$ after simplification due to its 2-neighbor $z$.
+Splitting on $u$ gives subproblems $\langle G-u,k-1\rangle$ and $\langle G-N[u],k-\deg(u)\rangle$ with drops $(0.5,1),(2,5)$ directly; the former has drop $(0.5,2)$ after simplification due to its 2-vertex $z$.
 Case II: $\boldsymbol{u}$ has a blocker $\boldsymbol{x}$ with $\boldsymbol{\deg(x)\geq 4}$.
@@
 Then the result follows from Proposition 23, noting that $[(1,4),(1,4)]$ dominates $[(1,3),(1,5)]$.
-Case III: $\boldsymbol{u}$ has a blocker $\boldsymbol{x}$ with $\boldsymbol{R(u,x)\geq 2}$.
-We apply (B) to vertices $u,x$.
+Case III: $\boldsymbol{u}$ has a blocker $\boldsymbol{x}$ with $\boldsymbol{S(u,x)\geq 2}$.
+We apply (B-Block) to vertices $u,x$.
 Subproblem $\langle G-\{u,x\},k-2\rangle$ has drop $(1,4)$ after simplification, and by Proposition 20 subproblem $\langle G-N[x],k-\deg(x)\rangle$ has drop $(1,4)$.
 We get branch-seq $[(1,4),(1,4)]$, which dominates $[(1,3),(1,5)]$.
-Case IV: $\boldsymbol{G-N[u]}$ has some non-singleton min-set $\boldsymbol{I}$.
+Case IV: $\boldsymbol{G-N[u]}$ has a non-singleton min-set $\boldsymbol{I}$.
 Let $x\in I$.
 Necessarily $\deg(x)=3$, else it would be covered in Case II.
-We apply (B) to $u,x$; subproblem $\langle G-\{u,x\},k-2\rangle$ has drop $(1,3)$ after simplification due to subquadratic vertex $z$.
+We apply (B-Block) to $u,x$; subproblem $\langle G-\{u,x\},k-2\rangle$ has drop $(1,3)$ after simplification due to subquadratic vertex $z$.
 By Proposition 20 subproblem $\langle G-N[x],k-\deg(x)\rangle$ has drop $(1,5)$ after simplification.
@@
 Suppose that no vertices in the graph are covered by any of the previous cases.
-Let $U$ denote the non-empty set of vertices $u$ with $\deg(u)\geq 5$ and $u$ having a 3-neighbor.
+Let $U\neq\emptyset$ denote the set of vertices $u$ with $\deg(u)\geq 5$ and $u$ having a 3-neighbor.
 We claim that every vertex $u\in U$ has degree exactly 5.
 For, suppose $\deg(u)\geq 6$, and since we are not in Case I we have $\operatorname{shad}(N[u])\leq 4-\deg(u)\leq-2$.
-Then necessarily a min-set of $G-N[u]$ would have size at least two, and it would be covered in Case IV.
+Then necessarily a min-set of GN[u]G-N[u] would have size at least two, and it would be covered in Case IV.
 So every $u\in U$ has degree exactly 5, and is blocked by a singleton 3-vertex $x$, i.e.
@@
 For, suppose that $x$ has two subquartic neighbors $z_{1},z_{2}$; necessarily $z_{1}\not\sim z_{2}$ since $G$ is simplified and $x$ has degree 3.
-Then $R(u,x)\geq 2$ by Observation 16, which would have been covered in Case III.
+Then $S(u,x)\geq 2$ by Observation 12, which would have been covered in Case III.
 On the other hand, we claim that each vertex $u\in U$ has at most one neighbor in $X$.
 For, suppose $u$ is linked to $x$ and $u$ has two 3-neighbors $x_{1},x_{2}\in X$, which must be non-adjacent since $G$ is simplified.
-Then Observation 16 gives $R(u,x)\geq 2$ due to subquadratic vertices $x_{1},x_{2}$.
+Then Observation 12 gives $S(u,x)\geq 2$ due to subquadratic vertices $x_{1},x_{2}$.
 This would have been covered in Case III.
@@
 Hence $|U|\geq 2|X|$, and so by the pigeonhole principle, there must be some vertex $x\in X$ which is linked to two vertices $u_{1},u_{2}\in U$.
-We apply (B) to $u_{1},u_{2},x$, getting subproblems $\langle G-\{u_{1},u_{2},x\},k-3\rangle$ and $\langle G-N[x],k-\deg(x)\rangle$.
+We apply (B-Block) to $u_{1},u_{2},x$, getting subproblems $\langle G-\{u_{1},u_{2},x\},k-3\rangle$ and $\langle G-N[x],k-\deg(x)\rangle$.
 By Proposition 20, the latter subproblem has drop $(1,4)$.
@@
 Proposition 25.
-Suppose a vertex $u$ has $\deg(u)\geq 4$ and $\operatorname{shad}(N[u])\leq 4-\deg(u)$. Then $G$ has available branch-seq $[(1,3),(1,5)]$ or $[(0.5,2),(2,5)]$.
+Suppose a vertex $u$ has $\operatorname{shad}(N[u])\leq 4-\deg(u)\leq 0$. Then $G$ has available branch-seq $[(1,3),(1,5)]$ or $[(0.5,2),(2,5)]$.
 Proof.
-Let $I$ be a min-set of $G-N[u]$.
-By Observation 18, there is a vertex $x\in I$ which shares some neighbor $t$ with $u$.
-If $\deg(x)\geq 4$, then we apply Proposition 23, noting that $[(1,4),(1,4)]$ dominates $[(1,3),(1,5)]$.
+Let $I$ be a min-set of GN[u]G-N[u].
+There must be some vertex $x\in I$ which shares some neighbor $t$ with $u$, as otherwise we would have $\operatorname{surp}_{G}(I)=\operatorname{surp}_{G-N[u]}(I)\leq 0$, contradicting that $G$ is simplified.
+If $\deg(x)\geq 4$, we apply Proposition 23, noting that $[(1,4),(1,4)]$ dominates $[(1,3),(1,5)]$.
 If $\deg(x)=3$ and $\deg(t)\geq 5$, we apply Lemma 24 to $t$.
-If $\deg(x)=3,\deg(t)\leq 4,|I|\geq 2$, we apply (B) to $u,x$; subproblem $\langle G-\{u,x\},k-2\rangle$ has drop $(1,3)$ after simplification (due to subquadratic vertex $t$) and by Proposition 19 subproblem $\langle G-N[x],k-\deg(x)\rangle$ has drop $(1,5)$.
+If $\deg(x)=3,\deg(t)\leq 4,|I|\geq 2$, we apply (B-Block) to $u,x$; subproblem $\langle G-\{u,x\},k-2\rangle$ has drop $(1,3)$ after simplification (due to subquadratic vertex $t$) and by Proposition 20 subproblem $\langle G-N[x],k-\deg(x)\rangle$ has drop $(1,5)$.
 So finally take $I=\{x\}$ where $\deg(x)=3$ and $x$ shares two neighbors $t_{1},t_{2}$ with $u$.
@@
 If either $t_{1}$ or $t_{2}$ has degree 5 or more, we can apply Lemma 24.
-If they are both subquartic, then we apply (B) to $u,x$; by Proposition 19, subproblem $\langle G-N[x],k-\deg(x)\rangle$ has drop $(1,4)$ and by Observation 16 there holds $R(u,x)\geq 2$ and subproblem $\langle G-\{u,x\},k-2\rangle$ has drop $(1,4)$ after simplification.
+If they are both subquartic, then we apply (B-Block) to $u,x$; by Proposition 20, subproblem $\langle G-N[x],k-\deg(x)\rangle$ has drop $(1,4)$ and by Observation 12 the subproblem $\langle G-\{u,x\},k-2\rangle$ has drop $(1,4)$ after simplification.
 Here, $[(1,4),(1,4)]$ dominates $[(1,3),(1,5)]$.
@@
 Lemma 26.
-Suppose vertex $u$ has $\deg(u)\geq 5$ and $\operatorname{shad}(N[u])\leq 5-\deg(u)$.
+Suppose vertex $u$ has $\operatorname{shad}(N[u])\leq 5-\deg(u)$.
   • If $\deg(u)=5$, then $G$ has available branch-seq $[(1,3),(1,4)]$ or $[(0.5,2),(2,5)]$.
@@
 Proof.
-Let $I$ be a min-set of $G-N[u]$ of smallest size.
+Let $I$ be an inclusion-wise-minimal min-set II of GN[u]G-N[u].
 If any vertex in $I$ has degree at least 4, the result follows from Proposition 23 (noting that $[(1,4),(1,4)]$ dominates $[(1,3),(1,5)]$.
 So suppose that all vertices in $I$ have degree 3.
-By Observation 18, we may select some $x\in I$ with $\operatorname{codeg}(x,u)\geq 1$.
-For this vertex $x$, define $Z=N(x)\cap N(u)$ and $Y=N_{G-N[u]}(x)$, where $|Z|+|Y|=\deg(x)=3$ and $|Z|\geq 1$.
+There must be some vertex $x\in I$ with $\operatorname{codeg}(x,u)\geq 1$ as otherwise we would have $\operatorname{surp}_{G}(I)=\operatorname{surp}_{G-N[u]}(I)\leq 0$.
+For this vertex $x$, define $Z=N(x)\cap N(u)$ and $Y=N(x)\setminus N(u)$, where $|Z|+|Y|=\deg(x)=3$ and $|Z|\geq 1$.
 If any vertex $z\in Z$ has degree at least 5, we apply Lemma 24 to $z$ with its 3-neighbor $x$.
@@
 The vertices in $Z$ must be independent, as otherwise $G$ would have a 3-triangle with $x$.
-If $\operatorname{codeg}(z,z^{\prime})\geq 3$ for any vertices $z,z^{\prime}\in Z$, then $\operatorname{shad}(N[z])\leq 0$ (since $G-N[z]$ has a 1-vertex $z^{\prime}$), and we can apply Proposition 25.
+If $\operatorname{codeg}(z,z^{\prime})\geq 3$ for any vertices $z,z^{\prime}\in Z$, then $\operatorname{shad}(N[z])\leq 0$ (due to vertex $z^{\prime}$), and we can apply Proposition 25.
 So we suppose the vertices in $Z$ share no neighbors besides $u$ and $x$.
-In this case, in the graph $G-\{u,x\}$, the vertices in $Z$ become subquadratic and share no common neighbors.
-By Proposition 12, we thus have $R(u,x)\geq|Z|$.
-We next claim that if $|I|\geq 2$, then $R(u,x)\geq|Y|$.
-For, every vertex $y\in Y$ must have a neighbor $\sigma(y)\in I\setminus\{x\}$, as otherwise we would have $\operatorname{surp}_{G-N[u]}(I^{\prime})\leq\operatorname{surp}_{G-N[u]}(I)$ for non-empty ind-set $I^{\prime}=I\setminus\{x\}$, contradicting minimality of $I$.
+In this case, in the graph $G-\{u,x\}$, the vertices in $Z$ have degree two and share no common neighbors.
+By Proposition 11, we thus have $S(u,x)\geq|Z|$.
+Now suppose that $|I|=1$, i.e.
+$I$ consists of a single vertex $x$ with $\deg_{G-N[u]}(x)\leq 6-\deg(u)$.
+In particular, if $\deg(u)=5$, then $|Z|\geq 2$ and if $\deg(u)\geq 6$ then $|Z|\geq 3$.
+In this case, when we apply (B-Block) to $u,x$, then subproblem $\langle G-N[x],k-\deg(x)\rangle$ has drop $(1,3)$ directly and subproblem $\langle G-\{u,x\},k-2\rangle$ has drop $(1,4)$ or $(1,5)$ after simplification, for $\deg(u)=5$ and $\deg(u)=6$ respectively.
+This shows the claimed bound.
+So suppose that $|I|\geq 2$.
+Each vertex $y\in Y$ must have a neighbor $\sigma(y)\in I\setminus\{x\}$, as otherwise we would have $\operatorname{surp}_{G-N[u]}(I^{\prime})\leq\operatorname{surp}_{G-N[u]}(I)$ for non-empty indset $I^{\prime}=I\setminus\{x\}$, contradicting minimality of $I$.
 The vertices $\sigma(y):y\in Y$ must be distinct, as a common vertex $x^{\prime}=\sigma(y)=\sigma(y^{\prime})$ would have $\operatorname{codeg}(x,x^{\prime})\geq 2$ which we have already ruled out.
-So, in the graph $G-N[x]$, each vertex $\sigma(y)\in I$ loses neighbor $y$.
-By Observation 15, we have $\operatorname{shad}(N[x])\geq 3-\deg(x)=0$.
-So $G-N[x]$ has $|Y|\leq 2$ non-adjacent subquadratic vertices.
-Observation 16 implies that $R(N[x])\geq|Y|$.
-At this point, our strategy is to apply (B) to $u,x$.
-If $|I|=1$, then either $\deg(u)=5,|Z|\geq 2$ or $\deg(u)\geq 6,|Z|=3$.
-Subproblem $\langle G-N[x],k-\deg(x)\rangle$ has drop $(1,3)$ directly and subproblem $\langle G-\{u,x\},k-2\rangle$ has drop $(1,2+|Z|)$ after simplification; that is, it has drop $(1,4)$ if $\deg(u)=5$ and drop $(1,5)$ otherwise.
-Otherwise, if $|I|\geq 2$, then subproblem $\langle G-\{u,x\},k-2\rangle$ has drop $(1,2+|Z|)$ after simplification and subproblem $\langle G-N[x],k-\deg(x)\rangle$ has drop $(1,3+|Y|)$ after simplification.
-Since $|Z|+|Y|=3$ and $|Z|\geq 1$, this always gives drops $(1,3),(1,5)$ or $(1,4),(1,4)$.
+So, in the graph GN[x]G-N[x], each vertex $\sigma(y)\in I$ loses neighbor $y$.
+By Observation 9, we have $\operatorname{shad}(N[x])\geq 3-\deg(x)=0$.
+So GN[x]G-N[x] has $|Y|\leq 2$ non-adjacent subquadratic vertices.
+Proposition 10 implies that $S(N[x])\geq|Y|$.
+Finally, our strategy is to apply (B-Block) to $u,x$.
+Subproblem $\langle G-\{u,x\},k-2\rangle$ has drop $(1,2+|Z|)$ after simplification and subproblem $\langle G-N[x],k-\deg(x)\rangle$ has drop $(1,3+|Y|)$ after simplification.
+Since $|Z|+|Y|=3$ and $|Z|\geq 1$, this always gives $(1,3),(1,5)$ or $(1,4),(1,4)$.
 ∎
@@
 4 A simple branching algorithm
-We now describe our first branching algorithm for vertex cover.
-As discussed earlier, the measure is a piecewise-linear function of $k$ and $\mu$.
-It is more convenient to describe it in terms of multiple self-contained “mini-algorithms,” with linear measure functions.
-As a starting point, we can describe the first algorithm:
-Simplify $G$.
-If $\operatorname{maxdeg}(G)\leq 3$, then run either the algorithm of Theorem 2 or the MaxIS-3 algorithm, whichever is cheaper, and return.
-Otherwise, apply Section 5 to an arbitrary vertex of degree at least $4$, and run Branch4Simple on the two resulting subproblems.
-Algorithm 1 Function Branch4Simple$(G,k)$
+We now describe our first branching algorithm for vertex cover, where the measure is a piecewise-linear function of $k$ and $\mu$.
+It is more convenient to describe this in terms of multiple self-contained “mini-algorithms,” with linear measure functions.
+As a starting point, we can describe the first mini-algorithm:
+1Simplify $G$.
+2If $\operatorname{maxdeg}(G)\leq 3$, then run either the algorithm of Theorem 1 or the MaxIS-3 algorithm, whichever is cheaper, and return.
+Otherwise, apply Figure 6 to an arbitrary vertex of degree at least $4$, and run Branch4Simple on the two resulting subproblems.
+Algorithm 1 Function Branch4Simple(G,k)(G,k)
 To clarify, despite the name Branch4Simple, the graph is allowed to have vertices of either higher or lower degrees.
@@
 Simplifying the graph in Line 1 can only reduce $\phi(G)=a\mu+bk$.
-If Section 5 is used, it gives a branch-seq $B$ with drops dominated by $[(1,3),(1,5)]$ or $[(0.5,1),(1.5,4)]$.
+If Figure 6 is used, it gives a branch-seq $B$ with drops dominated by $[(1,3),(1,5)]$ or $[(0.5,1),(1.5,4)]$.
 To satisfy Eq.
@@
 solving the problem when the graph has maximum degree $3$.
-The inequality (5), which governs this case, should be completely tight: there should be a “triple point” with respect to Branch4Simple, the MaxIS-3 algorithm, and the algorithm of Theorem 2.
+The inequality (4), which governs this case, should exhibit a “triple point” with respect to Branch4Simple, the MaxIS-3 algorithm, and the algorithm of Theorem 1.
 That is, for chosen parameters $a,b$, we should have
-$\mu\log(2.3146)=2(k-\mu)\log(1.083506)=a\mu+bk$
+$\mu\log(2.3146)=2(k-\mu)\log(1.083506)=a\mu+bk.$
 This allows us to determine $b$ in terms of $a$.
@@
 This same reasoning will be used for parameters in all our algorithms.
-For example, we can define algorithms for branching on degree 5 and degree 6 vertices.
-Simplify $G$ If $\operatorname{maxdeg}(G)\leq 4$, then run either the MaxIS-4 algorithm or Branch4Simple Otherwise, apply Section 5 to an arbitrary vertex of degree at least $5$.
-Algorithm 2 Function Branch5Simple$(G,k)$
-Simplify $G$ If $\operatorname{maxdeg}(G)\leq 5$, then run either the MaxIS-5 algorithm or Branch5Simple Otherwise, apply Section 5 to an arbitrary vertex of degree at least $6$.
-Algorithm 3 Function Branch6Simple$(G,k)$
-Along similar lines, we immediately obtain the results:
+Likewise, we define algorithms for branching on degree 5 and degree 6 vertices.
+1Simplify $G$
+2If $\operatorname{maxdeg}(G)\leq 4$, then run either the MaxIS-4 algorithm or Branch4Simple
+Otherwise, apply Figure 6 to an arbitrary vertex of degree at least $5$.
+Algorithm 2 Function Branch5Simple(G,k)(G,k)
+1Simplify $G$
+2If $\operatorname{maxdeg}(G)\leq 5$, then run either the MaxIS-5 algorithm or Branch5Simple
+Otherwise, apply Figure 6 to an arbitrary vertex of degree at least $6$.
+Algorithm 3 Function Branch6Simple(G,k)(G,k)
 Proposition 28.
@@
 Branch6Simple has measure $a\mu+bk$ for $a=0.20199,b=0.160637$.
-Combined with the MaxIS-5 algorithm using Lemma 10, Branch5Simple immediately gives an algorithm for degree-5 graphs with runtime $O^{*}(1.2491^{k})$.
 The final algorithm is very similar, but we only keep track of runtime in terms of $k$, not $\mu$ or $n$.
-If $\operatorname{maxdeg}(G)\leq 6$, then use Lemma 10 with algorithms of Branch6Simple and MaxIS-6 Otherwise, split on an arbitrary vertex of degree at least $7$.
-Algorithm 4 Function Branch7Simple$(G,k)$
+1If $\operatorname{maxdeg}(G)\leq 6$, then use Lemma 15 with algorithms of Branch6Simple and MaxIS-6
+Otherwise, split on an arbitrary vertex of degree at least $7$.
+Algorithm 4 Function Branch7Simple(G,k)(G,k)
 Theorem 29.
-Algorithm Branch7Simple$(G,k)$ runs in time $O^{*}(1.2575^{k})$.
+Algorithm Branch7Simple(G,k)(G,k) runs in time $O^{*}(1.2575^{k})$.
 Proof.
-If we split on a vertex of degree at least $7$, then we generate subproblems with vertex covers of size at most $k-1$ and $k-7$.
+If we split on a vertex of degree at least $7$, then we generate subproblems with vertex covers of size at most k1k-1 and k7k-7.
 These have cost $1.2575^{k-1}$ and $1.2575^{k-7}$ by induction.
-Since $11.2575^{-1}+1.2575^{-7}<$ 1, the overall cost is $O^{*}(1.2575^{k})$.
+Since $1.2575^{-1}+1.2575^{-7}<1$, the overall cost is $O^{*}(1.2575^{k})$.
 Otherwise the two algorithms in question have measure $a\mu+bk$ and $cn$ respectively; where $a=0.20199,b=0.160637,c=\log(1.1893)$; by applying Lemma 15, we get a combined algorithm with cost $O^{*}(1.2575^{k})$ as desired.
@@
 In the next sections, we consider more elaborate graph structures and branching rules.
-This improves the runtime to $O^{*}(1.25284^{k})$, along with improved algorithms for lower-degree graphs.
+This improves the runtime to $O^{*}(1.25284^{k})$, along with algorithms for lower-degree graphs.
 The outline is similar to Section 4: we develop a series of branching algorithms targeted to graphs with different degrees, building our way up from degree-4 graphs to degree-7 graphs.
 Here we highlight some of the main changes in our analysis.
-Multiway branching.
+5.1 Multiway branching
 So far, we have only analyzed two-way branching rules.
-The advanced algorithms will branch on multiple vertices at a time.
-Specifically, we say that we split on a sequence of vertices $x_{1},\dots,x_{\ell}$ if we first split on $x_{1}$; in the subproblem $\langle G-x_{1},k-1\rangle$ we split on $x_{2}$; in the subproblem $\langle G-\{x_{1},x_{2}\},k-2\rangle$ we split on $x_{3}$; and so on.
-Assuming that $x_{1}\dots,x_{\ell}$ are independent, this generates $\ell+1$ subproblems $\langle G-N[x_{1}],k-\deg(x_{1})\rangle,\langle G-x_{1}-N[x_{2}],k-1-\deg(x_{2})\rangle,\langle G-\{x_{1},x_{2}\}-N[x_{3}],k-2-\deg(x_{3})\rangle,\dots,\langle G-\{x_{1},\dots,x_{\ell-1}\}-N[x_{\ell}],k-(\ell-1)-\deg(x_{\ell})\rangle,\langle G-\{x_{1},\dots,x_{\ell}\},k-\ell\rangle$.
-We emphasize that the subproblem $G-N[x_{1}]$ does not have any further branching, such as guessing the status of $x_{2}$.
+The advanced algorithms will use multiway branching; for example, we may split on a given vertex $u$, and then apply further branching steps to the subproblems GuG-u and/or GN[u]G-N[u].
+A common tactic, which we refer to as splitting on a sequence of vertices $x_{1},\dots,x_{\ell}$, is to first split on $x_{1}$; in the subproblem $\langle G-x_{1},k-1\rangle$ then split on $x_{2}$; in the subproblem $\langle G-\{x_{1},x_{2}\},k-2\rangle$ then split on $x_{3}$, and so on.
+When $x_{1}\dots,x_{\ell}$ are independent, this generates subproblems $G_{1}=\langle G-N[x_{1}],k-\deg(x_{1})\rangle,G_{2}=\langle G-x_{1}-N[x_{2}],k-1-\deg(x_{2})\rangle,\dots,G_{\ell}=\langle G-\{x_{1},\dots,x_{\ell-1}\}-N[x_{\ell}],k-\ell+1-\deg(x_{\ell})\rangle,G_{0}=\langle G-\{x_{1},\dots,x_{\ell}\},k$-$\ell\rangle$.
+We emphasize that $G_{1}$ does not have further branching, such as guessing the status of $x_{2}$.
 Splitting on $x_{1},x_{2}$ is just a 3-way branch.
-Miscellaneous branching rules. We occasionally use some other branching rules; some of these can be applied to graphs which are not simplified. The following rules are used repeatedly:
-(A1): Branch on 2-vertex.
-For a 2-vertex $v$ with neighbors $x$, $y$, Lemma 5 implies there is either a good cover omitting $v$, or omitting both $x,y$.
-Thus, we can branch on subproblems $\langle G-N[v],k-|N(v)|\rangle$ and $\langle G-N[x,y],k-|N(x,y)|\rangle$.
-(A2): Branch on a funnel.
-For a funnel $u$ with out-neighbor $x$ and any other neighbor $v$ of $u$, we branch on whether $v\in C$ or whether $\{v,x\}\cap C=\emptyset$; this generates subproblems $\langle G-v,k-1\rangle$ and $\langle G-N[v,x],k-|N(v,x)|\rangle$.
-(A3): Branch on shared neighbors.
-For vertices $u,v$ branch on whether $\{u,v\}\subseteq C$, or $N(u)\cap N(v)\subseteq C$; this generates subproblems $\langle G-\{u,v\},k-2\rangle$ and $\langle G-N(u)\cap N(v),k-\operatorname{codeg}(u,v)\rangle$.
-We use a few specialized branching rules in various places; these will be discussed as needed.
-Creation of high-degree vertices and surplus-two indsets.
-One common tactic in our analysis is to branch in a targeted way to allow us the use a particularly profitable “follow-up” branching rule in the resulting graph.
-Two of the most common situations are to use preprocessing rules P2 or P3 in order to create new high-degree vertices, or branching to create a large surplus-two indset for which we can use brancing rules in Section 5.1.
-We say graph $G$ is $r$-vulnerable (or $r$-vul for short) if either (i) $S(G)\geq 1$ or (ii) $G$ has an indset with surplus at most two and size at least $r$.
+A related tactic is to branch in a targeted way to create large surplus-two indsets, or so that preprocessing rules (P2) or (P3) create new high-degree vertices.
+The branching rules in Section 3.1 can then be applied, which are highly profitable.
+We record a few examples of this analysis here.
 Proposition 30.
@@
 (ii) a funnel $u$ with out-neighbor $x$ and a vertex $y\in N(u)\setminus N(x)$.
-In either case, let $r=\deg(x)+\deg(y)-\operatorname{codeg}(x,y)-1$.
-If $r\geq 5$, then $G$ has available branch-seq $[(0,2)]$, $[(1,4),(1,6)]$ or $[(0.5,2),(2,6)]$.
+In either case, define $r=\deg(x)+\deg(y)-\operatorname{codeg}(x,y)-1.$
+If $r\geq 5$ then $G$ has available branch-seq $[(0,2)]$, $[(1,4),(1,6)]$ or $[(0.5,2),(2,6)]$.
 If $r\geq 6$ then $G$ has available branch-seq $[(0,2)]$, $[(1,4),(1,6)]$, $[(0.5,3),(2,6)]$, or $[(0.5,2),(2.5,r+1)]$.
@@
 In case (i), we apply (P2) to merge $x,y$ into a new $r$-vertex $z$.
-In case (ii), we apply (P3); if $\operatorname{codeg}(u,x)\geq 0$ this directly gives drop $[(0,2)]$.
+In case (ii), we apply (P3); if $\operatorname{codeg}(u,x)\geq 0$ this directly gives drop $(0,2)$.
 Otherwise, it produces an $r$-vertex $y$.
 If the resulting graph $G^{\prime}$ can be simplified further (aside from isolated vertices), then we get net drop $(0,2)$.
-Otherwise, we apply Section 5 to $G^{\prime}$.
+Otherwise, we apply Figure 6 to $G^{\prime}$.
 ∎
 Proposition 31.
-Suppose $\operatorname{mindeg}(G)\geq 1$.
+Suppose $\operatorname{mindeg}(G)\geq 1$. We say graph $G$ is $r$-vulnerable (or $r$-vul for short) if either (i) $S(G)\geq 1$ or (ii) $G$ has an indset with surplus at most two and size at least $r$.
 If $G$ is $2$-vulnerable it has available branch-seq $[(0,1)]$, $[(1,4),(1,4)]$, or $[(0.5,3),(2,5)]$.
@@
 Proof.
-If $G$ is not simplified, then $S(G)\geq 1$ giving branch-seq $(0,1)$.
-Otherwise, let $I$ be a surplus-two min-set of size $r$.
-If $r\geq 4$ then splitting on $I$ gives subproblems $G-I,G-N[I]$ with drops $(1,r)$ and $(1,r+2)$ respectively.
-If $r=2$ or $r=3$, we apply Lemma 22 or Lemma 21 respectively.
-∎
-For Propositions 30 and 31, we use the following notation for $\operatorname{val}_{a,b}(B)$ of the given branch-seq $B$.
+If $G$ is not simplified, then $S(G)\geq 1$ giving branch-seq $(0,1)$. Otherwise, we can apply Lemma 22 or Lemma 21 or just apply (B-Crit) depending on the size of the indset. ∎
+Notation for branching rules.
+For the remainder of this paper, when analyzing branching subproblems, we usually omit the solution size $k^{\prime}$.
+It should be determined from context: a subproblem $G^{\prime}=G-X-N[Y]$ is principal with excess $|Y|$ and has $k^{\prime}=k-|X|-|N(Y)|$.
+(For this notation, we are careful to ensure that $Y$ is an indset and that $X\not\sim Y$, to avoid double-counting.)
+We use the following important compositional property: if we generate subproblems $i=1,\dots,t$ each with drop $(\Delta\mu_{i},\Delta k_{i})$, and then apply a branching rule with branch-seq $B^{\prime}_{i}$ to the subproblem $i$, then the resulting branching rule $B$ overall has
+$\operatorname{val}_{a,b}(B)=\sum_{i=1}^{t}e^{-a\Delta\mu_{i}-b\Delta k_{i}}\operatorname{val}_{a,b}(B_{i}^{\prime}).$
+For Propositions 30 and 31, we use the following notation for the $\operatorname{val}(B^{\prime}_{i})$:
 […]
 […]
-Notational convention: omitting the value $k$.
-For the remainder of this paper, when analyzing branching subproblems, we will usually omit the solution size $k^{\prime}$.
-It should be determined from context, from the following mnemonic: for a subproblem $G-\{x_{1},\dots,x_{\ell}\}-N[y]$, we take $k^{\prime}=k-\ell-\deg(y)$.
-In using this notation, we must be careful to ensure that $\{x_{1},\dots,x_{\ell}\}\not\sim y$, as otherwise some $x_{i}$ may be double-counted.
+Thus, for instance, if we have a subproblem with drop $(\Delta\mu,\Delta k)$ and then we apply Proposition 30, its overall contribution to the Branching Inequality, would be given by $e^{-a\Delta\mu-b\Delta k}\psi_{r}.$
+5.2 Branching on funnels and other structures
+If the graph contains a funnel, the simplest tactic is to apply (P3).
+However this creates additional edges, which can make the graph too complicated to analyze further.
+It is often better to branch on the funnel itself.
+There are two possible branching rules for this situation:
+(B-Fun1) For a funnel $u$ with out-neighbor $x$, branch on whether $u\notin C$ or $x\notin C$, generating subproblems GN[u]G-N[u] and GN[x]G-N[x].
+[^7] Recall our notational convention that these subproblems have $k^{\prime}=k-\deg(u)$ and $k^{\prime}=k-\deg(x)$ respectively.
+(B-Fun2) For a funnel $u$ with out-neighbor $x$ and any other neighbor $v$ of $u$, branch on whether $v\in C$ or whether $\{v,x\}\cap C=\emptyset$; this generates subproblems GvG-v and GN[v,x]G-N[v,x].
+Note that, when we apply (B-Fun2), the vertex $u$ still has a funnel in the subproblem GvG-v (of size reduced by one); we can continue to apply (P3) to it, or use other funnel reductions.
+Consider a degree-two vertex $u$ with neighbors $x,y$.
+This can be regarded as a funnel, where $x$ is an out-neighbor and the vertex $\{y\}$ forms a (trivial) clique.
+If we apply (B-Fun2), we get subproblems GyG-y and GN[x,y]G-N[x,y].
+In the former subproblem, $u$ now has degree one; we can apply (P1) to remove $u$ and its neighbor $x$ and get subproblem $G-\{x,y\}=G-N[u]$.
+Since this comes up so frequently, we summarize this branching strategy as follows:
+(B-Pair) For a 2-vertex $u$ with neighbors $x,y$ branch on subproblems $G-\{x,y\}$ and GN[x,y]G-N[x,y].
+We occasionally use some other specialized branching rules; most of these will be discussed in the specific situations where they are needed.
+There is one branching rule worth further discussion, as it comes up in a number of contexts:
+(B-Shared) Suppose vertices $u,v$ share neighbors $x_{1},\dots,x_{\ell}$.
+Branch on whether $\{u,v\}\subseteq C$, or $\{x_{1},\dots,x_{\ell}\}\subseteq C$; this generates subproblems $G-\{u,v\}$ and $G-\{x_{1},\dots,x_{\ell}\}$.
+Note that, in applying (B-Shared), it is allowed to have $\operatorname{codeg}(u,v)>\ell$ strictly.
+5.3 Tracking degree-3 vertices
+In the algorithms we have encountered so far, we only use two parameters of the graph: the values $k$ and $\mu$.
+This will also be the case for most of our advanced algorithms.
+For the degree-4 algorithm, however, we will need to track a few additional pieces of information.
+The most important is whether the graph has any degree-3 vertices (the precise number of them is not tracked).
+These would lead to more profitable branching rules; as a rule of thumb, having a degree-3 vertex is almost as good as simplifying the graph.
+For the degree-4 algorithm, we extend our notations to track 3-vertices.
+We write $S(G)\geq\ell^{*}$ if either (i) $S(G)=\ell$ and the simplified graph $\langle G^{\prime},k$-$\ell\rangle$ has a 3-vertex, or (ii) $S(G)\geq\ell+1$.
+We say that a subproblem $\langle G^{\prime},k^{\prime}\rangle$ has drop $(\Delta\mu,\Delta k^{*})$ if $\mu^{\prime}\leq\mu-\Delta\mu$ and either (i) $G^{\prime}$ has a subcubic vertex and $k^{\prime}=k-\Delta k$ or (ii) $k^{\prime}<k-\Delta k$ strictly.
+Here we record a few observations about simplification and degree-3 vertices.
+Proposition 32.
+Let $s\geq 1$. If $G$ has at least $2(s+1)$ subcubic vertices, and it has a critical-set $I$ with $\operatorname{surp}_{G}(I)=1,|I|\geq s$, then $S(G)\geq s^{*}$.
+Proof.
+Apply (P2) to $I$. If $|I|>s$, this gives $S(G)\geq s+1$ directly. Otherwise, all subcubic vertices outside $N[I]$ remain subcubic, leaving at least 2(s+1)|N[I]|=12(s+1)-|N[I]|=1 subcubic vertices. ∎
+Proposition 33.
+If $G$ has no 4-cycles, has at least 7 subcubic vertices, and has two non-adjacent 2-vertices, then $S(G)\geq 2^{*}$.
+Proof.
+Let $u,v$ be non-adjacent 2-vertices.
+Apply (P2) to $v$; then all the subcubic vertices in $G$ outside from $N[v]$ remain subcubic in the resulting graph $G^{\prime}$.
+So $G^{\prime}$ has at least 4 subcubic vertices.
+Also, since $G$ has no 4-cycle, we have $\deg_{G^{\prime}}(u)=2$.
+By Proposition 32, this implies $S(G^{\prime})\geq 1^{*}$ and hence $S(G)\geq 2^{*}$.
+∎
+Lemma 34.
+Suppose that $\operatorname{maxdeg}(G)\leq 4$, and let $G^{\prime}=\langle G-(X\cup Y),k-|X|\rangle$ be a principal subproblem.
+If all components of $G$ have size at least $5|X|+1$, then $S(G^{\prime})$ has drop $(\Delta\mu,|X|^{*})$.
+Proof.
+There must be some vertex $z$ at distance two from $X$, since otherwise the component(s) containing $X$ would have size at most $5|X|$.
+So suppose $z$ has a neighbor $y\sim X$.
+Here $\deg_{G-X}(y)\geq 1$ (due to neighbor $z$) and also $\deg_{G-X}(y)\leq 3$ (it loses some neighbor in $X$).
+So GXG-X has a subcubic vertex.
+Removing isolated vertices does not affect this.
+∎
+Proposition 35.
+Let $r\geq 2$ be an arbitrary constant.
+Suppose that $G$ has at least 4 subcubic vertices and $\operatorname{maxdeg}(G)\leq 4$ and $\operatorname{minsurp}(G)\geq 1$ and $G$ has an indset $I$ with $|I|\geq r$ and $\operatorname{surp}_{G}(I)\leq 2$.
+Then $G$ has available branch-seq $(0,1^{*})]$ or $(1,r^{*}),(1,(r+2)^{*})]$.
+Proof.
+If $G$ has any components of size at most $5(r+2)$, we solve them exactly in constant time and remove them from $G$; since $\operatorname{minsurp}(G)\geq 1$, this reduces $k$ by at least two and hence gives drop $[(0,2)]$.
+If $\operatorname{surp}_{G}(I)=1$, then Proposition 32 with $s=1$ gives $S(G)\geq 1^{*}$.
+Otherwise, we apply (B-Crit); by Lemma 34 the subproblems have drops $(1,r^{*})$ and $(1,(r+2)^{*})$.
+∎
+Along the lines of Proposition 31, we say that $G$ is $r^{*}$-vul if either $S(G)\geq 1^{*}$ or the preconditions of Proposition 35 hold.
 6 Branching on 4-vertices
@@
 At a high level, the improvement over the baseline algorithm Branch4Simple comes from tracking the presence of 3-vertices.
-If a 4-vertex $u$ has a 3-neighbor $v$, then we can split on $u$ and then, in the branch $G-u$, the degree of $v$ gets reduced to two.
+If a 4-vertex $u$ has a 3-neighbor $v$, then we can split on $u$ and then, in the branch GuG-u, the degree of $v$ gets reduced to two.
 We then use (P2) to reduce $k$ by one – a “bonus” reduction.
@@
 We sketch an outline of the algorithm here; many additional details will be provided later.
-Apply branching rules to ensure $G$ has nice structure, e.g.
-simplified and no 4-cycles.
-(We will call this 4-base-position (4BP) later.) If $V_{3}\neq\emptyset$ and $V_{3}\not\sim V_{4}$, then solve vertex cover exactly on $V_{3}$.
-Split on a 4-vertex with a 3-neighbor, chosen according to tie-breaking rules to be described later.
-Branch on a triangle.
+1
+2Apply misc. branching rules to ensure $G$ has nice structure, e.g. simplified and no 4-cycles.
+3If $V_{3}\neq\emptyset$ and $V_{3}\not\sim V_{4}$, then solve vertex cover exactly on $V_{3}$.
+4Split on a 4-vertex with a 3-neighbor, chosen according to tie-breaking rules to be described later.
+5Branch on a triangle.
 Branch on a vertex $u$ according to tie-breaking rule to be described later.
-Algorithm 5 Function Branch4$(G,k)$
+Algorithm 5 Function Branch4(G,k)(G,k)
 Theorem 36.
@@
 Via Lemma 15, in combination with the MaxIS-4 algorithm, this will give our algorithm with runtime $O^{*}(1.21131^{k})$ for degree-4 graphs.
-Theorem 32 is very length to prove.
-One major complication is that it requires us to use a measure for $G$ tracking additional information beyond just the values $\mu$ and $k$.
-Specifically, our strategy is to show that Branch4 has the measure of the form
-$\phi(G)=a\mu+bk-\operatorname{bonus}(G);$
-where we define function $\operatorname{bonus}(G)$ as follows.
-First, we define a special 4-vertex to be a 4-vertex with at least two 3-neighbors, and we say $G$ is special if it has a special 4-vertex.
-We then define:
-$\operatorname{bonus}(G)$ =$=\begin{cases}\beta&\text{if $\operatorname{mindeg}(G)=3$ and $\operatorname{maxdeg}(G)=4$ and $G$ is special}\\ \alpha&\text{if $G$ has a subcubic vertex and the previous case does not hold}\\ 0&\text{otherwise}\end{cases}$
+Theorem 36 is very lengthy to prove.
+Our strategy is to show that Branch4 has the measure of the form
+$\phi(G)=a\mu+bk-\operatorname{bonus}(G).$
+Let us define a special 4-vertex to be a 4-vertex with at least two 3-neighbors, and we say $G$ is special if it has a special 4-vertex. We then define:
+[…]
 for parameters $\alpha=0.03894,\beta=0.05478$
@@
 $0\leq\alpha\leq b\leq\beta\leq 2b$ (5)
-Motivated by Eq.
-(6), we write $S(G)\geq\ell^{*}$ if either (i) $S(G)\geq\ell+1$, or (ii) $S(G)=\ell$ and there is a simplified graph $\langle G^{\prime},k-\ell\rangle$ with a 3-vertex.
-We also write $R(X)\geq\ell^{*}$ if $S(G-X)\geq\ell^{*}$, etc.
-Similarly, if we transform a graph $\langle G,k\rangle$ into a graph $\langle G^{\prime},k^{\prime}\rangle$, we say $G^{\prime}$ has “extended drop” $(\Delta\mu,\Delta k^{*})$ if $\mu^{\prime}\leq\mu-\Delta\mu$ and either (i) $G^{\prime}$ has a subcubic vertex and $k^{\prime}=k-\Delta k$ or (ii) $k^{\prime}<k-\Delta k$ strictly.
-To explain, note that since $\alpha\leq b$, we have $\phi(G^{\prime})\leq\phi(G)-a\Delta\mu-b\Delta k-\alpha+\operatorname{bonus}(G)$.
-So the contribution of $G^{\prime}$ to the sum in Eq.
-(3) is at most $e^{-a\Delta\mu-b\Delta k-\alpha}$.
-8.1 Some observations about simplification and degree-3 vertices
-We record a few general observations about simplification in graphs.
-Proposition 33.
-Let $s\geq 1$. If $G$ has at least $2(s+1)$ subcubic vertices, and it has a critical-set $I$ with $\operatorname{surp}(I)=1,|I|\geq s$, then $S(G)\geq s^{*}$.
-Proof.
-We apply (P2) to $I$. If $|I|>s$, this gives $S(G)\geq s+1$ directly. Otherwise, all subcubic vertices outside $N[I]$ remain subcubic, leaving at least $12(s+1)-|N[I]|=$ 1 subcubic vertices.
-Proposition 34.
-Suppose $G$ has no 4-cycles, has at least 7 subcubic vertices, and has two non-adjacent 2-vertices. Then $S(G)\geq 2^{*}$.
-Proof.
-Let $u,v$ be non-adjacent 2-vertices.
-If we apply (P2) to $v$, then all the subcubic vertices in $G$ outside from $N[v]$ remain subcubic in the resulting graph $G^{\prime}$.
-Since $G$ has no 4-cycle, we have $\deg_{G^{\prime}}(u)=2$.
-So $G^{\prime}$ has at least 4 subcubic vertices and has a 2-vertex $u$.
-By Proposition 33, this implies $S(G^{\prime})\geq 1^{*}$ and hence $S(G)\geq 2^{*}$.
-∎
-Lemma 35.
-Let $X$ be a vertex set. If $\operatorname{maxdeg}(G)\leq 4$ and $G$ has no components of size at most $8|X|$, then $G-X$ has a subcubic vertex.
-Proof.
-There must be some vertex $z$ at distance two from $X$, since otherwise the component(s) containing $X$ would have size at most $(4+\operatorname{maxdeg}(G))|X|\leq 8|X|$.
-So suppose $z$ has a neighbor $y\sim X$.
-Here $\deg_{G-X}(y)\geq 1$ (due to neighbor $z$) and also $\deg_{G-X}(y)\leq 3$ (it loses some neighbor in $X$).
-Hence, $G-X$ has a subcubic vertex.
-∎
-Proposition 36.
-Let $r\geq 2$ be an arbitrary constant.
-Suppose that $G$ has at least 4 subcubic vertices and $\operatorname{maxdeg}(G)\leq 4$ and $\operatorname{minsurp}(G)\geq 1$ and $G$ has an indset $I$ with $|I|\geq r$ and $\operatorname{surp}_{G}(I)\leq 2$.
-Then $G$ has available branch-seq $[(0,1^{*})]$ or $[(1,r^{*}),(1,(r+2)^{*})]$.
-Proof.
-If $G$ has any components of size less than $8(r+2)$, we solve them exactly in $O(1)$ time and remove them from $G$; since $\operatorname{minsurp}(G)\geq 1$, this reduces $k$ by at least two and hence gives drop $[(0,2)]$.
-If $\operatorname{surp}_{G}(I)=1$, then Proposition 33 with $s=1$ gives $S(G)\geq 1^{*}$.
-Otherwise, we split on min-set $I$; if $|I|>r$ strictly, the subproblem $G-I$ and $G-N[I]$ have drop $(1,(r+1))$ and $(1,(r+3))$ directly.
-If $|I|=r$, these subproblems have drops $(1,r)$ and $(1,r+2)$ and, by Lemma 35, both of them have subcubic vertices.
-We say that $G$ is $r^{*}$-vul if either $S(G)\geq 1^{*}$ or the preconditions of Proposition 36 hold, and we correspondingly define
-$\gamma^{*}_{r}=\max\{e^{-b-\alpha},e^{-a-rb-\alpha}+e^{-a-(r+2)b-\alpha}\}$
+To show that a branching rule generating subproblems $G_{1},\dots,G_{t}$ respects the measure $\phi$, we using a slight variant of the branching inequality:
+$\sum_{i=1}^{t}e^{\phi(G_{i})-a\mu-bk}\leq e^{-\operatorname{bonus}(G)}$ (6)
+As a consequence of Eq. (5) and Proposition 35, we have the following observation (which motivates our notation for tracking 3-vertices as we have described earlier):
+Observation 37.
+If $S(G)\geq\ell^{*}$, then the simplified graph $G^{\prime}$ has $\phi(G^{\prime})\leq\phi(G)-b\Delta k-\alpha$.
+If subproblem has drop $(\Delta\mu,\Delta k^{*})$, then its contribution to Eq. (6) is at most $e^{-a\Delta\mu-b\Delta k-\alpha}$.
+If subproblem has drop $(\Delta\mu,\Delta k)$ and is $r^{*}$-vul, its contribution to Eq.
+(6) is at most $e^{-a\Delta\mu-b\Delta k}\gamma^{*}_{r}$ where we define $\gamma^{*}_{r}=\max\{e^{-b-\alpha},e^{-a-rb-\alpha}+e^{-a-(r+2)b-\alpha}\}$.
 6.1 Regularizing the graph and basic analysis
-We will first ensure that $G$ has nice properties, before the “heavy-duty” branching rules. Specifically, we say $G$ is in 4-base position (or 4BP for short) if the following conditions hold:
+We will first ensure that $G$ has a number of nice properties:
   • $G$ is simplified.
@@
   • $G$ has no 4-cycle
-  • $G$ has no connected-components of size larger than some constant $K$.
-Note: Let us explain the last clause in the definition of 4BP, and the choice of constant $K$ (which we will not specify explicitly).
-In light of Lemma 35, our branching algorithms will generate principal subproblems $G-(U_{0}\cup U_{1})$ where $\Delta k=|U_{1}|$ is lower-bounded by some constant value $c$.
-Such subproblems may be produced directly (e.g.
-by guessing a vertex is in the cover), or may be obtained by applying (P1) to a principal subproblem $G-X$.
-If the constant $K$ is large enough (in particular, larger than $8c$), then any such principal subproblem $G-(U_{0}\cup U_{1})$ with drop $(\Delta\mu,c)$ automatically has drop $(\Delta\mu,c^{*})$ “for free”.
-For example, this principle was used in Proposition 36; we will use it implicitly through the remainder of the algorithm analysis.
+  • $G$ has no connected-components of size smaller than some constant $K$.
+Let us explain the last clause here.
+Our branching algorithms will generate principal subproblems GXG-X with some given drop $(\Delta\mu,\Delta k)$ a constant $\Delta k$.
+If $K>5\Delta k+1$, then by Lemma 34, such subproblem automatically has drop $(\Delta\mu,\Delta k^{*})$ “for free”.
+We choose $K$ to be larger than all such constants needed for the analysis; we do not specify it explicitly.
 Proposition 38.
-If $\operatorname{maxdeg}(G)\leq 4$ and $\operatorname{minsurp}(G)\geq 2$, but $G$ is not in 4BP, then it has available branch-seq $[(0,2)]$, $[(1,3),(1,5)]$, $[(0.5,2),(2,5)]$, or $[(1,3^{*}),(1,3^{*})]$
+If $\operatorname{maxdeg}(G)\leq 4$ and $\operatorname{minsurp}(G)\geq 2$, but $G$ does not satisfy the Line-1 conditions, then it has available branch-seq $[(0,2)]$, $[(1,3),(1,5)]$, $[(0.5,2),(2,5)]$, or $(1,3^{*}),(1,3^{*})]$
 Proof.
-If $G$ has a connected-component $X$ of constant size, we may solve vertex cover on it exactly in constant time, and update $k^{\prime}=k-|C_{X}|$ where $C_{X}$ is a good cover of $X$.
-Since $\operatorname{minsurp}(G)\geq 2$, we have $|C_{X}|\geq 2$, and so removing this component gives drop $[(0,2)]$.
-If $G$ has a funnel $u$ with out-neighbor $x$, then we branch on subproblems $G-N[u],G-N[x]$; these both have drop $(1,3^{*})$ by Lemma 35.
-Since $\operatorname{minsurp}(G)\geq 2$, rules (P1) and (P2) cannot be applied; so for the remainder of the analysis we may suppose that $G$ is simplified.
+If $G$ has a connected-component $X$ of size $O(1)$, we solve it exactly and update $k^{\prime}=k-|C_{X}|$ where $C_{X}$ is a good cover of $X$.
+Since $\operatorname{minsurp}(G)\geq 2$, we have $|C_{X}|\geq 2$ and this gives drop $[(0,2)]$.
+If $G$ has a funnel $u$ with out-neighbor $x$, then we apply (B-Fun1), generating subproblems GN[u],GN[x]G-N[u],G-N[x].
+Since $\operatorname{minsurp}(G)\geq 2$, these subproblems have drop $(1,3)$ directly; by Lemma 34 they indeed both have drop $(1,3^{*})$.
+Otherwise, rules (P1) and (P2) cannot be applied; so for the remainder of the analysis we may suppose that $G$ is simplified.
 If $G$ has a blocked 4-vertex, then we apply Proposition 25.
-Finally, if $G$ has a 4-cycle $u_{1},u_{2},u_{3},u_{4}$, we apply (A3) to get subproblems $G-\{u_{1},u_{3}\}$ and $G-N(u_{1},u_{3})$.
-We claim these subproblems both have drops $(1,3^{*})$ after simplification.
-Let us first show this holds for subproblem $G^{\prime}=G-\{u_{1},u_{3}\}$, which has drop $(1,2)$ directly.
-If $u_{2}\not\sim u_{4}$, then Observation 16 immediately gives $S(G^{\prime})\geq 2$.
-So suppose $u_{2}\sim u_{4}$; since $G$ is simplified, this is only possible if $u_{1},u_{2},u_{3},u_{4}$ all have degree 4 (else they would have a 3-triangle).
-Then $u_{1}\not\sim u_{3}$ as otherwise $u_{1}$ would have a funnel with its neighbors $u_{2},u_{3},u_{4}$.
-So $G^{\prime}$ has at least 4 subcubic vertices (the neighbors of $u_{1}$) and it has a 2-vertex $u_{2}$, so Proposition 33 with $s=1$ gives $S(G^{\prime})\geq 1^{*}$.
-We next argue that $G^{\prime\prime}=G-N(u_{1},u_{3})$ has drop $(1,3^{*})$.
-If $\operatorname{codeg}(u_{1},u_{3})\geq 3$, then $G^{\prime\prime}$ has drop $(1,3^{*})$ directly by Lemma 35.
-Otherwise, if $\operatorname{codeg}(u_{1},u_{3})=2$, then $G^{\prime\prime}=G-\{u_{2},u_{4}\}$; this has drop $(1,3^{*})$ by precisely the same reasoning as for the graph $G^{\prime}=G-\{u_{1},u_{3}\}$.
+Finally, if $G$ has a 4-cycle $u_{1},u_{2},u_{3},u_{4}$, we consider two cases.
+First, suppose there is a “cross-edge”, say $u_{1}\sim u_{3}$.
+We must have $\deg(u_{1})=\deg(u_{3})=4$, else they would have a funnel.
+Let $v_{1},v_{3}$ be the other neighbors of $u_{1},u_{3}$.
+We split on $v_{1}$; subproblem $G-N[v_{1}]$ has drop $(1,3^{*})$ by Lemma 34.
+Subproblem $G-v_{1}$ has drop $(0.5,1)$ directly, and has a kite involving vertex $u_{1}$; by Lemma 14, after applying (P3), we get net drop $(1,3)$ and the resulting graph has the 3-vertex $u_{3}$.
+Otherwise, suppose that $u_{1}\not\sim u_{3}$ and $u_{2}\not\sim u_{4}$.
+We apply (B-Shared) to get subproblems $G-\{u_{1},u_{3}\}$ and $G-\{u_{2},u_{4}\}$.
+These both have two non-adjacent subquadratic nodes ($u_{2},u_{4}$ and $u_{1},u_{3}$ respectively), so by Observation 12 they both have drop $(1,4)$.
 ∎
 Proposition 39.
-If $G$ is in 4BP and $u$ is a 4-vertex, then $G-N[u]$ has at least 8 subcubic vertices.
+After Line 1, for a 4-vertex $u$, the graph GN[u]G-N[u] has at least 8 subcubic vertices.
 Proof.
-Suppose $u$ has 3-neighbors $v_{1},\dots,v_{s}$ and 4-neighbors $v_{s+1},\dots,v_{s}$ We have $\operatorname{codeg}(u,v_{i})=0$ for $i\leq s$, since otherwise $v_{i}$ would have a 3-triangle.
+Suppose $u$ has 3-neighbors $v_{1},\dots,v_{s}$ and 4-neighbors $v_{s+1},\dots,v_{s}$ .
+We have $\operatorname{codeg}(u,v_{i})=0$ for $i\leq s$, since otherwise $v_{i}$ would have a 3-triangle.
 Likewise, we have $\operatorname{codeg}(u,v_{i})\leq 1$ for $i>s$, as otherwise $G$ would have a 4-cycle.
@@
 ∎
-We now analyze the easier steps of the algorithm line-by-line. Note that, starting at Line 4, the graph $G$ is 4-regular. Lines 3 and 5, which are more involved, will be considered later.
+We now analyze the easier steps of the algorithm line-by-line. Observe that, starting at Line 4, the graph $G$ is 4-regular. Lines 3 and 5, which are more involved, will be considered later.
 Line 1. We may remove any isolated vertices, as this can only increase $\operatorname{bonus}(G)$.
@@
 If $G$ is not simplified, then rules (P1)–(P3) give drop $[(0,1)]$.
-If $\operatorname{maxdeg}(G)\geq 5$, then we apply Section 5; otherwise, we apply Proposition 37.
-We need:
+If $\operatorname{maxdeg}(G)\geq 5$, then we apply Figure 6; otherwise, we apply Proposition 38.
+In order to satisfy Eq.
+(6), we need:
 $\max\{e^{-b},e^{-0.5a-b}+e^{-2a-5b},e^{-a-3b-\alpha}+e^{-a-3b-\alpha},e^{-a-3b}+e^{-a-5b}\}$ $\leq 1$ (4-1)
@@
 If $\operatorname{shad}(N[u])\leq 4-\deg(u)$, then we apply Lemma 26.
-If $\operatorname{shad}(N[u])\geq 5-\deg(u)$, then we split on $u$; subproblem $G-N[u]$ has drop $(2,5)$, while subproblem $G-u$ retains the subcubic vertex in the original graph and so has drop $(0.5,1^{*})$.
+If $\operatorname{shad}(N[u])\geq 5-\deg(u)$, then we split on $u$; subproblem GN[u]G-N[u] has drop $(2,5)$, while subproblem GuG-u retains the subcubic vertex in the original graph and so has drop $(0.5,1^{*})$.
 Over all cases, we need:
@@
 Let $N(u)=\{v_{1},\dots,v_{4}\}$ where $v_{1}\sim v_{2}$.
-Since $G$ has no 4-cycles, $v_{1}\not\sim\{v_{3},v_{4}\}$.
+Since $G$ has no 4-cycles, $\{v_{1},v_{2}\}\not\sim\{v_{3},v_{4}\}$.
 We split on $v_{3}$, getting subproblems $G-N[v_{3}],G-v_{3}$ with drops $(1.5,4^{*})$ and $(0.5,1^{*})$.
@@
 6.2 Line 3: Branching on a 4-vertex with a 3-neighbor
-Suppose Line 3 selects a 4-vertex $u$, with 3-neighbors $v_{1},\dots,v_{s}$ and $4$-neighbors $v_{s+1},\dots,v_{4}$.
-Let $x_{1},x_{2}$ be the other two neighbors of $v_{1}$, sorted so $\deg(x_{1})\geq\deg(x_{2})$.
-Since $G$ has no 4-cycles, and the 3-vertices $v_{1},\dots,v_{s}$ have no triangles, all vertices $u,v_{i},x_{i}$ are distinct, and $\{x_{1},x_{2}\}\not\sim\{u,v_{2},v_{3},v_{4}\}$.
-Note that there may be multiple ways of ordering vertices in this way; e.g., we can permute the vertices $v_{1},\dots,v_{s}$ or $v_{s+1},\dots,v_{4}$, or if if $\deg(x_{1})=\deg(x_{2})$ we can swap $x_{1}/x_{2}$.
-Let us fix some ordering and consider the graphs $G_{0}=G-N[u]$ and $G_{1}=G-u$ produced by splitting on $u$.
-There are two strategies we can use for further branching.
-First, we can apply (P2) to the graph $G_{1}$, contracting its 2-vertices $v_{1},\dots,v_{s}$ into vertices $z_{1},\dots,z_{s}$; we let $G_{1}^{\prime}$ denote the resulting graph, and in this case we overall generate two subproblems $G_{0},G_{1}^{\prime}$.
-Second, we can apply (A1) to the 2-vertex $v_{1}$ in $G-u$, generating further subproblems $G_{2}=G-N[v_{1}],G_{3}=G-v_{1}-N[x_{1},x_{2}]$; so in this case we overall generate three subproblems $G_{0},G_{2},G_{3}$.
-Throughout, we write $r=\deg(x_{1})+\deg(x_{2})-2=\deg_{G_{1}^{\prime}}(z_{1})$.
+Suppose Line 3 selects a 4-vertex $u$, with 3 neighbors $v_{1},\dots,v_{s}$ and 4-neighbors $v_{s+1},\dots,v_{4}$.
+We let $v=v_{1}$ be a chosen 3-vertex, and let $x,y$ be the other two neighbors of $v_{1}$, sorted so $\deg(x)\geq\deg(y)$.
+Since $G$ has no 4-cycles, and the 3-vertex $v$ has no triangles, all vertices $u,v_{i},x_{i},y_{i}$ are distinct, and $\{x,y\}\not\sim\{u,v_{2},v_{3},v_{4}\}$.
+See Figure 8.
+Figure 8: The vertex $u$ and its associated neighbor vertices. The degrees of the nodes are shown.
+Note that there may be multiple ways of ordering vertices in this way; e.g., we may permute $v_{1},\dots,v_{s}$, and if $\deg(x)=\deg(y)$ we can swap $x/y$.
+Let us fix some ordering.
+As a starting point, our strategy is to split on $u$, producing subproblems $G_{0}=G-N[u]$ and GuG-u.
+There are two main further branching strategies we can use on GuG-u:
+  • Apply (P2), contracting its 2-vertices $v_{1},\dots,v_{s}$ into vertices $z_{1},\dots,z_{s}$; we denote the resulting graph by $G_{1}$. Overall we generate two subproblems $G_{0},G_{1}$.
+  • Apply (B-Pair) to the 2-vertex $v$, generating further subproblems $G_{2}=G-N[v],G_{3}=G-u-N[x,y]$. Overall we generate three subproblems $G_{0},G_{2},G_{3}$.
+Throughout Section 6.2 we write
+$r$ $=\deg_{G_{1}}(z_{1})=\deg(x)+\deg(y)-2$
+$s$ =# $=\text{\# of 3-neighbors of }u$
 Proposition 40.
@@
 For the first result, if $s\leq 3$, then $G-N[v_{1}]$ has a 3-vertex $v_{4}$ and a 2-vertex $v_{2}\not\sim v_{4}$; applying (P2) to $v_{2}$ reduces $k$ by one, and still leaves a subcubic vertex $v_{4}$.
-If $s=4$, then $G-N[v_{1}]$ has non-adjacent 2-vertices $v_{2},v_{3}$, so $S(G_{2})\geq 2$ by Observation 11.
+If $s=4$, then $G-N[v_{1}]$ has non-adjacent 2-vertices $v_{2},v_{3}$, so $S(G_{2})\geq 2$ by Proposition 10.
 For the next two results, recall from Proposition 39 that $G_{0}$ has at least 8 subcubic vertices.
-If $r=4$, then $\deg_{G_{0}}(x_{1})=\deg_{G_{0}}(x_{2})=2$ and $x_{1}\not\sim x_{2}$, so by Proposition 34 we have $S(G_{0})\geq 2^{*}$.
-Similarly, if $r=5$, then $\deg_{G_{0}}(x_{2})=2$ so by Proposition 33 we have $S(G_{0})\geq 1^{*}$.
+If $r=4$, then $\deg_{G_{0}}(x)=\deg_{G_{0}}(y)=2$ and $x\not\sim y$, so by Proposition 33 we have $S(G_{0})\geq 2^{*}$.
+Similarly, if $r=5$, then $\deg_{G_{0}}(y)=2$ so by Proposition 32 we have $S(G_{0})\geq 1^{*}$.
 ∎
 Proposition 41.
-Suppose $r\geq 5$. Then $G_{3}$ is a principal subproblem with excess two and $\Delta k=r+2$ and $\operatorname{minsurp}(G_{3})\geq 4-r$. Furthermore, the following bounds all hold:
-  • If $r=5$ and $\operatorname{minsurp}(G_{3})=-1$, then $R(N[x_{1}])\geq 2^{*}$.
-  • If $r=5$ and $\operatorname{minsurp}(G_{3})=-1$ and $s=1$, then $R(N[x_{1}])\geq 3^{*}$.
-  • If $r=6$ and $\operatorname{minsurp}(G_{3})=-2$, then $R(N[x_{1}])\geq 3^{*}$.
-  • If $r=6$ and $\operatorname{minsurp}(G_{3})=-1$, then $G-N[x_{1}]$ is $3^{*}$-vul.
+If $r\geq 5$, then $G_{3}$ is a principal subproblem with excess two and $\Delta k=r+2$ and $\operatorname{minsurp}(G_{3})\geq 4-r$. Furthermore, the following bounds all hold:
+  • If $r=5$ and $\operatorname{minsurp}(G_{3})=-1$, then $S(N[x])\geq 2^{*}$.
+  • If $r=5$ and $\operatorname{minsurp}(G_{3})=-1$ and $s=1$, then $S(N[x])\geq 3^{*}$.
+  • If $r=6$ and $\operatorname{minsurp}(G_{3})=-2$, then $S(N[x])\geq 3^{*}$.
+  • If $r=6$ and $\operatorname{minsurp}(G_{3})=-1$, then GN[x]G-N[x] is $3^{*}$-vul.
 Proof.
-Let $H=G-u-N[x_{1},x_{2}]$.
-Since $G$ has no 4-cycles, we have $|N[x_{1},x_{2}]|=r+1$ and so $H$ is principal with excess two and $\Delta k=r+2$.
-Since $\deg(x_{1})\geq\deg(x_{2})$, we have $\deg(x_{1})=4$ and $\deg(x_{2})=r-2$.
-Note that if $s=1$, then the sole 3-neighbor of $u$ gets removed from $H$, and so $\operatorname{mindeg}(H)\geq 1$.
-Also note that, by Proposition 38, the graph $G-N[x_{1}]$ has at least 8 subcubic vertices, and since $G$ is in 4BP we have $\operatorname{shad}(N[x_{1}])\geq 1$.
-Let $I$ be a min-set in $H$ and $\ell=\operatorname{surp}_{H}(I)$.
+Here $\deg(x)=4$ and $\deg(y)=r-2$.
+Since $G$ has no 4-cycles, we have $|N[x,y]|=r+1$ and so $G_{3}$ is principal with excess two and $\Delta k=r+2$.
+If $s=1$, then the sole 3-neighbor of $u$ gets removed from $G_{3}$, and so $\operatorname{mindeg}(G_{3})\geq 1$.
+Also note that, by Proposition 39, the graph GN[x]G-N[x] has at least 8 subcubic vertices and $\operatorname{shad}(N[x])\geq 1$.
+Let $I$ be a min-set in $G_{3}$ and $\ell=\operatorname{surp}_{G_{3}}(I)$.
 If $\ell\leq-2$ or $s=1$ then necessarily $|I|\geq 2$.
-Then $\operatorname{surp}_{G-N[x_{1}]}(I\cup\{x_{2}\})\leq\ell+(\deg_{G}(x_{2})-1)=\ell+r-3$.
-Since $G$ is in 4BP, we have $\operatorname{shad}(N[x_{1}])\geq 1$ so $\ell\geq 4-r$.
-Furthermore, if $\ell=4-r$ exactly, then $\operatorname{surp}_{G-N[x_{1}]}(I\cup\{x_{2}\})=1$, and then Proposition 33 applied to indset $I\cup\{x_{2}\}$ gives $R(N[x_{1}])\geq 2^{*}$ for $|I|=1$ and $R(N[x_{1}])\geq 3^{*}$ for $|I|\geq 2$.
+Then $\operatorname{surp}_{G-N[x]}(I\cup\{y\})\leq\ell+(\deg_{G}(y)-1)=\ell+r-3$.
+We have $\operatorname{shad}(N[x])\geq 1$ so $\ell\geq 4-r$.
+Furthermore, if $\ell=4-r$ exactly, then $\operatorname{surp}_{G-N[x]}(I\cup\{y\})=1$, and then Proposition 32 applied to indset $I\cup\{y\}$ gives $S(N[x])\geq 2^{*}$ for $|I|=1$ and $S(N[x])\geq 3^{*}$ for $|I|\geq 2$.
 Finally, suppose $r=6$ and $\ell=-1$.
-If $|I|=1$, then $H$ has an isolated vertex $t$; since $G$ has no 4-cycles, this is only possible if $\deg_{G}(t)=3$ and $t\sim u$ and $\operatorname{codeg}(t,x_{1})=\operatorname{codeg}(t,x_{2})=1$.
-In this case, $\deg_{G-N[x_{1}]}(t)=2$ and by Proposition 33 we have $R(N[x_{1}])\geq 1^{*}$.
-Otherwise, we have $\operatorname{surp}_{G-N[x_{1}]}(I\cup\{x_{2}\})\leq 2$, which by Proposition 36 implies that $G-N[x_{1}]$ is $3^{*}$-vul.
+If $|I|\geq 2$, then $\operatorname{surp}_{G-N[x]}(I\cup\{y\})\leq 2$, which by Proposition 35 implies that GN[x]G-N[x] is $3^{*}$-vul.
+Otherwise, suppose $|I|=1$, i.e.
+$G_{3}$ has an isolated vertex $t$; since $G$ has no 4-cycles, this is only possible if $\deg_{G}(t)=3$ and $t\sim u$ and $\operatorname{codeg}(t,x)=\operatorname{codeg}(t,y)=1$.
+Then $\deg_{G-N[x]}(t)=2$ and by Proposition 32 we have $S(N[x])\geq 1^{*}$.
 ∎
@@
 Case I: $\boldsymbol{s\geq 2}$ and $\boldsymbol{S(G_{0})\geq 2^{*}}$.
-We generate subproblems $G_{0},G_{1}^{\prime}$.
-If $s\in\{2,3\}$, then $G_{1}^{\prime}$ has drop $(0.5,3)$ or $(0.5,4)$ and also has a 3-vertex $v_{4}$; if $s=4$, then $G_{1}^{\prime}$ has drop $(0.5,5)$ directly.
-In any case, $G_{1}^{\prime}$ has drop $(0.5,3^{*})$.
+We generate subproblems $G_{0},G_{1}$.
+If $s\in\{2,3\}$, then $G_{1}$ has drop $(0.5,3)$ and has a 3-vertex $v_{4}$; if $s=4$, then $G_{1}$ has drop $(0.5,5)$ directly.
+In any case, $G_{1}$ has drop $(0.5,3^{*})$.
 We then simply need:
@@
 By Proposition 40, we have $S(G_{2})\geq 1^{*}$; also, we must have $r\geq 5$, since if $r=4$ then $S(G_{0})\geq 2^{*}$ and it would covered in Case I.
-In particular, $\deg(x_{1})=4$.
-We claim that $G-N[x_{1}]$ is not $3^{*}$-vul.
-For, suppose it were, and consider $u^{\prime}=x_{1}$ and its 3-neighbor $v_{1}^{\prime}=v_{1}$ and its two neighbors $x_{1}^{\prime}=u,x_{2}^{\prime}=x_{2}$.
-We have $r^{\prime}=\deg(x_{1}^{\prime})+\deg(x_{2}^{\prime})-2=4+\deg(x_{2})=r$.
-If $\operatorname{shad}(u^{\prime},N[x_{1}^{\prime},x_{2}^{\prime}])\leq 4-r^{\prime}$, then Proposition 40 applied to $u^{\prime}$ would give $R(N[x_{1}^{\prime}])=S(G_{0})\geq 2^{*}$ so $u$ would be covered in Case I.
-On the other hand, if $\operatorname{shad}(u^{\prime},N[x_{1}^{\prime},x_{2}^{\prime}])=\operatorname{minsurp}(G_{3}^{\prime})\geq 5-r^{\prime}$, then vertex $u^{\prime}$, with $v_{1}^{\prime},x_{1}^{\prime},x_{2}^{\prime}$, would be covered by Case II.
-We claim next that $r=6$.
-For, suppose $r=5$, so $S(G_{0})\geq 1^{*}$.
-If $\operatorname{minsurp}(G_{3})\geq 0$, then we would be in Case II.
-On the other hand, if $\operatorname{minsurp}(G_{3})=-1$, then Proposition 40 gives $R(N[x_{1}])\geq 2^{*}$ and in particular $G-N[x_{1}]$ would be $3^{*}$-vul.
-Now our branching rule is to generate subproblems $G_{0},G_{2},G_{3}$.
-By Proposition 40, we have $\operatorname{minsurp}(G_{3})\geq 0$ since $G-N[x_{1}]$ is not $3^{*}$-vul.
-So $G_{2}$ has drop $(3,8^{*})$ and we need:
+In particular, $\deg(x)=4$.
+We claim that GN[x]G-N[x] is not $3^{*}$-vul.
+For, suppose it were, and consider $u^{\prime}=x$ and its 3-neighbor $v^{\prime}=v$ and its two neighbors $x^{\prime}=u,y^{\prime}=y$.
+We have $r^{\prime}=\deg(x^{\prime})+\deg(y^{\prime})-2=r$.
+If $\operatorname{shad}(u^{\prime},N[x^{\prime},y^{\prime}])\leq 4-r^{\prime}$, then Proposition 41 applied to $u^{\prime}$ would give $S(N[x^{\prime}])=S(G_{0})\geq 2^{*}$ so $u$ would be covered in Case I.
+On the other hand, if $\operatorname{shad}(u^{\prime},N[x^{\prime},y^{\prime}])=\operatorname{minsurp}(G_{3}^{\prime})\geq 5-r^{\prime}$, then vertex $u^{\prime}$, with $v^{\prime},x^{\prime},y^{\prime}$, would be covered by Case II.
+Now, if $\operatorname{minsurp}(G_{3})\leq-1$, then by Proposition 41, the graph GN[x]G-N[x] would be $3^{*}$-vul, which we have already ruled out.
+So $\operatorname{minsurp}(G_{3})\geq 0$.
+Consequently, if $r=5$ then observe that $S(G_{0})\geq 1^{*}$ and so we would be in Case II.
+So $r=6$, and our branching rule is to generate subproblems $G_{0},G_{2},G_{3}$, where $G_{2}$ has drop $(3,8^{*})$ and we need:
 $e^{-1.5a-4b-\alpha}+e^{-a-4b-\alpha}+e^{-3a-8b-\alpha}\leq e^{-\beta}$ (4-8)
 Note: henceforth, we know the graph has no special 4-vertex (it would be covered in Case III or earlier). In particular, we have $s=1$ and $\operatorname{bonus}(G)=\alpha$.
-Case IV: $\boldsymbol{S(G_{0})\geq 3^{*}}$.
-Here, we generate subproblems $G_{0},G_{1}^{\prime}$.
-Since $s=1$, the graph $G_{1}^{\prime}$ has a subcubic vertex $v_{2}$ and so it has drop $(0.5,2^{*})$.
-We need:
+Case IV: $\boldsymbol{S(G_{0})\geq 3^{*}}$. We generate subproblems $G_{0},G_{1}$. Since $s=1$, the graph $G_{1}$ has a subcubic vertex $v_{2}$ and so it has drop $(0.5,2^{*})$. We need:
 $e^{-1.5a-7b-\alpha}+e^{-0.5a-2b-\alpha}\leq e^{-\alpha}$ (4-9)
 Case V: $\boldsymbol{r\geq 5}$ and $\boldsymbol{G_{0}}$ is $\boldsymbol{3^{*}}$-vul.
-Here we generate subproblems $G_{0},G_{2},G_{3}$, where $G_{2}$ has drop $(1,3^{*})$.
-We have $\operatorname{minsurp}(G_{3})\geq 5-r$, as otherwise Proposition 40 would give $R(N[x_{1}])\geq 3^{*}$ (bearing in mind that $s=1$); then $u^{\prime}=x_{1}$ would be covered in Case IV.
+We generate subproblems $G_{0},G_{2},G_{3}$, where $G_{2}$ has drop $(1,3^{*})$.
+We have $\operatorname{minsurp}(G_{3})\geq 5-r$, as otherwise Proposition 41 would give $S(N[x])\geq 3^{*}$ (bearing in mind that $s=1$); then $u^{\prime}=x$ would be covered in Case IV.
 So subproblem $G_{3}$ has drop $(2.5,7^{*})$.
@@
 $e^{-1.5a-4b}\gamma_{3}^{*}+e^{-a-3b-\alpha}+e^{-2.5a-7b-\alpha}\leq e^{-\alpha}$ (4-10)
-Case VI: $\boldsymbol{r=6}$.
+Case VI: $\boldsymbol{r\geq 5}$. We must have specifically $r=6$, as if $r=5$ then by Proposition 41 we would have $S(G_{0})\geq 1^{*}$ putting us into Case V.
 Again we generate subproblems $G_{0},G_{2},G_{3}$, where $G_{2}$ has drop $(1,3^{*})$ as before.
-If $G-N[x_{1}]$ is $3^{*}$-vul, then vertex $u^{\prime}=x_{1}$ would be covered in Case V earlier (it has a 3-neighbor $v_{1}$ with a 4-neighbor $u$).
-So assume $G-N[x_{1}]$ is not $3^{*}$-vul; by Proposition 40, this implies $\operatorname{minsurp}(G_{3})\geq 0$ and so $G_{3}$ has drop $(3,8^{*})$.
+If GN[x]G-N[x] is $3^{*}$-vul, then vertex $u^{\prime}=x$ would be covered in Case V earlier.
+So assume GN[x]G-N[x] is not $3^{*}$-vul; by Proposition 41, this implies $\operatorname{minsurp}(G_{3})\geq 0$ and so $G_{3}$ has drop $(3,8^{*})$.
 We need:
 $e^{-1.5a-4b-\alpha}+e^{-a-3b-\alpha}+e^{-3a-8b-\alpha}\leq e^{-\alpha}$ (4-11)
-Case VII: $\boldsymbol{r=4}$.
-Here, $G$ cannot have any 3-vertex $v$ with two 4-neighbors $y_{1},y_{2}$, as then vertex $u^{\prime}=y_{1}$ would have $r^{\prime}\geq 5$ and be covered in previous cases.
+Case VII: Everything else.
+Here, $G$ cannot have any 3-vertex with two 4-neighbors, as any such 4-neighbor $u^{\prime}$ would have $r^{\prime}\geq 5$ and be covered in Case VI.
 So we assume that all 3-vertices in the graph have at least two 3-neighbors.
-In particular, we have $x_{1}\sim t_{1}$ and $x_{2}\sim t_{2}$ for 3-vertices $t_{1},t_{2}$ distinct from $v_{1}$.
-Then in $G_{1}^{\prime}$, the merged vertex $z_{1}$ has degree 4 and has 3-neighbors $t_{1},t_{2}$, that is, it is a special vertex (see Figure 6).
-Since $\operatorname{mindeg}(G_{1}^{\prime})\geq 3$ and $\operatorname{maxdeg}(G_{1}^{\prime})\leq 4$, we get $\operatorname{bonus}(G_{1}^{\prime})=\beta$.
+In particular $r=4$, and $x\sim x^{\prime}$ and $y\sim y^{\prime}$ for 3-vertices $x^{\prime},y^{\prime}$ distinct from $v$.
+We must have $x^{\prime}\neq y^{\prime}$ since $G$ has no 4-cycles.
+The merged vertex $z_{1}$ in $G_{1}$ has degree 4 with 3-neighbors $x^{\prime},y^{\prime}$, that is, it is a special vertex (see Figure 9).
+Since $\operatorname{mindeg}(G_{1})\geq 3$ and $\operatorname{maxdeg}(G_{1})\leq 4$, we get $\operatorname{bonus}(G_{1})=\beta$.
 Also, $S(G_{0})\geq 2^{*}$ since $r=4$.
@@
 $e^{-1.5a-6b-\alpha}+e^{-0.5a-2b-\beta}\leq e^{-\alpha}$ (4-12)
-Figure 6: We show the degrees of each vertex.
-In $G-u$, the 2-vertex $v_{1}$ has neighbors $x_{1},x_{2}$, which get merged into $z_{1}$.
-The new vertex has 4 neighbors in total and at least two 3-neighbors.
+Figure 9: We show the degrees of each vertex. In GuG-u, the 2-vertex $v$ has neighbors $x,y$, which get merged into $z_{1}$. The new vertex has 4 neighbors in total and at least two 3-neighbors.
 6.3 Line 5: Branching on a vertex in a 4-regular graph
-We select a vertex $u$ and an ordering of its neighbors $v_{1},v_{2},v_{3},v_{4}$; our general strategy is to split on $v_{1},v_{2}$.
-The graph $H=G-\{v_{1},v_{2}\}-N[v_{3},v_{4}]$ plays a central role in the analysis.
-We let $I$ be a min-set of $H$; for brevity, we define $A=\{v_{1},v_{2}\}$ in this section.
+We select a vertex $u$ and an ordering of its neighbors $v_{1},v_{2},v_{3},v_{4}$; our general strategy is to split on $v_{1},v_{2}$ so as to create a 2-vertex in the resulting graph $G-\{v_{1},v_{2}\}$.
+To analyze this process, we use the following notations in Section 6.3:
+$H=G-\{v_{1},v_{2}\}-N[v_{3},v_{4}],\qquad A=\{v_{1},v_{2}\}$
+$I=\text{some chosen min-set of $H$},\qquad r=|I|$
+The graph $H$ and its properties are critical to this analysis. We begin with a few observations.
 Proposition 42.
-Suppose $G$ has a 6-cycle $x_{1},\dots,x_{6}$. Then $R(x_{1},N[x_{4}])\geq 2^{*}$; furthermore, either $\operatorname{codeg}(x_{1},x_{4})=1$ or $x_{4}$ is a special 4-vertex in $G-N[x_{1}]$.
+Suppose $G$ has a 6-cycle $x_{1},\dots,x_{6}$. Then $S(x_{1},N[x_{4}])\geq 2^{*}$; furthermore, either $\operatorname{codeg}(x_{1},x_{4})=1$ or $x_{4}$ is a special 4-vertex in $G-N[x_{1}]$.
 Proof.
-At this point, since $\operatorname{girth}(G)\geq 5$ and $G$ is 4-regular, each vertex $v$ has twelve vertices at distance two which all have degree 3 in $G-N[v]$.
-The graph $G^{\prime}=G-x_{1}-N[x_{4}]$ has two 2-vertices $x_{2},x_{6}$.
-We have $x_{2}\not\sim x_{6}$, as otherwise would have a triangle $x_{1},x_{2},x_{6}$.
-Since $G^{\prime}$ has at least 11 subcubic vertices, Proposition 34 gives $S(G^{\prime})\geq 2^{*}$.
+Since $\operatorname{girth}(G)\geq 5$, the graph $G^{\prime}=G-x_{1}-N[x_{4}]$ has two non-adjacent 2-vertices $x_{2},x_{6}$.
+Since $G$ is 4-regular, $G^{\prime}$ has at least 11 subcubic (the 12 vertices at distance two to $x_{4}$, minus $x_{1}$).
+So Proposition 33 gives $S(G^{\prime})\geq 2^{*}$.
 Now suppose $\operatorname{codeg}(x_{1},x_{4})=0$.
@@
 We have $\operatorname{mindeg}(H)\geq 1$. Moreover, for any vertex $x\in H$, there holds:
-  • If $\deg_{H}(x)=1$, then $x\sim A$ and $R(u,N[x])\geq 2^{*}$.
-  • If $\deg_{H}(x)=2$ and $x\not\sim A$, then $x$ is a special 4-vertex in $G-N[u]$.
+  • If $\deg_{H}(x)=1$, then $x\sim A$ and $S(u,N[x])\geq 2^{*}$.
+  • If $\deg_{H}(x)=2$, then either $x\sim A$ or $x$ is a special 4-vertex in GN[u]G-N[u].
 Proof.
 Because $\operatorname{girth}(G)\geq 5$, we have $\operatorname{codeg}(x,v_{3})\leq 1,\operatorname{codeg}(x,v_{4})\leq 1$ and $|N(x)\cap A|\leq 1$.
-If $\deg_{H}(x)=1$, then necessarily $x$ has a neighbor $v_{i}\in A$ and a neighbor $r_{3}\sim v_{3}$ and a neighbor $r_{4}\sim v_{4}$ and thus $u,v_{3},r_{3},x,r_{4},v_{4}$ is a 6-cycle; by Proposition 41 we have $R(u,N[x])\geq 2^{*}$.
+If $\deg_{H}(x)=1$, then necessarily $x$ has a neighbor $v_{i}\in A$ and a neighbor $r_{3}\sim v_{3}$ and a neighbor $r_{4}\sim v_{4}$ and thus $u,v_{3},r_{3},x,r_{4},v_{4}$ is a 6-cycle; by Proposition 42 we have $S(u,N[x])\geq 2^{*}$.
 Likewise, if $\deg_{H}(x)=2$ and $x\not\sim A$, then $x$ has neighbors $r_{3}\sim v_{3},r_{4}\sim v_{4}$, and then $G$ has a 6-cycle $u,v_{3},r_{3},x,r_{4},v_{4}$.
 Furthermore, $x\not\sim A$ by hypothesis and $x\not\sim\{v_{3},v_{4}\}$ since $x\in H$.
-Hence $\operatorname{codeg}(u,x)=0$ and so by Proposition 41 the vertex $x$ is special in $G-N[u]$.
+Hence $\operatorname{codeg}(u,x)=0$ and so by Proposition 42 the vertex $x$ is special in GN[u]G-N[u].
 ∎
 Proposition 44.
-If $\operatorname{minsurp}(H)<0$ and $\deg_{H}(x)\leq 2$ for all $x\in I$, then $G-N[u]$ is special.
+If $\operatorname{minsurp}(H)<0$ and $\deg_{H}(x)\leq 2$ for all $x\in I$, then GN[u]G-N[u] is special.
 Proof.
@@
 Note that $y\not\sim\{v_{3},v_{4}\}$ since $y\in H$ and $y\not\sim\{v_{1},v_{2}\}$ since then $G$ would have a 3-cycle $v_{1},x_{1},y$ or $v_{2},x_{2},y$.
-So $\operatorname{codeg}(u,y)=0$ and hence $G-N[u]$ is special by Proposition 41.
+So $\operatorname{codeg}(u,y)=0$ and hence GN[u]G-N[u] is special by Proposition 42.
 ∎
-For our tie-breaking rule, we assume that if we are in a given case that ordering of neighbors $v_{1}^{\prime},\dots,v_{4}^{\prime}$ would fall into a previous case.
+We now describe the branching strategy, in particular, how to order $v_{1},\dots v_{4}$.
+For our tie-breaking rule, we assume that if we are in a given case that no other ordering of neighbors $v_{1}^{\prime},\dots,v_{4}^{\prime}$ would fall into a previous case.
 Case I: $\boldsymbol{\operatorname{shad}(N[v_{3}])\leq 1}$.
-Let $J$ be a min-set of $G-N[v_{3}]$; since $\operatorname{mindeg}(G-N[v_{3}])\geq 3$, necessarily $|J|\geq 2$, and since $G$ is in 4BP we have $\operatorname{surp}_{G-N[v_{3}]}(J)=1$.
+Let $J$ be a min-set of $G-N[v_{3}]$; since $\operatorname{mindeg}(G-N[v_{3}])\geq 3$, necessarily $|J|\geq 2$, and since we are after Line 1 we have $\operatorname{surp}_{G-N[v_{3}]}(J)=1$.
 Here, we split on $v_{3}$.
-By Proposition 38, $G-N[v_{3}]$ has at least 8 subcubic vertices, so $R(N[v_{3}])\geq 2^{*}$ by Proposition 33.
+By Proposition 39, $G-N[v_{3}]$ has at least 8 subcubic vertices, so $S(N[v_{3}])\geq 2^{*}$ by Proposition 32.
 Overall, we need:
@@
 Case II: $\boldsymbol{\operatorname{minsurp}(H)\leq-2}$.
-Since $\operatorname{mindeg}(H)\geq 1$, we must have $|I|\geq 3$.
+Since $\operatorname{mindeg}(H)\geq 1$, we must have $r\geq 3$.
 Let $J=I\cup\{v_{4}\}$.
@@
 Since we are not in Case I, we have $\operatorname{shad}(N[v_{3}])\geq 2$ and so $J$ is a min-set in $G-N[v_{3}]$.
-Here again we split on $v_{3}$ and then split on $J$ in the subproblem $G-N[v_{3}]$, generating subproblems $G-v_{3},G-N[v_{3}]-J,G-N[v_{3},J]$.
-Since we are not in Case II, we have $\operatorname{shad}(N[v_{3}],J)\geq\operatorname{shad}(N[v_{3}])-|J|\geq 2-|J|$ and likewise $\operatorname{shad}(N[v_{3},J])\geq\operatorname{shad}(N[v_{3}]-|N[J]|\geq 2-|N[J]|$.
+Here we split on $v_{3}$ and in the subproblem $G-N[v_{3}]$, we apply (B-Crit) to $J$.
+Overall, this generates subproblems $G-v_{3},G-N[v_{3}]-J,G-N[v_{3},J]$.
+Since we are not in Case II, we have $\operatorname{shad}(N[v_{3}],J)\geq\operatorname{shad}(N[v_{3}])-|J|\geq 2-|J|$ and likewise $\operatorname{shad}(N[v_{3},J])\geq\operatorname{shad}(N[v_{3}])-|N[J]|\geq 2-|N[J]|$.
 So the latter two subproblems have drops $(2.5,8^{*}),(2.5,10^{*})$ respectively.
@@
 $e^{-0.5a-b-\alpha}+e^{-2.5a-8b-\alpha}+e^{-2.5a-10b-\alpha}\leq 1$ (4-14)
-Note: In all subsequent cases, our strategy is to split on $v_{1},v_{2}$, and then apply (A1) to vertex $u$ in subproblem $G-\{v_{1},v_{2}\}$.
-This gives four subproblems $G_{1}=G-N[u],G_{2}=G-N[v_{1}],G_{3}=G-v_{1}-N[v_{2}],H$.
-Here, $G_{1},G_{2},G_{3}$ have drop $(1.5,4^{*}),(1.5,4^{*}),(2,5^{*})$ directly.
+Note: In all subsequent cases, our strategy is to split on $v_{1},v_{2}$, and then apply (B-Pair) to vertex $u$ in subproblem $G-\{v_{1},v_{2}\}$.
+This gives four subproblems $G_{1}=G-N[u],G_{2}=G-N[v_{1}],G_{3}=G-v_{1}-N[v_{2}],H$ where $G_{1},G_{2},G_{3}$ have drop $(1.5,4^{*}),(1.5,4^{*}),(2,5^{*})$ directly.
 Case III: $\boldsymbol{\operatorname{minsurp}(H)=0}$. Here $H$ has drop $(3.5,9^{*})$, so we need:
 $e^{-1.5a-4b-\alpha}+e^{-1.5a-4b-\alpha}+e^{-2a-5b-\alpha}+e^{-3.5a-9b-\alpha}\leq 1$ (4-15)
-Note: For the remainder of the analysis, we assume $\operatorname{minsurp}(H)=-1$, and we let $r=|I|$; we have $r\geq 2$ since $\operatorname{mindeg}(H)\geq 1$.
+Note: For the remainder of the analysis, we assume $\operatorname{minsurp}(H)=-1$.
+We have $r\geq 2$ since $\operatorname{mindeg}(H)\geq 1$.
 For $r\leq O(1)$, the graph $H$ has drop $(3,(8+r)^{*})$ after applying (P1).
@@
 $e^{-1.5a-4b-\alpha}+e^{-1.5a-4b-\alpha}+e^{-2a-7b-\alpha}+e^{-3a-12b-\alpha}\leq 1$ (4-16)
-Next suppose that $r=3$. Clearly $\deg_{H}(x)\leq 2$ for all $x\in I$, so $G-N[u]$ is special by Proposition 43. Since in this case $H$ has drop $(3,11^{*})$, we then need:
+Next suppose that $r=3$ and $H$ has drop $(3,11^{*})$. For $x\in I$, we have $\deg_{H}(x)\leq|N(I)|=r-1=2$, so GN[u]G-N[u] is special by Proposition 44. We then need:
 $e^{-1.5a-4b-\beta}+e^{-1.5a-4b-\alpha}+e^{-2a-7b-\alpha}+e^{-3a-11b-\alpha}\leq 1$ (4-17)
-Finally suppose $r=2$.
-Again, $G-N[u]$ is special by Proposition 43.
+Finally suppose $r=2$ and $H$ has drop $(3,10^{*})$.
+Again, GN[u]G-N[u] is special by Proposition 44.
 Let $I=\{x_{1},x_{2}\},N_{H}(I)=\{y\}$.
@@
 Now consider the 6-cycle $u,v_{1},x_{1},y,x_{2},v_{2}$.
-Here, $x_{2}$ has neighbors $v_{2},y,r_{3}\sim y_{3},r_{4}\sim y_{4}$.
+Let us enumerate $N(x_{2})=\{v_{2},y,r_{3},r_{4}\}$ for vertices $r_{3}\sim v_{3},r_{4}\sim v_{4}$.
 We must have $\operatorname{codeg}(v_{1},x_{2})=0$; for, if $v_{1}\sim y$, then $G$ would have a 3-cycle $v_{1},x_{1},y$, and if $v_{1}\sim v_{2}$, then $G$ would have a 3-cycle $u,v_{1},v_{2}$, and if $v_{1}\sim r_{i}$ for $i\in\{3,4\}$, then $G$ would have a 4-cycle $r_{i},v_{i},u,v_{1}$.
-So by Proposition 41, the graph $G_{2}=G-N[v_{1}]$ is also special.
-Accordingly, in this case we need:
+So by Proposition 42, the graph $G_{2}=G-N[v_{1}]$ is also special. Accordingly, in this case we need:
 $e^{-1.5a-4b-\beta}+e^{-1.5a-4b-\beta}+e^{-2a-7b-\alpha}+e^{-3a-10b-\alpha}\leq 1$ (4-18)
 Case V: $\boldsymbol{G_{3}}$ is $\boldsymbol{3^{*}}$-vul.
-If $\deg_{H}(x)=1$ for any $x\in I$, then by Proposition 42, $u,x$ share a neighbor $v\in A$ and $R(u,N[x])\geq 2^{*}$.
+If $\deg_{H}(x)=1$ for any $x\in I$, then by Proposition 43, $u,x$ share a neighbor $v\in A$ and $S(u,N[x])\geq 2^{*}$.
 But then vertex $u^{\prime}=v$ and its two neighbors $v_{1}^{\prime}=u,v_{2}^{\prime}=x$ would be covered in Case IV, or an earlier case.
 Accordingly, we must have $\deg_{H}(x)\geq 2$ for all $x\in I$.
-At this point, consider counting the two-element subsets of $N_{H}(I)$.
+Now consider counting the two-element subsets of $N_{H}(I)$.
 Each vertex $x\in I$ contributes $\binom{\deg_{H}(x)}{2}$ such subsets and since $G$ has no four-cycles these are all distinct.
 So $\sum_{x\in I}\binom{\deg_{H}(x)}{2}\leq\binom{|N_{H}(I)|}{2}=\binom{r-1}{2}$.
-Since $\deg_{H}(x)\geq 2$ for all $x\in I$, this implies that $r\geq 5$ and, furthermore, if $r=5$ then $\deg_{H}(x)=2$ for all $x\in I$; in the latter case, by Proposition 43 the graph $G-N[u]$ is special.
-So depending on whether $r=5$ or $r\geq 6$, we need respectively:
+Since $\deg_{H}(x)\geq 2$ for all $x\in I$, this implies that $r\geq 5$ and, furthermore, if $r=5$ then $\deg_{H}(x)=2$ for all $x\in I$; in the latter case, by Proposition 44 the graph GN[u]G-N[u] is special.
+So for $r=5$ or $r\geq 6$, we need respectively:
 $e^{-1.5a-4b-\beta}+e^{-1.5a-4b-\alpha}+e^{-2a-5b}\gamma_{3}^{*}+e^{-3a-13b-\alpha}\leq 1$ (4-19)
@@
 Case VI: Everything else.
-Since $G$ is in 4BP, we have $\operatorname{shad}(v_{1},N[v_{3}])\geq\operatorname{shad}(N[v_{3}])-1\geq 0$.
+Since we are not in Case I, we have $\operatorname{shad}(v_{1},N[v_{3}])\geq\operatorname{shad}(N[v_{3}])-1\geq 1$.
 On the other hand, $\operatorname{surp}_{G-v_{1}-N[v_{3}]}(I\cup\{v_{4}\})\leq\operatorname{minsurp}(H)+1+(\deg_{G-N[v_{3}]}(v_{4})-1)=2$.
 So $G-v_{1}-N[v_{3}]$ is $3^{*}$-vul via indset $I\cup\{v_{4}\}$; it would be covered by Case V.
-So Case VI is impossible.
+Case VI is impossible.
 7 Branching on 5-vertices
+Our strategy for the degree-5 algorithm depends on the relative density of 5-vertices.
+When 5-vertices are very common, we can find closely-clustered groups of them for more advanced branching rules; the denser the 5-vertices, the more favorable the branching.
+We will actually develop three separate degree-5 algorithms with different “intensity” levels.
+There is a major difference between the degree-4 and degree-5 algorithms.
+In the former, branching on a single vertex can create new 2-vertices.
+This cannot be maintained in the degree-5 case: at least two 5-vertices are jointly needed for the desired simplifications.
+The process is not self-sustaining.
+On the positive side, it means that we do not need to keep track of 3-vertices; as far as our analysis goes, they will be created anew at each branching step.
 We will show for the following main result in this section:
 Theorem 45.
-Branch5$(G,k)$ has measure $a_{5}\mu+b_{5}k$ for $a_{5}=0.379406,b_{5}=0.097471$.
+Branch5(G,k)(G,k) has measure $a_{5}\mu+b_{5}k$ for $a_{5}=0.379406,b_{5}=0.097471$.
 By Lemma 15, Theorem 45 combined with the MaxIS-5 algorithm immediately gives the algorithm with runtime $O^{*}(1.24394^{k})$ for degree-5 graphs.
-There is a major difference between the degree-4 and degree-5 settings.
-In the former, we can ensure that the graph has many 3-vertices, and so branching on a single 4-vertex can create new 2-vertices.
-This structure cannot be maintained in the degree-5 case: at least two 5-vertices are jointly needed for the desired simplifications.
-The process is not self-sustaining.
-Our strategy depends on the relative density of 5-vertices.
-When 5-vertices are very common, we can find closely-clustered groups of them for more advanced branching rules; the denser the 5-vertices, the more favorable the branching.
-We will actually develop three separate degree-5 algorithms with different “intensity” levels.
 The first algorithm is very simple and is targeted to the regime where 5-vertices are rare:
-Simplify $G$ If $\operatorname{maxdeg}(G)\leq 4$, then run either the MaxIS-4 algorithm or Branch4 Apply Section 5 to an arbitrary vertex of degree at least $5$.
-Algorithm 6 Function Branch5-1$(G,k)$
+1Simplify $G$
+2If $\operatorname{maxdeg}(G)\leq 4$, then run either the MaxIS-4 algorithm or Branch4
+Apply Figure 6 to an arbitrary vertex of degree at least $5$.
+Algorithm 6 Function Branch5-1(G,k)(G,k)
 Proposition 46.
@@
 Proof.
-Given the potential branch-seqs from Section 5, we need the following constraints:
+Given the potential branch-seqs from Figure 6, we need the following constraints:
 $\max\{e^{-a-3b}+e^{-a-5b},e^{-0.5a-b}+e^{-2a-5b}\}\leq 1$ (5-1)
@@
 The latter constraint can be routinely checked by LP algorithms. ∎
-The next algorithms Branch5-2 and Branch5-3 are targeted to medium and high density of 5-vertices respectively. They share some branching rules, so we describe them in a single listing.
-1 Apply miscellaneous branching rules to ensure $G$ has nice structure.
-(We call this 5-base-position (5BP).) Branch on non-adjacent 5-vertices $u,v$ with $R(u,v)\geq 1$ and $R(u,N[v])\geq 1$.
-Branch on vertices $u,v$ where $\deg(u)=5,\deg_{G-u}(v)\geq 4$, and $G-\{u,v\}$ has a kite.
-Branch on a 4-vertex with two non-adjacent 5-neighbors Branch on a 5-vertex $u$ with neighbors $v_{1},v_{2},v_{3},v_{4},v_{5}$ where $v_{1}\sim v_{2}$ and $\deg(v_{1})=\deg(v_{3})=\deg(v_{4})=\deg(v_{5})=5$.
-Branch on a $5$-vertex with five $5$-neighbors.
-if $L=2$ then
-2        Branch on a 5-vertex $u$ where $G-N[u]$ is 3-vul.
-Branch on non-adjacent 5-vertices $u,v$ with $R(u,v)\geq 1$.
-Branch on a $4$-vertex which has a triangle involving a 5-vertex.
-Branch on a 5-vertex with fewer than three 4-neighbors.
-Run the cheaper of Branch5-$(L-1)$$(G,k)$ or the MaxIS-5 algorithm
-Algorithm 7 Function Branch5-$L$$(G,k)$ for $L=2,3$
+The next algorithms Branch5-2 and Branch5-3 are targeted to medium and high density of 5-vertices respectively. They share many branching rules, so we describe them in a single listing.
+1
+2Apply miscellaneous two-way branching rules to ensure $G$ has nice structure.
+3Branch on non-adjacent 5-vertices $u,v$ with $S(u,v)\geq 1$ and $S(u,N[v])\geq 1$.
+4Branch on vertices $u,v$ where $\deg(u)=5,\deg_{G-u}(v)\geq 4$, and $G-\{u,v\}$ has a kite.
+5Branch on a 4-vertex with two non-adjacent 5-neighbors
+6Branch on a 5-vertex $u$ with neighbors $v_{1},v_{2},v_{3},v_{4},v_{5}$ where $v_{1}\sim v_{2}$ and $\deg(v_{1})=\deg(v_{3})=\deg(v_{4})=\deg(v_{5})=5$.
+7Branch on a 55-vertex with five 55-neighbors.
+8if $L=2$ then
+9   Branch on a 5-vertex $u$ where GN[u]G-N[u] is 3-vul.
+10  Branch on non-adjacent 5-vertices $u,v$ with $S(u,v)\geq 1$.
+11  Branch on a 44-vertex which has a triangle involving a 5-vertex.
+12  Branch on a 5-vertex with fewer than three 4-neighbors.
+Run the cheaper of Branch5-(L1)(L-1)$(G,k)$ or the MaxIS-5 algorithm
+Algorithm 7 Function Branch5-LL(G,k)(G,k) for $L=2,3$
 We will show the following results for these algorithms:
 Theorem 47.
-Branch5-2$(G,k)$ has measure $a_{52}\mu+b_{52}k$ for $a_{52}=0.437086,b_{52}=0.080751$.
-Branch5-3$(G,k)$ has measure $a_{53}\mu+b_{53}k$ for $a_{53}=0.379406,b_{53}=0.097471$.
+Branch5-2(G,k)(G,k) has measure $a_{52}\mu+b_{52}k$ for $a_{52}=0.437086,b_{52}=0.080751$.
+Branch5-3(G,k)(G,k) has measure $a_{53}\mu+b_{53}k$ for $a_{53}=0.379406,b_{53}=0.097471$.
 Thus, algorithm Branch5-3 will achieve the claim of Theorem 45.
 7.1 Regularizing the graph and structural properties
-We say that $G$ is in 5-base position (or 5BP for short) if the following conditions hold:
+In Line 1, we will apply a series of branching rules to ensure the following conditions:
   • $G$ is simplified.
   • $\operatorname{maxdeg}(G)\leq 5$.
-  • Every 5-vertex $u$ has $\operatorname{shad}(N[u])\geq 1$ and $R(u)=0$.
+  • Every 5-vertex $u$ has $\operatorname{shad}(N[u])\geq 1$ and $S(u)=0$.
   • Every 4-vertex $u$ has $\operatorname{shad}(N[u])\geq 1$.
-  • There are no vertices $u,v$ with $\operatorname{codeg}(u,v)\geq 2$ and $R(u,v)+R(N(u)\cap N(v))+\operatorname{codeg}(u,v)\geq 5$.
-Note that, if $G$ is in 5BP, then any 5-vertex $u$ only has neighbors of degree 4 and 5 (else we would have $R(u)\geq 1$).
-Unlike in the degree-3 case, we cannot ensure that $G$ has no four-cycles.
+  • There are no vertices $u,v$ with $\operatorname{codeg}(u,v)\geq 2$ and $S(u,v)+S(N(u)\cap N(v))+\operatorname{codeg}(u,v)\geq 5$.
+Note that, when these conditions hold, then no 5-vertex $u$ can have subcubic neighbors (else we would have $S(u)\geq 1$).
+Unlike in the degree-4 case, we cannot ensure that $G$ has no four-cycles.
 However, we can show that the four-cycles are heavily restricted.
 Proposition 48.
-If $G$ is in 5BP, then any pair of non-adjacent vertices $x,y$ have $\operatorname{codeg}(x,y)\leq 2$.
+After Line 1, any pair of non-adjacent vertices $x,y$ have $\operatorname{codeg}(x,y)\leq 2$.
 Proof.
@@
 If $\deg(x)=3$, then $\{x,y\}$ would be an indset with surplus one, contradicting that $G$ is simplified.
-If $\deg(y)=4$, then $\operatorname{shad}(N[y])\leq 0$ (due to vertex $x$), contradicting that $G$ is in 5BP.
+If $\deg(y)=4$, then $\operatorname{shad}(N[y])\leq 0$ (due to vertex $x$), and we would branch at Line 1.
 So we suppose $\deg(x)=\deg(y)=5$.
 Again, if $\operatorname{codeg}(x,y)\geq 4$ then $\operatorname{shad}(N[y])\leq 0$ due to vertex $x$, so we suppose that $\operatorname{codeg}(x,y)=3$ exactly.
-Then $R(N(x)\cap N(y))\geq 2$ due to non-adjacent 2-vertices $x,y$.
-So $R(x,y)+R(N(x)\cap N(y))+\operatorname{codeg}(x,y)\geq 5$ contradicting that $G$ is in 5BP.
+Then $S(N(x)\cap N(y))\geq 2$ due to non-adjacent 2-vertices $x,y$.
+So $S(x,y)+S(N(x)\cap N(y))+\operatorname{codeg}(x,y)\geq 5$ and we would have branched at Line 1.
 ∎
 Proposition 49.
-If $G$ is in 5BP, then any 4-cycle $x_{1},x_{2},x_{3},x_{4}$ satisfies the following properties:
+After Line 1, any 4-cycle $x_{1},x_{2},x_{3},x_{4}$ satisfies the following properties:
   • All the vertices $x_{i}$ have degree at least four.
-  • If $\deg(x_{1})=4$ then $x_{1}\not\sim x_{3}$
+  • If $x_{1}\sim x_{3}$ then $\deg(x_{1})=\deg(x_{3})=5$.
   • At least two of the vertices $x_{i}$ have degree five.
@@
 By Proposition 48, this means that $x_{2},x_{4}$ are the only shared neighbors of $x_{1},x_{3}$.
-If $\deg(x_{3})=4$, then we would have $\operatorname{shad}(N[x_{3}])\leq 0$ (due to 1-vertex $x_{1}$), contradicting that $G$ is simplified.
+If $\deg(x_{3})=4$, then $\operatorname{shad}(N[x_{3}])\leq 0$ (due to 1-vertex $x_{1}$), contradicting that $G$ is simplified.
 So, suppose that $\deg(x_{3})=5$.
 Since $x_{2},x_{4}$ are neighbors of 5-vertex $x_{3}$ and 3-vertex $x_{1}$, they must have degree 4 (as 5-vertices cannot be neighbors of 3-vertices).
-At this point, observe that $R(x_{1},x_{3})\geq 2$ (due to non-adjacent 2-vertices $x_{2},x_{4}$) and $R(N(x_{1})\cap N(x_{3}))=R(x_{2},x_{4})\geq 1$ (due to 1-vertex $x_{1}$), contradicting that $G$ is in 5BP.
-For the second result, suppose $\deg(x_{1})=4$ and $x_{1}\sim x_{3}$ and $\deg(x_{2})\geq\deg(x_{4})$; we have already shown that $\deg(x_{4})\geq 4$.
+At this point, observe that $S(x_{1},x_{3})\geq 2$ (due to non-adjacent 2-vertices $x_{2},x_{4}$) and $S(N(x_{1})\cap N(x_{3}))=S(x_{2},x_{4})\geq 1$ (due to 1-vertex $x_{1}$); we would branch at Line 1.
+For the second result, suppose without loss of generality that $\deg(x_{1})=4$ and $x_{1}\sim x_{3}$ and $\deg(x_{2})\geq\deg(x_{4})$; we have already shown that $\deg(x_{4})\geq 4$.
 We have $x_{2}\not\sim x_{4}$, as otherwise $x_{1}$ has a funnel (due to clique $x_{2},x_{3},x_{4}$).
-If $\deg(x_{2})=5$, then $R(x_{2})\geq 1$ (due to 3-triangle $x_{1},x_{3},x_{4}$).
-Otherwise, if $\deg(x_{2})=\deg(x_{4})=4$, then $R(N(x_{1})\cap N(x_{3}))\geq 1$ (due to subquadratic vertex $x_{2}$) and $R(x_{1},x_{3})\geq 2$ (due to non-adjacent 2-vertices $x_{2},x_{4}$); this contradicts that $G$ is in 5BP.
+If $\deg(x_{2})=5$, then $S(x_{2})\geq 1$ (due to 3-triangle $x_{1},x_{3},x_{4}$).
+Otherwise, if $\deg(x_{2})=\deg(x_{4})=4$, then $S(N(x_{1})\cap N(x_{3}))\geq 1$ (due to subquadratic vertex $x_{2}$) and $S(x_{1},x_{3})\geq 2$ (due to non-adjacent 2-vertices $x_{2},x_{4}$); we would branch at Line 1.
 For the final result, suppose that $\deg(x_{1})=\deg(x_{2})=\deg(x_{3})=4$.
@@
 So $G-\{x_{2},x_{4}\}$ has two non-adjacent subquadratic vertices $x_{1},x_{3}$.
-Thus $R(N(x_{1})\cap N(x_{3}))=R(x_{2},x_{4})\geq 2$ by Observation 16, and $R(x_{1},x_{3})\geq 1$ due to subquadratic vertex $x_{2}$.
-This contradicts that $G$ is in 5BP.
+Thus $S(N(x_{1})\cap N(x_{3}))=S(x_{2},x_{4})\geq 2$ by Observation 12, and $S(x_{1},x_{3})\geq 1$ due to subquadratic vertex $x_{2}$.
+We would branch at Line 1.
 ∎
 Proposition 50.
-Suppose $G$ is in 5BP.
-If a 4-vertex $u$ has a 5-neighbor, then the neighbors of $u$ can be enumerated as $v_{1},v_{2},v_{3},v_{4}$, where $\deg(v_{1})=5$ and $\{v_{3},v_{4}\}\not\sim\{v_{1},v_{2},v_{3},v_{4}\}$.
-Moreover, after Line 4, we may further suppose that $v_{3},v_{4}$ are subquartic.
+After Line 1, the neighbors of any 4-vertex $u$ can be enumerated as $v_{1},\dots,v_{4}$ to satisfy one of the following three conditions:
+  • (I)
+    $v_{1},\dots,v_{4}$ are all subquartic.
+  • (II)
+    $\deg(v_{1})=5$ and $v_{3},v_{4}$ are subquartic, and $G[N(u)]$ has exactly one edge $(v_{1},v_{2})$.
+  • (III)
+    $\deg(v_{1})=5$, and all the vertices $v_{1},\dots,v_{4}$ are independent.
+Moreover, after Line 4, in case (III) we may further suppose that $v_{2},v_{3},v_{4}$ are subquartic.
 Proof.
-Let $v_{1}$ be a 5-neighbor of $u$.
+Suppose $u$ has a 5-neighbor $v_{1}$.
 By Proposition 49, we must have $\operatorname{codeg}(u,v_{1})\leq 1$.
-Also, if there are any edges among the other neighbors of $u$, then $R(v_{1})\geq 1$ due to the resulting 3-triangle on $u$.
-Finally, if either $v_{3}$ or $v_{4}$ has degree 5, then $u$ would have two non-adjacent 5-neighbors and we would branch at Line 4.
+Also, if there are any edges among the other neighbors of $u$, then $S(v_{1})\geq 1$ due to the resulting 3-triangle on $u$.
+In Case II, if say $\deg(v_{3})=5$, then we would have $S(v_{3})\geq 1$ (due to the 3-triangle on $u,v_{1},v_{2}$), and we would have applied the Line-1 branching rules.
+In Case III, if say $\deg(v_{2})=5$, then $u$ would have two non-adjacent 5-neighbors and we would branch at Line 4.
 ∎
@@
 Suppose $x\sim u$ and $u,x$ share neighbors $y_{1},y_{2}$; let $v_{1},v_{2}$ be the two remaining neighbors of $u$, where $\deg(v_{1})=5$.
-Since we did not branch on $v_{1},v_{2}$ at Line 3, we must have $\deg(v_{2})=4,v_{1}\sim v_{2}$ and $\deg(y_{3})=\deg(y_{2})=\deg(x)=5$.
-By Proposition 49, we have $\{y_{1},y_{2},x\}\not\sim v_{2}$.
-Then $y_{1}\not\sim v_{1}$, as otherwise $G-\{y_{2},x\}$ would have a kite, where $\deg_{G-y_{2}}(v_{2})=\deg_{G}(v_{2})=4$, and we would branch at Line 3.
-So $R(v_{1},y_{1})\geq 1$ (due to the 3-triangle $u,x,y_{2}$) and $R(v_{1},N[y_{1}])\geq 1$ (due to 2-vertex $v_{2}$), and we would branch on non-adjacent 5-vertices $v_{1},y_{1}$ at Line 2.
+Note that $G-\{v_{1},v_{2}\}$ contains a kite $u,x,y_{1},y_{2}$.
+Since we did not branch on $v_{1},v_{2}$ at Line 3, we must have $\deg(v_{2})=4,v_{1}\sim v_{2}$.
+Since $u$ has at least four 5-neighbors, this implies that $\deg(y_{1})=\deg(y_{2})=\deg(x)=5$.
+The 4-vertex $v_{2}$ must be in Case II of Proposition 50.
+So $\{y_{1},y_{2},x\}\not\sim v_{2}$.
+Then $y_{1}\not\sim v_{1}$, as otherwise $G-\{y_{2},v_{2}\}$ would have a kite, where $\deg_{G-y_{2}}(v_{2})=4$, and we would branch at Line 3.
+So $S(v_{1},y_{1})\geq 1$ (due to the 3-triangle $u,x,y_{2}$) and $S(v_{1},N[y_{1}])\geq 1$ (due to 2-vertex $v_{2}$), and we would branch on non-adjacent 5-vertices $v_{1},y_{1}$ at Line 2.
 ∎
@@
 Proof.
-Let $A_{51},A_{5+}$ be the set of 5-vertices with exactly one 4-neighbor or more than one 4-neighbor, respectively, and let $B$ be the set of 4-vertices with exactly one 5-neighbor.
+Let $A_{51},A_{5+}$ be the set of 5-vertices with exactly one 4-neighbor or more than one 4-neighbor respectively, and let $B$ be the set of 4-vertices with exactly one 5-neighbor.
 Note that $A_{51},A_{5+}$ partition $V_{5}$ at this point; for, if a 5-vertex $u$ has five independent 5-neighbors, we would branch on it at Line 6; if it has a triangle in its neighborhood, we would branch on $u$ at Line 5.
-We claim that, for any vertex $u\in A_{51}$ with 4-neighbor $x$, we have $x\in B$.
-For, suppose that $x$ has a second 5-neighbor $v$; we must have $u\sim v$ else we would branch on $x$ at Line 4.
-Then $u$ has three other 5-neighbors $v_{3},v_{4},v_{5}$; we would branch on $u$ at Line 5.
-By counting edges between $B$ and $A_{51}$, we thus see that $|B|\geq|A_{51}|$.
-Next let us count the edges between $V_{4}$ and $V_{5}$.
-Clearly there are least $|A_{51}|+2|A_{5+}|$ such edges.
-Furthermore, each vertex $v\in B$ has exactly one 5-neighbor, and by Proposition 49 any 4-vertex $v$ has at most two 5-neighbors.
+We claim that, for any vertex in $A_{51}$ its sole 4-neighbor must be in $B$, and hence $|B|\geq|A_{51}|$.
+For, suppose that $u\in A_{51}$ has a neighbor $x\in B$ and $x$ has a second 5-neighbor $v$; if $u\not\sim v$ we would branch on $x$ at Line 4 and if $u\sim v$ we would branch on $u$ at Line 5.
+There are at least $|A_{51}|+2|A_{5+}|$ edges between $V_{4}$ and $V_{5}$.
+Conversely, each vertex $v\in B$ has exactly one 5-neighbor, and by Proposition 50 any 4-vertex $v$ has at most two 5-neighbors.
 So $|B|+2(|V_{4}|-|B|)\leq|A_{51}|+2|A_{5+}|$.
-Combining this inequality with $|B|\geq|A_{51}|$, we conclude that $|V_{4}|=|B|+(|V_{4}|-|B|)\geq|A_{51}|+|A_{5+}|=|V_{5}|.$
-Observation 52.
+Since $|B|\geq|A_{51}|$, we conclude that $|V_{4}|=|B|+(|V_{4}|-|B|)\geq|A_{51}|+|A_{5+}|=|V_{5}|.$
+Observation 53.
 After Line 8 of algorithm Branch5-2, any 5-vertex $x$ has $\operatorname{shad}(N[x])\geq 2$, and has no 3-vertex within distance two.
@@
 Proof.
-If $\operatorname{shad}(N[x])\leq 1$, then since $G$ is in 5BP we would have $R(N[x])\geq 1$ and we would branch on $x$ at Line 2.
+If $\operatorname{shad}(N[x])\leq 1$, then after Line 1 we would have $S(N[x])\geq 1$ and we would branch on $x$ at Line 8.
 So $\deg_{G-N[x]}(y)\geq 3$, i.e.
 $\operatorname{codeg}(x,y)\leq\deg(y)-3$.
-We have already seen that $x$ has no 3-neighbor since $G$ is in 5BP; this also shows it cannot have any 3-vertex at distance two.
+We have already seen that $x$ has no 3-neighbor; this also shows it cannot have any 3-vertex at distance two.
 ∎
@@
 For, consider a vertex $t\in G^{\prime}$.
-If $\deg_{G}(t)=3$, then by Observation 52 we have $\operatorname{codeg}(x,t)=0$ and by Proposition 48 we have $\operatorname{codeg}(y,t)\leq 1$.
+If $\deg_{G}(t)=3$, then by Observation 53 we have $\operatorname{codeg}(x,t)=0$ and by Proposition 49 we have $\operatorname{codeg}(y,t)\leq 1$.
 So $\deg_{G^{\prime}}(t)\geq 2$.
@@
 So again $\deg_{G^{\prime}}(t)\geq 2$.
-Next suppose $\deg_{G}(t)=4$; by Proposition 47 and Observation 52 we have $\operatorname{codeg}(y,t)\leq 2$ and $\operatorname{codeg}(x,t)\leq 1$.
+Next suppose $\deg_{G}(t)=4$; by Proposition 48 and Observation 53 we have $\operatorname{codeg}(y,t)\leq 2$ and $\operatorname{codeg}(x,t)\leq 1$.
 To get $\deg_{G^{\prime}}(t)=1$, vertex $t$ must share two neighbors $z_{1},z_{2}$ with $y$ and share a distinct neighbor $z_{3}$ with $x$.
 By Proposition 49 we must have $\deg(z_{1})=\deg(z_{2})=5$.
-If $z_{1}\not\sim z_{2}$, then $R(z_{1},z_{2})\geq 1$ (due to subquadratic vertex $y$), and we would branch on $z_{1},z_{2}$ at Line 9.
-If $z_{1}\sim z_{2}$, then in the graph $G-N[x]$, vertex $t$ loses its neighbor $z_{3}$, and so it has a 3-triangle $z_{1},z_{2}$; in particular, $R(N[x])\geq 1$ and we would branch on $x$ at Line 8.
+If $z_{1}\not\sim z_{2}$, then $S(z_{1},z_{2})\geq 1$ (due to subquadratic vertex $y$), and we would branch on $z_{1},z_{2}$ at Line 9.
+If $z_{1}\sim z_{2}$, then in the graph GN[x]G-N[x], vertex $t$ loses its neighbor $z_{3}$, and so it has a 3-triangle $z_{1},z_{2}$; in particular, $S(N[x])\geq 1$ and we would branch on $x$ at Line 8.
 Thus, in all cases, $\deg_{G^{\prime}}(t)\geq 2$.
@@
 Then $\operatorname{surp}_{G-N[x]}(I\cup\{y\})\leq 0+\deg_{G-N[x]}(y)-1=2$.
-So $G-N[x]$ would be 3-vul and we would branch on $x$ at Line 8.
+So GN[x]G-N[x] would be 3-vul and we would branch on $x$ at Line 8.
 ∎
-Observation 54.
-After Line 10 of algorithm Branch5-2, any 4-vertex has at most one 5-neighbor and is not part of any 4-cycle.
+Observation 55.
+After Line 10 of algorithm Branch5-2, any 4-vertex $v$ has at most one 5-neighbor and is not part of any 4-cycle.
 Proof.
-Consider a 4-vertex $v$.
-If $v$ has two 5-neighbors we would branch at Line 6 or Line 10, depending if they were adjacent or not.
+If $v$ has two 5-neighbors we would branch at Line 4 or Line 10, depending if they were adjacent or not.
 Now consider a four-cycle $v,x_{1},x_{2},x_{3}$.
-By Observation 52, we have $\deg(x_{2})\leq 4$.
-Thus, by Proposition 48, both $x_{1}$ and $x_{3}$ have degree 5; that is, $v$ has two 5-neighbors, which is what we have showed is impossible.
+By Observation 53, we have $\deg(x_{2})\leq 4$.
+Thus, by Proposition 49, we have $\deg(x_{1})=\deg(x_{3})=5$; that is, $v$ has two 5-neighbors, which is what we have showed is impossible.
+∎
 7.2 Basic branching rules
 We analyze it line by line; Lines 6 and 11, which are more involved, will be handled later.
-Line 1: We can first simplify $G$.
-If a vertex $u$ has degree at least 6, we apply Section 5.
+Line 1: We first simplify $G$.
+If a vertex has degree at least 6, we apply Figure 6.
 If a 5-vertex $u$ has $\operatorname{shad}(N[u])\leq 0$ we apply Lemma 26.
-Otherwise if a 5-vertex $u$ has $R(u)\geq 1$, then we split on $u$ giving branch-seq $[(0.5,2),(2,5)]$.
+Otherwise if a 5-vertex $u$ has $S(u)\geq 1$, then we split on $u$ giving branch-seq $[(0.5,2),(2,5)]$.
 If a 4-vertex $u$ has $\operatorname{shad}(N[u])\leq 0$, then we apply Proposition 25.
-If vertices $u,v$ have $R(u,v)+R(N(u)\cap N(v)))+\operatorname{codeg}(u,v)\geq 5$, we apply (A3).
+If vertices $u,v$ have $S(u,v)+S(N(u)\cap N(v)))+\operatorname{codeg}(u,v)\geq 5$, we apply (B-Shared).
 Over all cases, we get branch-seq $[(1,3),(1,4)]$, $[(1,2),(1,5)]$, $[(0.5,2),(2,5)]$, or $[(0.5,1),(2.5,6)]$. Accordingly, we need:
@@
 Line 2: We split on $u,v$, obtaining subproblems $G_{1}=G-u,G_{2}=G-u-N[v],G_{3}=G-\{u,v\}$.
-Since $G$ is in 5BP, they have drop $(2,5),(2.5,6),(1,2)$ directly.
+They have drop $(2,5),(2.5,6),(1,2)$ directly.
 After simplification, $G_{2}$ has drop $(2.5,7)$ and $G_{3}$ has drop $(1,3)$.
@@
 Line 3: We split on $u,v$ and apply (P3) to $G-\{u,v\}$.
-Since $G$ is in 5BP, subproblems $G-N[u]$ and $G-u-N[v]$ both have drops $(2,5)$.
+Since we are after Line 1, subproblems GN[u]G-N[u] and GuN[v]G-u-N[v] both have drops $(2,5)$.
 Also, $\operatorname{shad}(u,v)\geq\operatorname{shad}(u)-1\geq 1$ so by Lemma 14 the subproblem $G-\{u,v\}$ has drop $(1.5,4)$ after simplification.
@@
 $e^{-2a-5b}+e^{-2a-5b}+e^{-1.5a-4b}\leq 1$ (5-5)
-Line 4: Consider a 4-vertex $u$ with non-adjacent 5-neighbors $y_{1},y_{2}$, and let $x_{1},x_{2}$ be the other two neighbors of $u$ where $\deg(x_{1})\leq\deg(x_{2})$.
-Our plan is to split on $y_{1},y_{2}$, generating subproblems $G_{1}=G-N[y_{1}],G_{2}=G-y_{1}-N[y_{2}],G_{3}=G-\{y_{1},y_{2}\}$.
-Since $G$ is in 5BP, the subproblems have drops $(2,5),(2.5,6),(1,2)$ directly.
-By Proposition 49, the neighbors of $u$ must be independent.
-If $\deg(x_{1})\leq 3$, then $R(y_{1},y_{2})\geq 1$ (due to 2-vertex $u$) and $R(y_{1},N[y_{2}])\geq 1$ (due to 2-vertex $x_{1}$); in this case we would branch on $y_{1},y_{2}$ at Line 2.
-So we may assume that $\deg(x_{1})\geq 4$.
+Line 4: Consider a 4-vertex $u$ with non-adjacent 5-neighbors $y_{1},y_{2}$.
+This must be Case III of Proposition 50, where all neighbors of $u$ are independent; let $x_{1},x_{2}$ be the other two neighbors of $u$, sorted so that $\deg(x_{1})\leq\deg(x_{2})$.
+We split on $y_{1},y_{2}$, generating subproblems $G_{1}=G-N[y_{1}],G_{2}=G-y_{1}-N[y_{2}],G_{3}=G-\{y_{1},y_{2}\}$.
+Since we are after Line 1, the subproblems have drops $(2,5),(2.5,6),(1,2)$ directly.
+If $\deg(x_{1})=3$, then $S(y_{1},y_{2})\geq 1$ (due to 2-vertex $u$) and $S(y_{1},N[y_{2}])\geq 1$ (due to 2-vertex $x_{1}$); in this case we would branch on $y_{1},y_{2}$ at Line 2.
+So suppose that $\deg(x_{1})\geq 4$.
 We can apply Proposition 30 to 2-vertex $u$ in $G_{3}$ with $r=\deg(x_{1})+\deg(x_{2})-\operatorname{codeg}(x_{1},x_{2})-1$.
@@
 $e^{-2a-5b}+e^{-2.5a-6b}+e^{-a-2b}\psi_{6}\leq 1$ (5-6)
-Line 5: Vertices $v_{3},v_{4},v_{5}$ cannot form a clique (else $R(v_{1})\geq 1$ due to the funnel $u$), so assume without loss of generality $v_{3}\not\sim v_{4}$.
+Line 5: Vertices $v_{3},v_{4},v_{5}$ cannot form a clique (else $S(v_{1})\geq 1$ due to the funnel $u$), so assume without loss of generality $v_{3}\not\sim v_{4}$.
 Then we split on $v_{3},v_{4}$, generating subproblems $G_{1}=G-N[v_{3}],G_{2}=G-v_{3}-N[v_{4}],G_{3}=G-\{v_{3},v_{4}\}$, which have drops $(2,5),(2.5,6),(1,2)$ directly.
@@
 So $v_{1}\not\sim v_{5}$ and $\deg_{G_{3}}(v_{1})=5,\deg_{G_{3}}(v_{5})\geq 4$ and, by Proposition 48, we have $\operatorname{codeg}(v_{1},v_{5})\leq 2$.
-So $\deg_{G_{3}}(v_{1})+\deg_{G_{3}}(v_{5})-\operatorname{codeg}_{G_{3}}(v_{1},v_{5})-1\geq 5+4-2-1=6$.
+So $\deg_{G_{3}}(v_{1})+\deg_{G_{3}}(v_{5})-\operatorname{codeg}_{G_{3}}(v_{1},v_{5})-1\geq 6$.
 Overall we need
@@
 which is just a repetition of (5-6).
-Line 8: Since $G$ is in 5BP, subproblem $G-N[u]$ has drop $(2,5)$ directly, and subproblem $G-u$ has drop $(0.5,1)$. We apply Proposition 31 to $G-N[u]$, so we need:
+Line 8: Since we are after Line 1, subproblem GN[u]G-N[u] has drop $(2,5)$ directly, and subproblem GuG-u has drop $(0.5,1)$. We apply Proposition 31 to GN[u]G-N[u], so we need:
 $e^{-2a-5b}\gamma_{3}+e^{-0.5a-b}\leq 1\qquad\text{(when $L=2$)}$ (5-7)
-Line 9: We get subproblems $G-N[u],G-u-N[v],G-\{u,v\}$. Since $G$ is in 5BP, these have drops $(2,5),(2.5,6),(1,2)$ directly, and the latter has drop $(1,3)$ after simplification. We need:
+Line 9: We get subproblems $G-N[u],G-u-N[v],G-\{u,v\}$. Since we are after Line 1, these have drops $(2,5),(2.5,6),(1,2)$ directly, and the latter has drop $(1,3)$ after simplification. We need:
 $e^{-2a-5b}+e^{-2.5a-6b}+e^{-a-3b}\leq 1\qquad\text{(when $L=2$)}$ (5-8)
-Line 10: Suppose we are branching on a 4-vertex $u$ with neighbors $x_{1},x_{2},y_{1},y_{2}$ where $\deg(x_{1})=5$ and $x_{1}\sim x_{2}$.
-By Proposition 49, $u$ has no other edges in its neighborhood and $y_{1},y_{2}$ are subquartic.
-Since $x_{2},y_{1},y_{2}$ are within distance two of 5-vertex $x_{1}$ we have $\deg(y_{1})=\deg(y_{2})=4$ and so by Observation 52 we have $\operatorname{codeg}(x_{1},y_{2})=\operatorname{codeg}(x_{1},y_{1})=1$.
-Here, we split on $y_{1}$, and then apply (A2) to $x_{1}$ in subproblem $G-y_{1}$.
+Line 10: Let us branch on a 4-vertex $u$ with neighbors $x_{1},x_{2},y_{1},y_{2}$ where $\deg(x_{1})=5$ and $x_{1}\sim x_{2}$.
+This is Case II of Proposition 50: $u$ has no other edges in its neighborhood and $y_{1},y_{2}$ are subquartic.
+Since $x_{2},y_{1},y_{2}$ are within distance two of 5-vertex $x_{1}$ we have $\deg(y_{1})=\deg(y_{2})=4$ and so by Observation 53 we have $\operatorname{codeg}(x_{1},y_{2})=\operatorname{codeg}(x_{1},y_{1})=1$.
+Here, we split on $y_{1}$, and then apply (B-Fun2) to $x_{1}$ in subproblem $G-y_{1}$.
 This generates subproblems $G_{1}=G-N[y_{1}],G_{2}=G-y_{1}-N[y_{2},x_{1}]$, and $G_{3}=G-\{y_{1},x_{1}\}$.
-Since $G$ is in 5BP, subproblem $G_{1}$ has drop $(1.5,4)$ directly.
-Also, by Lemma 53, we have $\operatorname{minsurp}(G_{2})\geq\operatorname{shad}(N[y_{2},x_{1}])-1\geq 0$.
-So $G_{2}$ has drop $(3.5,9)$.
+Subproblem $G_{1}$ has drop $(1.5,4)$.
+By Lemma 54, $\operatorname{minsurp}(G_{2})\geq\operatorname{shad}(N[y_{2},x_{1}])-1\geq 0$, so $G_{2}$ has drop $(3.5,9)$.
 Subproblem $G_{3}$ has drop $(1,2)$ directly, and has a 2-vertex $u$ so it has drop $(1,3)$ after simplification.
@@
 which can be checked routinely by LP algorithms.
-For Branch5-2, each 5-vertex has at least three 4-neighbors (else we would branch at Line 12), and each 4-vertex has at most one 5-neighbor. So $n_{4}\geq 3n_{5}$, and we need
+For Branch5-2, each 5-vertex has at least three 4-neighbors (else we would branch at Line 11), and each 4-vertex has at most one 5-neighbor. So $n_{4}\geq 3n_{5}$, and we need
 $\min\{a_{51}\mu+b_{51}k,2(k-\mu)\cdot(w_{4}\cdot 3/4+1/4)\log 1.17366\}\leq a\mu+bk$ (5-11)
@@
 7.3 Line 6: A 5-vertex with five independent 5-neighbors
-Our basic plan is to choose an ordering of the neighbors of $u$ as $v_{1},\dots,v_{5}$, and then split on $v_{1},v_{2},v_{3}$ to create a 2-vertex $u$.
-The graph $H=G-\{v_{1},v_{2},v_{3}\}-N[v_{4},v_{5}]$ is central to this analysis; note that $\operatorname{codeg}(v_{4},v_{5})\in\{1,2\}$ by Proposition 47.
-Let $I$ be a min-set of $H$, and define parameters
-$s=|N(I)|,\qquad\ell=\operatorname{minsurp}(H),\qquad d_{45}=\operatorname{codeg}(v_{4},v_{5})$
-For this section, we define $A=\{v_{1},v_{2},v_{3}\}$ and $B=\{v_{4},v_{5}\}$, so that $H=G-(A\cup N[B])$.
+Our basic plan is to choose an ordering of the neighbors of $u$ as $v_{1},\dots,v_{5}$, and then split on $v_{1},v_{2},v_{3}$ to create a 2-vertex $u$. For Section 7.3, we use the following notations:
+$A=\{v_{1},v_{2},v_{3}\},\qquad B=\{v_{4},v_{5}\},\qquad d_{45}=\operatorname{codeg}(v_{4},v_{5})$
+$H=G-\{v_{1},v_{2},v_{3}\}-N[v_{4},v_{5}]=G-A-N[B]$
+$I=\text{a chosen min-set of $H$},\qquad s=|N(I)|,\qquad\ell=\operatorname{minsurp}(H)=\operatorname{surp}_{H}(I)$
+The properties of the graph $H$ are central to this analysis; note that $d_{45}\in\{1,2\}$ by Proposition 48. We begin with a few observations.
 Proposition 56.
-If there is any vertex $x\in H$ with $\deg_{H}(x)\leq 5-\deg_{G}(x)$, then $u$ has neighbors $v_{i},v_{j}$ with $R(v_{i},N[v_{j}])\geq 1$.
+If there is any vertex $x\in H$ with $\deg_{H}(x)\leq 5-\deg_{G}(x)$, then $u$ has neighbors $v_{i},v_{j}$ with $S(v_{i},N[v_{j}])\geq 1$.
 Proof.
-If $\deg_{G}(x)=3$, then $x\not\sim A$; to get $\deg_{H}(x)\leq 2$, we must have $\operatorname{codeg}(x,v_{j})\geq 1$ for $v_{j}\in B$, so $R(v_{1},N[v_{j}])\geq 1$.
-If $\deg_{G}(x)=4$, then $|N(x)\cap A|\leq 1$ (else we would branch on $x$ at Line 3); so to get $\deg_{H}(x)\leq 1$, we have either $\operatorname{codeg}(x,v_{j})\geq 2$ for $v_{j}\in B$, or $|N(x)\cap A|\geq 1$.
-If $\operatorname{codeg}(x,v_{j})\geq 2$ for $v_{j}\in B$, then $R(v_{1},N[v_{j}])\geq 1$ (due to subquadratic vertex $x$), while if $\operatorname{codeg}(x,v_{4})=\operatorname{codeg}(x,v_{5})=1,x\sim v_{i}\in A$, then $R(v_{i},N[v_{4}])\geq 1$ (due to subquadratic vertex $x$).
+If $\deg_{G}(x)=3$, then $x\not\sim A$; to get $\deg_{H}(x)\leq 2$, we must have $\operatorname{codeg}(x,v_{j})\geq 1$ for $v_{j}\in B$, so $S(v_{1},N[v_{j}])\geq 1$.
+If $\deg_{G}(x)=4$, then $|N(x)\cap A|\leq 1$ (else we would branch on $x$ at Line 4); so to get $\deg_{H}(x)\leq 1$, we have either $\operatorname{codeg}(x,v_{j})\geq 2$ for $v_{j}\in B$, or $x\sim A$.
+In the former case, we have $S(v_{1},N[v_{j}])\geq 1$ (due to subquadratic vertex $x$).
+In the latter case, if $\operatorname{codeg}(x,v_{4})=\operatorname{codeg}(x,v_{5})=1,x\sim v_{i}\in A$, then $S(v_{i},N[v_{4}])\geq 1$ (due to subquadratic vertex $x$).
 Finally, suppose $\deg_{G}(x)=5$ and $x$ is isolated in $H$.
 By Proposition 48 we have $\operatorname{codeg}(x,u)\leq 2$ and $\operatorname{codeg}(x,v_{4})\leq 2$ and $\operatorname{codeg}(x,v_{5})\leq 2$.
-If $\operatorname{codeg}(x,v_{4})=\operatorname{codeg}(x,v_{5})=2$, then $x$ has a neighbor $v_{i}\in A$, so $R(v_{i},N[v_{4}])\geq 1$ (due to subquadratic vertex $x$).
-Otherwise, if $\operatorname{codeg}(x,v_{j})=2$ for $v_{j}\in B$ and $x\sim v_{i}\in A$, then $R(v_{i},N[v_{j}])\geq 1$ (due to subquadratic vertex $x$) ∎
+If $\operatorname{codeg}(x,v_{4})=\operatorname{codeg}(x,v_{5})=2$, then $x$ has a neighbor $v_{i}\in A$, so $S(v_{i},N[v_{4}])\geq 1$ (due to subquadratic vertex $x$).
+Otherwise, if $\operatorname{codeg}(x,v_{j})=2$ for $v_{j}\in B$ and $x\sim v_{i}\in A$, then $S(v_{i},N[v_{j}])\geq 1$ (due to subquadratic vertex $x$) ∎
 Proposition 57.
 If $s=1$ and $\ell=-1$, then one of the following holds:
-  • There are distinct neighbors $v_{i},v_{j}$ where $G-v_{i}-N[v_{j}]$ is 3-vul.
-  • $d_{45}=1$ and there are neighbors $v_{i},v_{i^{\prime}},v_{j}$ with $R(v_{i},v_{i^{\prime}},N[v_{j}])\geq 1$.
-  • $d_{45}=2$ and the graph $G-N[u]$ has a 5-vertex with a 3-neighbor.
+  • There are neighbors $v_{i},v_{j}$ where $G-v_{i}-N[v_{j}]$ is 3-vul.
+  • $d_{45}=1$ and there are neighbors $v_{i},v_{i^{\prime}},v_{j}$ with $S(v_{i},v_{i^{\prime}},N[v_{j}])\geq 1$.
+  • $d_{45}=2$ and the graph GN[u]G-N[u] has a 5-vertex with a 3-neighbor.
 Proof.
@@
 If $I\not\sim A$, then $\operatorname{surp}_{G-N[v_{4},v_{5}]}(I)=-1$, and hence $\operatorname{surp}_{G-v_{1}-N[v_{4}]}(I\cup\{v_{5}\})\leq 2$, and $G-v_{1}-N[v_{4}]$ is 3-vul via in the indset $J=I\cup\{v_{5}\}$.
-If some $x\in I$ has exactly one neighbor $v_{i}\in A$, then $\operatorname{codeg}(x,v_{j})\geq 2$ for some $v_{j}\in B$ and $R(v_{i},N[v_{j}])\geq 1$ (due to subquadratic vertex $x$).
-If $d_{45}=1$, the only remaining case is that there is some $x\in I$ with two neighbors $v_{i},v_{i^{\prime}}$ in $A$.
-Then $\operatorname{codeg}(x,v_{j})\geq 1$ for some $v_{j}\in B$, and so $R(v_{i},v_{i^{\prime}},N[v_{j}])\geq 1$ as claimed.
-Finally suppose $d_{45}=2$.
+If some $x\in I$ has exactly one neighbor $v_{i}\in A$, then $\operatorname{codeg}(x,v_{j})\geq 2$ for some $v_{j}\in B$ and $S(v_{i},N[v_{j}])\geq 1$ (due to subquadratic vertex $x$).
+Similarly, if there is some $x\in I$ with two neighbors $v_{i},v_{i^{\prime}}$ in $A$, then $\operatorname{codeg}(x,v_{j})\geq 1$ for some $v_{j}\in B$, and so $S(v_{i},v_{i^{\prime}},N[v_{j}])\geq 1$.
+The only remaining case if $d_{45}=2$ and every $x\in I$ has at least two neighbors in $A$.
 If $|N(I)\cap A|<3$, say without loss of generality $v_{1}\not\sim I$, then $\operatorname{surp}_{G-\{v_{2},v_{3}\}-N[v_{4},v_{5}]}(I)=-1$ and hence $\operatorname{surp}_{G-v_{2}-N[v_{4}]}(I\cup\{v_{5}\})\leq 2$; so $G-v_{2}-N[v_{4}]$ is 3-vul (via indset $I\cup\{v_{5}\})$.
-So let us suppose that every vertex in $A$ has at least one neighbor in $I$, and no $x\in I$ has exactly one neighbor in $A$.
-Without loss of generality, we may suppose that $v_{1}\sim x_{1}$ and $v_{2}\sim x_{2}$, and $v_{3}\sim x_{1}$ and $v_{3}\sim x_{2}$, with no other edges between $A$ and $I$.
-Now $v_{3}$ has four 5-neighbors $u,x_{1},x_{2},y$ and $y,v_{3}$ share neighbors $x_{1},x_{2}$.
-So by Proposition 50 we have $y\not\sim v_{3}$.
-If $y\sim v_{i}$ for $v_{i}\in\{v_{1},v_{2}\}$, then $R(v_{i},N[v_{3}])\geq 1$ due to subquadratic vertex $y$.
-Otherwise, if $y\not\sim A$, then in the graph $G-N[u]$, vertices $x_{1},x_{2}$ remain with degree 3 (losing neighbors $v_{1},v_{2},v_{3}$), while $y$ remains with degree 5.
+So we may suppose without loss of generality that the edges between $A$ and $I$ are precisely $(v_{1},x_{1}),(v_{2},x_{2}),(v_{3},x_{1}),(v_{3},x_{2})$.
+Now $v_{3}$ has three 5-neighbors $u,x_{1},x_{2}$ and $y,v_{3}$ share neighbors $x_{1},x_{2}$.
+So by Proposition 51 we have $y\not\sim v_{3}$ (else $v_{3}$ would have four 5-neighbors and its neighborhood would be a non-matching).
+If $y\sim v_{i}$ for $v_{i}\in\{v_{1},v_{2}\}$, then $S(v_{i},N[v_{3}])\geq 1$ due to subquadratic vertex $y$.
+Otherwise, if $y\not\sim A$, then in the graph GN[u]G-N[u], vertices $x_{1},x_{2}$ remain with degree 3 (losing neighbors $v_{1},v_{2},v_{3}$), while $y$ remains with degree 5.
 ∎
-We now analyze the subproblems. When describing each case, we assume that no other vertex ordering could go into an earlier case.
-Case I: $\boldsymbol{\operatorname{shad}(v_{1},v_{2},N[v_{3}])\leq-1.}$ Let $G^{\prime}=G-\{v_{1},v_{2}\}-N[v_{3}]$.
+We now analyze the branching rule and subproblems. When describing each case, we assume that no other vertex ordering could go into an earlier case.
+Case I: $\boldsymbol{\operatorname{shad}(v_{1},v_{2},N[v_{3}])\leq-1.}$ Let $G^{\prime}=G-\{v_{1},v_{2}\}-N[v_{3}]$ and let $J$ be a min-set of $G^{\prime}$.
 We claim that $\operatorname{mindeg}(G^{\prime})\geq 1$.
@@
 Finally, if $\deg(x)=3$, then $x\not\sim\{v_{1},v_{2}\}$ and $\deg_{G^{\prime}}(x)\geq 1$.
-So $\operatorname{surp}_{G^{\prime}}(J)\leq-1$ for an indset $J$, where necessarily $|J|\geq 2$.
-Then $\operatorname{surp}_{G-N[v_{3}]}(J)\leq 1$; since $G$ is in 5BP, this implies $R(N[v_{3}])\geq|J|\geq 2$.
-Our strategy is to split on $v_{3}$, generating subproblems $G-v_{3},G-N[v_{3}]$; these have drops $(0.5,1),(2,7)$ after simplification.
+So necessarily $|J|\geq 2$.
+Then $\operatorname{surp}_{G-N[v_{3}]}(J)\leq 1$, which implies that $S(N[v_{3}])\geq|J|\geq 2$.
+Our strategy is to split on $v_{3}$, generating subproblems $G-v_{3},G-N[v_{3}]$ with drops $(0.5,1),(2,7)$ after simplification.
 We need:
 $e^{-0.5a-b}+e^{-2a-7b}\leq 1$ (5-12)
-Case II: $\boldsymbol{R(v_{1},v_{2},N[v_{3}])\geq 3.}$ Split on $v_{1},v_{2},v_{3}$, generating subproblems $G-N[v_{1}],G-v_{1}-N[v_{2}],G-\{v_{1},v_{2}\}-N[v_{3}],G-\{v_{1},v_{2},v_{3}\}$.
+Case II: $\boldsymbol{S(v_{1},v_{2},N[v_{3}])\geq 3.}$ We split on $v_{1},v_{2},v_{3}$, generating subproblems $G-N[v_{1}],G-v_{1}-N[v_{2}],G-\{v_{1},v_{2}\}-N[v_{3}],G-\{v_{1},v_{2},v_{3}\}$.
 Since we are not in Case I, the first three have drops $(2,5),(2.5,6),(3,7)$ directly.
-Since $G$ is in 5BP, subproblem $G-\{v_{1},v_{2},v_{3}\}$ has drop $(1.5,3)$ directly; we apply Proposition 30 to its 2-vertex $u$ where $\deg(v_{4})+\deg(v_{5})-d_{45}-1\geq 7$.
+Note that $\operatorname{shad}(v_{1},v_{2},v_{3})\geq\operatorname{shad}(v_{1})-2\geq 0$.
+So subproblem $G-\{v_{1},v_{2},v_{3}\}$ has drop $(1.5,3)$ directly, and we can apply Proposition 30 to its 2-vertex $u$ where $\deg(v_{4})+\deg(v_{5})-d_{45}-1\geq 7$.
 So we need:
@@
 Then $\operatorname{surp}_{G-\{v_{1},v_{2}\}-N[v_{4}]}(I\cup\{v_{5}\})\leq 1$.
-So either $\operatorname{shad}(v_{1},v_{2},N[v_{4}])\leq-1$ (which would put us into Case I with alternate vertex ordering $v_{1},v_{2},v_{4}$) or, by Lemma 13 we have $R(v_{1},v_{2},N[v_{4}])\geq|I\cup\{v_{5}\}|\geq 3$, putting us into Case II.
-Our strategy in all subsequent cases is to split on $v_{1},v_{2},v_{3}$, generating subproblems $G_{1}=G-N[v_{1}],G_{2}=G-v_{1}-N[v_{2}],G_{3}=G-\{v_{1},v_{2}\}-N[v_{3}]$, and $G-\{v_{1},v_{2},v_{3}\}$.
-We then apply (A1) to the 2-vertex $u$ in $G-\{v_{1},v_{2},v_{3}\}$, obtaining subproblems $G_{0}=G-N[u]$ and $H=G-\{v_{1},v_{2},v_{3}\}-N[v_{4},v_{5}]$.
+So either $\operatorname{shad}(v_{1},v_{2},N[v_{4}])\leq-1$ (which would put us into Case I with alternate vertex ordering $v_{1},v_{2},v_{4}$) or, by Lemma 13 we have $S(v_{1},v_{2},N[v_{4}])\geq|I\cup\{v_{5}\}|\geq 3$, putting us into Case II.
+Our strategy in all subsequent cases is to split on $v_{1},v_{2},v_{3}$, generating subproblems $G_{1}=G-N[v_{1}],G_{2}=G-v_{1}-N[v_{2}],G_{3}=G-\{v_{1},v_{2}\}-N[v_{3}]$, and $G-\{v_{1},v_{2},v_{3}\}$; we then apply (B-Pair) to the 2-vertex $u$ in $G-\{v_{1},v_{2},v_{3}\}$, to get subproblems $G_{0}=G-N[u]$ and $H=G-\{v_{1},v_{2},v_{3}\}-N[v_{4},v_{5}]$.
 Since we are not in Case I, subproblems $G_{0},G_{1},G_{2},G_{3}$ have drops $(2,5),(2,5),(2.5,6)$ and $(3,7)$ respectively, and the final subproblem $H$ has drop
-$(5.5-d_{45}/2,13-d_{45})\qquad\text{ or }\qquad(5.5-(d_{45}+\ell)/2,13-d_{45}+s)$
-where the latter is obtained if we apply P1 to $I$.
+$(5.5-d_{45}/2,\ \ \ \ \ \ \ \ 13-d_{45})$ if $\ell\geq 0$
+$(5.5-d_{45}/2+\ell/2,13-d_{45}+s)$ if $\ell<0$ (after applying P1)
 Case III: $\boldsymbol{G_{3}}$ is 2-vul, and either $\boldsymbol{G_{1}}$ or $\boldsymbol{G_{2}}$ is 3-vul.
@@
 $e^{-2a-5b}+\max\{e^{-2a-5b}\gamma_{3}+e^{-2.5a-6b},e^{-2a-5b}+e^{-2.5a-6b}\gamma_{3}\}+e^{-3a-7b}+e^{-4.5a-11b}\leq 1$ (5-15)
-Note: After Case IV, we must have $s\geq 1$; for, if $s=0$, then $H$ has an isolated vertex; by Proposition 55 there are then neighbors $v_{i},v_{j}$ with $R(v_{i},N[v_{j}])\geq 1$, putting us into Case IV.
+Note: After Case IV, we must have $s\geq 1$; for, if $s=0$, then $H$ has an isolated vertex; by Proposition 56 there are then neighbors $v_{i},v_{j}$ with $S(v_{i},N[v_{j}])\geq 1$, putting us into Case IV.
 Case V: $\boldsymbol{G_{3}}$ is 4-vul.
-We know $\ell\geq-3+d_{45}$.
-First, suppose $d_{45}=2,\ell=-1,s=1$ exactly.
-So $H$ has drop $(4,12)$.
+Recall that $\ell\geq-3+d_{45}$.
+First, suppose $d_{45}=2,\ell=-1,s=1$, so $H$ has drop $(4,12)$.
 By Proposition 57, there are either vertices $v_{i},v_{j}$ where $G-v_{i}-N[v_{j}]$ is 3-vul (putting us into Case IV), or $G_{0}$ has a 5-vertex with a 3-neighbor.
@@
 $\qquad\qquad+e^{-2a-5b}+e^{-2.5a-6b}+e^{-3a-7b}\gamma_{4}+e^{-4a-12b}\leq 1$ (5-16)
-Otherwise, if $\ell\geq-2+d_{45}$ or $s\geq 2$ or $d_{45}=1$, then $H$ has drop $(4.5,11)$ or $(4,13)$, and we need:
+Otherwise, if $\ell\geq-2+d_{45}$ or $s\geq 2$ or $d_{45}=1$, then $H$ has drop $(4.5,11)$ or $(4,13)$; we need:
 $e^{-2a-5b}+e^{-2a-5b}+e^{-2.5a-6b}+e^{-3a-7b}\gamma_{4}+\max\{e^{-4a-13b},e^{-4.5a-11b}\}\leq 1$ (5-17)
@@
 If $\ell\leq-2$, then $\operatorname{surp}_{G-\{v_{1},v_{2}\}-N[v_{4}]}(I\cup\{v_{5}\})\leq 2$ and $|I|=s-\ell\geq 3$, so $G-\{v_{1},v_{2}\}-N[v_{4}]$ is 4-vul; this would put us into Case V.
-If $s=1,\ell=-1$, then by Proposition 56, there are either neighbors $v_{i},v_{i^{\prime}},v_{j}$ with $R(v_{i},v_{i^{\prime}},N[v_{j}])\geq 1$ (putting us into Case V) or $G-v_{i}-N[v_{j}]$ is 3-vul (putting us into Case IV).
-So either we haver $\ell\geq 0$ or we have $\ell=-1,s\geq 2$, and thus $H$ has drop $(4.5,14)$ or $(5,12)$.
+If $s=1,\ell=-1$, then by Proposition 57, there are either neighbors $v_{i},v_{i^{\prime}},v_{j}$ with $S(v_{i},v_{i^{\prime}},N[v_{j}])\geq 1$ (putting us into Case V) or $G-v_{i}-N[v_{j}]$ is 3-vul (putting us into Case IV).
+So either $\ell\geq 0$ or $\ell=-1,s\geq 2$, and thus $H$ has drop $(5,12)$ or $(4.5,14)$ respectively.
 We need:
@@
 7.4 Line 11: A 5-vertex with fewer than three 4-neighbors
-We suppose we are in Branch5-2 here.
-Let $U$ be the non-empty set of 5-vertices which have fewer than three 4-neighbors.
+We suppose we are in Branch5-2.
+Let $U$ be the non-empty set of 5-vertices with fewer than three 4-neighbors.
 For any vertex $u\in U$, define the induced neighborhood $H_{u}=G[N(u)]$.
@@
 Theorem 58.
-There is some vertex $u\in U$ and an enumeration of its neighbors $v_{1},\dots,v_{5}$ to satisfy one of the following three conditions:
+There is a vertex $u\in U$ and an enumeration of its neighbors $v_{1},\dots,v_{5}$ to satisfy one of the following three conditions:
   1. (I)
@@
   2. (II)
-    $H_{u}$ is a matching and $v_{1}\sim v_{2},\deg(v_{1})=\deg(v_{2})=\deg(v_{3})=5,\deg(v_{5})=4,\deg_{G-v_{3}}(v_{4})=4$ and $\operatorname{codeg}(v_{1},v_{4})=\operatorname{codeg}(v_{2},v_{4})=\operatorname{codeg}(v_{1},v_{5})=\operatorname{codeg}(v_{2},v_{5})=\operatorname{codeg}(v_{3},v_{4})=1$.
+    $H_{u}$ is a matching with an edge $(v_{1},v_{2})$ and possibly an edge $(v_{3},v_{4})$, where $\deg(v_{1})=\deg(v_{2})=\deg(v_{3})=5,\deg(v_{5})=4,\deg_{G-v_{3}}(v_{4})=4$ and $\operatorname{codeg}(v_{1},v_{4})=\operatorname{codeg}(v_{2},v_{4})=\operatorname{codeg}(v_{1},v_{5})=\operatorname{codeg}(v_{2},v_{5})=\operatorname{codeg}(v_{3},v_{4})=1$.
   3. (III)
-    $v_{1}\sim v_{2},v_{1}\sim v_{3}$, and $\deg(v_{1})=5,\deg(v_{4})=\deg(v_{5})=4$ and $\{v_{4},v_{5}\}\not\sim v_{1}$.
+    $v_{1}$ has two neighbors $v_{2},v_{3}$, where $\deg(v_{1})=5,\deg(v_{4})=\deg(v_{5})=4$ and $\{v_{4},v_{5}\}\not\sim v_{1}$.
   4. (IV)
-    $v_{1}\sim v_{2},v_{3}\sim v_{4},\deg(v_{1})=\deg(v_{2})=\deg(v_{3})=\deg(v_{4})=5$.
+    $H_{u}$ is a matching with edges $(v_{1},v_{2})$ and $(v_{3},v_{4})$, where $\deg(v_{1})=\deg(v_{2})=\deg(v_{3})=\deg(v_{4})=5$.
 Furthermore, there are 5-vertices $x_{13},x_{24},y_{2},y_{4}$ and 4-vertices $z_{2},z_{4}$, all distinct from each other and from $N[u]$, such that $v_{2},x_{24},y_{2}$ and $v_{4},x_{24},y_{4}$ are both triangles in $G$ and $x_{2}\sim x_{24}\sim x_{4}$ and $x_{13}\not\sim\{x_{24},v_{2},v_{4}\}$ and $v_{2}\sim z_{2},v_{4}\sim z_{4}$.
 Proof.
-If $H_{u}$ has no edges, it falls into the first category (choose $v_{1},v_{2},v_{3}$ to be 5-neighbors of $u$).
-Suppose $H_{u}$ is not a matching, i.e.
-there is a neighbor $v_{1}$ with $\operatorname{codeg}(v_{1},u)\geq 2$.
-Let $v_{2},v_{3}$ be two shared neighbors of $v_{1},u$ and let $v_{4},v_{5}$ be any other neighbors of $u$.
+If $H_{u}$ has no edges, it falls into the Case I (choose $v_{1},v_{2},v_{3}$ to be 5-neighbors of $u$).
+Suppose $H_{u}$ is not a matching.
+Let $v_{1}$ be a neighbor which shares two other neighbors $v_{2},v_{3}$ with $u$.
+Let $v_{4},v_{5}$ be the other neighbors of $u$.
 Note that $G-\{v_{4},v_{5}\}$ has a kite.
-Since 4-vertices are isolated in $H_{u}$, the only possible structure where we would not use the branching rule at Line 3 must be that $\deg(v_{1})=5,\deg(v_{4})=\deg(v_{5})=4$; since 4-vertices are isolated in $H_{u}$ necessarily $\{v_{1},v_{2},v_{3}\}\not\sim\{v_{4},v_{5}\}$.
-This gives Case III.
+Since 4-vertices are isolated in $H_{u}$, the only possible structure where we would not use the branching rule at Line 3 is $\deg(v_{1})=5,\deg(v_{4})=\deg(v_{5})=4$; since 4-vertices are isolated in $H_{u}$ necessarily $\{v_{1},v_{2},v_{3}\}\not\sim\{v_{4},v_{5}\}$.
+This is Case III.
 Suppose $H_{u}$ is a matching with exactly one edge $(v_{1},v_{2})$, and an additional 5-vertex $v_{3}$.
-If it has any other 5-neighbor $t$, then $R(v_{3},t)\geq 1$ (due to 3-triangle $u,v_{1},v_{2}$) and we would branch at Line 9.
+If $u$ has any other 5-neighbor $t$, then $S(v_{3},t)\geq 1$ (due to 3-triangle $u,v_{1},v_{2}$) and we would branch at Line 9.
 So $u$ must have two 4-neighbors $v_{4},v_{5}$.
-Then Observation 54 implies $\operatorname{codeg}(v_{1},v_{4})=\operatorname{codeg}(v_{2},v_{4})=\operatorname{codeg}(v_{1},v_{5})=\operatorname{codeg}(v_{2},v_{5})=\operatorname{codeg}(v_{3},v_{4})=1$.
-This gives Case II.
+Then Observation 55 implies $\operatorname{codeg}(v_{1},v_{4})=\operatorname{codeg}(v_{2},v_{4})=\operatorname{codeg}(v_{1},v_{5})=\operatorname{codeg}(v_{2},v_{5})=\operatorname{codeg}(v_{3},v_{4})=1$.
+This is Case II.
 Thus, for the remainder of the proof, we suppose that, for every vertex $u\in U$, the graph $H_{u}$ is a matching with exactly two edges, where the one unmatched neighbor has degree 4.
 Consider some $u\in U$, where $H_{u}$ has edges $(v_{1},v_{2}),(v_{3},v_{4})$ and has a 4-vertex $v_{5}$.
-By Observation 52 we automatically have $\operatorname{codeg}(v_{1},v_{5})=\operatorname{codeg}(v_{2},v_{5})=1$.
+By Observation 53 we automatically have $\operatorname{codeg}(v_{1},v_{5})=\operatorname{codeg}(v_{2},v_{5})=1$.
 We claim that $\operatorname{codeg}(v_{3},v_{4})=1$.
@@
 Now consider forming a bipartite graph $J$ with left-nodes $v_{1},v_{2}$ and right-nodes $v_{3},v_{4}$, with an edge $(v_{i},v_{j})$ if $\operatorname{codeg}_{G}(v_{i},v_{j})=2$.
-If any node has degree zero in $J$, we may suppose without loss of generality that $\deg_{J}(v_{4})=0$ in which case $\operatorname{codeg}(v_{1},v_{4})=\operatorname{codeg}(v_{2},v_{4})=1$ giving us Case II.
+If any node has degree zero in $J$, we may suppose without loss of generality it is $v_{4}$, in which case case $\operatorname{codeg}(v_{1},v_{4})=\operatorname{codeg}(v_{2},v_{4})=1$ giving us Case II.
 Otherwise, if every node in $J$ has at least one neighbor, then by Hall’s theorem $J$ has a perfect matching; by swapping vertices as needed we assume this matching is $(v_{1},v_{3}),(v_{2},v_{4})$, that is, $v_{1},v_{3}$ share a neighbor $x_{13}\neq u$ and $v_{2},v_{4}$ share a neighbor $x_{24}\neq u$.
@@
 Clearly $u\not\sim\{x_{13},x_{24}\}$.
-By Observation 52, we have $\operatorname{codeg}(x_{13},u)\leq 2$ and $\operatorname{codeg}(x_{24},u)\leq 2$, so $x_{13}\not\sim\{v_{2},v_{4}\}$ and $x_{24}\not\sim\{v_{1},v_{3}\}$.
+By Observation 53, we have $\operatorname{codeg}(x_{13},u)\leq 2$ and $\operatorname{codeg}(x_{24},u)\leq 2$, so $x_{13}\not\sim\{v_{2},v_{4}\}$ and $x_{24}\not\sim\{v_{1},v_{3}\}$.
 In particular, $x_{13}\neq x_{24}$.
@@
 By our hypothesis on the structure of $U$, it must be that $v_{2}$ has an edge $(x_{24},y_{2})$ in its neighborhood and has a 4-neighbor $z_{2}$, and likewise $v_{4}$ has an edge $(x_{24},y_{4})$ in its neighborhood and has a 4-neighbor $z_{4}$.
-By Observation 52, we have $\operatorname{codeg}(x_{24},u)\leq 2$, and so $y_{2},y_{4}\notin N[u]$.
+By Observation 53, we have $\operatorname{codeg}(x_{24},u)\leq 2$, and so $y_{2},y_{4}\notin N[u]$.
 Also $y_{2}\neq y_{4}$, as otherwise the vertex $y=y_{2}=y_{4}$ would have $\operatorname{codeg}(v_{2},v_{4})\geq 3$, sharing neighbors $u,x_{24},y$.
@@
 Split on $v_{1},v_{2},v_{3}$, getting subproblems $G_{1}=G-N[v_{1}],G_{2}=G-v_{1}-N[v_{2}],G_{3}=G-\{v_{1},v_{2}\}-N[v_{3}],G_{4}=G-\{v_{1},v_{2},v_{3}\}$.
-By Observation 52, these have drops $(2,5),(2.5,6),(3,7),(1,3)$ directly.
-In $G_{4}$, we apply Proposition 30 to 2-vertex $u$, where by Observation 52 and Observation 54 we have $r=\deg_{G_{4}}(v_{4})+\deg_{G_{4}}(v_{5})-\operatorname{codeg}_{G_{4}}(v_{4},v_{5})-1\geq 6$ irrespective of the degrees of $v_{4},v_{5}$.
+By Observation 53, these have drops $(2,5),(2.5,6),(3,7),(1,3)$ directly.
+In $G_{4}$, we apply Proposition 30 to 2-vertex $u$, where by Observation 53 and Observation 55 we have $r=\deg_{G_{4}}(v_{4})+\deg_{G_{4}}(v_{5})-\operatorname{codeg}_{G_{4}}(v_{4},v_{5})-1\geq 6$ irrespective of the degrees of $v_{4},v_{5}$.
 We need:
@@
 Case IIa: $\boldsymbol{\operatorname{shad}(v_{3},N[v_{1},v_{4}])\geq 1}$.
-Split on $v_{3},v_{5}$ and apply (A2) to vertex $v_{1}$ in $G-\{v_{3},v_{5}\}$.
+Split on $v_{3},v_{5}$ and apply (B-Fun2) to vertex $v_{1}$ in $G-\{v_{3},v_{5}\}$.
 This generates subproblems $G_{1}=G-N[v_{3}],G_{2}=G-v_{3}-N[v_{5}],G_{3}=G-\{v_{3},v_{5}\}-N[v_{1},v_{4}],G_{4}=G-\{v_{1},v_{3},v_{5}\}$.
-Since $G$ is in 5BP, and $\deg_{G}(v_{3})=5,\deg_{G}(v_{5})=4,v_{3}\not\sim v_{5}$, subproblems $G_{1},G_{2}$ both have drop $(2,5)$ directly.
-Case IIa gives $\operatorname{minsurp}(G_{3})\geq 0$; since $\operatorname{codeg}(v_{1},v_{4})=1$, it has drop $(4,10)$.
+Since we are after Line 1, and $\deg_{G}(v_{3})=5,\deg_{G}(v_{5})=4,v_{3}\not\sim v_{5}$, the subproblems $G_{1},G_{2}$ both have drop $(2,5)$ directly.
+Since $\operatorname{codeg}(v_{1},v_{4})=1$, and since we are in Case IIa with $\operatorname{minsurp}(G_{3})\geq 0$, the subproblem $G_{3}$ has drop $(4,10)$.
 Subproblem $G_{4}$ has drop $(1.5,3)$ directly, and we also apply Proposition 30 to 2-vertex $u$ in $G_{4}$, noting that $\deg_{G_{4}}(v_{2})+\deg_{G_{4}}(v_{4})-1-\operatorname{codeg}_{G_{4}}(v_{2},v_{4})=4+4-1-1=6$.
@@
 Case IIb: $\boldsymbol{\operatorname{shad}(v_{3},N[v_{1},v_{4}])\leq 0}$.
-Split on $v_{3},v_{4}$, then apply (A2) to vertex $v_{1}$ in $G-\{v_{3},v_{4}\}$.
+Split on $v_{3},v_{4}$, then apply (B-Fun2) to vertex $v_{1}$ in $G-\{v_{3},v_{4}\}$.
 We get subproblems $G_{1}=G-N[v_{3}],G_{2}=G-v_{3}-N[v_{4}],G_{3}=G-\{v_{3},v_{4}\}-N[v_{1},v_{5}],G_{4}=G-\{v_{1},v_{3},v_{4}\}$.
-Since $G$ is in 5BP, subproblems $G_{1},G_{2}$ both have drop $(2,5)$.
+Subproblems $G_{1},G_{2}$ both have drop $(2,5)$.
 In subproblem $G_{2}$, consider vertex $v_{1}$; it has degree 4 since $\operatorname{codeg}(v_{1},v_{4})=1$ and $v_{1}\not\sim v_{3}$.
 Case IIb gives $\operatorname{shad}_{G_{2}}(v_{1})=\operatorname{shad}(v_{3},N[v_{1},v_{4}])\leq 0$, that is, $v_{1}$ is a blocked 4-vertex in $G_{2}$.
-Also, $\operatorname{minsurp}(G_{2})\geq\operatorname{shad}(N[v_{5}])-1\geq 0$.
+Also, $\operatorname{minsurp}(G_{2})\geq\operatorname{shad}(N[v_{4}])-1\geq 0$.
 So, either $S(G_{2})\geq 1$ or we can apply Proposition 25 to $G_{2}$.
@@
 Case IV: Split on the independent 5-vertices $u,x_{13},x_{24}$, generating subproblems $G_{1}=G-N[u],G_{2}=G-u-N[x_{13}],G_{3}=G-\{u,x_{13}\}-N[x_{24}],G_{4}=G-\{x_{14},x_{23},u\}$.
-Note that $\operatorname{minsurp}(G_{3})\geq\operatorname{shad}(N[x_{24}])-2\geq 0$ and $\operatorname{minsurp}(G_{4})\geq\operatorname{shad}(x_{14})-2\geq 0$.
+By Observation 53, we have $\operatorname{minsurp}(G_{2})\geq\operatorname{shad}(N[x_{13}])-1\geq 1$ and $\operatorname{minsurp}(G_{3})\geq\operatorname{shad}(N[x_{24}])-2\geq 0$ and $\operatorname{minsurp}(G_{4})\geq\operatorname{shad}(x_{14})-2\geq 0$.
 So these have drops $(2,5),(2.5,6),(3,7),(1.5,3)$ directly.
-Consider graph $G_{2}$.
-By Observation 52, we have $\operatorname{minsurp}(G_{2})\geq\operatorname{shad}(N[x_{13}])-1\geq 1$, in particular $\operatorname{mindeg}(G_{2})\geq 2$.
-Also note that $v_{4}\not\sim v_{4}$ and both $z_{2},z_{4}$ remain in the graph $G_{2}$, since $z_{2},z_{4}$ are 4-vertices which already have a 5-neighbors $v_{2}$ or $v_{4}$ respectively.
+Note that both $z_{2},z_{4}$ remain in the graph $G_{2}$, since $z_{2},z_{4}$ are 4-vertices which already have a 5-neighbors $v_{2}$ or $v_{4}$ respectively.
 We claim that $S(G_{2})\geq 2$; there are a few different cases depending on $|N(x_{13})\cap\{y_{2},y_{4}\}|$.
 First, suppose $x_{13}\sim y_{2}$ and $x_{13}\sim y_{4}$; then, in the graph $G_{2}$, vertices $v_{2}$ and $v_{4}$ have degree two (losing neighbors $v_{1},u,y_{2}$ and $v_{3},u,y_{4}$ respectively).
-So by Observation 11 we immediately have $S(G_{2})\geq 2$.
+So by Proposition 10 we immediately have $S(G_{2})\geq 2$.
 Next, suppose $|N(x_{13})\cap\{y_{2},y_{4}\}|=1$, say without loss of generality $x_{13}\sim y_{2},x_{13}\not\sim y_{4}$.
 In this case, we apply (P2) to the 2-vertex $v_{2}$ in $G_{2}$; in the resulting graph $G_{2}^{\prime}$, the vertices $z_{2},x_{24}$ get merged into a new vertex $z^{\prime}$.
-In $G_{2}^{\prime}$, the vertex $v_{4}$ has neighbors $z_{4},z^{\prime},y_{4}$ where $z^{\prime}\sim y_{4}$ in $G_{2}^{\prime}$ since $x_{24}\sim y_{4}$ in $G_{2}$.
-So we can apply (P3) to 3-triangle $v_{2}^{\prime}$, yielding $S(G_{2}^{\prime})\geq 1$ and hence $S(G_{2})\geq 2$.
+We then have $N_{G_{2}^{\prime}}(v_{4})=\{z_{4},z^{\prime},y_{4}\}$ where $z^{\prime}\sim y_{4}$, since $x_{24}\sim y_{4}$ in $G_{2}$.
+So we can apply (P3) to the 3-triangle on $v_{4}$, yielding $S(G_{2}^{\prime})\geq 1$ and hence $S(G_{2})\geq 2$.
 Finally suppose $x_{13}\not\sim\{y_{2},y_{4}\}$; then in $G_{2}$, the vertex $v_{2}$ has degree 3 with a triangle $v_{1},y_{2}$.
@@
 So $S(G_{2}^{\prime})\geq 1$ and $S(G_{2})\geq 2$.
-Consider graph $G_{3}$; here, vertex $v_{1}$ loses neighbors $u,x_{13},v_{2}$ and likewise $v_{3}$ loses neighbors $u,x_{13},v_{4}$).
-Moreover $\operatorname{minsurp}(G_{3})\geq\operatorname{minsurp}(G_{24})-2\geq 0$.
-Thus, by Observation 11 we have $S(G_{3})\geq 2$.
+Consider graph $G_{3}$; here, vertex $v_{1}$ loses neighbors $u,x_{13},v_{2}$ and likewise $v_{3}$ loses neighbors $u,x_{13},v_{4}$. Thus, by Proposition 10 we have $S(G_{3})\geq 2$.
 So $G_{2}$ has drop $(2.5,8)$ after simplification and $G_{3}$ has drop $(3,9)$ after simplification. We need:
@@
 Theorem 59.
-Branch6$(G,k)$ has measure $a_{6}\mu+b_{6}k$ for $a_{6}=0.166944,b_{6}=0.166214$.
+Branch6(G,k)(G,k) has measure $a_{6}\mu+b_{6}k$ for $a_{6}=0.166944,b_{6}=0.166214$.
 Combined with Lemma 15, this gives the $O^{*}(1.25214^{k})$ algorithm for degree-6 graphs. The first algorithm is very simple, and is targeted toward the regime where degree-6 vertices are sparse.
-Simplify $G$ If $\operatorname{maxdeg}(G)\leq 5$, then run the cheaper of Branch5-3 or the MaxIS-5 algorithm. Apply Section 5 to a vertex of degree at least $6$
-Algorithm 8 Function Branch6-1$(G,k)$
+1Simplify $G$
+2If $\operatorname{maxdeg}(G)\leq 5$, then run the cheaper of Branch5-3 or the MaxIS-5 algorithm.
+3Apply Figure 6 to a vertex of degree at least 66
+Algorithm 8 Function Branch6-1(G,k)(G,k)
 Proposition 60.
@@
 Proof.
-Given the potential branch-seqs from Section 5, we have the following constraints:
+Given the potential branch-seqs from Figure 6, we have the following constraints:
 $\max\{e^{-a-3b}+e^{-a-5b},e^{-0.5a-2b}+e^{-2a-5b},e^{-0.5a-b}+e^{-2.5a-6b}\}\leq 1$ (6-1)
@@
 The more advanced algorithms try to branch on multiple 6-vertices in order to get bonus simplifications. These algorithms differ only in a few small steps, so we we describe them jointly.
-1 Apply miscellaneous branching rules to ensure $G$ has nice structure e.g.
-max degree 6.
-Split on a 6-vertex $u$ where $G-N[u]$ is 2-vul.
-Branch on vertices $u,v$ with $\deg_{G}(u)=6,\deg_{G-u}(v)\geq 5$, and either $R(u,v)\geq 1$ or $\operatorname{shad}(u,N[v])\leq 0$.
-Branch on non-adjacent 6-vertices $x_{1},x_{2},x_{3}$ with $R(x_{1},x_{2},x_{3})\geq 1$.
-Branch on a 6-vertex $u$ with six 6-neighbors, or five 6-neighbors and one 5-neighbor.
-if $L=2$ then
-2        Branch on non-adjacent 6-vertices $x_{1},x_{2}$ with $R(x_{1},N[x_{2}])\geq 1$.
-Branch on a 5-vertex $u$ which has two 6-neighbors and three 5-neighbors,            UNLESS each of these 6-neighbors itself has at least four 5-neighbors.
-Run the cheaper of Branch6-$(L-1)(G,k)$ or the MaxIS algorithm
-Algorithm 9 Function Branch6-$L$$(G,k)$ for $L=2,3$
+1
+2Apply miscellaneous branching rules to ensure $G$ has nice structure.
+3Split on a 6-vertex $u$ where GN[u]G-N[u] is 2-vul.
+4Branch on vertices $u,v$ with $\deg_{G}(u)=6,\deg_{G-u}(v)\geq 5$ and $S(u,v)+S(u,N[v])\geq 1$.
+5Branch on non-adjacent 6-vertices $x_{1},x_{2},x_{3}$ with $S(x_{1},x_{2},x_{3})\geq 1$.
+6Branch on a 6-vertex $u$ with six 6-neighbors, or five 6-neighbors and one 5-neighbor.
+7if $L=2$ then
+8
+9  Branch on non-adjacent 6-vertices $x_{1},x_{2}$ with $S(x_{1},N[x_{2}])\geq 1$.
+10  Branch on a 5-vertex $u$ which has two 6-neighbors and three 5-neighbors, UNLESS each of these 6-neighbors itself has at least four 5-neighbors.
+11
+12Run the cheaper of Branch6-(L1)(G,k)(L-1)(G,k) or the MaxIS algorithm
+Algorithm 9 Function Branch6-LL(G,k)(G,k) for $L=2,3$
 We will show the following:
@@
 8.1 Regularizing the graph and structural properties
-There is a grab bag of isolated local obstructions that we will remove. We say that $G$ is in 6-base position (or 6BP for short) if the following conditions hold:
+In Line 1, we will ensure that the following conditions hold:
   • $G$ is simplified.
   • $\operatorname{maxdeg}(G)\leq 6$.
-  • Every 6-vertex $u$ has $\operatorname{shad}(N[u])\geq 2$ and $R(u)=0$.
-  • Every 5-vertex $u$ has $\operatorname{shad}(N[u])\geq 1$ and $R(u)=0$.
-  • There are no vertices $u,v$ with $\operatorname{codeg}(u,v)\geq 2$ and $R(u,v)+R(N(u)\cap N(v))+\operatorname{codeg}(u,v)\geq 5$.
-Observation 61.
-  • If $G$ is in 5BP, then any 6-vertex $u$ and vertex $v\not\sim u$ have $\operatorname{codeg}(u,v)\leq\deg(v)-3$.
-  • If $G$ is in 5BP, then any 5-vertex $u$ and vertex $v\not\sim u$ have $\operatorname{codeg}(u,v)\leq 2$.
+  • Every 6-vertex $u$ has $\operatorname{shad}(N[u])\geq 2$ and $S(u)=0$.
+  • Every 5-vertex $u$ has $\operatorname{shad}(N[u])\geq 1$ and $S(u)=0$.
+  • There are no vertices $u,v$ with $\operatorname{codeg}(u,v)\geq 2$ and $S(u,v)+S(N(u)\cap N(v))+\operatorname{codeg}(u,v)\geq 5$.
+Observation 62.
+  • After Line 1, any 6-vertex $u$ and vertex $v\not\sim u$ have $\operatorname{codeg}(u,v)\leq\deg(v)-3$.
+  • After Line 1, any 5-vertex $u$ and vertex $v\not\sim u$ have $\operatorname{codeg}(u,v)\leq 2$.
   • After Line 3, any vertices $u,v$ with $\deg(u)=6$ and $\deg_{G-u}(v)\geq 5$ have $\operatorname{shad}(u,v)\geq 2$ and $\operatorname{shad}(u,N[v])\geq 1$.
@@
 In the second case, we may suppose that also $\deg(v)\leq 5$ (else it is covered by the first case); then the analysis is the same as Proposition 48.
-In the third case, if $\operatorname{shad}(u,v)\leq 1$ or $\operatorname{shad}(u,N[v])\leq 0$, we would branch at Line 3.
+In the third case, if $\operatorname{shad}(u,v)\leq 1$ or $\operatorname{shad}(u,N[v])\leq 0$, then we would have $S(u,v)\geq 1$ or $S(u,N[v])\geq 1$ (note that $\operatorname{shad}(u,N[v])\geq\operatorname{shad}(N[v])-1\geq 0$), and we would branch at Line 3.
 In the final case, we have $\operatorname{shad}(x_{1},x_{2},x_{3})\geq\operatorname{shad}(x_{1},x_{2})-1\geq 1$.
@@
 Let $v_{4},v_{5}$ be the other two neighbors of $u$.
-They must have degree at least 4, else $R(u)\geq 1$ contradicting that $G$ is in 6BP.
-If $v_{1},v_{2},v_{3}$ are independent, then $R(v_{1},v_{2},v_{3})\geq 1$ due to 2-vertex $u$, and we would apply Line 4 to $v_{1},v_{2},v_{3}$.
+They must have degree at least 4, else $S(u)\geq 1$ and we would have branched at Line 1.
+If $v_{1},v_{2},v_{3}$ are independent, then $S(v_{1},v_{2},v_{3})\geq 1$ due to 2-vertex $u$, and we would apply Line 4 to $v_{1},v_{2},v_{3}$.
 So we may suppose without loss of generality that $v_{1}\sim v_{2}$.
 Now suppose for contradiction that $\deg(v_{4})\geq 5$.
-If $v_{3}\sim v_{4}$, then $R(v_{1},v_{2})\geq 1$ due to 3-triangle $u,v_{3},v_{4}$, and we would branch on $v_{1},v_{2}$ at Line 3.
-Otherwise, if $v_{3}\not\sim v_{4}$, then $R(v_{3},v_{4})\geq 1$ due to 3-triangle $u,v_{1},v_{2}$, and we would branch on $v_{3},v_{4}$ at Line 3.
+If $v_{3}\sim v_{4}$, then $S(v_{1},v_{2})\geq 1$ due to 3-triangle $u,v_{3},v_{4}$, and we would branch on $v_{1},v_{2}$ at Line 3.
+Otherwise, if $v_{3}\not\sim v_{4}$, then $S(v_{3},v_{4})\geq 1$ due to 3-triangle $u,v_{1},v_{2}$, and we would branch on $v_{3},v_{4}$ at Line 3.
 ∎
@@
 Also let $A_{52}^{\prime}\subseteq A_{52}$ be the set of 5-vertices with two 6-neighbors and at least one 4-neighbor.
-By double-counting edges between $V_{5}$ and $V_{6}$, we have:
+By counting edges between $V_{5}$ and $V_{6}$, we have:
 $|A_{51}|+2|A_{52}|+3|A_{53}|=|A_{61}|+2|A_{62}|+3|A_{63}|+4|A_{64}|+5|A_{65}|+6|A_{66}|$ (7)
@@
 Now consider the edges between $V_{4}$ and $V_{5}\cup V_{6}$.
-Each vertex in $A_{60}\cup A_{61}$ contributes one such edge, each vertex in $A_{52}^{\prime}$ contributes at least one edge, each vertex in $A_{53}$ contributes at least two edges by Proposition 62, and each vertex in $V_{4}$ has at most four such edges.
+Each vertex in $A_{60}\cup A_{61}$ contributes one such edge, each vertex in $A_{52}^{\prime}$ contributes at least one edge, each vertex in $A_{53}$ contributes at least two edges, and each vertex in $V_{4}$ has at most four such edges.
 So we get:
@@
 If a vertex $u$ of degree 5 or 6 has $\operatorname{shad}(N[u])\leq 5-\deg(u)$, we apply Lemma 26.
-If a vertex $u$ of degree 5 or 6 has $R(u)\geq 1$, we split on $u$ with branch-seq $[(0.5,2),(2,5)]$.
-If a vertex has degree at least 7, we apply Section 5.
-If there are vertices $u,v$ with $R(u,v)+R(N(u)\cap N(v))+\operatorname{codeg}(u,v)\geq 5$, we apply (A3).
+If a vertex $u$ of degree 5 or 6 has $S(u)\geq 1$, we split on $u$ with branch-seq $[(0.5,2),(2,5)]$.
+If a vertex has degree at least 7, we apply Figure 6.
+If there are vertices $u,v$ with $S(u,v)+S(N(u)\cap N(v))+\operatorname{codeg}(u,v)\geq 5$, we apply (B-Shared).
 Over all cases, we get branch-seq $[(1,3),(1,4)]$, $[(1,2),(1,5)]$, or $[(0.5,2),(2,5)]$. We need:
 $\max\{e^{-a-3b}+e^{-a-4b},e^{-a-2b}+e^{-a-5b},e^{-0.5a-2b}+e^{-2a-5b}\}\leq 1$ (6-3)
-Line 2. Since $G$ is in 6BP, subproblems $G-u$ and $G-N[u]$ have drop $(0.5,1)$ and $(2.5,6)$ directly, and we can apply Proposition 31 to the latter problem. We need:
+Line 2. After Line 1, subproblems GuG-u and GN[u]G-N[u] have drop $(0.5,1)$ and $(2.5,6)$ directly, and we can apply Proposition 31 to the latter problem. We need:
 $e^{-0.5a-b}+e^{-2.5a-6b}\gamma_{2}\leq 1$ (6-4)
-Line 3. Split on $u$, and in subproblem $G-u$, either apply Lemma 26 or split on $v$. Since $G$ is in 6BP, subproblems $G-N[u],G-u$ have drops $(2.5,6),(0.5,1)$ directly and are simplified. We need:
+Line 3. Split on $u$, and in subproblem GuG-u, either apply Lemma 26 or split on $v$. After Line 1, subproblems GN[u],GuG-N[u],G-u have drops $(2.5,6),(0.5,1)$ directly and are simplified. We need:
 $e^{-2.5a-6b}+e^{-0.5a-b}\max\{e^{-b},e^{-a-3b}+e^{-a-4b},e^{-0.5a-2b}+e^{-2a-5b}\}\leq 1$ (6-5)
@@
 Split on $x_{1},x_{2},x_{3}$, getting subproblems $G_{1}=G-N[x_{1}],G_{2}=G-x_{1}-N[x_{2}],G_{3}=G-\{x_{1},x_{2}\}-N[x_{3}],G_{4}=G-\{x_{1},x_{2},x_{3}\}$.
-By Observation 61 we have $\operatorname{minsurp}(G_{1})\geq 2,\operatorname{minsurp}(G_{2})\geq 1,\operatorname{minsurp}(G_{3})\geq\operatorname{shad}(N[x_{3}])-2\geq 0$ and $\operatorname{minsurp}(G_{4})\geq\operatorname{shad}(x_{1},x_{2})-1\geq 1$.
+By Observation 62 we have $\operatorname{minsurp}(G_{1})\geq 2,\operatorname{minsurp}(G_{2})\geq 1,\operatorname{minsurp}(G_{3})\geq\operatorname{shad}(N[x_{3}])-2\geq 0$ and $\operatorname{minsurp}(G_{4})\geq\operatorname{shad}(x_{1},x_{2})-1\geq 1$.
 Because $S(G_{4})\geq 1$, these subproblems $G_{1},G_{2},G_{3},G_{4}$ have drops $(2.5,6),(3,7),(3.5,8),(2,5)$.
@@
 Line 7.
-Since $G$ is in 6BP, subproblems $G-N[x_{1}]$ and $G-x_{1}-N[x_{2}]$ and $G-\{x_{1},x_{2}\}$ have drops $(2.5,6),(3,7),(1,2)$ directly, and $G-x_{1}-N[x_{2}]$ has drop $(3,8)$ after simplification.
+After Line 1, subproblems $G-N[x_{1}]$ and $G-x_{1}-N[x_{2}]$ and $G-\{x_{1},x_{2}\}$ have drops $(2.5,6),(3,7),(1,2)$ directly, and $G-x_{1}-N[x_{2}]$ has drop $(3,8)$ after simplification.
 We need:
@@
 Proof.
-First observe that $H$ cannot have any triangle $v_{1},v_{2},v_{3}$; for, in that case, we would have $R(v_{4},v_{5})\geq 1$ where $v_{4},v_{5}$ are any other 6-neighbors of $u$; we would branch on $v_{4},v_{5}$ at Line 3.
+First observe that $H$ cannot have any triangle $v_{1},v_{2},v_{3}$; for, in that case, we would have $S(v_{4},v_{5})\geq 1$ where $v_{4},v_{5}$ are any other 6-neighbors of $u$; we would branch on $v_{4},v_{5}$ at Line 3.
 If $H$ has no edges, then let $v_{1},\dots,v_{5}$ be any five 6-neighbors of $u$, and we get Case I.
-If $H$ is a matching with exactly one edge $(v_{1},v_{2})$, with three other 6-neighbors $v_{3},v_{4},v_{5}$, we would have $R(v_{3},v_{4},v_{5})\geq 1$ due to the 3-triangle $u,v_{1},v_{2}$ and we would branch Line 4.
+If $H$ is a matching with exactly one edge $(v_{1},v_{2})$, with three other 6-neighbors $v_{3},v_{4},v_{5}$, we would have $S(v_{3},v_{4},v_{5})\geq 1$ due to the 3-triangle $u,v_{1},v_{2}$ and we would branch Line 4.
 If $H$ is a matching with exactly two edges, suppose without loss of generality they are $(v_{1},v_{2}),(v_{3},v_{4})$ where $\deg(v_{4})=6$.
-If $\deg(v_{5})=\deg(v_{6})=6$, then again have $R(v_{4},v_{5},v_{6})\geq 1$ due to 3-triangle $u,v_{1},v_{2}$, and we would branch at Line 4.
+If $\deg(v_{5})=\deg(v_{6})=6$, then again have $S(v_{4},v_{5},v_{6})\geq 1$ due to 3-triangle $u,v_{1},v_{2}$, and we would branch at Line 4.
 So suppose without loss of generality that $\deg(v_{5})=5$, in which case the ordering $v_{1},\dots,v_{6}$ satisfies Case II.
 If $H$ is a matching with exactly three edges, we can arrange them as $(v_{1},v_{2}),(v_{3},v_{4}),(v_{5},v_{6})$, where all vertices except $v_{2}$ have degree 6. This also satisfies Case II.
-Next suppose $H$ is not a matching.
-First, suppose some $v_{4}\in H$ has neighbors $v_{5},v_{6}\in H$ where one of the vertices $v_{4},v_{5},v_{6}$ has degree 5.
-Since $H$ has no triangle, we order the other vertices as $v_{1},v_{2},v_{3}$ where $v_{2}\not\sim v_{3}$, satisfying Case III.
-Finally, suppose $v_{4}\in H$ has neighbors $v_{5},v_{6}\in H$ where $\deg(v_{4})=\deg(v_{5})=\deg(v_{6})=6$.
+Next suppose $H$ is not a matching, and some vertex $v_{4}\in H$ has neighbors $v_{5},v_{6}\in H$ where the other neighbors $v_{1},v_{2},v_{3}$ all have degree 6.
+Since $H$ has no triangle, we order these vertices so that $v_{2}\not\sim v_{3}$, satisfying Case III.
+Finally, suppose $H$ is not a matching and some $v_{4}\in H$ has neighbors $v_{5},v_{6}\in H$ where $\deg(v_{4})=\deg(v_{5})=\deg(v_{6})=6$.
 The induced subgraph $H^{\prime}=G[\{v_{1},v_{2},v_{3}\}]$ must be a matching, as otherwise it would be covered by the argument in the preceding paragraph.
@@
 Split on $v_{1},\dots,v_{4}$, generating subproblems $G_{1}=G-N[v_{1}],G_{2}=G-v_{1}-N[v_{2}],G_{3}=G-\{v_{1},v_{2}\}-N[v_{3}],G_{4}=G-\{v_{1},v_{2},v_{3}\}-N[v_{4}],G_{5}=G-\{v_{1},v_{2},v_{3},v_{4}\}$.
-By Observation 61, we have $\operatorname{minsurp}(G_{1})\geq 2$, $\operatorname{minsurp}(G_{2})\geq 1$, $\operatorname{minsurp}(G_{3})\geq\operatorname{shad}(N[v_{3}])-2\geq 0$, $\operatorname{minsurp}(G_{4})\geq\operatorname{shad}(N[v_{4}])-3\geq-1$ and $\operatorname{minsurp}(G_{5})\geq\operatorname{shad}(v_{1},v_{2},v_{3})-1\geq 1$.
+By Observation 62, we have $\operatorname{minsurp}(G_{1})\geq 2$, $\operatorname{minsurp}(G_{2})\geq 1$, $\operatorname{minsurp}(G_{3})\geq\operatorname{shad}(N[v_{3}])-2\geq 0$, $\operatorname{minsurp}(G_{4})\geq\operatorname{shad}(N[v_{4}])-3\geq-1$ and $\operatorname{minsurp}(G_{5})\geq\operatorname{shad}(v_{1},v_{2},v_{3})-1\geq 1$.
 So $G_{1},G_{2},G_{3},G_{4},G_{5}$ have drop $(2.5,6),(3,7),(3.5,8),(3.5,9),(2,4)$ directly.
@@
 $e^{-2.5a-6b}+e^{-3a-7b}+e^{-3.5a-8b}\gamma_{3}+e^{-3.5a-9b}+e^{-2a-5b}\leq 1$ (6-10)
-Case Ib: Split on $v_{1},v_{2},v_{3},v_{4}$, and apply (A1) to vertex $u$ in the subproblem $G-\{v_{1},v_{2},v_{3},v_{4}\}$.
+Case Ib: Split on $v_{1},v_{2},v_{3},v_{4}$, and apply (B-Pair) to vertex $u$ in the subproblem $G-\{v_{1},v_{2},v_{3},v_{4}\}$.
 This generates subproblems $G_{1}=G-N[v_{1}],G_{2}=G-v_{1}-N[v_{2}],G_{3}=G-\{v_{1},v_{2}\}-N[v_{3}],G_{4}=G-\{v_{1},v_{2},v_{3}\}-N[v_{4}],G_{0}=G-N[u]$ and $G_{5}=G-\{v_{1},v_{2},v_{3},v_{4}\}-N[v_{5},v_{6}]$.
-Since $G$ is 6BP, $G_{0},G_{1},G_{2},G_{3}$ have drops $(2.5,6),(2.5,6),(3,7),(3.5,8)$ directly.
+Since we are after Line 1, subproblems $G_{0},G_{1},G_{2},G_{3}$ have drops $(2.5,6),(2.5,6),(3,7),(3.5,8)$ directly.
 We must have $\operatorname{minsurp}(G_{4})\geq 0$, as otherwise $\operatorname{shad}(v_{1},v_{2},N[v_{4}])\leq 0$, which would put us into Case Ia (with swapping the order of $v_{3}/v_{4}$).
 So $G_{4}$ has drop $(4,9)$.
-For subproblem $G_{5}$, let $d=\deg_{G-N[v_{5}]}(v_{6})$; by Observation 61 we have $d\geq 3$.
+For subproblem $G_{5}$, let $d=\deg_{G-N[v_{5}]}(v_{6})$; by Observation 62 we have $d\geq 3$.
 We claim that $\operatorname{minsurp}(G_{5})\geq 2-d$.
 For, consider an indset $I$ with $\operatorname{surp}_{G_{5}}(I)\leq 1-d\leq-2$; necessarily $|I|\geq 2$.
-Then $\operatorname{surp}_{G-\{v_{1},v_{2}\}-N[v_{5}]}(I\cup\{v_{6}\})\leq(1-d)+(d-1)+1=2$.
+Then $\operatorname{surp}_{G-\{v_{1},v_{2}\}-N[v_{5}]}(I\cup\{v_{6}\})\leq(1-d)+(d-1)+2=2$.
 So $G-\{v_{1},v_{2}\}-N[v_{5}]$ is 3-vul, putting us into Case Ia with alternate vertex ordering $v_{1}^{\prime}=v_{1},v_{2}^{\prime}=v_{2},v_{3}^{\prime}=v_{5}$.
@@
 Split on $v_{6},v_{1},v_{5}$, obtaining subproblems $G_{1}=G-N[v_{6}],G_{2}=G-v_{6}-N[v_{1}],G_{3}=G-\{v_{1},v_{6}\}-N[v_{5}],G_{4}=G-\{v_{1},v_{6},v_{5}\}$.
-By Observation 61, we have $\operatorname{minsurp}(G_{1})\geq 2,\operatorname{minsurp}(G_{2})\geq 1,\operatorname{minsurp}(G_{3})\geq\operatorname{shad}(v_{6},N[v_{5}])-1\geq 0,\operatorname{minsurp}(G_{4})\geq\operatorname{shad}(v_{1},v_{6})-1\geq 1$.
+By Observation 62, we have $\operatorname{minsurp}(G_{1})\geq 2,\operatorname{minsurp}(G_{2})\geq 1,\operatorname{minsurp}(G_{3})\geq\operatorname{shad}(v_{6},N[v_{5}])-1\geq 0,\operatorname{minsurp}(G_{4})\geq\operatorname{shad}(v_{1},v_{6})-1\geq 1$.
 So $G_{1},G_{2},G_{3},G_{4}$ have drops $(2.5,6),(3,7),(3,7),(1.5,3)$ directly.
@@
 $e^{-2.5a-6b}+e^{-3a-7b}\gamma_{3}+e^{-3a-7b}+e^{-1.5a-4b}\leq 1$ (6-12)
-Case IIb: Split on $v_{6},v_{4},v_{3}$ and apply (A2) to vertex $v_{1}$ in subproblem $G-\{v_{6},v_{4},v_{3}\}$.
+Case IIb: Split on $v_{6},v_{4},v_{3}$ and apply (B-Fun2) to vertex $v_{1}$ in subproblem $G-\{v_{6},v_{4},v_{3}\}$.
 Overall, this generates subproblems $G_{1}=G-N[v_{6}],G_{2}=G-v_{6}-N[v_{4}],G_{3}=G-v_{6}-N[v_{3}],G_{4}=G-\{v_{3},v_{4},v_{6},v_{1}\}$ and $G_{5}=G-\{v_{3},v_{4},v_{6}\}-N[v_{1},v_{5}]$.
-By Observation 61, we have $\operatorname{minsurp}(G_{1})\geq 2$ and $\operatorname{minsurp}(G_{2})\geq 1$ and $\operatorname{minsurp}(G_{3})\geq 1$ and $\operatorname{minsurp}(G_{4})\geq\operatorname{shad}(v_{4},v_{6})-2\geq 0$.
+By Observation 62, we have $\operatorname{minsurp}(G_{1})\geq 2$ and $\operatorname{minsurp}(G_{2})\geq 1$ and $\operatorname{minsurp}(G_{3})\geq 1$ and $\operatorname{minsurp}(G_{4})\geq\operatorname{shad}(v_{4},v_{6})-2\geq 0$.
 So the subproblems have drops $(2.5,6),(3,7),(3,7),(2,4)$ directly; subproblem $G_{4}$ has a 2-vertex $u$, so it has drop $(2,5)$ after simplification.
 For subproblem $G_{5}$, let $d=\deg_{G-v_{6}-N[v_{1}]}(v_{5})$.
-We have $d\geq 3$, as otherwise $R(v_{6},N[v_{1}])\geq 1$ and it would be covered in Case IIa.
+We have $d\geq 3$, as otherwise $S(v_{6},N[v_{1}])\geq 1$ and it would be covered in Case IIa.
 We must have $\operatorname{minsurp}(G_{5})\geq 2-d$; for, if we have an indset $I$ with $\operatorname{surp}_{G_{5}}(I)\leq 1-d\leq-2$, where necessarily $|I|\geq 2$, then $\operatorname{surp}_{G-v_{6}-N[v_{1}]}(I\cup\{v_{5}\})\leq 2$ and $G-v_{6}-N[v_{1}]$ is 3-vul; it would be covered in Case IIa.
@@
 Split on $v_{1},v_{2},v_{3}$, obtaining subproblems $G_{1}=G-N[v_{1}],G_{2}=G-v_{1}-N[v_{2}],G_{3}=G-\{v_{1},v_{2}\}-N[v_{3}],G_{4}=G-\{v_{1},v_{2},v_{3}\}$.
-By Observation 61, we have $\operatorname{minsurp}(G_{1})\geq 2$ and $\operatorname{minsurp}(G_{2})\geq 1$ and $\operatorname{minsurp}(G_{3})\geq\operatorname{shad}(v_{1},N[v_{3}])-1\geq 0$ and $\operatorname{minsurp}(G_{4})\geq\operatorname{shad}(v_{1},v_{2})-1\geq 1$.
+By Observation 62, we have $\operatorname{minsurp}(G_{1})\geq 2$ and $\operatorname{minsurp}(G_{2})\geq 1$ and $\operatorname{minsurp}(G_{3})\geq\operatorname{shad}(v_{1},N[v_{3}])-1\geq 0$ and $\operatorname{minsurp}(G_{4})\geq\operatorname{shad}(v_{1},v_{2})-1\geq 1$.
 So the subproblems have drops $(2.5,6),(2.5,6),(3,7),(1.5,3)$ directly; furthermore, $G_{4}$ has a kite $u,v_{4},v_{5},v_{6}$ and so it has net drop $(2,5)$ by Lemma 14.
@@
 8.4 Line 8: Branching on a 5-vertex with two 6-neighbors
-In the algorithm Branch6-2, we consider a 5-vertex $u$ with 6-neighbors $y_{1},y_{2}$ and 5-neighbors $v_{1},v_{2},v_{3}$, where either $y_{1}$ or $y_{2}$ has at most three 5-neighbors.
-See Figure 9 for an illustration.
+In the algorithm Branch6-2, we consider a 5-vertex $u$ with 6-neighbors $y_{1},y_{2}$ and 5-neighbors $v_{1},v_{2},v_{3}$, where either $y_{1}$ or $y_{2}$ has at most three 5-neighbors. See Figure 12.
 Figure 12: The neighborhood of vertex $u$. The edge $(y_{1},y_{2})$ may or may not be present.
@@
 Proof.
-  • Vertices $v_{1},v_{2},v_{3}$ are independent, as otherwise $R(y_{1},y_{2})\geq 1$ (due to 3-triangle $u$) and we would branch at Line 3.
+  • Vertices $v_{1},v_{2},v_{3}$ are independent, as otherwise $S(y_{1},y_{2})\geq 1$ (due to 3-triangle $u$) and we would branch at Line 3.
 There must be some pair $y_{i},v_{j}$ with $y_{i}\not\sim v_{j}$ (else both $y_{1},y_{2}$ would have four 5-neighbors $u,v_{1},v_{2},v_{3}$); say without loss of generality that $y_{1}\not\sim v_{1}$.
-We have $y_{2}\not\sim\{v_{1},v_{2},v_{3}\}$, else $R(y_{1},v_{1})\geq 1$ due to the 3-triangle involving $u,y_{2},v_{j}$, and we would apply Line 3 to $y_{1},v_{1}$.
-This in turn implies that $y_{1}\not\sim\{v_{2},v_{3}\}$, else $R(y_{2},v_{1})\geq 1$ due to the resulting 3-triangle and we would apply Line 3 to $y_{2},v_{1}$.
+We have $y_{2}\not\sim\{v_{1},v_{2},v_{3}\}$, else $S(y_{1},v_{1})\geq 1$ due to the 3-triangle involving $u,y_{2},v_{j}$, and we would apply Line 3 to $y_{1},v_{1}$.
+This in turn implies that $y_{1}\not\sim\{v_{2},v_{3}\}$, else $S(y_{2},v_{1})\geq 1$ due to the resulting 3-triangle and we would apply Line 3 to $y_{2},v_{1}$.
   • Say without loss of generality that $y_{1},v_{1}$ share a vertex $x\neq u$.
 Since none of the neighbors of $u$ are adjacent to $v_{1}$, we have $x\not\sim u$.
-If $\deg(x)\leq 4$, then $R(y_{1},v_{1})\geq 1$ (due to subquadratic vertex $x$), and we would branch on $y_{1},v_{1}$ at Line 3.
-If $\deg(x)=6$ and $x\sim y_{2}$, then $R(N[x])\geq 1$ (due to vertex $u$ losing neighbors $y_{1},v_{1},y_{2}$) and we would branch at Line 2.
-If $\deg(x)=6$ and $x\not\sim y_{2}$, then $R(y_{2},N[x])\geq 1$ (again due to vertex $u$ losing neighbors $y_{1},v_{1},y_{2}$), and we would branch on $y_{2},x$ at Line 7.
+If $\deg(x)\leq 4$, then $S(y_{1},v_{1})\geq 1$ (due to subquadratic vertex $x$), and we would branch on $y_{1},v_{1}$ at Line 3.
+If $\deg(x)=6$ and $x\sim y_{2}$, then $S(N[x])\geq 1$ (due to vertex $u$ losing neighbors $y_{1},v_{1},y_{2}$) and we would branch at Line 2.
+If $\deg(x)=6$ and $x\not\sim y_{2}$, then $S(y_{2},N[x])\geq 1$ (again due to vertex $u$ losing neighbors $y_{1},v_{1},y_{2}$), and we would branch on $y_{2},x$ at Line 7.
   • Suppose $y_{i}$ shares neighbor $t_{ij}\neq u$ with $v_{j}$ for $j=1,2,3$; as we have just seen, $\deg(t_{ij})=5$ and $t_{ij}\not\sim u$ for all $j$.
-Vertices $t_{i1},t_{i2},t_{i3}$ must be distinct; for, if $t_{ij}=t_{ij^{\prime}}=t$, then $\operatorname{codeg}(t,u)\geq 3$ (sharing vertices $v_{j},v_{j^{\prime}},y_{i}$), contradicting Observation 61.
+Vertices $t_{i1},t_{i2},t_{i3}$ must be distinct; for, if $t_{ij}=t_{ij^{\prime}}=t$, then $\operatorname{codeg}(t,u)\geq 3$ (sharing vertices $v_{j},v_{j^{\prime}},y_{i}$), contradicting Observation 62.
 So $y_{i}$ has four distinct 5-neighbors $t_{i1},t_{i2},t_{i3},u$.
@@
 For the graph $G^{\prime}_{ij}=G-\{v_{1},v_{2},v_{3}\}-N[y_{i},v_{j}]$, we have $\operatorname{minsurp}(G^{\prime}_{ij})\geq-3+\operatorname{codeg}(y_{i},v_{j})$.
-Moreover, if $\operatorname{codeg}(y_{i},v_{j})=1$ and $\operatorname{minsurp}(G^{\prime}_{ij})=-2$, then either $S(G^{\prime}_{ij})\geq 1$ or $R(v_{j^{\prime}},N[v_{j}])\geq 1$ for some $j^{\prime}\in\{1,2,3\}$.
+Moreover, if $\operatorname{codeg}(y_{i},v_{j})=1$ and $\operatorname{minsurp}(G^{\prime}_{ij})=-2$, then either $S(G^{\prime}_{ij})\geq 1$ or $S(v_{j^{\prime}},N[v_{j}])\geq 1$ for some $j^{\prime}\in\{1,2,3\}$.
 Proof.
@@
 If $|I|\geq 3$ then $S(G^{\prime})\geq 1$ as desired, so suppose $|I|=2$.
-By Observation 61, we have $\operatorname{codeg}(v_{2},v_{3})\leq 2$, so there must be some vertex $x\in I$ and $v_{j}\in\{v_{2},v_{3}\}$ with $x\not\sim v_{j}$; say without loss of generality $x\not\sim v_{3}$ and $x$ is isolated in $G^{\prime}$.
+By Observation 62, we have $\operatorname{codeg}(v_{2},v_{3})\leq 2$, so there must be some vertex $x\in I$ and $v_{j}\in\{v_{2},v_{3}\}$ with $x\not\sim v_{j}$; say without loss of generality $x\not\sim v_{3}$ and $x$ is isolated in $G^{\prime}$.
 Let $d=\deg_{G}(x)$.
-By Observation 61 we have $\operatorname{codeg}(x,y_{1})\leq d-3$ and $\operatorname{codeg}(x,v_{1})\leq 2$.
-The only way for $x$ to be isolated in $G^{\prime}$ is to have $x\sim v_{2},\operatorname{codeg}(x,v_{1})=2,\operatorname{codeg}(x,y_{1})=d-3$.
-If $d\leq 5$, we thus have $R(v_{2},N[v_{1}])\geq 1$ due to subquadratic vertex $x$.
-If $d=6$, then by Proposition 66 we have $x\not\sim y_{2}$ (else the 6-vertex $x$ would be shared by $y_{2},v_{2}$); in this case, $G-N[x]$ has a 3-vertex $y_{1}$ with a triangle $y_{1},y_{2},u$, and we would branch on $x$ at Line 2.
+By Observation 62 we have $\operatorname{codeg}(x,y_{1})\leq d-3$ and $\operatorname{codeg}(x,v_{1})\leq 2$.
+So the only way for $x$ to be isolated in $G^{\prime}$ is to have $x\sim v_{2},\operatorname{codeg}(x,y_{1})=d-3,\operatorname{codeg}(x,v_{1})=2$.
+If $d\leq 5$, we thus have $S(v_{2},N[v_{1}])\geq 1$ due to subquadratic vertex $x$.
+If $d=6$, then by Proposition 66 we have $x\not\sim y_{2}$ (else the 6-vertex $x$ would be shared by $y_{2},v_{2}$); in this case, GN[x]G-N[x] has a 3-vertex $y_{1}$ with a triangle $y_{1},y_{2},u$, and we would branch on $x$ at Line 2.
 ∎
@@
 Split on $y_{1},y_{2},v_{1}$, getting subproblems $G_{1}=G-N[y_{1}],G_{2}=G-y_{1}-N[y_{2}],G_{3}=G-\{y_{1},y_{2}\}-N[v_{1}],G_{4}=G-\{y_{1},y_{2},v_{1}\}$.
-By Observation 61, we have $\operatorname{minsurp}(G_{1})\geq 2$ and $\operatorname{minsurp}(G_{2})\geq 1$ and $\operatorname{minsurp}(G_{3})\geq\operatorname{shad}(y_{1},v_{1})-1\geq 1$ and $\operatorname{minsurp}(G_{4})\geq\operatorname{shad}(y_{1},y_{2})-1\geq 1$.
+By Observation 62, we have $\operatorname{minsurp}(G_{1})\geq 2$ and $\operatorname{minsurp}(G_{2})\geq 1$ and $\operatorname{minsurp}(G_{3})\geq\operatorname{shad}(y_{1},v_{1})-1\geq 1$ and $\operatorname{minsurp}(G_{4})\geq\operatorname{shad}(y_{1},y_{2})-1\geq 1$.
 So $G_{1},G_{2},G_{3},G_{4}$ have drops $(2.5,6),(3,7),(3,7),(1.5,3)$ directly.
-We also apply Proposition 30 to 2-vertex $u$ in $G_{4}$, where $\deg(v_{2})+\deg(v_{3})-1-\operatorname{codeg}(v_{2},v_{3})\geq 7$ by Observation 61.
+We also apply Proposition 30 to 2-vertex $u$ in $G_{4}$, where $\deg(v_{2})+\deg(v_{3})-1-\operatorname{codeg}(v_{2},v_{3})\geq 7$ by Observation 62.
 So we need:
@@
 Case II: $\boldsymbol{y_{1}\sim y_{2}}$.
-We choose some enumeration $v_{1},v_{2},v_{3}$, and split on $v_{1},v_{2}$; in the subproblem $G-\{v_{1},v_{2}\}$, we apply (A2) to $u,y_{1}$.
+We choose some enumeration $v_{1},v_{2},v_{3}$, and split on $v_{1},v_{2}$; in the subproblem $G-\{v_{1},v_{2}\}$, we apply (B-Fun2) to $u,y_{1}$.
 This generates subproblems $G_{1}=G-N[v_{1}],G_{2}=G-v_{1}-N[v_{2}],G_{3}=G\{v_{1},v_{2}\}-N[v_{3},y_{1}],G_{4}=G-\{v_{1},v_{2},y_{1}\}$.
-By Observation 61, $G_{1},G_{2},G_{4}$ have drops $(2,5),(2.5,6),(1.5,3)$ directly; furthermore, we can apply Proposition 30 to the 2-vertex $u$ in $G_{4}$, where $r=\deg(v_{1})+(\deg(y_{2})-1)-\operatorname{codeg}(y_{2},v_{1})-1\geq 5+5-2-1=7$ by Observation 61.
-Proposition 66 and Observation 61 give $\operatorname{minsurp}(G_{3})\geq-3+\operatorname{codeg}(y_{1},v_{3})$ and $\operatorname{codeg}(y_{1},v_{3})\leq 2$.
+By Observation 62, $G_{1},G_{2},G_{4}$ have drops $(2,5),(2.5,6),(1.5,3)$ directly; furthermore, we can apply Proposition 30 to the 2-vertex $u$ in $G_{4}$, where $r=\deg(v_{1})+(\deg(y_{2})-1)-\operatorname{codeg}(y_{2},v_{1})-1\geq 5+5-2-1=7$ by Observation 62.
+Proposition 67 and Observation 62 give $\operatorname{minsurp}(G_{3})\geq-3+\operatorname{codeg}(y_{1},v_{3})$ and $\operatorname{codeg}(y_{1},v_{3})\leq 2$.
 In light of Proposition 66, there are two cases to select the ordering:
-Case IIa: $\boldsymbol{R(v_{1},N[v_{2}])\geq 1.}$ In this case $G_{2}$ has drop $(2.5,7)$ after simplification. Since $G_{3}$ has drop $(4,11)$, we need:
+Case IIa: $\boldsymbol{S(v_{1},N[v_{2}])\geq 1.}$ In this case $G_{2}$ has drop $(2.5,7)$ after simplification. Since $G_{3}$ has drop $(4,11)$, we need:
 $e^{-2a-5b}+e^{-2.5a-7b}+e^{-4a-11b}+e^{-1.5a-3b}\psi_{7}\leq 1$ (6-16)
-Case IIb: $\boldsymbol{\operatorname{codeg}(y_{1},v_{3})=1.}$ By Proposition 66, either $R(v_{j},N[v_{3}])\geq 1$ for some $j$ (which would put us into Case IIa, with some other vertex ordering), or $S(G_{3})\geq 1$, or $\operatorname{minsurp}(G_{3})\geq-1$.
+Case IIb: $\boldsymbol{\operatorname{codeg}(y_{1},v_{3})=1.}$ By Proposition 67, either $S(v_{j},N[v_{3}])\geq 1$ for some $j$ (which would put us into Case IIa, with some other vertex ordering), or $S(G_{3})\geq 1$, or $\operatorname{minsurp}(G_{3})\geq-1$.
 In the latter two cases, $G_{3}$ has drop $(4,13)$ or $(4.5,12)$.
@@
 We first describe the degree-7 branching algorithm.
-Simplify $G$ Branch on a vertex $u$ with degree larger than 7 Split on a 7-vertex $u$ with $\operatorname{shad}(N[u])\geq 0$.
-Branch on a 7-vertex $u$ which has a blocker $x$ with $\deg(x)\geq 5$.
-Branch on a vertex $x$ which is a blocker of three 7-vertices $u_{1},u_{2},u_{3}$.
-Run either Branch6-3 or the MaxIS-7 algorithm.
-Algorithm 10 Function Branch7$(G,k)$
+1Simplify $G$
+2Branch on a vertex $u$ with degree larger than 7
+3Split on a 7-vertex $u$ with $\operatorname{shad}(N[u])\geq 0$.
+4Branch on a 7-vertex $u$ which has a blocker $x$ with $\deg(x)\geq 5$.
+5Branch on a vertex $x$ which is a blocker of three 7-vertices $u_{1},u_{2},u_{3}$.
+6Run either Branch6-3 or the MaxIS-7 algorithm.
+Algorithm 10 Function Branch7(G,k)(G,k)
 Theorem 68.
@@
 To show this, let us examine the algorithm line by line.
-Line 2. Here, we apply Section 5; we need:
+Line 2. Here, we apply Figure 6; we need:
 $\max\{e^{-a-3b}+e^{-a-4b},e^{-0.5a-2b}+e^{-2a-5b},e^{-0.5a-b}+e^{-2.5a-8b}\}\leq 1$ (7-1)
-Line 3. In this case $G-N[u]$ has drop $(3,7)$. So we need:
+Line 3. In this case GN[u]G-N[u] has drop $(3,7)$. So we need:
 $e^{-0.5a-b}+e^{-3a-7b}\leq 1$ (7-2)
-Line 4. Apply (B) to $u,x$, getting subproblems $G-N[x],G-\{u,x\}$ with drops $(1,5)$ and $(1,2)$ respectively. We need:
+Line 4. Apply (B-Block) to $u,x$, getting subproblems $G-N[x],G-\{u,x\}$ with drops $(1,5)$ and $(1,2)$ respectively. We need:
 $e^{-a-2b}+e^{-a-5b}\leq 1$ (7-3)
 Line 5.
-Apply (B) to vertices $u_{1},u_{2},u_{3},x$.
-The resulting subproblems $G-\{u_{1},u_{2},u_{3},x\}$ and $G-N[x]$ have drops $(1,4),(1,3)$ directly; the branch-seq $[(1,3),(1,4)]$ is already covered in (7-1).
+Apply (B-Block) to $u_{1},u_{2},u_{3},x$.
+The resulting subproblems $G-\{u_{1},u_{2},u_{3},x\}$ and GN[x]G-N[x] have drops $(1,4),(1,3)$ directly; the branch-seq $[(1,3),(1,4)]$ is already covered in (7-1).
 Line 6.
@@
 To finish up, we use a simple degree-8 branching algorithm.
-If $\operatorname{maxdeg}(G)\leq 7$, then use Lemma 10 with the algorithms Branch7 or MaxIS-7. Otherwise, split on a vertex of degree at least $8$.
-Algorithm 11 Function Branch8$(G,k)$
+1If $\operatorname{maxdeg}(G)\leq 7$, then use Lemma 15 with the algorithms Branch7 or MaxIS-7.
+Otherwise, split on a vertex of degree at least $8$.
+Algorithm 11 Function Branch8(G,k)(G,k)
 Theorem 69.
@@
 Proof.
-Splitting on a vertex of degree $r\geq 8$ generates subproblems with vertex covers of size at most $k-1$ and $k-8$.
+Splitting on a vertex of degree $r\geq 8$ generates subproblems with vertex covers of size at most k1k-1 and k8k-8.
 These have cost $1.25284^{k-1}$ and $1.25284^{k-8}$ by induction.
-Since $11.25284^{-1}+1.25284^{-8}<$ 1, the overall cost is $O^{*}(1.25284^{k})$.
+Since $1.25284^{-1}+1.25284^{-8}<1$, the overall cost is $O^{*}(1.25284^{k})$.
 Otherwise, Lemma 15 (with parameters $a_{7},b_{7}$ and $c=\log(1.19698)$) gives the desired algorithm with runtime $O^{*}(1.25284^{k})$.
@@
 Appendix A Proofs for basic results
-Proof of Proposition 4.
-Consider any basic solution $\theta\in\{0,\tfrac{1}{2},1\}^{V}$ to $\operatorname{LPVC}(G)$.
+Proof of Proposition 4 and Proposition 5.
+Consider the linear program defined by $\operatorname{LPVC}(G)$ and the additional constraints $\theta(v)=0:v\in I$ and $\theta(v)\geq 1/2:v\notin X$, and let its optimum value be denoted by $\lambda_{I,X}$.
+We claim that any indset $J$ of minimum surplus with $I\subseteq J\subseteq X$ corresponds to an optimal basic solution to this linear program, and has $\lambda_{I,X}=\tfrac{1}{2}(n+\operatorname{surp}_{G}(J))$.
+This will show Proposition 5 (since the linear program can be solved in polynomial time), and it will also show Proposition 4 (since $\lambda(G)=\lambda_{\emptyset,V}=\operatorname{minsurp}^{0}(G)$).
+First, we claim that any such basic solution $\theta\in[0,1]^{V}$ is half-integral.
+For, suppose not, and let $A,B$ denote the vertices $v$ with $\theta(v)\in(0,1/2)$ and $\theta(v)\in(1/2,1)$ respectively.
+It can be seen that modifying $\theta(v):v\in A$ by $\pm\epsilon$ and $\theta(v):v\in B$ by $\mp\epsilon$ preserves feasibility.
+Thus, in a basic solution, we must have $A=B=\emptyset$.
+Now consider any such basic solution $\theta\in\{0,\tfrac{1}{2},1\}^{V}$.
 Let $I_{0}=\theta^{-1}(0)$ and $I_{1}=\theta^{-1}(1)$; note that $I_{0}$ is an indset (possibly empty) and $N(I_{0})\subseteq I_{1}$.
-So $\sum_{v}\theta(v)=|I_{1}|+\tfrac{1}{2}(n-|I_{0}|-|I_{1}|)\geq(n+|N(I_{0})|-|I_{0}|)/2=(n+\operatorname{surp}(I_{0}))/2\geq(n+\operatorname{minsurp}^{-}(G))/2$.
-On the other hand, if $I_{0}$ is a min-set of $G$, then setting $\theta(v)=1$ for $v\in N(I_{0})$ and $\theta(v)=0$ for $v\in I_{0}$ and $\theta(v)=1/2$ for all other vertices $v$ gives a fractional vertex cover with $\sum_{v}\theta(v)=(n+\operatorname{minsurp}^{-}(G))/2$.
+So $\sum_{v}\theta(v)=|I_{1}|+\tfrac{1}{2}(n-|I_{0}|-|I_{1}|)\leq(n+|N(I_{0})|-|I_{0}|)/2$.
+On the other hand, for any min-set $J$ with $I\subseteq J\subseteq X$, then setting $\theta(v)=1$ for $v\in N(J)$ and $\theta(v)=0$ for $v\in J$ and $\theta(v)=1/2$ for all other vertices $v$ gives a fractional solution cover with $\sum_{v}\theta(v)=(n+\operatorname{surp}_{G}(J))/2$.
 ∎
@@
 For, suppose that $I\cap C=J$ where $\emptyset\neq J\subsetneq I$.
-In this case, we must also have $N(I\setminus J)\subseteq C$, so overall $|C\cap N[I]|\geq|J|+|N(I\setminus J)|=|J|+|I\setminus J|+\operatorname{surp}(I\setminus J)$; since $I\setminus J$ is a non-empty subset of $I$, we have $\operatorname{surp}(I\setminus J)\geq\operatorname{surp}(I)$, hence this is at least $|J|+|I\setminus J|+\operatorname{surp}(I)=|I|+\operatorname{surp}(I)\geq N(I)$.
+In this case, we must also have $N(I\setminus J)\subseteq C$, so overall $|C\cap N[I]|\geq|J|+|N(I\setminus J)|=|J|+|I\setminus J|+\operatorname{surp}(I\setminus J)$; since $I\setminus J$ is a non-empty subset of II, we have $\operatorname{surp}(I\setminus J)\geq\operatorname{surp}(I)$, hence this is at least $|J|+|I\setminus J|+\operatorname{surp}(I)=|I|+\operatorname{surp}(I)\geq N(I)$.
 Thus, replacing $C\cap N[I]$ by $N(I)$ would give a vertex cover $C^{\prime}$ with $C^{\prime}\cap I=\emptyset$ and $|C^{\prime}|\leq|C|$.
@@
 If $C\cap J\neq\emptyset$, this would be a good cover with $I\cap C=\emptyset$.
-∎
-Proof of Proposition 6.
-Let $J$ be an indset with $\operatorname{surp}_{G-N[I]}(J)=\operatorname{minsurp}^{-}(G-N[I])$.
-Then $J\cup I$ is a non-empty indset of $G$ with surplus $\operatorname{surp}_{G-N[I]}(J)+\operatorname{surp}_{G}(I)=\operatorname{minsurp}^{-}(G-N[I])+\operatorname{surp}_{G}(I)$; in particular, if $\operatorname{minsurp}^{-}(G-N[I])+\operatorname{surp}_{G}(I)=\operatorname{minsurp}(G)$, then $I$ is a min-set of $G$.
-On ther other hand, if $I$ is contained in a min-set $J$ of $G$, then $\operatorname{surp}_{G-N[I]}(J\setminus I)=\operatorname{surp}_{G}(J)+|I|-|N(I)|=\operatorname{surp}_{G}(J)-\operatorname{surp}_{G}(I)$.
-So $\operatorname{minsurp}^{-}(G-N[I])\leq\operatorname{minsurp}(G)-\operatorname{surp}_{G}(I)$.
-∎
-Proof of Proposition 7.
-By applying Proposition 6 to indsets $I=\{x\}$, we get $\operatorname{minsurp}(G)=\min_{x\in V}\bigl{(}\operatorname{minsurp}^{-}(G-N[x])+\deg(x)-1\bigr{)}$.
-We can use Proposition 4 to compute each value $\operatorname{minsurp}^{-}(G-N[x])$ in polynomial time (by solving the LP).
-This gives an algorithm to compute $\operatorname{minsurp}(G)$.
-Likewise, given any vertex-set $X$, we can use Proposition 6 to determine if $X$ is contained in a min-set of $G$ and, if so, we can use Proposition 4 to find a vertex set $J$ attaining $\operatorname{surp}_{G-N[X]}(J)=\operatorname{minsurp}^{-}(G-N[X])$; then $J\cup X$ is a min-set of $G$ containing $X$.
 ∎
@@
 ∎
-Proof of Lemma 9.
-Consider a kite $u,x,y,z$.
-We can view the application of (P3) to funnel $u$ as a two-part process.
-First, we remove the shared neighbor $y\in N(u)\cap N(x)$; then, we merge vertices $z,x$ into a new vertex $u^{\prime}$.
-In the first step, we obtain a graph $G^{\prime\prime}=G-y$ with $k^{\prime\prime}=k-1$ and $n^{\prime\prime}=n-1$.
-So, by Proposition 4, we have $\lambda(G^{\prime\prime})=\lambda(G)-1/2$ and hence $\mu(G^{\prime\prime})\leq\mu(G)-1/2$ and $k^{\prime\prime}=k-1$.
-By Proposition 8, the second step gives $\mu(G^{\prime})\leq\mu(G^{\prime\prime})$ and $k^{\prime}=k^{\prime\prime}-1$.
+Proof of Observation 9.
+Here $\operatorname{minsurp}(G)\geq 2$.
+If $J$ is a min-set of GXG-X, then $2\leq\operatorname{surp}_{G}(J)\leq\operatorname{surp}_{G-X}(J)+|X|$.
+So $\operatorname{surp}_{G-X}(J)\geq 2-|X|$.
+Similarly, if $J$ is a min-set of GN[I]G-N[I], then $2\leq\operatorname{surp}_{G}(I\cup J)=\operatorname{surp}_{G-N[I]}(J)+\operatorname{surp}_{G}(I)$.
 ∎
-Proof of Lemma 10.
-First, exhaustively apply rule (P1) to $G$; the resulting graph $G^{\prime}$ has $n^{\prime}\leq n$ and $k^{\prime}\leq k$, and $\operatorname{minsurp}(G^{\prime})\geq 0$.
-In particular, $\mu(G^{\prime})=k^{\prime}-\lambda(G^{\prime})=k^{\prime}-n^{\prime}/2$.
-Note that $G^{\prime}\in\mathcal{G}$ since (P1) just deletes vertices.
-Now dovetail the two algorithms for $G^{\prime}$, running both simultaneously and returning the output of whichever terminates first.
-This has runtime $O^{*}(\min\{e^{cn^{\prime}},e^{a(k^{\prime}-n^{\prime}/2)+bk^{\prime}}\})$.
-For fixed $k^{\prime}$, this is maximized at $n^{\prime}=\frac{2k^{\prime}(a+b)}{a+2c}$, at which point it takes on value $e^{dk^{\prime}}\leq e^{dk}$.
+Proof of Proposition 10.
+If $\deg(x)=\deg(y)=1$ and $\operatorname{codeg}(x,y)=0$, then $\{x,y\}$ is a min-set with $\operatorname{surp}(\{x,y\})=0$, and applying P1 reduces $k$ by two.
+Otherwise, if $\deg(x)\geq 2$, then we can apply (P2) to $x$, and vertex $y$ remains subquadratic, and we can follow up by applying (P1) or (P2) again to $y$.
+Note that if $\operatorname{minsurp}(G)\geq 0$, then at least one of the conditions (i) or (ii) must hold.
 ∎
+Proof of Proposition 11.
+We show it by induction on $\ell$.
+The base case $\ell=0$ is vacuous.
+For the induction step, we apply (P2) to $x_{\ell}$, forming a graph $G^{\prime}$ where the neighbors $y,z$ get contracted into a single new vertex $t$.
+By hypothesis, $y,z$ are distinct from $x_{1},\dots,x_{\ell}$.
+We claim that $\operatorname{dist}_{G^{\prime}}(x_{i},x_{j})\geq 3$ and $\deg_{G^{\prime}}(x_{i})\geq 2$ for any pair $i<j<\ell$.
+For, clearly $x_{i}\not\sim x_{j}$ in $G^{\prime}$.
+If $x_{i},x_{j}$ share a neighbor in $G^{\prime}$, it must be vertex $t$ as no other vertices were modified.
+This is only possible if $\{x_{i},x_{j}\}\sim\{y,z\}$; but this contradicts our hypothesis that $\operatorname{codeg}(x_{i},x_{\ell})=\operatorname{codeg}(x_{j},x_{\ell})=0$.
+Finally, suppose that $\deg_{G^{\prime}}(x_{i})\leq 1$.
+This is only possible if the two neighbors of $x_{i}$ got merged together, i.e.
+$x_{i}\sim y$ and $x_{i}\sim z$.
+Again, this contradicts that $\operatorname{codeg}(x_{i},x_{\ell})=0$.
+By induction hypothesis applied to $G^{\prime}$, we have $S(G^{\prime})\geq\ell-1$ and hence $S(G)\geq\ell$. ∎